A Complete Bibliography of Publications in Communications in Statistics: Simulation

and Computation: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

18 May 2024 Version 1.06

Title word cross-reference

(k, l) [RSG21]. 1 [EHF23]. 2 [EHF23, NN20, OP20]. 2 × 2 [CH23]. 3 [EHF23, HCHK23, OP20]. [0, ∞) [Som22]. ^{1} [GZZ22]. _{2} [LWW22]. _{2.5} [GZ21b]. _{RS} [CKCG21]. A [MPD20]. α [GDWY22, Moh22, MEA24, ZKW24]. X̄ [Hua22, MMMG21, MMMCS22b, SNKA24, CKT^{+}20, HCTZ21, NKSY22]. β [GWW^{+}23]. C_{p} [AA23b]. C_{pk} [PDO^{+}20]. C_{pm} [SCK23, TdSCdSdCP22]. D [GKJ23, LFMMRH22, PMM23, UIC23, GBHA20, GBHA21]. ℓ_{1} [EJKL23]. ℓ_{p} [BDK^{+}23]. ℓ_{q} [BDK^{+}23]. ϵ [DP22]. F [Lue23, SX22]. G [AAHR20, YW24]. GM(1, N) [CLL23, CLLX22]. I [CLJK21]. I^{2} [HBM^{+}20]. INARMA(p, q) [BA22b]. INGARCH(1, 1) [MB23d]. K′ [Wan23c, LW21, TF22b, ARA23, AALR23, FY21, GPP21, GK23, MAMA20, SI23a, SR23, TKR24, Wan23c]. K^{−} [ZYSS22]. L [Hut21, SAA^{+}23]. L^{p}(p > 1) [Xia22]. L_{1} [GPGMK22, EJKL22]. L_{2} [Zhe21, Zhe24]. L_{p} [SD22]. M [MA23a, Pol20, RYA20, SCK21, SCA23, GBHA20, KS23b, TKR24]. n [BQCP23, TKR24, ZYSS22]. P [UY23, Cav24, CL22b, DR23, DL20, ZMNL23]. ϕ [DCY22, DC24]. PINAR(1) [BS22a]. Q [dSLHMB23]. r [SR23, TKR24]. ρ [Bar20a]. S_{B} [MM20]. S_{n} [KA22, SS23]. sin^{k}(x) [MS22]. SINAR(1, 1) [SP23]. t [ACKL23, BA22a, DS24, HEK22, KVJH23, MKSH22, NIN20, Pan20c, RBN^{+}22, SKM^{+}23, XS23]. T^{2} [AMK^{+}21a, AF22b, ENA^{+}23, KM22, MLR^{+}22, SSBA22]. U [GK20]. φ [LY20]. W [NCC^{+}22, SMM22]. X [PM20].

-class [FY21]. -complexity [DP22]. -control [PM20]. -copulas [GBHA21]. -designs [KVJH23]. -distributed [TF22b]. -distribution [XS23]. -divergence [MEA24]. -estimation [DA21]. -estimator [MA23a, RYA20, SCK21, SCA23]. -estimators [Hut21]. -expectation [GWW^{+}23]. -formula [YW24]. -inflated [RSG21]. -Loss [Zhe21, Zhe24]. -matrices [dSLHMB23]. -means [Wan23c, LW21]. -mixing [DCY22, DC24, GDWY22, LY20]. -modulated [DL20]. -moments [SAA^{+}23]. -multivariate [SI23a]. -of- [SMM22]. -optimal [GKJ23, LFMMRH22, MPD20, PMM23, UIC23]. -out-of [TKR24]. -out-of- [TKR24, ZYSS22]. -penalized [GPGMK22]. -player [HCHK23]. -regression [Pol20]. -sample [GBHA20]. -Score [DS24]. -spline [CLJK21]. -stable [Moh22, ZKW24]. -stage [NN20]. -statistics [GK20]. -test [Lue23]. -transformed [Pan20c]. -value [CL22b, ZMNL23]. -values [Cav24, ZMNL23, UY23]. -Variation [DR23]. -within- [TKR24].

//doi.org/10.1080/03610918.2020.1812649 [Kri22].

1 [BC21, JD23b]. 19 [SM23b].

3117 [Ano24].

4398 [AE23c]. 46 [AE23c]. 48 [Ano24].

AANA [LZSC23]. AB/BA [ZL20]. abrupt [MMMG21]. absence [EH22, GABJ23]. Absolute [ZXW23]. absolutely [RK20]. accelerated [ABPK21, Bar20a, CSL20, Çet23, GAS20, HB22a, LY24, NDWE24, XZS22, ZLSC20]. accelerating [SU23a]. Acceptance [Ger20, BNA23, DR22, uASS22, SDG22]. acceptance-rejection [SDG22]. accordance [LX23]. Accounting [ZL22a]. accuracy [AJIAA22, AJI24, AT21, CL22b, Kol24, LWX23, PY20, WDZ22]. Accurate [BRTW24]. across [De 20]. Active [WKcIC24, LZ24]. acyclic [HPZ^{+}21]. adaLASSO [EGVdR21]. adapted [AF22b, dSdSTC21]. Adaptive [ESES20, FY21, WWT22, WMXL24, YYZ23, AH20, AJI24, Ali22, ASM22, FF23, GI20, GJQ23, JBJ23, KS23a, KB21, Lev21, LYY21, Mur23, NN20, PP23, SK24, STdL21, Som22, WKT22, WG23, Yad22, YDHW21, YSG23, ZA23, dSC23]. additive [AS21b, LE23, LMP20, PHN22, PSL23, PBLM22, SW23, SKV20a, ZG22a, ZFG22, ZGL22, ZLF23]. address [ARAR24, AA22c]. Addressing [MAAA21, Yad22]. Adjusted [TNN20, GABJ23, RAN23, TIK24, TAA22]. adjustment [CSE^{+}21, Hep21b, PSL23, AE23a]. Advances [FP20]. AFANS [AM24]. after [Wu22]. against [AG21, OY24, OP22]. age [NG22, ZlBVGW^{+}23]. age-specific [ZlBVGW^{+}23]. agent [YH21a]. aggregated [LW23]. aggregating [MV22a]. aggregation [TK22b]. agreement [TDD20]. agroforestry [BVJ^{+}22]. AIDS [TWTT21]. air [KKO21, TPC^{+}23]. aircraft [LRF20]. Ait [ZS23]. Ait-Sahalia-type [ZS23]. alarm [TF22b]. Alex [Bek23]. algorithm [AJI24, ASM21, BLXL24, BFS22, CCH20, CHL22a, CL22c, CAA22, DKKA22, ESLK22, EHF23, Els24, GW20, HYSC21, KTZ20, KJB23, LLV23, LW21, LWZ^{+}23, MS22, MW20, NN22, PR22, RM20, SWG24, SZ23, SMO21, TY22, Tel22, TWTT20, UIC23, VVNNT23, VPT21, Wan23c, ZZL23]. algorithm-based [LWZ^{+}23]. algorithmic [MPD20]. Algorithms [JPH21, BV22, Che22, ES22, NS22, NIN20]. allocate [JBJ23]. allocation [AM21, BFS22, CMMV21, LZ24]. allowing [CBRL22]. AlM [Bek23]. Almon [MA23a]. Almost [AQYA22, KL22, VW21]. alpha [TD23]. alphas [TD23]. alternating [GW20]. Alternative [KTK22, PLBP24, SSCT23, SM23b, Tak23, ZKW24, AF22b, Bap22, DS24, Fuk22, HZ23, SKV20a, SF23, YW24, AM24]. alternatives [Akd22, AG21, BKM23, Lue23]. American [MFS21, WZX23]. among [Fuk22, PM23]. amplitude [BLXL24]. Analysis [AKE23, AMP23, LY24, RPT23, RftADNI22, RSG21, SVN24, AKG23, ALB22, ALZ24, ANP20, Avc21, BYK24, BBD^{+}24, BH23a, BQCP23, Bia22, BS21, CSL20, CLY23, Cha23b, CKK21, CDB^{+}21, CSG22, CKCG21, DA22, DSO^{+}23, DF22, FN20, FLXY22, Gao23, GGMM22, GW20, HHvR22, HAT^{+}22, Har21, KOT20, KKK23, LR23, LJ22a, LSC23, LB23a, LS22b, Moh22, MFPP21, OM21, PM23, Pan20a, PRH21, Pol22, RM20, RY21, SBR^{+}23, SS23, Seo22, SMB22, SIK24, SLME23, SBZG22, TY24, TKR24, WP22, WCC22, WMS22, XZS22, YZF22, Yan23, Yas22, YWT22, ZA21, ZL22a, ZNKN22, ZWW23, ZX22b, ZL20, dOMD^{+}21, dSRC22]. analysis-based [FN20]. Analytic [LPP21]. analytics [ML22]. analyze [RSG22]. analyzing [LYZ^{+}21, YP23]. Anderson [BK23]. animal [PHN22]. anisotropic [Wan23b]. announcement [LWS23]. anomalies [ZSM23]. ANOVA [Cav24, Fig23, Lue23]. Anticipated [LX22]. APP [CWWT24]. Application [ASM21, KBYS22, RH23, TPK22, ULN21, YH21a, Abd18, AT23, AMD22, AQAN22, AW22, AM22b, Baz22, Bis22, CL22a, CSX20, CS22b, CLLX22, CLL23, CDB^{+}21, DA22, DPS22, FQKA23, GZ21b, HAT^{+}22, HORA22, HYSC21, IMR21, LDA23, LHY^{+}23, LLZ22, LZZZ23, eLhS23, LB23b, Lin22, LBH23, LWW22, MRM22, PASB21, RBRAa21, SKM^{+}23, SPG^{+}22, SS21a, SM23b, SF24, SBB24, fSLC21, SKJ20, SHN23, TWTT21, TL23b, VPOC24, XWZL22, XZJ^{+}21, ZQYW22, ZL22b]. Applications [Alt20b, LR23, ZL23c, AAF23, AAHR20, AK22c, AE23c, BAEk21, BSNT21, Bek24, BE20, Cha23b, CS22c, GNE21, GMC23, HHvR22, JKH20, KF24, KKH17, KS20c, KR22, KH23, KS23d, KE22, LX22, LB22, LRF20, RLP23, Som22, WW23, YHWW23, Yas22, ZB22]. Applied [MSL21, AA20d, BFAR20, BFS22, NJ22, RCSS22, Squ22, Ver21]. Applying [BF23, GAS20, TED24]. Appraisal [ALMA23]. approach [AJIAA22, AES21, Alg20, Alt20a, Ami21, AK23b, Avc21, AA20d, BB24b, BKIR23, BAL24, BDD^{+}22, CYS21, CLL20, CsSL22, CY24, CII22, De 20, DP23, DCL^{+}24, GI20, Gai21, GAS20, GSG22, HI22, HBS22, HB23, HMRN22, Hut21, KS23a, KT23, KA23, LCC22, LWZ^{+}23, LF24, LK24, Mac24, MPD20, Meg23, MRB24, Nad21, PG23, PDK23, PHS21, PASB21, RAA22, RH23, RM20, RBKA23, Sad24, SFG21, SAA^{+}23, ULN21, YC22, ZA23, ZLXT23, ZKW24]. Approaches [Wes20, BBK23, KR23, LSC23, PV20b, SSCT23, YWL^{+}23]. Approximate [HK20, ABC^{+}23, HCHK23, YW20]. approximated [ÇKK23]. approximating [HB23]. approximation [AY21b, BYK24, BRTW24, CL22b, Lav21, Lee20, LZ24, RAA23, TK23, iTS23]. Approximations [NZ21, BBM24, BDG23, CWWT24, GBHA21, Hom20, LZSC23]. arbitrarily [Hor22]. ARCH} [OB21, BGM20]. area [AESY22, AC23b, FHM22, IC21, MW20, MSS20, PG23]. area-based [PG23]. area-level [FHM22]. Areal [BJ20]. arising [BAEk21]. arm [WH24]. array [Ger20]. Artificial [KE22, CAA22]. asian} [CHL22b]. ask [LWC22]. aspects [AALR23, FP20]. Assessing [Lee20, MCJ22, WH24, ZBA23, AR22, CP22]. assessment [ANP20, CS22a, MB23a]. asset [AM21, MFS21]. asset-allocation [AM21]. associated [AZAZ23, Büy23, DYC22, GABG22, GBHA20, HZAF22, LMFMA23, ZQYW22, ZWC^{+}23]. association [Cha23b, CH23, DK22b, WK23]. assuming [LLLY23b, OY23]. assumption [Par20]. assumptions [CLY23, DYC22]. assurance [BJU20]. asymmetric [AGI23, Hep21a, MZZA24, Zhe21]. asymmetry [CQ22, GZ22b, ZG22a]. Asymptotic [AE23b, BPN22, Ega24, Esk21, LZSC23, MB24, Mil20, BS21, GDWY22, MVLMA22, SB24, dCSF24, TKF^{+}20, WLC^{+}22]. asymptotically [DRB20, DYC22]. asymptotics [GW22, MHW24, dXLY20a, dXGLY20, dXLY20b]. atmospheric [KKO21]. attri [Bal22]. attri-vari [Bal22]. attribute [ABPK21, RSS22, WZ21a, WCRM23]. attribute-variable [WZ21a]. attributed [FAT23, MASA22, TWSW23]. attributes [BQCP23, MS20]. Augmentation [AM24]. augmented [AGA20]. Australian [WC23a]. auto [BNA23, El 24]. auto-correlated [BNA23, El 24]. autocorrelated [FAT23, SMM22, SCK23, WHS23]. autocorrelation [KA23, RAG21]. automated [EGVdR21]. Automatic [AT21, Ciu22, LLV23, VPT21, Tel22, VVNNT23]. Autometrics [EGVdR21]. Autoregressive [GKM20, AGI23, ASA23, AE21, CWLH23, GNE21, GO22, LLZ22, LY22, LW22b, MWNSB23, NIN20, OR23, PBC23, PB23, SB24, SSZ23, WWYX21, WHW21, ZLF23, ZWCY24, ZKW24]. auxiliary [AS21a, AS22c, ASK23, BPP20, CD23, HB22b, JuAH20, NKSY22, uARS22, uAJHD22, uAA23, SNKA24, SIG23, SVM20, SVK22, TPK22]. Availability [KCS24, SS21a]. available [BF23]. Average [FZY24, ASA23, AK20, ACKL22, ACKL23, CLS22, CTK^{+}22, HGWZ22, HZZ^{+}24, KNSA23, LLZ22, uARS22, uAA23, OR23, OP20, PBC23, Tak23, WHS23, dSC23]. average-variance [uAA23]. averaged [TL23b]. averages [uASS22, QTH24]. averaging [HD22, HCZ21, YOD^{+}23, dSdSTC21]. averse [LZ24]. axially [VWZ21]. axis [RP23].

B [JKLK24, YZ21]. B-spline [YZ21]. B-splines [JKLK24]. BA [ZL20]. background [CS22b]. backward [LX22]. Balakrishnan [TS22]. balanced [AATD21, ARA23, ASA24, BB24a, BVJ^{+}22, DSS23, KVJH23, KBR^{+}23, MPD20, RBT^{+}22]. balancing [HAT^{+}22]. band [DLRZ21]. banded [JM20]. bandwidth [AHB23, ESES20, Sol21, Som22, ZZ22a]. Banerjee [PM20]. bankruptcy [BBK23]. barrier [LG21]. bars [BCC24]. Bartlett [Pic21]. base [YWT22]. based [ASZN21, ABHF22, AMK^{+}21a, AY21a, AS22b, Aki24, AGI23, AJIAA22, ALZ24, AHB23, Ali24, AR22, AJ22, ASA24, AAGP23, AA22b, ABPK21, ABJ21b, AA20d, ASP22, BBM24, Bal22, BNA23, BD23, BC21, BMG22, BM21, CWWT24, CKK21, CLS22, CS22a, CWLH23, CW23, CS22b, CMMV21, CTK^{+}22, CCM20, CNL21, CAA22, DKKA22, De 20, DSS23, DMP23, DA21, DWLS23, DCL^{+}24, El 24, Ere20, FN20, FLXY22, FBL22, FAT23, FL23, GK22a, GK22b, GWZC22, GDWY22, GSG22, HI22, HR22, Hep21b, HCTZ21, HE21, HE24, HS23, Hut22, IM20, IGA24, IS20, KR21, KF24, KJB22, KA24, KL23, KJB23, KR23, KG22, KS22b, LSC23, LL23, LLV23, LZSC23, LWZ^{+}23, LY24, LES21, LS22a, LLL22, Lon23, LK24, MG21, MWTR23, MNL21, MZZA24, MSL21, MÖ23, MQ21, MLR^{+}22, Mil20, MB21, MEA24, MBBS22]. based [MR22, MW20, MK22, NdSA20, NN21, NKSY22, uASS22, PHY20, Pak23, PG23, PD21, PHS21, RSM22, RYA20, RAJ21, RSG22, RM23, RCSS22, SI20, STdL21, SJ23, SBTP20, SU23a, SANH^{+}22, SAA^{+}23, SIK24, Sha22b, SL21, SMWJ20, SS21b, SLME23, SB21b, SM20, SZJG21, SKS21, SN23, SCK22, SSSR23, SCK20, SCK21, SCA23, SLE20, SHN23, TWSW23, TS22, TWTT20, TP22a, VDN22, WS20, WWYX21, WWT22, WZY22, WJZ^{+}22, WW23, WMJ21, Wu21, XXH22, XWL22, XH21, YYZ23, YÖT20, YW24, YL22a, Yoo23, YZ21, ZZFG20, ZGC21, ZG22a, ZNKN22, ZBA23, ZB20, ZfSRJ22, ZWCY24, ZX22b, ZLG23, ZHJ22, ZLSC20, ZBZS20, ZB24, ZKW24, dCdAB23, vO21]. based-estimators [DCL^{+}24]. baseline [BBM24, DSL20, PBLM22]. basketball [SS21b, SZS20, TGS22]. Basu [dOMA21]. batch [PK22]. bathtub [BMG22]. bathtub-shaped [BMG22]. Bayes [AC23b, AMK21b, CW23, Esk21, RY22, SU23a, YW20, ZGC21]. Bayesian [ASZN21, AKG23, ALZ24, ASA23, AA20b, AA20c, AM22a, Alh23, ARAR24, Ami21, AJ22, ALVQ22, AAGP23, Avc21, AA20d, BJ20, BC21, BC23a, BR24, CSL20, CKK21, CLY21, DA22, FW23, FLPB20, Gao23, GABJ23, GGB20, GK23, GJK21, Han20, Han21, Han22, HvZ21, HCLF23, dAIIS20, IT21, KSP23, KS23c, KK22b, KR23, KBYS22, KP21, Lev21, LE23, LLLY23b, MNN21, MdN22, Meg23, MZZF21, Min22, Min23, OLOK22, OY23, PC22, PCWL22, PMK23b, RAA22, RdAIS^{+}20, Seo22, Sha23, SIK24, STH24, SMB23, Som22, dSS22, SLE20, ÜFR22, VDN22, WP22, WWT22, XS23, Yad22, YBC22, YLX^{+}22, Yan23, YWT22, YOD^{+}23, ZLXT23, ZB20, ZWW23, ZLSC20, dSLHMB23]. bee [CAA22]. behavior [WJB21]. behavioral [ML22]. behaviors [EFD21]. Beijing [LWS23]. Bell [AAM23]. Ben [PM20]. benchmark [MG24]. bending [Squ22]. benefits [BH23b]. Benford [Kaz23]. Bent [ZZZ23]. Bent-cable [ZZZ23]. Berk [HM21]. Berkson [WCF21]. Bernoulli [CWWT24, Pho24]. Bernstein [DY22a, HS23]. Berry [DYC22, KL22]. best [Ami21, CZ22, Hut22, Ver21]. Beta [AT23, AMA22, CNL21, KP21, MAB23, PBC23, WW23, dSPdSB23, CWWT24, TIK24]. Beta-Binomial [TIK24]. Better [IR21]. between [AK23a, AK20, AE21, Büy23, CST21, DuA23, FS23, FSGMM23, Gül22, HQZ^{+}23, Lui22, MNN21, MAF23, MPS21, OP20, PK22, PAT22, SF23, UYMK22, WS22, YC22, ZZ21]. between-batch [PK22]. between-within [SF23]. Beyond [CHL22a]. Bi [SS21a, RSS22, WLZ23]. bi-attribute [RSS22]. Bi-dimensional [SS21a]. Bi-level [WLZ23]. Bias [KAVR20, MB23a, Tei22, WN23, AÇ23a, BBD^{+}24, BZ22a, BR21, GPS23, GCF^{+}20, HZAF22, MMdOC23, SCHS22, Sha22a, ZPW21, Zha24]. bias-corrected [Sha22a]. bias-corrections [MMdOC23]. biased [AAF23, AFJ23, AA22c, CZS21]. biasing [ÖT20, RAA22, SCK20]. bid [LWC22]. bidimensional [GTW22]. bifractional [KL22]. big [FF23, XH21]. bilateral [QLG23]. BiMM [SWC^{+}20]. bimodal [ROL^{+}20]. binary [AESY22, BJA22, DKKA22, FS22, HCLF23, KFM23, LFMMRH22, LKR^{+}23, Lue23, MASA22, OBN22, SI20, SMB23, SWC^{+}20]. BINMA [SKJ20]. binomial [AF22a, AKG23, Arn23, AKR23, ESLK22, GPP21, HE24, Jod23, KS20c, LB23b, MB23c, MB23d, PR22, PZ24, SK22a, SYPV23, UYMK22, WWYX21, ZBB23, ZWY20, TIK24]. binormal [GM24a]. biochemical [HAT^{+}22]. biological [ZCH23]. biomarkers [AESY22, JBJ23]. biosimilar [PK22]. Birnbaum [BA22a, CLZ23a, KJB22, LMFMA23, LBH23, Sha23]. birth [RdAIS^{+}20]. birthdates [PRH21]. Bivariate [Ere20, KH23, MPS21, AK22c, ASP22, Bap20, Bap22, GAS22, GMC23, LSTE24, MLR^{+}22, MEA24, NAKK20, NM21, PV22, RJJ23, RHG22, RK20, SAK24, SSBA22, SZ21, SLME23, TK23, WLT22, dXLY20a, dXGLY20, dOMA21]. blank [SKV20b]. blinded [TF22a]. block [Bey21, BJVV21, KS23b, KB23, MPD20, MBDK21, SD22, Yan23]. block-sum [KS23b]. blocked [ZK20b]. blockmodel [ESLK22]. blocks [Akd22]. BLUPs [Büy23, Gül22]. bond [SM24]. bone [LBH23]. boosting [LYP22, ZQY22]. Bootstrap [BCP21, FH22, ÖA23, PS23, YGL22, YWL^{+}23, ZZ22a, Bey21, BRA20, FL23, Hut21, Hut22, Hwa23, LSC23, MBDK21, MRM23, MSS20, MMVBE23, PP23, PDO^{+}20, PV20b, Sol21, WDZ22, ZMNL23]. bootstrap-based [Hut22]. bootstrapped [AS23]. Bootstrapping [CS20, XW20]. bootstraps [IHS22]. border [DSO^{+}23]. both [CW24]. bottleneck [ML22]. bound [DYC22]. Boundary [FM23, NZ21]. Boundary-free [FM23]. Bounded [GE23, dALFB21]. bounds [KL22, XSS22]. bracelets [DG23b]. Bradley [SDG22]. branching [RY21]. breaking [DK22a]. breaks [De 20]. bridge [Alh23]. broader [HZ22]. Brownian [DPS22, KL22]. BSDEs [Xia22]. BTPs [Bia22]. bubbles [LWS23]. Budget [ZRC23]. Burr [AS22b, LS23, MCJ22, WCKC23]. Bursa [NN22]. business [ML22].

cable [ZZZ23]. calculate [ZMNL23]. calculation [YL22a]. calculations [CFM21]. calibrated [Mag20]. Calibration [AS22c, CD23, LWP24, SSS20, MF22, SAA^{+}23, SBB23, ZXW23, ZX22a]. cancer [HAT^{+}22, RBRAa21, YW24, ZA21, ZB22]. canonical [GW20, Pol22]. capability [AK22a, ABJ21b, Bal22, PG22, PDO^{+}20, RAJ21, SCK23]. capacity [ZYSS22]. capture [Kal21]. Capturing [BCR^{+}21]. carbon [LWS23]. cardia [YW24]. cards [SKV20b]. care [SCE^{+}20]. Carlo [AQAN22, CL22b, DK22b, FPAA21, KKZ21, KSP23, KOT20, uIKKH22, Kur21c, LPNP20, Squ22, UU23, Uya22, WMS22, ZL20]. CARMA [STH24]. case [BS22a, BKIR23, CSE^{+}21, FS23, GKZ^{+}22, KOT20, MNL21, MAMA20, NN22, ONE22, Pan20a, PD22, PSL23, SB20, SPG^{+}22, SCE^{+}20, SMB22, SANH^{+}22, SW23, SRO20, SU23b, SK23, dOMA21]. case- [MAMA20]. case-cohort [Pan20a, PD22, PSL23, SW23]. case-mix [CSE^{+}21, SCE^{+}20]. CAT [AW22]. categorical [AMK^{+}21a, JL24, PM23]. category [Avc21, HLT23]. Cauchy [KAVR20, KAVdSFG22]. Causal [ZCG24, KS23a, YW24]. causes [CSL20]. caution [SLJ23]. CDF [Ras20]. censored [ASZN21, AHAMP23, AZAZ23, Alh23, ARAB21, AJ22, AMK21b, AHG20, ASP22, BKM23, CLY21, CsSL22, CW23, DWLS23, GKS24, HB22a, KJB23, KKK23, KG22, Lee20, LTB23, Lon23, LHG24, MG21, MWTR23, MK21, MAMA20, MB21, MBBS22, NDWE24, PHY20, Pak23, PC22, PCWL22, PV20a, POC23, RBRAa21, SBTP20, Sha22b, SLSZ22, SVN24, TY22, TRF^{+}22, WG23, WCC22, Wu21, WLZ23, ZA21, ZWW23, ZB24]. censoring [AKE23, AB22, BR24, BB24b, BMG22, CSL20, CKK21, DSS23, GAS20, GK23, KB21, KP21, LTR21, LS23, MJ20, Mis20, PMK20, PMK23b, Seo22, Tei22, Wan22, WLT22, WL22, WC23b, WCKC23, YBC22, YSG23]. censorship [FZZ20]. center [Bek23]. centered [HQN22, ONS22]. central [RP23, TD23]. centrality [ZSM23]. Centroid [SWG24]. CEV [ARAB21]. CEV-Hybrid [ARAB21]. chain [HE21, HE24, KSP23, MKS22, SS21b, SLE20]. Change [TY24, BSNT21, DR23, KL23, LW23, LLLY23b, MMMG21, QYC22, WTN22, Wu22, ZT20, ZHJ22]. change-point [BSNT21, WTN22]. change-points [LLLY23b, ZT20]. changed [Wu22]. changes [BVJV22]. changing [TPvdH23]. chaos [BCDA21]. character [SKV20b]. characteristic [AESY22, GM24a]. characteristics [CSE^{+}21, CS22a, MK21, WQW^{+}22, Wij23, YBC22]. characterization [KBB24, Mil20]. Characterizing [Büy23]. charged [EHF23]. chart [ANR^{+}23, ASS22, AMK^{+}21a, AK20, ACKL22, AK22b, ACKL23, ASM22, ARJ20, ABPK21, CKL24, CLS22, CW24, CKT^{+}20, CTK^{+}22, Cos23, ENA^{+}23, GK22a, HU20, HEK22, HORA22, HCTZ21, JuAH20, JuAH23, JL24, KM22, KNSA23, KST^{+}22, KL21, LK21a, LKHC23, LTJC23, dALFB21, MMMG21, MLR^{+}22, MKSK23, uARS22, uAJHD22, uAA23, RAN23, SK24, SAF21, SAA21, SKM^{+}23, STdL21, SB21b, TAS24, TAA22, WZ21a, WHS23, ZHJ22]. charts [AH20, ARAB21, EAJ20, FN20, GK22b, GM24b, HB22b, HQZ^{+}23, HZZ^{+}24, Hua22, Hua23, HMRN22, KA24, KBR21, KS20a, MMMCS22a, MÖ23, MKSH22, NKSY22, PM20, SNKA24, SSBA22, SN22b, WCRM23]. checkerboard [GBHA21]. checking [ZG22a]. chemical [AQAN22]. children [IM20]. China [CLL23]. Chinese [LZQ23]. chirp [BLXL24]. Choice [AK23a, ALVQ22, LW22a]. Cholesky [Lu23, vO21]. Cholesky-based [vO21]. Choosing [Ver21]. CIR [MFS21, TMZL21]. circular [AATD21, ARA23, BZ22a, KBR^{+}23, RBT^{+}22, SRF^{+}23, TA24]. claim [TK22b]. claims [HB23, TYLX22]. class [AF22a, ATYK20, AA22c, ASK23, BVJ^{+}22, BZ22b, CHL22a, Che22, CCM20, FY21, HZ22, Hwa23, KS23d, Ruz22, SNSE22, STH24, SK22c, SBB23, SR23, TBK23, WKSeF22, WYY22, ZLG23]. classes [AK23a, NG22]. Classical [BC23a, BR24, Sha23, HB23]. Classification [AM24, JD23a, LX23, BD23, BQCP23, Cha23a, CY24, DKKA22, DP22, IAG21, IGA24, KTK22, MS20, ÖDS24, Wan23c, YÖT20, ZC22, ZCH23]. classifications [dOVQP21]. classifiers [DKKA22, DMP23]. classifying [PK21]. clean [CLL23]. cleaned [LB23a]. clinical [HHOC23, JBJ23, ZL23c]. close [ZLG23]. Closed [Lev21, HK20, LWC22, PHN22]. Closed-form [Lev21, LWC22]. closeness [AR22, MR22, WWTM22]. CLS [MB24]. CLT [KL22]. cluster [BM21, CDB^{+}21, KS20c, dOMD^{+}21]. cluster-based [BM21]. clustered [AS22a, DCL^{+}24, FJ24, LSTE24, Sha22a, SWC^{+}20, SF23, Yoo23]. Clustering [JH23, AY21a, BJA22, CGS21, CCH20, CII22, CQ22, CAA22, HYSC21, LLV23, LH20, REMM22, Tel22, TPvdH23, VVNNT23, VPT21, ZBA23, ZfSRJ22]. clusters [BDD^{+}22, DSO^{+}23, FBNRG21, Tel22, ZNKN22]. clutter [TF22b]. Co [BJA22, LWW22]. Co-clustering [BJA22]. Cochran [HMV^{+}22]. COD [ÜFR22]. code [SLP22]. codifference [GTW22]. coefficient [AS21a, CWLH23, CKL22, DK22b, ESS22, GWZC22, GZZ22, GP20, HCZ21, HZZ^{+}24, HZY22, IR21, KJ20, LTJC23, LZ22, LW22b, Mos22, RuA22, SGG22, Sha22a, SVM20, SBZG22, TNN20, WKT22, WH24, YXYX22, ZZFG20, ZG22a, Zha22a, ZXW23, ZL23b, ZYF24, ZZWH20, ZZ22a, ZXF22, ZXZL22]. coefficient-based [ZG22a]. coefficients [AE23b, EAT22, PLBP24, SF24, TL23b]. cognitive [ML22]. coherent [LZ24]. cohort [Pan20a, PD22, PSL23, SW23]. Coincidence [Pic21]. cointegrating [AÖ20b]. cointegration [Hep21a, uIKKH22]. collinearity [KFM23, SR23]. colony [CAA22]. COM [AK22b, BBM24, SKJ20]. COM-Poisson [AK22b, SKJ20]. combination [REN22, SU21, ZMNL23]. combinations [De 20, YH21a, Yad22]. combine [CL21]. Combined [Bal22, Hua20, Cha23a, GM24b, QLG23, WZ21a]. Combining [AESY22, CGS21, KLM24, ÖA22, IC21]. Comment [GRR24]. Comments [CL22a]. Common [Kal21, KJ20, KLM24, KTK22, TNN20, Wu22]. commonly [Lui22, Uya22]. commonly-used [Lui22]. communication [SMWJ20]. Communications [AE23c, Kri22, Ano24]. community [TWSW23]. compact [WJB21]. company [BS21, WDZ22]. Comparative [AR22, EJKL22, SSS23, AE21, BA22a, Bay21, CLZ23b, FSGMM23, KGR21, LWP24, MFPP21, ÖA23, SMB23]. compare [CYS21, RM23]. Comparing [AGI23, AG21, CSG22, dAIIS20, NS22, SRN^{+}22, SS20b, Wil22, Dog22, FO22, MHW24, SG21, SS20a, ZB22]. Comparison [Aga23, Akd22, AK23c, Cav24, CLL20, CH23, CCBN22, GM24a, KR23, MÖ23, NBC23, NHZ22, ONE22, PBLM22, PL22, SM23a, UR22, AFJ23, AKS^{+}21, BS22b, CMMV21, FS23, GLT20, HHH23, IM20, IRA21, JM20, KSP23, uIKKH22, LSC23, Lui22, MSL21, MPS21, SCHS22, SS20a, SMMO22, TPvdH23, UY23, Uya22, WL20, WC23b, ZSM23, ZlBVGW^{+}23, dSRC22]. Comparisons [AGA20, ÖÇ23, BDL23, HM21, KLM24, WWW24, Wu21]. Competing [MN17, AK22c, AHG20, CSL20, CSA23, LY24, LTB23, LHG24, PR22, WLT22, WL22]. Complete [LHY^{+}23, ASZN21, ASP22, BKM23, YHWW23]. complex [BCR^{+}21, HAA^{+}23, Ors24, PK21, PHS21, RH23, SLP22]. complexity [DP22, Els24]. component [AE21, Çet23, FN20, Kur21b, Moh22, PMM23, TII22, YGL22, ZNKN22]. components [LMP20, SRF^{+}23, VLvdW22, YW20]. Composite [JGL22, MMR22, DA22, HHY21, LMP20, MW20, MV22b, OY24, Pan20b, SK22c, TD23, TWTT21]. compositional [GGMM22, MZ23]. compound [dCSF24, ZK20b]. compromised [ASK23]. Computation [GLT20, Kri22, ABC^{+}23, AE23c, Bar20a, KAVdSFG22, SMO21, Ano24]. Computational [AALR23, BFAR20, HE24, AK23b, CdCG22, DG22, FP20, GSG22, YÖT20]. computationally [CsSL22, LW22a, dCSF24, SY22]. computer [Jod23, LFPL24, LWX23]. Computing [Nag21a, SVK22, SZ23, Wil24]. concentration [MG24]. concepts [ZNKN22]. conceptual [KÖ21a]. concomitant [AA22b]. concomitant-based [AA22b]. concordance [TL23b]. condition [FPAA21, Jhw24]. Conditional [KL23, SP23, AALR23, BCP21, Baz22, CLC22, DKL22, ESES20, GMC23, HCTZ21, KBR21, LE23, WK23, dXLY20a, XWL22]. conditionally [AGI23, Mil21, YH21b]. conditions [LW22b]. conduct [dOVQP21]. Confidence [CST21, Cur23, DG23a, DP20, Kri22, QLG23, SPV22, UYMK22, dCCM^{+}22, BCP21, CTCL22, DLRZ21, IRS20, IR21, LLL22, MNN21, MCJ22, MMVBE23, PDO^{+}20, RLP23, RM23, SI23a, SYPV23, TNN20, WWTM22, Wil24, ZX22a]. conformable [XWZL22]. conformal [TED24]. confounders [DNWG22, WGlTL23]. conic [YÖT20]. conjugate [PR22]. consideration [VSJ^{+}24]. considerations [BK23, RAN23]. considering [MASA22]. consistency [DC24]. consistent [MNL21, PMK23a, TYLX22, YL22b, ZL22b]. Constant [KS23b, LG21, NDWE24]. constant-stress [NDWE24]. constrained [CHL22a, Pan20a, ZRC23, ZBZS20]. constraints [DY22a, MQ21, PD22, RVV22]. construct [dSC23]. constructed [JHR21, SSBA22]. Constructing [GGS23, Kaz23]. Construction [Bey21, KBR^{+}23, LFPL24, MPD20, dSC23, ARA23, BGJ^{+}23, UIC23, VJV^{+}23]. consumption [CLL23]. contagion [CDB^{+}21]. containing [GAS22]. contaminated [Dog22, SFG21]. contamination [AM22b]. context [DKL22, LLLY23b]. contingency [AS22d, HLT23, KJK20, WK23]. continual [UKK20, ZWY21, ZWMY23]. continuous [AESY22, AMK^{+}21a, CBRL22, GABJ23, GGB20, KLM24, LFPL24, LK24, RSG21, RS24, RK20, Sha22a, ZG22a]. Control [EAJ20, KBR21, ASS22, AMK^{+}21a, ACKL22, AK22b, ACKL23, ASM22, ARJ20, ABPK21, Ava21, BQCP23, CW24, DC20, ENA^{+}23, FN20, GK22a, GK22b, HORA22, HZZ^{+}24, HE21, Hua22, HMRN22, JuAH20, JuAH23, JL24, KA24, KM22, KFM23, KST^{+}22, KL21, LTJC23, dALFB21, MMMG21, MÖ23, MLR^{+}22, MKSK23, uARS22, uAJHD22, uAA23, PM20, RAN23, RYA20, SK24, SAA21, STdL21, SSBA22, SMO21, SS20b, SN22b, TAA22, WZ21a, WHS23, WWW24, Wu21, Wu22, WCRM23, Yas22, ZL23a, ZHJ22]. controlled [SRN^{+}22]. Controlling [LAZ21]. convergence [GBHA20, LHY^{+}23, YHWW23, ZQYW22]. conversion [KBR^{+}23, RBT^{+}22]. converting [Mag20]. coordinates [GGMM22, JKLK24, KS20b]. Copula [ADK21, QL21, SIK24, BH23b, GAS20, GBHA20, HD20, HE21, HE24, HS23, KCS24, KBR21, LY24, LES21, LSTE24, LK24, MEA24, RSG22, RS24, SLE20, dXLY20a, dXGLY20, YC22]. copula-based [HE21, HE24, LES21, RSG22, SLE20]. copulas [Bek24, CS22a, GBHA21, SSBA22, JHR21]. corporate [WZX23]. correct [Aub23, CNL21]. Corrected [GCF^{+}20, Özk21, Sha22a, Tei22]. Corrected-Hill [GCF^{+}20]. Correcting [BR21]. Correction [Ano24, AE23c, AÇ23a, BZ22a, HZAF22, Hep21a, ZPW21]. corrections [MMdOC23]. correlated [AA22a, AS21b, BNA23, El 24, KFM23, KBB24, LES21, LY23, Meg23, PL22, RSG22, Yad22, YTL22]. Correlation [CZP21, ZG22a, AE23b, Bap20, Bap22, DK22b, DS24, GWZC22, GZZ22, GW20, Kol24, MS22, OR23, PLBP24, Pol22, Sha22a, SLJ23, TL23b, UR22, Wes20, ZXW23, ZL23b, ZYF24, vO21]. Cost [AMA^{+}23, LJ22a, KS23c, MYVD21, PM20, SA22]. cost-effectiveness [KS23c]. costs [KS23c, ZC22, ZC23]. count [AK23a, AS22a, AKR23, BLR23, HD20, Hom20, MRB24, MV22a, OLOK22, RSG22, RS24, SF23, TII22, TW24, TEO20, Yoo23]. counterpart [Wan23a]. counting [GBHA20, LZ22, LTN23, PAT22, SM23b]. counts [MPS21, TK22b, dSdSTC21]. course [ML22, VBB23]. covariance [ALB22, AM21, BZ22b, GZ21a, Jhw24, JM20, Kal22, KA22, MX24, MRM23, RAG21, SS23, SANH^{+}22, TCJ23, iTS23, YP23, ZK20b]. Covariate [GABJ23, AHAMP23, BYK24, HAT^{+}22, MAMA20, PSL23, RHG22].

Covariate-adjusted [GABJ23]. covariates [DSL20, Kur23, LE23, MZ23, MZZF21, MGS23, NCC^{+}22, PLS23, SLSZ22, SBZG22, WHW21, WCF21, XS22, ZXZL22]. Coverage [IRS20]. covering [ZL23c]. COVID [SM23b]. COVID-19 [SM23b]. Cox [AJ22, BYK24, CMA21, Pan20a, PD22, SSCT23, SI20, ULN21, WSZ23, ZGC21]. Cpk} [Cos23]. Cramér [XSS22, ZZFG20]. credit [CHL22b, GHW20]. cricket [BT20]. crisis [NN22]. Criteria [Hom20, CCM20, Kim21, SLJ23, Wes20]. criterion [AA23b, BSNT21, DRB20, HK22, HCZ21, KÖ21a, MGS22, Mis20]. critical [ABC^{+}23, KCS24]. Cross [GTW22, LY22, Alg20, Pol22, RHG22, SYPV23, ZCH23]. Cross-codifference [GTW22]. cross-language [Pol22]. cross-product [SYPV23]. cross-ratio [RHG22]. cross-validation [Alg20, ZCH23]. crossing [HLHC22, KJB23, NZ21]. crossover [KS22b, Yan23, ZL20]. Crosswise [AE23a]. crucial [Spa22]. cum [KGF20]. cumulative [CP23, CHC20, GPS23, LDA23]. cure [BBM24, LBH23, PR22, RdAIS^{+}20, SL21, WLZ23]. current [HHH23, SPS22, WFLS24, XZS22]. curse [ÜFR22]. curtailed [CLY21]. Curve [Ors24, AESY22, ANP20, CdCG22, HHvR22, LK21b, PY20, Wan23a, YZF22]. curves [CHC20, GM24a, HLHC22]. customer [LJ22a, YWT22]. customer-base [YWT22]. customized [KS20a]. CUSUM [AH20, DR23, GDWY22, GM24b, HB22b, KL21, LL23, MMMCS22a, RYA20, ZL23a]. CUSUM-GWMA [MMMCS22a]. cutoff [Tak23].

D [Ava21, EHF23, OP20]. D-Trace [Ava21]. daily [SMB22]. Danish [MNL21]. Darling [BK23]. Data [VWZ21, Abd18, AS22a, AMD22, AMK^{+}21a, AY21a, AHAMP23, AKG23, AA20b, Alh23, ARAB21, AM21, AR22, AQAN22, AJ22, ALVQ22, AMK21b, AÖ20b, AM22b, AE23c, AHG20, ASP22, BJ20, BLR23, Bek23, BJA22, BYS24, BSAS22, BKPS23, Bis22, BDD^{+}22, Bul24, CL22a, Cha23a, CLL20, CCH20, CZP21, CsSL22, CY24, Ciu22, CZS21, DP22, DP20, DG22, DA22, DuA23, DWLS23, DCL^{+}24, DPS22, El 24, EE23, ENA^{+}23, FW23, FF23, Fan22, FJ24, FZM21, FM23, FZZ20, Fig23, Fuk22, GI20, GZZ24, GAS22, GY23, GGB20, Gho23, GGMM22, GKS24, GGS23, GP20, GZ21b, HMSRM23, HHvR22, HD20, HHOC23, HHH23, HCZ21, HZY22, HLT23, HYSC21, IAG21, IC21, KF24, KS22a, KKH17, KJB23, KG22, KS22b, LW23, LLV23, LE23, LH20, LYY21, LY24, LB23a, dALFB21, LES21, LSTE24, LBH23, LTN23]. data [LYZ^{+}21, LTB23, Lon23, LRF20, Lu23, LK24, MZ22a, MWTR23, MT21, MT23, MAMA20, MB21, MA23b, MWNSB23, MRB24, MBBS22, MV22a, Mos22, Nad21, Nak21, Nak22, NDWE24, NAG21b, OLOK22, Ors24, ÖDS24, PM23, PHY20, Pak23, PG23, PC22, PCWL22, Pap20, PK21, PHS21, POC23, PV20b, QLG23, RPT23, RftADNI22, RBRAa21, RH23, RSG22, SAA21, SSCT23, SBTP20, SCHS22, SMB22, SFG21, SANH^{+}22, SIK24, Sha22b, STH24, SLME23, SKK21, SU23b, Spa22, SCK23, SVN24, SF23, SHN23, TED24, TII22, TY21, TK22b, TY22, TPvdH23, TCJ23, TWTT20, TRF^{+}22, TL23b, TEO20, UM22, UR22, UU23, VPOC24, VBB23, VVNNT23, Wan22, WHW21, WL22, WWTM22, WFLS24, WJB21, WCC22, WS22, WLZ23, XZS22, XH21, YSH23, YZ23, YZF22, YP23, Yoo23, ZA21, ZGL22, ZZL23, ZB20, ZZWH20, ZWW23, ZCH23]. data [ZQY22, ZL20, dSRC22]. database [Ver21]. datasets [Cha23b, CL22c, JGL22]. Davis [Aki24]. Daya [PM20]. DEA [ÜFR22]. deal [KS22a, SKK21, SK22c]. Dealing [KHK22, MRB24]. death [KKK23]. debiased [KS23a]. debit [GHW20]. deciles [SB21b]. decision [MM22, NG22, PMK20, SWC^{+}20]. decomposition [AJIAA22, AJI24, BYS24, DP23, Lu23]. Decrease [EL21]. decreasing [KS23d, PM20]. deeper [ML22]. default [CHL22b, fSLC21]. defective [HG22]. degradation [CSDT20, Hu23, KBYS22, SBZG22, WWT22, ZX22b]. degree [IMR21]. degrees [SF23, TEO20]. Delaunay [NAJ22]. delay [SMWJ20, SMO21]. delay-based [SMWJ20]. delayed [TYLX22, YH21a]. deletion [IHS22]. delta [MNN21]. delta-lognormal [MNN21]. dengue [HYSC21]. denoising [Lev21, TP22b]. densities [QL21]. Density [MRM23, Aki24, BNN22, BRTW24, Bar20a, BCP21, BZ22a, CZS21, DC24, GJQ23, HK22, dAIIS20, LJC^{+}20, MEA24, QL21, Sha20, SD22, Som22, WP22, WN21b, ZCW23]. density-hazard [BNN22]. departure [AS22d]. dependence [ADK21, CDB^{+}21, DBS23, RH23]. dependency [YC22, ZKW24]. dependent [AHAMP23, AK22c, AMA^{+}23, AÖ20b, ARJ20, ABPK21, ABJ21a, ABJ21b, BCR^{+}21, CSL20, CsSL22, JHR21, LY24, LZ22, MAMA20, Med22, NG22, PV20b, RAJ21, SM23b, Sha20, SM20, TKF^{+}20, dOVQP21, WLT22, Wil22, XZS22, ZD22, ZMNL23]. depth [DMP23]. derivation [FZY24]. derivative [MKK22]. derivatives [YDHW21]. derived [UYMK22]. description [MT23]. Design [MNL21, PHS21, SPG^{+}22, WCRM23, ABJ21a, CP22, CKT^{+}20, FLXY22, HQN22, Hua23, HMRN22, KJB22, KGF20, KL21, KK22b, KB23, KBB24, LKHC23, LZZZ23, LWX23, PMM23, PSL23, PM20, RAN23, RGA21, REN22, SY22, TD23, WCKC23, Yad22, Yan23, ZWY21, ZD22]. Design-based [PHS21]. Design-consistent [MNL21]. designed [WZ21a]. Designing [ABPK21, BJU20, MS20, Els24, HORA22, SAF21]. designs [AATD21, ARA23, AGA20, AJ22, AAA20, BB24a, BJVV21, BVJV22, BGJ^{+}23, BVJ^{+}22, BDK^{+}23, CTK^{+}22, FTC21, GJK21, GKJ23, HAA^{+}23, KVJH23, KS23b, KFEL23, KBR^{+}23, KS22b, LFPL24, LFMMRH22, Lui22, MPD20, NN20, NKSY22, OY24, RBT^{+}22, TdSCdSdCP22, Tak23, UIC23, VSJ^{+}24, VJV^{+}23]. desirability [SCHS22]. destructive [PR22, RdAIS^{+}20]. detect [DSO^{+}23, GK22a, GK22b]. Detecting [AE21, OP20, WZ21a, FBNRG21, Ham22, ZSM23]. Detection [BYS24, GZ22b, GZ22a, WS21a, WS24, ZGL21, AMK^{+}21a, BDD^{+}22, CMM23, DK22b, GABJ23, HCLF23, KFEL23, KA22, LLLY23b, MRB23, Mur23, NHZ22, SCK22, TWSW23, TK22a, UEA20, WTN22, Wu22, YXYX22]. detections [QYC22]. deteriorating [KFB23]. determinant [BZ22b, Kal22].

Determination [AS22d, ABJ21a, MY23, HHOC23, KS23c, SGG22]. determined [CQ22]. Determining [PJL21, BB24b, IM20]. deterministic [JGGDMRLD21]. detrending [GG22]. Developed [ÇKK23]. Developing [AS22a, BNA23, NN21, Spa22]. development [FS22, PK22, VPOC24]. Deviance [Kim21, MÖ23]. deviances [RM23]. deviation [BRTW24]. device [SKV20b, SSK22]. Dewma [ARAB21]. Dhar [dOMA21]. diagnosed [IM20]. diagnosis [ÖDS24, YW24]. diagnostic [ANP20, GABJ23, HBM^{+}20, HE21, KL21]. Diagnostics [SMG23, AÖ20a, KUA21, KUA23, MSNN^{+}24]. diagrams [Aga23]. dichotomous [LSC23, ZL20]. difference [CST21, MNN21, UYMK22, WS22]. differences [OP20, QLG23, RLP23]. Different [CSA23, SWG24, ARA23, AGI23, AM21, CYS21, CdCG22, DSL20, FS23, GLT20, Han20, Han21, Han22, HHY21, OY23, Pol20, ROL^{+}20, RAA23, SSS23, SYPV23, TK22b, UKK20]. differential [LX22, MZZF21]. diffusion [GHW20, LLLL22, LLLY23a, MFS21, SBZG22, TH23, WZX23, WMJ21]. diffusions [MF22]. Diggle [CLL20]. dimension [ABC^{+}23, AM24, CZP21, CW24, CQ22, Kol24, WZY22]. dimensional [AALR23, Bul24, CKCG21, DNWG22, Ega24, FY21, Fan22, GY23, IAG21, JM20, Kur21a, Kur21b, Kur23, LYY21, LXMT22, MT21, MX24, Nak21, Nak22, SS21a, SMB23, SLSZ22, TCJ23, VLvdW22, WX21, WKT22, WCF21, XS22, ZWTY23, ZCH23]. dimensionality [ÜFR22]. dimensions [HHY21]. directed [HPZ^{+}21, ZSM23]. direction [GW20]. directional [ADK21, BJVV21, DMP23, NAG21b]. Dirichlet [GKJ23, LH20, MVLMA22, WJB21]. discontinuity [Tak23]. discontinuous [AA20d]. discovery [BR21]. discrepancy [HQN22]. Discrete [Ano24, HZAF22, JS19, ADK21, AK23b, CSBdES22, DPS22, LR23, SS21b, TH23, VPT21, YL22a, Yoo23, ZlBVGW^{+}23]. discrete-time [SS21b]. discriminant [KS20b, SS23, WP22]. Discriminating [PAT22, PDK23]. Discrimination [UM22, MGS22, TP22b]. Discussion [ACKL22]. disease [ÖDS24]. disorder [IM20]. dispersed [TW24]. dispersion [AS21a, BLR23, CW24, ESS22, SVM20, Spa22, dOMD^{+}21]. dissimilar [CII22]. Distance [DMP23, NdSA20, AY21a, LLV23, LWP24]. Distance-based [DMP23, NdSA20]. distances [SRF^{+}23]. Distinguishing [GKZ^{+}22]. distortion [FZY24, GZZ24, dXLY20b, ZG21, ZG22a, ZFG22, ZGL22, ZXW23, ZL23b, ZCW23, Zha23, ZYF24, Zha24, ZZ22b]. Distributed [FF23, GS23, ÇKK23, CHL24, DSS23, DDK21, DuA23, HORA22, IRS20, MAAA21, MA23a, ÖT20, SSZ23, Spa22, TF22b]. Distribution [Alt20b, Ham22, MMMCS22a, TF22a, iTS23, Abd18, ASZN21, ABHF22, AHAMP23, AOA23, AZAZ23, AK22b, ARAB21, AK22c, Ami21, Ano24, Arn23, AMK21b, AAGP23, AA22b, ABPK21, AE23c, ASP22, BA22a, Bar20a, BAEk21, BKG22, Bee22, BC23b, BMG22, CL22a, CSDT20, CW23, CBRL22, CCBN22, Dog22, DWLS23, FK22, Fig23, GDWY22, GPP21, GABG22, GW22, GM22, GMC23, HG22, HR22, Hua22, IMR21, JAP21, JS19, JKH20, Jod23, KAVR20, KR21, Kal21, KLK21, KJB22, KTZ20, KKH17, KNSA23, KBR21, KB21, KS20c, KR22, KH23, KS23d, KP21, LH20, LPP21, LB23b, LB22, LX24, LTR21, LTB23, Lon23, LRF20, LY20, LHG24, MWTR23, MRM22, MK21, MSL21, MB24, MG24, MM20, MMdOC23, MBBS22, MMVBE23, MV22b, NAKK20, NBC23, NdSA20, NIN20, NM21, ONS22, ON21, OP22, PHN22, PBLM22]. distribution [PV22, PAS23, PASB21, PMK23a, PMK20, PMK23b, RL21, RBRAa21, Ras20, RSS22, Ric21, RTJ24, RC23, RK20, SAK24, Seo22, Sha23, SJ24, SIG23, Sha22b, TAS24, TS22, TY24, TRF^{+}22, TP22a, UY23, UKK20, VG22, WLT22, WTN22, WG23, WZW23, WQW^{+}22, Wij23, WCKC23, XS23, YBC22, YWLD20, YSG23, ZNLW20, ZWC^{+}23, ZX22b, ZLG23, ZBZS20, ZL20, ZB24, dOMA21, Biç20]. Distribution-free [Ham22, MMMCS22a, TP22a]. Distributions [GBHA20, AMA22, AMD22, AS22b, AGI23, AAHR20, ARJ21, BNN22, BJU20, BPN22, CKK21, CCBJ21, CLY21, Cur23, FN20, FBL22, Fuk22, Han22, HHOC23, Hom20, KAVdSFG22, Kaz23, Kim21, Kri22, KLM24, KGR21, LE23, LF24, Lin22, MZZA24, MJ20, MNN21, MB21, MFPP21, Pan20c, PDK23, PAT22, RYA20, RLP23, RY22, SU22, SK22b, SKG22, SPV22, Sul22b, Sul23, Sun23, TII22, TIK24, TJ20, UEA20, WZY^{+}21, WMXL24, ZB22]. disturbances [AS23, VH23]. divergence [AS22d, LK21b, MEA24, MK22, PHY20]. Diverse [DKKA22]. diversification [BH23b]. Dividend [LC22, LG21, TYLX22, ZC23]. DMTTF [BKM23]. DMUs [ÜFR22]. do [Gan22]. Does [BH23b]. doi.org [Kri22]. domain [LFPL24, ZX22a]. dose [NN20]. dose-response [NN20]. Double [LK21a, MMMCS23, ZWTY23, ANR^{+}23, AKG23, AS22c, AK20, ACKL22, HEK22, HZZ^{+}24, KS23a, KM22, LTJC23, LLLY23a, dALFB21, Lon23, MMMG21, MMMCS22b, YYZ23]. double-robust [YYZ23]. double-sampling [AKG23, LTJC23]. double/debiased [KS23a]. doubly [PIA22, Wu21]. down [LC22]. downside [LZQ23]. draw [LC22]. draw-down [LC22]. driven [BAL24, GWZ23, LC22, STH24]. Driving [LWW22]. dropout [CLL20]. drug [Yad22]. DSLRIG [SAD^{+}21]. Dual [fSLC21, AAGP23, HB22b, LFZ21]. Dual-frailty [fSLC21]. due [KS22a]. Dynamic [SLME23, dSdSTC21, DNWG22, HMSRM23, MPS21, RRA20, SMWJ20, TEM20, TAA22, YL23]. dynamics [DF22].

E-Bayesian [Han20, Han22, Han21]. E-MSE [Han21]. E-posterior [Han20, Han22]. each [Ver21]. early [PM20, SEMR23, TK22a, YW24]. EBLUP [MSS20]. ECF [Lin22]. ECF-WS [Lin22]. Economic [Hua23, KJB22, LKHC23, NKSY22, ABJ21a, LWW22, PM20, RAN23, RGA21, REN22]. Economic-statistical [LKHC23, NKSY22, REN22]. economical [DR22]. economically [WZ21a]. EDF [CSA23]. Education [SN22b]. educational [UM22]. Effect [ÖDS24, TL23a, BCC24, BJVV21, FJ24, FLXY22, GZ21a, Han22, Har21, Hua22, IM20, LWW22, ML22, MV22a, NSN21, SEMR23, SRN^{+}22, SRF^{+}23, SVK22, Tak23, XS23]. effective [BKIR23, SW23]. effectively [Els24]. effectiveness [ASA23, KS23c]. Effects [PK22, TEM20, AW22, BCR^{+}21, CHL24, CdCG22, DY22b, Daw23, EH22, GLT20, GKJ23, Gül22, Hua20, Lar20, LE23, LWZ^{+}23, MASA22, PHN22, Pol20, RHG22, SLW20, TWTT20, TWTT21, VJV^{+}23, WGlTL23, YGL22, YW20, YW24, ZBB23, ZK20a, ZL20]. Efficiency [ÖT20, AK23c, CKL22, LKR^{+}23, Mil20, PM23, SSBA22, XWL22]. Efficient [BS22a, BS23b, BM21, MYVD21, NIN20, RBT^{+}22, SB20, SU21, ANR^{+}23, BKIR23, BLXL24, BVJ^{+}22, CsSL22, Esk21, FLXY22, MS22, NS23, SVK22, SU23b, SY22, TD23, VVNNT23, YYZ23, ZHJ23]. EGARCH [SB22]. eigenvalue [Ava21, WS21b]. eigenvalues [iTS23]. elastic [AJIAA22, LYY21, Pol22]. elastic-net [LYY21]. election [SBB24]. electricity [SHN23, WC23a]. elementary [FLXY22]. elements [VPT21]. elicitation [SK22a]. elliptical [RY22]. elongated [Tel22]. emerging [TK22a]. emissions [LWW22]. emphasis [HBM^{+}20]. Empirical [Che21, HHH23, LK21b, LLL22, LY20, SN23, WS22, AJIAA22, AJI24, AFJ23, Bia22, CWLH23, CZ22, Esk21, Gho20, HS23, KF24, LWW22, MQ21, ÖA22, SKS21, TY21, UY23, WY23, WN21b, YYZ23, YWT22, ZZWH20, ZWW23, ZLG23, dCdAB23]. Empirical-likelihood-based [SN23]. Employing [TPC^{+}23]. Encompassing [TD23]. energy [CLL23, EHF23, YC22]. enhance [DP23]. enhanced [HU20]. Enhancing [Ver21, AJI24]. Ensemble [IAG21, LYP22, DKKA22, DK22b, IGA24, ÖDS24, ZCH23]. entrance [XXH22]. entropies [NYM20]. Entropy [MYS22, AFJ23, CP23, CZP21, CQ22, Mac24, Mis20, RJJ23, SSSR23]. envelope [YV20, ZHJ23]. environmental [dALFB21, LWS23]. epistemic [BCDA21]. equal [GWW^{+}23]. equal-tailed [GWW^{+}23]. equality [AG21, MJ20, PFM21, YSH23]. equation [Dog22, FW23, MZZF21, dSC23]. equations [AK23c, FAT23, LX22, LYZ^{+}21, REMO22, SLJ23, Wes20]. equilibrium [TEM20]. equivalence [DC20]. ergodicity [TMZL21]. ERNN [SI23b]. error [CR21, CYS21, CL23, DC20, EH22, EL21, GZ22c, Gha22, GB22, Han21, Hep21a, Hor22, HZY22, IRS20, JuAH23, LAZ21, MRB23, MSS20, MGS23, uARS22, uAJHD22, PK22, PB23, SCE^{+}20, Sha20, Squ22, TL23a, WX21, WN21a, WCF21, YA21, YRAB22, ZNLW20, ZGC21, ZLF23, ZZ22b]. error-in-variable [HZY22, WCF21]. error-in-variables [WX21]. errors [AE23b, BKIR23, DCY22, FZY24, GZ22b, GZ22a, GWZC22, GZZ22, GZZ24, KZG21, LY23, MAAA21, Meg23, MKSK23, MS20, PZ20, RGA21, Sha20, SVK22, SBB22, SBB23, WS20, WLY22, YDLX20, ZG21, ZFG22, ZGL22, ZL23b, ZYF24, Zha24, ZZ22b]. errors-in-variables [WS20]. Esseen [DYC22, KL22]. estimate [CdCG22, Gho20, SBB22, Tak23, ZBB23]. estimated [CKT^{+}20, CTK^{+}22, LK21a, MMMCS22b, SAF21, YHGA21]. estimates [DP23, Gho20, KAVR20, SRN^{+}22, Squ22, WN23]. Estimating [Bee22, Bis22, CHL24, CLZ23b, GBHA21, LTR21, LD23, MG24, TK22b, TY22, WGlTL23, YDHW21, YW24, AK23c, BLXL24, CWWT24, FTC21, FAT23, KLM^{+}22, eLhS23, LYZ^{+}21, NS23, Ras20, REMO22, SLJ23, TPK22, WKSeF22, Wes20, Wu22, ZZL23]. Estimation [AOA23, ALB22, AHB23, BKPS23, CWLH23, DWLS23, ESLK22, HPZ^{+}21, HUK23, HZY22, JHR21, KSAR21, KB21, LES21, Lon23, MRM22, MK21, MES21, MBBS22, NS20, NDWE24, PKA22, PZ20, RSM22, RHG22, SEMR23, SU22, SA22, Sha22a, Sha20, SKV20b, SKJ20, VLvdW22, WWYX21, YSG23, ZNLW20, ZGC21, ZL23b, ZYF24, ZXF22, AG20, ASZN21, AA23a, AS21a, AS22c, AAM23, AK23b, AC23b, ASA24, Ava21, AM22b, BYK24, BA22a, BB24a, BS22a, BS23b, BS22b, BKIR23, BPP20, BMG22, BRA20, BZ22b, CCBJ21, CLZ23a, CL23, CSX20, CS22b, CLLX22, CLJK21, CZS21, DRB20, De 20, DC24, DDK21, DNWG22, EAT22, ESES20, EN22, EE23, ENK^{+}22, Esk21, FHM22, FY21, FZY24, FLPB20, GNE21, GY23, GJQ23, GABJ23, GCF^{+}20, HD20, Han22, HMV^{+}22, HZ22, HGWZ22, HHY21, Hua22, dAIIS20, IC21, KZG21, KUA23, KN22]. estimation [Kol24, KBYS22, KVS22, KP21, KA22, LDA23, LPNP20, LJC^{+}20, LK21b, LW21, LYY21, LWZ^{+}23, LF24, LMP20, LY23, LWP24, LRF20, Lu23, MG21, MZ22b, MRB23, MGS22, MMdOC23, MZZF21, Mil21, MEA24, MRM23, MR22, Mos22, MFPP21, MSNN^{+}24, MYVD21, MSB23, MHW24, NAKK20, NBC23, NIN20, Nis21, ÖT20, ÖA22, PM23, PMM23, Pan20a, PV22, PZ24, PMK23a, QL21, RJJ23, RAA23, RAA22, RBN^{+}22, RY22, RVV22, SB20, SSGH22, SP23, SZ20, SCE^{+}20, SJ24, SIG23, SW23, SD22, SKV20a, SK22c, SPS22, SSK22, Sol21, SMMO22, SSSR23, SCK20, TKF^{+}20, TIK24, TYLX22, TF22a, Tel22, TÇA22, TH23, WY21, WZ21b, WZZY23, Wan23b, War23, WQW^{+}22, WMXL24, XS23, XWL22, XS22, YBC22, YZ23, YXYX22, YCD23, YRAB22, YWLD20, YTL22, ZA23, ZG22b, ZHJ23, ZCW23, Zha23, ZS23, ZWTY23, ZZ22b, ZLTP23, ZKW24]. estimation [dSLHMB23, DA21]. estimations [Han20, Han21, NYM20, SAK24, fSLC21]. estimator [AÖ20a, AA22a, Aki24, AAA22a, AAK^{+}24, AA20a, AK21, AQYA22, AAA22b, AAM23, AA22c, BZ22a, Bek23, Bou23, Bul23, CWW23, CD23, CW23, Daw23, GDWY22, Gha22, GB22, Hut22, IS20, IMR21, Kal22, KFM23, KGF20, LA23, Lin22, LW22a, LWC22, LD23, LAKA22, LAK24, MKK22, MA23a, MA23b, OY23, PN21, RYA20, Sal23, Som22, SCK23, SCK21, SCA23, TEM20, UOA23, VW21, VG22, Wan23a, YA21, YW20, ZK23]. Estimators [LFMMRH22, AF22a, AT23, AAF23, AKS^{+}21, AM21, AQAN22, ACP21, AMK21b, AA22b, ABB22, AA22c, ASK23, BS23a, BZ22b, CK20, ÇKK23, CCM20, DYC22, DCL^{+}24, DPS22, ESS22, FQKA23, GLT20, HZAF22, Hua20, Hut21, JM20, Kal21, KK22a, KR23, Kur21c, LXMT22, LFZ21, MAAA23, MB24, MAB23, MM20, MW20, MKS22, ÖK21, ÖÇ23, PKW22, PS23, QTH24, Ruz22, SNSE22, SZ21, SANH^{+}22, SAA^{+}23, SVM20, SU21, SK22c, SVK22, SBB22, SBB23, SU23b, Squ22, TdSCdSdCP22, Tei22, TBK23, TL23b, WLC^{+}22, WKSeF22, WDZ22]. European [SBR^{+}23]. evaluate [DSO^{+}23]. Evaluating [AE23a, PM23, PG22, DC20, Hom20]. Evaluation [AK22a, HBM^{+}20, TdSCdSdCP22, UEA20, CdCG22, CII22, Han21, KR21, LPP21, OBN22, SMWJ20, ZL20, ZYSS22]. Evaluations [MW20, CL21, SN22b]. evasion [ZfSRJ22]. event [AÇ23a, FJ24, HHOC23, HCLF23, LLLY23b, ONE22, RHG22, VBB23]. events [AK20, BDD^{+}22, DuA23, HQZ^{+}23, PRH21]. evidence [BT20, Kur21c, SR23]. evolution [Jhw24]. EWMA [ASM22, CKL24, CTK^{+}22, HEK22, HCTZ21, HQZ^{+}23, JuAH23, KNSA23, LKHC23, NN21, NKSY22, uAJHD22, uAAF23, PAS23, SAA21, SNKA24, TAS24, TAA22, WCRM23]. EWMA- [HEK22]. Exact [ABC^{+}23, GHW20, GWW^{+}23, LBL22, SN22a, ZBZS20, ZB24, AS23, GM22, Nag21a, SAF21]. examination [CsSL22]. example [BGM20, MMR22]. exceedance [Ere20]. excess [ALMA23]. exciting [BS23b, LZQ23]. existence [YÖT20]. exit [GHW20]. exp [AAHR20]. exp- [AAHR20]. expansions [BCDA21]. expectation [ESLK22, GWW^{+}23, SSCT23, dXLY20a]. expectation-maximization [SSCT23]. expectations [ZL22b]. expected [BRTW24, BDG23, Han21, SHN23]. expected-based [SHN23]. expectile [PLS23, SHN23]. experiences [VJV^{+}23]. experiment [PMM23, ZBZS20]. experimental [LFMMRH22, MMR22, UIC23, VJV^{+}23]. experimentation [BVJ^{+}22]. experiments [Els24, KB23, LFPL24, LAZ21, LWX23, RCSS22, SPG^{+}22, SCHS22]. experimentwise [LAZ21]. experts [YWLD20]. exploitation [SN22b]. exploratory [FBNRG21]. Exponential [BLR23, LL23, LRF20, AK23a, ASZN21, AHAMP23, AZAZ23, AR22, AK22c, AE23c, AE21, BKG22, CLY21, CTCL22, CBRL22, FK22, GKS24, GJK21, HHOC23, Jod23, KR21, KKH17, MWTR23, MRM22, MJ20, MKS22, PV22, PMK23a, PMK20, PMK23b, POCC22, RL21, RK20, SBTP20, TAS24, TY24, WL22, WKSeF22, Wu21, YCD23, YSG23, ZB22, MBBS22]. exponential-geometric [AZAZ23]. Exponential-Poisson [LRF20]. exponentiality [GZZ24, UU23]. Exponentially [uARS22, AK20, CTK^{+}22, DuA23, HZZ^{+}24, uASS22, QTH24, WHS23]. exponentiated [POCC22, RBRAa21, Ras20, WG23]. expression [HYSC21]. expressions [LXMT22]. Extended [AAHR20, Abd18, ASP22, CL22a, MRM22, POC23, RCSS22, Sha23, VPOC24, YCD23]. extension [SAK24, SDG22]. extensions [MdN22]. extracted [dSC23]. extreme [AOA23, BGM20, BRTW24, BE20, CW23, KGR21, Sun23, WN23]. extremely [ZQY22]. extremes [MdN22, YCD23]. Extropy [NJ22, SJ23, LX24]. eye [OP20]. Eyraud [dXLY20a, dXGLY20].

Facebook [TPC^{+}23]. facing [FO22]. factor [BH23a, Bia22, CSG22, LR23, MF22, MPS21, PM23]. factorial [BVJV22, LAZ21]. factors [ML22]. failure [BMG22, CSL20, Han22, KCS24, KS23d, NG22, RPT23, SLME23, SVN24, WCC22, XZS22, YZ23, ZWW23]. false [BR21, TF22b]. families [Bar20a, BAEk21]. family [AMA22, AMD22, AAHR20, BNN22, BFAR20, BPN22, BKM23, CKK21, CCBJ21, HPZ^{+}21, LL23, LXMT22, MG21, MCWL22, MZZA24, MB21, MKS22, PLBP24, SK22b, SAA^{+}23]. familywise [DC20, Hor22]. Farlie [dXLY20a, dXGLY20]. fast [Els24, LYP22, LW22a]. Fay [AC23b, MGS23]. Fay-Herriot [AC23b]. FDR [Wu22]. Feature [AM24, CGS21, CLC22, DS24, Fan22, IGA24, SX22, SLSZ22, ZCH23]. features [TWTT20, Zhe24]. featuring [WKcIC24]. FGM [Bek24]. fiber [HORA22]. Fiducial [WZY^{+}21, CP22, YDLX20]. field [KS23b]. fields [HI22, Lev21, Wan23b]. filter [Lev21]. Filtering [DL20]. filtration [ZL22b]. filtration-consistent [ZL22b]. financial [AM21, CDB^{+}21, Fuk22, NN22, WS21a, WS24, ZT20]. finders [Gan22]. finding [Tel22]. findings [ÖT20]. Finite [SCK23, TJ20, BRA20, DKL22, HUK23, KK22b, LY20, PKW22, QL21, REMO22, SNSE22, SS21b, Wan23a, WZW23, WQW^{+}22]. finite-sample [PKW22, Wan23a]. finite-state [SS21b]. fires [TBH22]. First [LJ22b, LZ22, BMG22, ÇKK23, Che22, FSGMM23, SSZ23, Sol21, TBH22]. first-failure [BMG22]. First-order [LZ22, ÇKK23, FSGMM23, SSZ23, Sol21]. Fisher [Fig23, HR22, JD23a, WCY24, ZMNL23]. Fit [LK21b, AS22a, ABHF22, BH20, CSA23, CMA21, Dog22, FM23, FS22, Gai21, KGR21, MG21, MQ21, Mil20, NM21, OP22, PHY20, Pak23, Pho24, Sul22a, Sul22b, WYY22, YL22b]. fitted [PAS23]. fitting [AS22d, LK21b, MW20, Ors24]. fixed [CHC20, Hua20, SN22a, ZBB23]. flexible [BCR^{+}21, LXMT22, ROL^{+}20, WHW21]. Focused [HCZ21]. following [TF22a]. forecast [AT21, Bey21, CL21, RM20]. Forecasting [AKR23, De 20, SI23b, EGVdR21, IRA21, TPC^{+}23, WZ21b]. forecasts [XZJ^{+}21]. Forest [SLW20, Cha23a, MNL21]. forests [CGS21, WSZ23]. form [Lev21, LWC22, dALNPdOF21, YZ23]. formula [Bar20a, HCHK23, SZJG21, YW24]. formulas [DL20]. Formulation [SBB22, RBN^{+}22]. four [WL20]. Fourier [dAIIS20, Zhe24]. fractal [Kol24]. fraction [HG22]. fractional [DPS22, LPP21, MF22, ONS22, SB24, XWZL22, XW20]. frailty [CSBdES22, CsSL22, KKK23, LSTE24, LBH23, PBLM22, PL22, fSLC21, ULN21, ZGC21]. frailty-copula [LSTE24]. frame [LFZ21]. framework [ARAR24, Kaz23, WHW21]. free [CLC22, DP22, FM23, Ham22, MMMCS22a, SLSZ22, TK22a, TP22a, XS22]. freedom [SF23]. frequency [AM21, CL21, STH24]. frontier [RRA20]. Full [LE23]. Function [NCC^{+}22, AS23, AOA23, AA22b, BRTW24, BCP21, BS22b, CLZ23a, CHC20, CZP21, DP23, Gho20, Hua23, Hut24, KSAR21, KE22, LDA23, Lav21, LY24, LY20, MYVD21, Nad21, PKA22, PZ20, RJJ23, Sal23, SS21a, SA22, SIG23, SD22, SSSR23, WW23, ZNLW20, ZLG23, dOMD^{+}21]. function-based [ZLG23]. function-on-function [BS22b]. Functional [BSAS22, Cha23a, WZY22, ZHW21, BJA22, BV22, CCH20, GP20, HGWZ22, KS20b, Mos22, NSN21, PB22, Sha20, YZ21, dSRC22]. functions [AMK21b, AA20d, Bar20a, DSL20, Han20, Han21, KEE20, LWP24, LX23, LD23, Ors24, Pic21, Pol20, RAA23, Sul23, WN21b, YDHW21, YGL22, YTL22]. further [ÖK20]. fused [HGWZ22]. future [BKG22, CM22, MMMCS23]. fuzzy [KEE20, Tür20]. gains [AM21]. galaxies [BCC24]. Galois [KS23b]. gambler [HCHK23]. games [LX22, SZS20, TGS22]. Gamma [MMdOC23, NAG21b, AAF23, AK22a, AA20a, AQYA22, BJU20, ÇKK23, CsSL22, Hu23, KLK21, Kur21c, LAKA22, LAK24, NBC23, OP22, PKA22, Ras20, RLP23, SAA21, SJ24, Som22, VG22, WZY^{+}21, YH21a, ZK20a, TIK24]. gamma-frailty [CsSL22]. Gamma-Poisson [TIK24]. Gamma-Uniform [MMdOC23]. gap [SL21]. GARCH [ALMA23, Alt20a, Bou23, De 20, DDK21, IT21, TKF^{+}20]. GARCH-TSLx [Alt20a]. GARCH-type [ALMA23]. Gaussian [AA23a, AAA22a, AAK^{+}24, AQAN22, BRTW24, Bia22, Bul23, CSDT20, DC23, DE22, GS23, GSG22, HD20, Lev21, LZQ23, Nag21a, RSG22, RS24, SU23a, TGS22, WCF21, ZBA23, ZfSRJ22]. GEE [XWL22]. gene [HPZ^{+}21, HYSC21, LXMT22]. General [ARA23, WN21a, ZZ22b, AÖ20a, Baz20, BKIR23, CLY23, FM23, Kaz23, LK24, LHG24, MG21, MSB23, SM23a, SBB23, TEM20, TBK23, Xia22, YZ23, Yao23, ZG21, ZL23b]. generalised [BJ20]. Generalization [Alt20b, GP20, AE23c, JKH20, KKH17, dOMA21]. generalizations [Jod23]. generalize [SRN^{+}22]. Generalized [AHAMP23, Bek24, BE20, CLZ23a, DK22a, FQKA23, GNE21, LBH23, LYZ^{+}21, Pan20a, PD22, SPS22, YDLX20, AAF23, AGI23, AK22c, ARJ21, AAGP23, AK23c, ARJ20, ABJ21a, AE23c, BAEk21, BS23a, BFAR20, CKK21, DA21, FK22, FHM22, FF23, FZM21, FAT23, GKM20, GM22, HG22, HK20, Hut24, KLK21, KKH17, KT23, KH23, LH20, LWW22, MK21, MKS22, NS22, Özk21, OP22, PM20, PV22, PMK23a, PMK23b, RAJ21, REMM22, RY22, REMO22, RCSS22, RK20, SSGH22, SBTP20, Seo22, Sha23, SI23a, SW23, SF23, SLJ23, TAS24, TY21, TNN20, VLvdW22, WLT22, WL22, WHS23, WKSeF22, WQW^{+}22, Wes20, XS23, YYZ23, YSG23, ZK23, ZWY20, ZD22, ZS23, ZWC^{+}23, ZB24]. generally [ACKL23, CLS22]. Generating [NCC^{+}22, EHF23, Gho20, MS22, Min23]. Generation [KTZ20, BVJV22, DG22, DG23b, Jod23, VWZ21]. generators [Xia22]. genes [CII22]. Genetic [ES22, BFS22, NS22, PHN22, TY22, VPT21]. Geometric [Alt20b, AZAZ23, Biç20, GMC23, LB22, NAKK20, PAT22, PIA22, RL21, YCD23, dOMA21]. geostatistical [DRB20, HD20]. GEV [NHZ22]. Gibbs [AA20c, Min22, OKO21]. Gini [MA23b, Pak23, SKS21]. given [Aki24]. glass [HORA22]. glaucoma [OP20]. GLD [AA23a]. global [ES22, FLXY22]. GLS [GG22]. GM [DA21, CSX20, CS22b]. GMDH [DKKA22]. GMDH-type [DKKA22]. GMWB [OO21]. goal [AES21]. Gompertz [CL22a, Abd18, AAGP23, GABG22, HG22, HHOC23, LTB23, LHG24]. good [EFD21]. Goodness [ABHF22, CMA21, KGR21, NM21, OP22, PHY20, Pak23, Pho24, AS22a, CSA23, FM23, FS22, Gai21, LK21b, MG21, MQ21, Mil20, Sul22a, Sul22b, WYY22, YL22b]. Goodness-of-fit [CMA21, KGR21, NM21, Pak23, LK21b, CSA23, FM23, FS22, Gai21, Mil20, Sul22a, Sul22b, YL22b]. GPAbin [NSlRGL23]. Gradient [Yas22, DKL22, LYP22, PR22, ZQY22]. gradually [KFB23]. graph [ESLK22, HvZ21]. Graphical [HHvR22, DC23, Kur21c]. graphs [HPZ^{+}21]. gray [CS22b, CLLX22, CLL23, WW23, XWZL22]. grey [CSX20]. gross [WLY22]. Group [SX22, CKT^{+}20, GK22a, GK22b, KJB22, MKSK23, SKM^{+}23, ULN21, WKT22, War23]. grouped [DP20]. groups [CM22, MAF23]. groupTesting [War23]. Groupwise [ZHJ23]. growth [HHvR22, KBYS22, LWW22, MD22]. growth-application [LWW22]. GSTAR [IRA21]. GSTAR-SUR [IRA21]. GSTAR-SUR/neural [IRA21]. GTL [VH23]. Guaranteed [HCTZ21]. Gumbel [dXLY20a, dXGLY20]. GWMA [HU20, MMMCS22a]. GWMA-CUSUM [MMMCS22a].

Haenszel [HMV^{+}22]. Half [SMB22, LB22, RBRAa21]. Half-spectral [SMB22]. Handling [PRH21, VBB23, BYS24]. Hanurav [Aub23]. Harrell [Aki24]. having [DSL20, SKG22]. hazard [AMK21b, BNN22, BD23, CLZ23a, PM20, Sal23, SK22b, SKG22, SR23]. hazards [BBM24, CMA21, FS23, PSL23, SW23, WFLS24, ZCG24, ZWW23]. head [OP20]. health [SCE^{+}20]. healthcare [RAN23]. heart [SVN24]. heavy [AMD22, GI20, GO22, QYC22, dXLY20a, dXGLY20]. heavy-tailed [GI20]. hedonic [DE22]. Heligman [SRO20]. heritability [VLvdW22]. Herriot [AC23b, MGS23]. Heston [MFS21]. heterogeneity [Har21]. heterogeneous [LWZ^{+}23, Nad21]. heteroscedastic [BKIR23, DCY22, MAAA21, TL23a, YH21b, ZA23]. heteroscedasticity [AGI23, ARAR24, Uya22, WWW24, Wu21, WCF21, ZGL21]. heteroskedastic [AS21b]. heteroskedasticity [AS23, ZLF23]. HEWMA [JuAH20]. Hidden [EN22, Gho23, HMSRM23, HW24]. Hierarchical [AY21a, AC23b, CFM21, CQ22, REMM22, ZLXT23]. High [MT21, AM21, AALR23, ABC^{+}23, AM24, Bul24, CZP21, CL21, CKCG21, Ega24, FY21, FS22, GY23, IAG21, JM20, Kur21a, Kur21b, Kur23, LYY21, LXMT22, LWX23, MX24, MCJ22, Nak21, Nak22, RSS22, STH24, SMB23, TCJ23, VLvdW22, WX21, WKT22, WCF21, XS22, ZCH23]. high-accuracy [LWX23]. high-dimension [CZP21]. High-dimensional [MT21, CKCG21, FY21, GY23, IAG21, JM20, Kur23, LYY21, LXMT22, MX24, Nak21, Nak22, SMB23, TCJ23, VLvdW22, WKT22, XS22, ZCH23]. high-frequency [AM21, CL21]. high-quality [RSS22]. Higher [MVLMA22, SN22b, Bap20, BPP20, Che22, ZLF23]. Higher-order [MVLMA22, ZLF23]. highest [Aki24]. highly [KS20a]. Hill [GCF^{+}20]. histogram [DC24]. histograms [DP20]. hitting [LJ22b]. HIV [SKG22]. Hjorth [YBC22]. Hollander [GM22]. homogeneity [ATYK20, GZY22, GKS24, Gök22, GSG22, LSC23, MX24, PD21]. homogeneous [XWZL22, YWL^{+}23]. Homogeneously [uAA23]. homoscedasticity [FO22]. Horvitz [IMR21]. Hoteling [ENA^{+}23]. Hotelling [MLR^{+}22, SSBA22]. house [DE22]. household [RH23]. HPD [LPNP20]. https [Kri22]. Hudson [AE23c]. human [VPOC24]. humidity [AMP23, KKO21]. hybrid [AKE23, CSL20, CLY21, DWLS23, IRA21, KT23, KB21, KA23, Lon23, MBBS22, NG22, uASS22, PMK23b, RBRAa21, RH23, SNSE22, SBTP20, Seo22, SI23b, WLT22, WL22, YSG23, ZB24, ARAB21]. hyper [KR22]. hyper-Poisson [KR22]. hyperbolic [ARJ21, GKM20, XS23]. hypercubes [EHF23]. hyperkinetic [IM20]. hypotheses [Bis22, Wij23, ZK20b]. Hypothesis [ZD22, GZZ24, NdSA20, PD21, SN23, TA24, YV20].

I/II [Yad22]. IBNR [TK22b]. ICBayes [PCWL22]. ideas [MMMCS23]. identically [EHF23]. identifiability [RBN^{+}22]. Identification [AZH^{+}22, SI20, Sun23, BH23a, CQ22, KB23, LMP20, TEM20]. Identifying [NAJ22, PY20, Wij23, ZT20]. ignorable [RSG21]. II [ASS22, AZAZ23, AMK21b, AB22, BR24, CKK21, CW23, DSS23, FAT23, GAS20, Gho23, GK23, HA24, KA24, KB21, KG22, KP21, LTR21, LS23, LHG24, MG21, MWTR23, MK21, MJ20, NDWE24, Pak23, PV20a, PMK23b, REN22, SBTP20, Seo22, Sha22b, TRF^{+}22, WG23, Yad22, YBC22, YHGA21, YSG23]. illness [KKK23]. illness-death [KKK23]. illustration [PR22]. image [Lev21]. images [CMM23, JH23, OP20]. imbalanced [CY24, Nak21, Nak22, Wan23c]. Impact [SCE^{+}20, UKK20, CSE^{+}21, GZY22, SBZG22]. impatient [LJ22a]. imperfect [BQCP23, KFB23]. implementation [Aub23, CKCG21, JPH21]. implied [BH20]. importance [NT20]. improve [SPG^{+}22]. Improved [CS22b, MAB23, MM20, MMdOC23, PDO^{+}20, RuA22, AAK^{+}24, BZ22b, Daw23, KVS22, MEA24]. Improvement [LPNP20, CSX20, ZK20a]. Improving [AJIAA22, CL22b, LW21, PY20, SN22b, Lev21, XWL22]. Imputation [SANH^{+}22, ASK23, BPP20, FH22, KS22a, LK24, SKK21, SSS23, Tak23]. imputations [NSlRGL23, SSCT23]. in-play [SS21b]. In}-type [CK20]. inadequate [CSE^{+}21]. INAR [WZ21b, BS23b, AKR23, HZ23, Lar20, LS22a, LZ22, SB20, SM23b, SM20]. incentives [WMJ21]. incidence [CHC20]. including [BBD^{+}24]. incomplete [HLT23, KJK20, KB23, MPD20, YSH23, Yan23]. Inconsistency [KM20]. incorporating [KCS24]. incorrect [CNL21]. increase [GAS22]. independence [Bul24, Fan22, HS23, KF24, KuIK20, MAF23, MEA24, YL22b]. independent [ACP21, Gho20, Kri22, KLM24, Min23, SPV22]. index [AA20b, ABJ21b, ARKA22, Bal22, BT20, CS20, El 24, GZ22a, GZ22c, Hwa23, JKLK24, JGL22, Kur23, MA23b, Nad21, NN21, PDO^{+}20, RAJ21, SJ24, Sha20, SCK23, VPOC24, WC23b, WCKC23, ZNLW20, ZGL21, ZHW21, ZFG22, ZXF22, ZXZL22, dSC23, Pak23]. Indian [SR23]. indicator [Özk21]. indicators [BH23a, MMR22]. indices [ANP20, OP20, RSM22, Sol21, SKS21]. indirect [Bee22]. individual [HBS22, LAZ21, SLME23]. individual-patient [SLME23]. individuals [JAP21]. induced [CSBdES22]. industry [YC22]. inequality [RSM22]. infant [SKG22, SR23]. infected [JAP21]. Inference [AB22, ASP22, BS23a, HR22, Hut21, KFB23, LS23, MJ20, RBRAa21, SBTP20, SJ24, Sha22b, WX21, WLT22, WL22, ZLF23, AZAZ23, ASA23, Alt20a, AAGP23, AHG20, BC21, BC23a, Bee22, Biç20, CHC20, DCY22, FH22, FK22, FP20, GABG22, Gho23, GK23, HK20, HLT23, IT21, KS23a, KJ20, LMFMA23, LW22b, LHG24, MWTR23, MYS22, MdN22, MB21, PD22, PSL23, PIA22, SB21a, SCE^{+}20, Sha23, SKS21, SF23, SLE20, WZY^{+}21, WG23, YDLX20, YGL22, YTL22, ZCG24, Zha22a, ZB20, ZWC^{+}23, ZBZS20, ZB24]. Inferences [AS22b, DSS23, KG22, MCWL22, SRO20, BR24]. inferential [LBH23]. inferiority [CP22, HHOC23, WH24]. inferring [DC23]. infinite [Che21, GTW22]. inflated [Ali24, ASM22, BLR23, CFM21, GZY22, HA24, KR22, Pho24, RSG21, RLP23, RSG22, RS24, RC23, SZ20, SN23, YL22a, ZBB23, ZQY22]. inflation [LTN23]. Influence [LAK24, MRB23, AÇ23a, LS22b, YW24]. Influential [KUA21, MAMA20, AZH^{+}22, KB23, SI20]. information [AS21a, AS22c, BSNT21, EE23, HB22b, HR22, HCZ21, HHY21, IMR21, JuAH20, Kim21, KK22a, NKSY22, uAJHD22, uAA23, Pol22, SNKA24, Seo22, SVM20, SVK22, TPK22, WCY24, XH21, ZX22a]. information-weighted [HHY21]. informative [Wan22]. INGARCH [PZ24]. inhomogeneous [LD23]. initial [FPAA21]. INLA [Ver21]. innovation [SM20, WMJ21, YH21b]. innovations [GKM20, HZ23, PB23, SSZ23, SKJ20, TKF^{+}20]. inputs [SLW20]. inspection [AMA^{+}23, Bal22, HB22a, RGA21, WZ21a]. inspections [BQCP23]. instrumental [ZZWH20, ZWTY23]. insurance [AMA22, AMD22, BNN22, BH23b, BS21, DA22, WDZ22]. integer [AES21, BA22c, LL23, RBKA23, SSZ23, WWYX21, OB21]. integer-valued [BA22c, LL23, SSZ23, WWYX21, OB21]. integral [TK23]. integrated [PM20, SNA20, YZF22]. integrating [LWX23]. integration [Bap20, EL21, XW20]. intensities [FSGMM23]. intensity [BC23a, CMM23, LD23, NG22, fSLC21]. intensive [dCSF24]. inter [TDD20, WC23a]. inter-rater [TDD20]. inter-relationships [WC23a]. interactive [LE23]. interarrival [LJ22b, PIA22]. intercept [AJ22]. interest [GHW20, LG21, MMVBE23, SM24, XZJ^{+}21, ZS23]. interesting [EFD21]. interpolation [HI22, WZZY23]. interpretable [SRF^{+}23]. interquantile [ACP21]. interquartile [SA22]. Intersection [AÖ20b]. Interval [ACP21, Aki24, AHAMP23, AHG20, BSAS22, CST21, CsSL22, CTCL22, FM23, GWZ23, HB22a, KM22, KKK23, LKHC23, LLV23, LPP21, MZ22b, MAMA20, MB21, NBC23, PC22, PD21, SZ20, SJ24, SI23a, SB21b, Tei22, VVNNT23, WCC22, Wil24, WC23b, WCKC23]. interval-censored [AHAMP23, CsSL22, PC22, WCC22]. interval-valued [BSAS22]. intervals [BCP21, CNL21, Cur23, DG23a, DP20, GWW^{+}23, IRS20, IR21, Kri22, LPNP20, MNN21, MCJ22, MKSH22, MMVBE23, PBC23, PDO^{+}20, QLG23, RLP23, RM23, SPV22, SYPV23, TNN20, UYMK22, WJB21, Xia22, ZX22a]. intervel [PCWL22]. intervel-censored [PCWL22]. intervened [KS20c]. interweaving [eLhS23]. intra [ATYK20]. intra-class [ATYK20]. intraclass [AE23b, PLBP24, Sha22a]. introduced [BBD^{+}24]. Intuitionistic [KEE20]. invariant [LHY^{+}23]. inventory [MNL21]. Inverse [Bul23, AAA22a, AAK^{+}24, AK22c, AQAN22, AA20d, CSDT20, GSG22, SU23a, TGS22, TRF^{+}22, WL20, MCJ22]. inverted [CKK21, MWTR23, YSG23]. Investigation [BCC24, GZ21a, PKW22, Kur21c, TA24]. investment [BH23b]. involving [BDG23]. Iranian [SMB22]. irregularly [Mos22]. Isolating [Wu22]. isomorphic [KFEL23]. issue [KS22a, SK22c]. Itô [SF24]. Italian [Bia22]. Item [Brz20, SBR^{+}23, BH23a, Mag20, TPK22, UM22, ZLXT23, ZZL23, dSLHMB23]. Iterative [SLP22, MW20, RH23]. Iteratively [Zha22b, Zhe21].

Jackknife [DCL^{+}24, SKS21, WY21, WLY22, UOA23, WY23]. James [AAA22b]. Jayakumar [Ano24]. Jeffreys [MK22]. Jensen [LK21b]. Jensen-Shannon [LK21b]. jmcm [YP23]. Johnson [MM20, PY20]. joining [Ors24]. Joint [MX24, uAAF23, WCC22, AKE23, AS22a, AB22, BR24, DSS23, GY23, GGB20, GK23, HPZ^{+}21, JuAH23, KKO21, KG22, LSTE24, RSG22, RS24, SMG23, YLX^{+}22, YP23]. jointly [VBB23]. Jones [HM21]. jump [LLLL22, LLLY23a, WZX23]. jump-diffusion [WZX23]. jumps [MFS21, WS21a, WS24, XS23, ZZ22a].

kaphom [ATYK20]. kappa [ATYK20]. Kenward [CLL20, SF23]. Kernel [EAT22, ZCW23, Ali22, AHB23, BZ22a, CLZ23a, FM23, GJQ23, HZAF22, PK21, Pol22, RJJ23, Sal23, Som22, WY23]. kernel-smoothed [FM23]. kernels [MZZA24]. Khan [AE23c]. Kibria [UOA23]. kinds [SBB23]. King [AE23c]. kink [MYS22]. knockoffs [GPGMK22]. known [AZH^{+}22, ASK23, CTK^{+}22, RSM22]. Kolmogorov [BK23, YZF22, ZZFG20]. kriging [ASM21]. Krylov [EJKL23]. Kumaraswamy [KB21, dALFB21, WLT22]. Kurtosis [ALMA23, BDG23, EHHS20, PFM21].

L [AE23c, Kri22]. label [CY24, HvZ21]. lactation [CdCG22]. lag [FPAA21, MAAA21, MA23a, ÖT20, SKG22]. Lambert [NCC^{+}22]. landmark [SCK22]. language [Pol22]. Laplace [Bou23, LBL22, ZBZS20, ZB24]. Laplacian [HvZ21, LH20]. Large [HW24, Bil23, CL22c, HB23, HvZ21, PV20b, STH24, YHWW23, ZC22, ZMNL23]. large-sample [PV20b]. Large-scale [HW24]. larger [SRN^{+}22]. largest [ABC^{+}23, CTCL22]. Lasso [Gan22, AA20c, AWZ24, FF23, HPZ^{+}21, WKT22, AJI24, AM22a, EGVdR21, ULN21, WJZ^{+}22]. LASSO/adaLASSO [EGVdR21]. late [SEMR23, Yad22]. late-onset [Yad22]. latency [RdAIS^{+}20]. latent [CHL22a, HK20, ZLXT23]. lattice [EFD21, MWNSB23]. law [CS22c, Kaz23, LLLY23b]. Laws [YHWW23]. lead [ZHJ22]. learning [DY22b, GZ21b, KS23a, LYP22, ÖDS24, RCSS22, SPG^{+}22, WKcIC24, SVN24]. Least [MT23, Che22, DRB20, KLM^{+}22, SP23, WY21, WLY22, WZZY23, ZA23, ZG22b, Zha22b, Zhe21, ZZ22b, AK23c]. Least-squares [MT23]. left [Alh23, MAMA20, MB21]. left-censored [Alh23]. left-truncated [MAMA20]. Legendre [LJC^{+}20]. Lehman [GM24a]. Lehmann [TRF^{+}22]. length [AFJ23, BC21, CTK^{+}22, FPAA21, HCTZ21, PAS23, WCRM23]. length-biased [AFJ23]. Leone [AMK21b, BR24, LB23b]. level [AHB23, AY21b, BVJV22, CL22b, Els24, FHM22, FS22, LAZ21, NS20, VJV^{+}23, WLZ23]. levels [TK22b]. Levene [BDL23]. Levene-style [BDL23]. leverage [XS23]. Leveraging [SAD^{+}21]. Lévy [JHR21, LC22, OO21, SM24, STH24]. Lévy-driven [STH24]. liabilities [WZX23]. life [ABPK21, BJU20, CSL20, Çet23, DR22, GAS20, HB22a, KJB22, LCC22, NDWE24, PZ20, SJ23, SBTP20, SZJG21, WDZ22, ZLSC20]. lifetime [Abd18, Ami21, BKG22, CL22a, KS23d, MRM22, PAT22, SAK24, SJ24, SLME23, WC23b, WCKC23]. lifetimes [DSS23, Wu21]. Likelihood [GK23, TWTT20, WTN22, AFJ23, BA22a, Bou23, CLL20, CWLH23, Che21, CS22c, DDK21, DKL22, DPS22, GNE21, Gök22, HD20, HK20, HHH23, JGL22, KAVR20, LK21b, LWC22, LLL22, LX23, LY20, LWW22, MQ21, MM20, MMdOC23, MRB24, OKO21, ÖA22, PZ24, SZ21, SKS21, SN23, TKF^{+}20, TY21, Tei22, WY23, WS22, YYZ23, YH21b, ZGC21, ZZWH20, ZBZS20, ZB24, dCdAB23]. Likelihood-based [TWTT20, LLL22, SKS21]. Lilliefors [Sul22a]. limit [GABJ23, XXH22]. limitations [AW22]. limits [TAA22]. Lin [Mil20, PHY20]. Lin-Wong [PHY20]. Lindley [Biç20, WZ21b, DSS23, DWLS23, GABG22, GW22, GK23, KG22, MSNN^{+}24, NDWE24, PBLM22, RPT23, Sha22b, SSZ23, TW24]. Linear [CW23, FBL22, GAS22, YW20, ZG21, AG20, AÖ20a, AA22a, ASS22, AZH^{+}22, AES21, AK21, AALR23, ARAR24, ALVQ22, BJ20, BBM24, BSNT21, BSAS22, BKIR23, Bil23, BAL24, Büy23, CMM23, CL23, CHL24, CCM20, Ciu22, DLRZ21, DCL^{+}24, EGVdR21, FHM22, FY21, FF23, FP20, GI20, GZ22b, GZ22a, GLT20, Gha22, GB22, GKJ23, Gül22, HBS22, HK20, HCZ21, Hut22, IHS22, IRS20, KAVdSFG22, KA24, KSAR21, KN22, KJK20, KÖ21a, KÖ21b, LHY^{+}23, Lee20, LMP20, LY23, MAAA23, MRB23, MYVD21, Nag21a, Nis21, ÖA22, ÖK20, Özk21, PR22, PV22, PLA22, RAG21, RAA23, RBN^{+}22, Ric21, RCSS22, RVV22, RBKA23, REN22, SM23a, SA22, SFG21, SF23, SCK21, SCA23, TY21, TÇA22, TL23a, Tür20, UOA23, VLvdW22, VH23, WX21, WCY24, WCF21, YA21, YRAB22, YL22b, YZ21, ZA23]. linear [ZK23, ZNLW20, ZGL21, ZD22, ZRC23, ZHJ23, ZLF23, ZYF24, ZZWH20, ZXF22, ZX22a, ZWTY23]. linear-time-invariant [LHY^{+}23]. link [GPS23, Nad21]. Linnik [Ano24, JS19]. Lipschitz [MF22]. Lipschitzian [SF24]. liquid [GHW20]. literature [MMMCS23]. Liu [AÖ20a, AAK^{+}24, AA20a, AA22c, Bul23, Gha22, KÖ21b, ÖK20, PS23, VW21]. Liu-ridge-type [AA22c]. Liu-type [AA20a, Bul23]. local [AALR23, LS22b, MB23a, Nis21, SB24, Tak23]. Locally [JKLK24, GJQ23]. location [ANR^{+}23, AG21, Bek23, CTCL22, CW24, Ham22, HMRN22, Hut21, Hut22, Kal21, MCWL22, MMMCS22a, MMMG21, MB21, PKW22, RYA20, STdL21, WLY22, WZW23, YWL^{+}23, dCCM^{+}22]. locations [NAJ22, ZB22]. Log [MSL21, YCD23, CS22c, FS23, GGMM22, KJK20, Kri22, KP21, Kur21c, LJC^{+}20, MZZA24, MAMA20, MB21, PDK23, SPV22, Wan23b]. log-density [LJC^{+}20]. Log-extended [YCD23]. log-gamma [Kur21c]. log-likelihood [CS22c]. log-location-scale [MB21]. log-logistic [PDK23]. Log-normal [MSL21, Kri22, MAMA20, PDK23, SPV22]. log-periodogram [Wan23b]. log-ratio [GGMM22]. log-symmetric [MZZA24]. Logarithmic [SYPV23, ÖÇ23, SBB22, ZXW23]. Logistic [Cha23b, AÇ23a, EJKL22, EJKL23, IAG21, IGA24, IS20, LB22, Mil21, NN20, ONE22, OBN22, PDK23, PJL21, RBRAa21, dSS22, SAD^{+}21, UYMK22, dOVQP21, VW21, Wan23c, WKcIC24, ZZ21]. logistic-geometric [LB22]. logit [DSL20, FQKA23, VPOC24]. logit-normal [VPOC24]. Lognormal [Sal23, Hua22, MNN21, PY20, MV22b]. Lomax [AB22]. long [AGI23, AS21b, BS23a, FZZ20, FL23, ROL^{+}20, TPC^{+}23, Wan23b]. long-memory [AGI23]. long-term [FZZ20, ROL^{+}20]. Longitudinal [NSN21, ALVQ22, BYK24, BYS24, GI20, GZ21a, GGB20, HMSRM23, HHvR22, HCZ21, Lu23, RftADNI22, RSG21, RS24, SWC^{+}20, TY21, TPvdH23, TWTT20, TWTT21, TL23b, WCC22, XWL22, YP23, YOD^{+}23, ZZWH20]. look [ZLG23]. Loss [Zhe21, AMD22, ARKA22, CZP21, Han20, Han21, Hua23, RAA23, SIK24, WDZ22, dXLY20a, dXGLY20, dXLY20b, Zhe24]. losses [AMA22, BNN22]. lot [El 24]. lots [Bal22]. low [CZP21, Els24, LWX23]. low-accuracy [LWX23]. low-sample-size [CZP21]. lower [XSS22]. Lowering [XSS22]. lowest [BC23b]. luck [GKZ^{+}22]. Lukman [UOA23]. LVQ [LW21].

M [AE23c, BC21, BC21, DA21, JD23b]. M/M/1 [BC21]. M/PH/1 [JD23b]. Machine [RCSS22, AWZ24, DY22b, Ega24, KS23a, Nak21, Nak22, SPG^{+}22, SVN24]. machines [FN20, MT21]. magnitude [TAS24]. maintenance [KFB23, NG22, PM20, SNA20, SPG^{+}22]. Makeham [CL22a, Abd18]. making [MM22]. Malaysia [NN22]. Mallows [AA23b, SDG22]. Mann [MMMCS22a]. Mansson [ÖT20]. Mantel [HMV^{+}22]. manufacturer’s [MY23]. manufacturer’s/seller [MY23]. manufacturing [SKM^{+}23]. MAR [CLL20]. Mardia [EHHS20]. margin [ZC22]. Marginal [KÖ21a, Mil21, PZ24, SL21, AKR23, CS22a, NSN21, PRH21, YZ23, ZGL21]. marginalized [OKO21]. marginals [ADK21, ZWY20]. margins [BF23]. marked [LZQ23]. markers [ANP20]. market [CL21, LWS23, SHN23, WS24, YC22, ZT20]. markets [DPS22, WC23a]. Markov [BCP21, EN22, HMSRM23, HW24, HE21, HE24, KSP23, KBR21, KHK22, Lev21, LES21, OKO21, SS21b, SLE20, TMZL21, Yas22, ZC23]. Markov-modulated [ZC23]. Markov-switching [KHK22, OKO21]. Markov-type [Yas22]. Markovian [ALZ24, BC23a]. marrow [LBH23]. Marshall [HG22]. martingale [Lee20, LPP21]. masked [CSL20, JH23, LY24, RPT23]. Masking [SCHS22, KB23]. massive [Ciu22, JGL22, MA23b]. matched [CP22, WWTM22]. matched-pairs [CP22]. matches [BT20, SS21b]. matching [KLK21, OBN22, ZL22a]. maternal [SKG22]. matrices [MS22, MX24, SANH^{+}22, dSLHMB23, vO21]. matrix [Ava21, AM22b, Jhw24, RAG21, TCJ23, iTS23, WS21b]. matters [Lav21]. max [Hua23, JuAH20, JuAH23, uAAF23, TAS24]. Max-EWMA [JuAH23, uAAF23, TAS24]. Max-HEWMA [JuAH20]. maxima [MBDK21, RGA21]. maximin [MGS22]. maximization [ESLK22, SSCT23]. maximizing [AESY22]. Maximum [AFJ23, BA22a, DPS22, HD20, Bou23, CLL20, DDK21, DKL22, GNE21, KAVR20, LK21b, LWC22, LX24, Mac24, MM20, MMdOC23, MX24, uAJHD22, SZ21, TKF^{+}20, Tei22, ZNKN22]. maximum-type [MX24]. Maxwell [POC23, SAF21]. MCEM [TWTT20]. MCMC [BB24b, eLhS23, OKO21]. MCMC-type [BB24b]. MCMC4Extremes [MdN22]. MD [TEM20]. Mean [KZG21, SSGH22, ANR^{+}23, AH20, AS22c, AK22b, ACKL23, BJU20, Baz20, Baz22, Bek23, BPP20, BKPS23, CKL24, CD23, DR23, DC20, DLRZ21, EL21, GK22a, Han21, HU20, HB22b, HEK22, HZ22, JuAH23, KGF20, KBR21, KTK22, MZ22b, MNN21, MKSH22, MRM23, MSS20, MYVD21, MSB23, NS23, uAAF23, Pap20, PFM21, PKA22, PZ20, RAG21, RLP23, RuA22, SNKA24, SZ20, SIG23, SANH^{+}22, SMM22, SMO21, SKV20b, SKV20a, SVK22, SBB22, SPS22, SSK22, TPK22, TA24, WZW23, WKSeF22, Wil24, Wu21, WS22, YSH23, YP23]. mean-covariance [YP23]. means [AC23b, Bar20a, Cur23, FO22, KVS22, LW21, MJ20, MNN21, MX24, MES21, MHW24, OY23, PD21, SN23, Wan23c]. measure [AS22d, Bap22, Bek23, ESS22, LK21b, LKR^{+}23, MK22, REMM22, SSK22, dXLY20b, ZG22a]. measured [VBB23]. measurement [Brz20, CR21, FZY24, GZZ24, Gha22, GB22, JuAH23, KZG21, MRB23, MKSK23, MGS23, uARS22, uAJHD22, PB23, PZ20, SCE^{+}20, SVK22, SBB22, TL23a, YDLX20, YA21, YRAB22, ZGC21, ZG21, ZFG22, ZGL22, ZL23b, ZCW23, ZYF24, Zha24, ZZ22b]. measurements [AATD21, ARA23, KBR^{+}23, RBT^{+}22]. measures [Akd22, CP23, Dog22, LAK24, MRB23, MAMA20, SRF^{+}23, SK23, TDD20, Tür20, UR22, WK23, Wil22, WDZ22, ZSM23, ZKW24]. measures-based [ZKW24]. Measuring [GWZC22, SLW20]. mechanical [MRM22]. mechanics [CHL22b]. mechanism [RSG21]. media [KKZ21]. Median [GPS23, Cav24, CST21, CTK^{+}22, GK22b, HCTZ21, HMRN22, KST^{+}22, MKSK23]. medians [Kri22, SPV22]. mediated [ZCG24]. mediation [CLY23]. mediator [PHS21]. medical [Cha23b]. Mellin [LLLY23a]. Memory [QTH24, AGI23, AS21b, BS23a, FL23, TPC^{+}23, Wan23b]. Memory-type [QTH24]. MEPDF [WN21b]. mesh [Ver21]. meta [Avc21, Har21, LSC23, SLME23]. meta-analysis [Avc21, Har21, LSC23, SLME23]. method [AA23a, Ali22, Cha23a, CLL23, DS24, DA21, EAT22, FLXY22, GPGMK22, GW20, EJKL23, HLHC22, IGA24, LW21, LS22a, LBH23, LWW22, Med22, MBDK21, MEA24, MK22, Ors24, ÖT20, ÖK20, Pan20c, PK21, PMK23a, SAF21, SDG22, SLP22, SD22, SZJG21, SWC^{+}20, SF23, SY22, Sul23, UKK20, UEA20, WP22, WS20, WY21, WZY22, WLY22, WZZY23, YZF22, YWT22, ZT20, ZWY21, ZWMY23, ZfSRJ22, ZMNL23, dSRC22]. methodology [Squ22, XWL22]. methods [AÖ20a, AE23a, AAF23, AÇ23a, AK23c, BYK24, BFAR20, BPP20, CHC20, CSE^{+}21, CHL22a, CS22b, CdCG22, CCBN22, CMMV21, CL22b, CLZ23b, EGVdR21, FQKA23, Fig23, EJKL22, HZAF22, HLT23, HE24, IM20, KSP23, KFEL23, KS22a, KLM^{+}22, Kur21a, Kur21b, MES21, NBC23, NHZ22, Nis21, OLOK22, OBN22, ÖA22, PM23, Ras20, SWG24, SM23a, SRN^{+}22, SKK21, SSS23, SMMO22, TIK24, TPvdH23, UY23, WL20, YÖT20, YOD^{+}23, ZlBVGW^{+}23, ZGC21]. metric [CSG22]. metrics [BBK23]. MEWMA [KA24, SSBA22]. Mexican [SRO20]. MGR [CKT^{+}20]. MI [RH23]. microarray [Bis22]. MIDAS [XZJ^{+}21]. Minimal [AATD21, ARA23, Ami21, KBR^{+}23, RBT^{+}22]. MiniMax [AOA23]. minimization [BDK^{+}23, JPH21]. minimizing [CCM20]. minimum [AM22b, BVJV22, BZ22b, EHF23, HZ22, Kal22]. Mis [SBZG22, Hu23]. Mis-specification [SBZG22, Hu23]. misclassification [dOVQP21]. misclassified [AKG23, ÜFR22, WFLS24]. Mises [Fig23, JD23a, BZ22a, MG24, ZZFG20]. missing [BKPS23, DY22b, Fan22, GAS22, HAA^{+}23, HLT23, KS22a, LE23, NSlRGL23, PLS23, PRH21, RSG21, SSCT23, SANH^{+}22, SKK21, SU23b, WS22, ZXZL22]. missingness [RftADNI22, VBB23]. misspecification [ASA24, GZY22, dSPdSB23]. misspecifying [ZL20]. mix [AMK^{+}21a, CSE^{+}21, SCE^{+}20, XWL22]. mix-GEE [XWL22]. Mixed [AES21, BB24a, DG22, AY21a, ALB22, AW22, ALVQ22, Büy23, CHL24, CdCG22, DPS22, Els24, FHM22, FJ24, FBL22, Gha22, GB22, GGB20, GKJ23, HK20, KÖ21a, KÖ21b, MMMCS22a, MRB23, NSN21, ÖK20, RftADNI22, RSG21, RBN^{+}22, RS24, RBKA23, SF23, TWTT20, TWTT21, VJV^{+}23, WHW21, WQW^{+}22, YA21, YRAB22, ZBB23]. mixed-effects [AW22, CdCG22]. mixed-integer [RBKA23]. mixed-level [Els24, VJV^{+}23]. mixing [DCY22, DC24, GDWY22, LY20, SB24]. Mixture [LZZZ23, BAEk21, BKG22, CLL20, CAA22, DA21, KK22b, LH20, LZQ23, LF24, LES21, LBH23, LKR^{+}23, Nad21, OB21, PMM23, QL21, RSS22, REMO22, RCSS22, TII22, TJ20, TY22, Wan22, WZW23, WJB21, WLZ23, YWLD20, ZfSRJ22]. mixtures [CKCG21, FBL22, KAVdSFG22, Kim21, LFMMRH22, MCJ22, TJ20, WYY22]. MNAR [CLL20]. modal [KSAR21, Lu23]. mode [AJIAA22, AJI24, CWW23, DP23, Ruz22]. Model [ASA24, ALVQ22, Bar20b, CLC22, DP22, DC23, HD22, HE21, KJB23, SLSZ22, TKF^{+}20, AG20, AÖ20a, AS22a, AA22a, AAA22a, AAF23, AAK^{+}24, AZH^{+}22, ALB22, AA20a, Alg20, AK21, AE23b, AK22c, ARAR24, Alt20a, AS22d, AQYA22, AAA22b, AAM23, AJ22, AC23b, AMA^{+}23, ABB22, AHG20, BYK24, BD23, BC21, BSNT21, BS22a, BKIR23, BLXL24, Bil23, Büy23, CSL20, CSBdES22, CWW23, Çet23, Cha23b, CLL20, CSDT20, CHL22a, CsSL22, CWLH23, CL23, CSX20, CS22b, CLLX22, CLL23, CBRL22, CCM20, Ciu22, CMA21, CNL21, CAA22, DE22, DG23a, DA22, DZLS22, DSL20, DYC22, DCY22, DY22a, Dog22, DNWG22, DCL^{+}24, EH22, EN22, EE23, FW23, FJ24, FQKA23, FZM21, GZ22b, GZ22a, GWZC22, GZZ22, GHW20, GZY22, Gao23, GJK21, GKJ23, GWZ23, Ham22, HCZ21, Hu23, HZ23, HE24, IRA21, IRS20, JAP21]. model [JKLK24, JHR21, JM20, JD23b, KSAR21, KL21, KKK23, KS23c, KK22b, KBYS22, KVS22, KÖ21a, Kur21c, Kur23, Lee20, LMFMA23, LZSC23, LZZZ23, eLhS23, LY24, LS22a, LSTE24, LW22a, LY22, LLL22, LBH23, LG21, LKR^{+}23, LAKA22, MWTR23, MNL21, MRM22, MA23a, MAAA23, MB23c, MB23d, MB23b, MB24, MAMA20, MFS21, MF22, MVLMA22, MD22, MES21, MRB24, MGS23, MSNN^{+}24, NN20, NS23, NDWE24, dALNPdOF21, NG22, NAG21b, OBN22, ÖT20, PR22, PBC23, Pan20a, PC22, PD22, PSL23, PB22, PZ24, PS23, PAS23, Pho24, PLA22, POCC22, RPT23, RRA20, ROL^{+}20, RAA23, RHG22, RftADNI22, RBN^{+}22, RCSS22, RBKA23, SB20, SNA20, SM24, SI20, SDG22, SM23a, Sha20, SW23, SL21, SI23b, SMWJ20, SS21b, SLME23, SM20, SRO20, SMG23, SLW20, SCK22, fSLC21, SCK20, SCK21, SKJ20, TYLX22, TS22, TY22, TW24, TPC^{+}23]. model [TMZL21, TL23b, UOA23, UKK20, ULN21, dOVQP21, VH23, WS20, WLC^{+}22, WW23, WSZ23, WCY24, WFLS24, WKSeF22, WMJ21, XXH22, XWZL22, XS23, dXLY20b, XZJ^{+}21, XZS22, XH21, YDLX20, YV20, YA21, YGL22, YOD^{+}23, Yoo23, ZA23, ZBB23, ZK23, ZCG24, ZHW21, Zha22a, ZBA23, ZRC23, ZZZ23, ZHJ23, ZLF23, ZS23, ZZL23, ZYF24, ZXF22, Zha22b, ZWW23, ZCH23, ZLTP23, ZC23, ZKW24, dSdSTC21, AE23a]. model* [ÖÇ23, YW20]. model-averaged [TL23b]. Model-based [ASA24, KJB23, CAA22, MNL21, SCK22, ZBA23]. Model-free [CLC22, DP22]. modeled [DPS22, OO21]. Modeling [BS21, CL21, LZQ23, PHN22, Sad24, SKG22, TGS22, YC22, dSRC22, AT21, BH23b, CLL23, CDB^{+}21, Dog22, DA21, DKL22, GI20, GY23, GGB20, HMSRM23, HG22, KCS24, KJK20, LW23, MWNSB23, MSNN^{+}24, Mur23, SWC^{+}20, SF23, TJ20, Tür20, Ver21, ZLXT23]. Modelling [AMA22, SZS20, AJI24, AA20d, GZ21a, WHW21, WC23a]. models [AK23a, ALMA23, AGI23, AJIAA22, AALR23, AW22, ALVQ22, AE21, BCDA21, BJ20, BBM24, BB24a, BLR23, BR24, BH20, BA22c, BA22b, BSAS22, Bia22, BAL24, BCR^{+}21, BE20, Brz20, Büy23, CR21, CST21, CFM21, CHL24, Che22, CdCG22, CSA23, DC23, De 20, DLRZ21, EGVdR21, ENK^{+}22, FHM22, FY21, FW23, FF23, FZZ20, FS22, FBL22, FP20, GZ22b, GZ22a, GWZC22, GZ22c, GZZ22, GNE21, GLT20, GO22, Gha22, GB22, GPS23, Gül22, HMSRM23, HK22, HHvR22, HD20, HB23, Hep21a, HK20, HW24, HCZ21, HE21, HHY21, HZY22, IAG21, IGA24, IT21, IRA21, Jhw24, JGL22, JGGDMRLD21, KAVdSFG22, KKO21, KZG21, KHK22, KÖ21a, KÖ21b, LE23, LLZ22, LZQ23, LES21, LMP20, LC22, LY23, Lu23, LW22b, Mag20, MAAA21, MRB23, MAB23, MGS22, MZZF21, MPS21, Mos22, Nad21, Nag21a]. models [NSN21, NIN20, NHZ22, ONE22, Ors24, OKO21, OB21, ÖA22, ÖK20, Özk21, PHN22, PBLM22, PL22, PV20a, PM20, RSG22, RS24, RVV22, SB22, SB21a, SB24, SU23a, STH24, SKV20a, SR23, SCA23, SLE20, TY21, TEM20, TF22a, TWTT20, TWTT21, TL23a, TEO20, UM22, VLvdW22, WP22, WWYX21, WX21, WKT22, WZW23, WKcIC24, WJB21, WCF21, WLZ23, XWL22, YZ23, YLX^{+}22, YXYX22, YP23, YRAB22, YWLD20, YL22a, YL22b, ZNLW20, ZZFG20, ZGL21, ZGC21, ZG21, ZD22, ZG22b, ZFG22, ZQYW22, ZCW23, Zha23, Zha24, ZZWH20, ZZ22a, ZXZL22, ZWTY23, ZWCY24, ZZ22b, ZK20b, dSdSTC21, dSLHMB23, dSC23, dSPdSB23, SVN24]. modes [GPP21, KCS24]. Modification [Ali22, TW24, HZ23, Med22]. Modifications [PP23]. Modified [BSNT21, CS22c, HQZ^{+}23, Kur21c, LAKA22, SAA21, Sul22a, VW21, WKT22, ALMA23, Alg20, ABPK21, AA22c, CLLX22, CKT^{+}20, DSL20, HZ22, HYSC21, JH23, KUA21, KR23, LA23, Lu23, LAK24, MRM22, MR22, NG22, PAS23, RYA20, SZ21, SSZ23, SSK22, TK22a]. modulated [DL20, ZC23]. moment [BPP20, Che21, Gho20, MQ21]. moments [AZAZ23, LPP21, LB22, LWW22, ONS22, SAA^{+}23, WN21a]. monitor [HBS22, HORA22, MMMCS22a, SMM22]. Monitoring [BNK21, CKL24, DuA23, ENA^{+}23, LTJC23, PG23, PV20a, RSS22, ANR^{+}23, ASS22, ACKL23, ASM22, CW24, EAJ20, FAT23, HA24, HQZ^{+}23, HZZ^{+}24, HMRN22, JuAH23, KA24, KST^{+}22, KS20a, dALFB21, MT21, MMMG21, MÖ23, MASA22, MMMCS22b, MMMCS23, uAAF23, PASB21, RAG21, RYA20, RuA22, SK24, SNKA24, SMM22, Spa22, TAS24, TAA22, TP22a, WHS23, YHGA21, YL23, ZL23a, ZHJ22]. monotone [CLJK21, YSH23]. monotonicity [RVV22, Xia22]. Monte [Squ22, AQAN22, CL22b, DK22b, FPAA21, KKZ21, KSP23, KOT20, uIKKH22, Kur21c, LPNP20, UU23, Uya22, WMS22, ZL20]. Morgenstern [dXLY20a, dXGLY20]. Moriña [GRR24]. mortality [SRO20, SR23]. Most [KuIK20, uIKKH22, Uya22]. motion [DPS22, KL22]. moves [TPK22]. Moving [KNSA23, OR23, AOA23, ASA23, AK20, ACKL22, ACKL23, CLS22, CTK^{+}22, HZZ^{+}24, LLZ22, NZ21, uARS22, uASS22, uAA23, PBC23, QTH24, WHS23, YCD23]. moving-average [LLZ22]. MSE [Han21]. much [SRN^{+}22]. Multi [BDD^{+}22, KS22b, RAN23, REN22, Avc21, Bey21, Bia22, CY24, EAJ20, Gai21, KB23, MFS21, ZYSS22]. multi-asset [MFS21]. multi-category [Avc21]. multi-factor [Bia22]. multi-label [CY24]. Multi-objective [BDD^{+}22, RAN23, REN22]. multi-performance [ZYSS22]. Multi-response [KS22b, KB23]. multi-sample [Gai21]. multi-stage [EAJ20]. multi-state [ZYSS22]. multi-step [Bey21]. Multicategory [ZC22]. multiclass [Ega24]. multicollinear [EE23]. multicollinearity [AA22c, Daw23, Sad24, SCK21, SCA23, ZZ21]. Multidimensional [KA22, UM22, ZLXT23, dSLHMB23]. multilevel [BF23, FS22, KSP23, LE23, PM23, ZLXT23]. multinomial [DSL20, Esk21, FQKA23, LW22a]. Multiple [DF22, Tak23, ASS22, AMA^{+}23, ARJ20, ABPK21, ABJ21a, ABJ21b, BBD^{+}24, BDD^{+}22, CH23, CM22, CSG22, GK20, HW24, KCS24, KJB23, KS22b, KB23, LCC22, LB23a, NSlRGL23, OY23, RAJ21, RLP23, SS20a, SK23, SLJ23, TWTT20, WHW21, WWW24, Wes20, Wu21, YDLX20, ZT20, ZHW21, ZRC23, ZLXT23, ZL23c]. multiple-index [ZHW21]. multiplex [TWSW23]. multiplicative [BZ22a, CL23, FZY24, GZZ24, HZAF22, ZXW23, ZCW23, Zha23, ZYF24, Zha24, ZZ22b]. multipliers [Gan22, GW20]. multiserver [BC23a]. multistage [BNK21, KN22]. Multivariate [EHHS20, GJQ23, HMSRM23, LMFMA23, WN21b, ASA23, ARJ21, BNK21, BCC24, Baz20, Baz22, CSBdES22, CWLH23, CW24, CKL22, Dog22, ENA^{+}23, FN20, Gao23, Gho23, HHvR22, HZAF22, HW24, HCLF23, IT21, JL24, JM20, KAVdSFG22, Kal21, KA24, KM22, KS20b, LYZ^{+}21, MCWL22, MVLMA22, MAF23, MV22a, MSS20, Mur23, NAG21b, ONS22, OR23, Pan20c, Par20, RAG21, SK24, SU22, SI23a, SK23, TCJ23, UEA20, WYY22, WMXL24, dXLY20b, YZ23, YHGA21, YL23, ZHJ23]. Multiwavelet [CZS21]. myopia [FW23].

Nadaraya [Ali22, AHB23, ESES20]. national [MNL21]. Navarro [GRR24]. NBA [SS21b, SZS20]. nearest [FH22, GS23, TED24, SWG24]. nearest-neighbor [GS23]. nearly [KBR^{+}23, MB24]. Necessary [WWTM22]. necklaces [DG23b]. need [dOVQP21]. negation [LX24]. negative [Arn23, AKR23, GPP21, KS20c, LC22, MB23c, MB23d, PR22, PZ24, WWYX21, ZBB23, ZWY20]. negatively [DYC22, ZQYW22]. neighbor [GS23]. neighbors [FH22, TED24]. neighbour [BJVV21]. Nelson [ZLTP23]. nerve [OP20]. Nested [ZCH23, EH22, MGS22]. nested-error [EH22]. net [AJIAA22, LYY21, Pol22]. network [BVJ^{+}22, DKKA22, IRA21, IMR21, SBR^{+}23]. networks [CMM23, CLZ23b, FAT23, GZ21b, HPZ^{+}21, KE22, Min22, Min23, MASA22, SMWJ20, SZ23, TWSW23, ZSM23]. Neumann [Squ22]. neural [CLZ23b, DKKA22, IRA21, KE22]. neutrosophic [ARKA22]. newly [TK22a]. Neyman [NAJ22]. NMBAR [YWL24]. NN [AALR23]. No [ALZ24]. No-U-Turn [ALZ24]. node [TWSW23]. node-attributed [TWSW23]. noise [AS21b, DDK21, ZK20a]. noises [SB24]. noisy [MZZF21]. nominal [UR22]. nomination [RGA21]. Non [Bar20b, MWNSB23, Sol21, AATD21, ARA23, ASZN21, Akd22, AHB23, ANP20, AS22d, CP22, DE22, DBS23, FN20, FO22, GE23, HHOC23, KKO21, KFEL23, KS22a, KJB23, KR23, LWP24, LK24, MF22, MSB23, PR22, Pan20c, RSG21, SA22, SI23b, SF24, SU21, SK22c, SBB22, SBB23, Wan22, WDZ22, WCF21, WH24, XWZL22]. non-Bayesian [ASZN21, KR23]. non-circular [AATD21, ARA23]. non-crossing [KJB23]. non-Gaussian [DE22, WCF21]. non-homogeneous [XWZL22]. non-ignorable [RSG21]. non-inferiority [CP22, HHOC23, WH24]. non-informative [Wan22]. non-isomorphic [KFEL23]. non-life [WDZ22]. non-linear [PR22, SA22]. non-Lipschitz [MF22]. non-Lipschitzian [SF24]. non-normal [FN20, LK24, Pan20c]. non-normality [FO22]. Non-Parametric [Bar20b, Sol21, Akd22, AHB23, ANP20, DBS23]. non-probability [LWP24]. non-response [KS22a, MSB23, SU21, SK22c]. non-sampling [SBB23]. non-sparse [AS22d]. Non-stationary [MWNSB23, DE22, GE23, KKO21, SI23b]. noncompliance [SEMR23]. noncontractual [YWT22]. Nonlinear [ARJ21, CL21, ZG22b, Zha23, AES21, BD23, CdCG22, EE23, GG22, Hep21a, Hep21b, KA23, Lu23, NSN21, RBKA23, Tür20, Wan23b, ZL22b]. nonlinearity [AE21]. Nonnegative [LYY21]. nonnormal [Cur23]. Nonparametric [BZ22a, CDB^{+}21, Fan22, GJK21, GK20, HvZ21, JL24, LDA23, LFZ21, SZ20, SK22a, SD22, SSSR23, YWT22, Zha24, ANR^{+}23, Ali22, AG21, AY21b, CLS22, DYC22, GZZ22, Gho20, HLHC22, KK22b, MKK22, Meg23, Ruz22, YL23, ZQYW22, ZLG23]. Nonparametrics [AM24]. nonrandom [AJ22]. nonresponse [KJK20]. nonsampled [KK22b]. nonstationary [AC23b]. Nordic [DPS22]. normal [AS23, BJU20, BAEk21, Baz20, Baz22, BFAR20, CWW23, CWWT24, CSDT20, Dog22, FN20, FZM21, FBL22, GZ21a, Gök22, HR22, HK20, HZ22, KAVR20, KAVdSFG22, Kri22, KTK22, LMFMA23, LF24, LES21, LTR21, LK24, MCWL22, MSL21, MAMA20, MES21, NZ21, ONS22, ON21, PHN22, Pan20c, PD21, PDK23, SI23a, SS20b, SS20a, SPV22, TJ20, TNN20, TK23, VPOC24, WWTM22, WZW23, WQW^{+}22, WYY22, WMXL24, YWL^{+}23, ZX22b]. Normality [Bil23, PV20b, Bay21, CYS23, EHHS20, FO22, MQ21, OR23, Par20, Sul22a, UY23, WMS22]. normalized [WW23]. normalizing [EHHS20]. normally [CHL24, IRS20, Spa22]. Note [SK22b, AW22, BJVV21, DLRZ21, GW22, Gho20, Jod23, Kri22, Mag20, RL21, SVM20, TMZL21, WLC^{+}22, Wil24, YV20, YTL22]. Notes [Lui22, ZL20]. novel [ASS22, ARAR24, CII22, EHF23, Els24, HBS22, KA23, RAA22, SAA^{+}23, WW23]. np} [LK21a]. NSGA [REN22]. NSGA-II [REN22]. nuisance [GZY22]. null [Bis22, uIKKH22]. number [JAP21, Tel22, ZMNL23]. numbers [Tür20, YHWW23]. numerator [AKG23]. Numerical [CKK21, KLM^{+}22, KBB24, WQW^{+}22, BV22, JL24, ÖT20, SZJG21].

object [Hor22]. Objective [Seo22, BDD^{+}22, KE22, RAN23, REN22]. objectives [ZL23c]. observation [AZH^{+}22, HAA^{+}23]. observational [KT23, YW24, ZL22a]. observations [CM22, FL23, HBS22, JD23a, KB23, MHW24, PV20a, QYC22, SI20, TH23]. obtained [NS22, Squ22]. occasion [KS22a, SKK21, SK22c, SPS22]. occurrence [CBRL22, LRF20]. occurring [SN22a]. OCtS [DS24]. odds [CP22]. old [AKS^{+}21]. Olkin [HG22]. Olson [ZLTP23]. Oman [AMP23]. oncology [Yad22]. One [Wu21, BD23, BPN22, Cav24, CKL24, HQZ^{+}23, KJK20, Mis20, RYA20, RAJ21, TCJ23]. one-sided [HQZ^{+}23, RAJ21]. one-step [Mis20, RYA20]. one-truncation [BPN22]. one-way [BD23, Cav24]. Online [XH21, BQCP23]. only [BF23]. onset [Yad22]. operating [AESY22, FTC21, GM24a]. operators [SM20]. optic [OP20]. optical [KKZ21]. optima [dCCM^{+}22]. Optimal [CLY21, CTCL22, CKT^{+}20, CTK^{+}22, El 24, HB22a, HMV^{+}22, Hor22, Mis20, SK23, AJ22, BB24b, BJVV21, BFS22, EAJ20, Els24, GJK21, GKJ23, HI22, LZ24, LFMMRH22, MPD20, MY23, Nak22, NN20, PMM23, PJL21, TYLX22, UIC23, YHGA21]. optimalities [HAA^{+}23]. optimality [BPP20]. OptimaRegion [dCCM^{+}22]. optimization [ASM21, BM21, CHL22a, CS22b, CCM20, ES22, IS20, SI20, YÖT20, dOMD^{+}21]. optimized [XWZL22]. Optimum [Hu23, PMM23, SBTP20, ZLSC20]. Option [GWZ23, LLLL22, MFS21, SM24]. optional [KZG21, LJ22a, NS20, WKSeF22]. options [LLLY23a, WZX23]. order [AZAZ23, ARJ21, AAA20, AAGP23, BRTW24, Bap20, Baz20, Baz22, BC23b, BPP20, ÇKK23, Che22, DRB20, FSGMM23, GPP21, GBHA20, HR22, LZ22, LB22, MVLMA22, MES21, MHW24, NJ22, SSZ23, Sol21, Sun23, VSJ^{+}24, dXLY20b, ZLF23, Zha23]. ordered [Akd22, AG21, Bey21, CMM23, Gao23, KTK22]. ordering [CM22]. ordinal [AA20b, GZ21a, GGB20, PG22, PG23, RftADNI22, RSG21, TDD20, WK23, ZZL23]. ordinary [MZZF21, NS22]. organizing [CCH20]. Orthogonal [ZWCY24, EH22, GGMM22, KFEL23, LDA23, OY24]. orthogonality [ZZWH20]. other [DP20, FS23, Lui22]. outcome [BT20, KS23a, LSC23, SK23, ZD22]. outcome-adaptive [KS23a]. outcome-dependent [ZD22]. outcomes [RHG22, Sha22a, SLSZ22, SWC^{+}20, TDD20, Yad22, YTL22]. Outlier [AMK^{+}21a, Mur23, ALMA23, KA22, OY24, UEA20, YXYX22]. outlier-modified [ALMA23]. outliers [BYS24, CYS23, DS24, FBNRG21, KA23, LBL22, MA23a, MAAA23, MRB23, MMVBE23, SRF^{+}23, SCK21, SCA23, Tel22, TEO20, WCF21, YÖT20, ZA23]. outlying [MHW24]. over-dispersed [TW24]. overdispersed [AKR23, BLR23, TL23b]. overlap [LLV23, VJV^{+}23]. overlapping [EAT22].

p [PAS23]. package [ATYK20, FBNRG21, JPH21, MdN22, MM22, MN17, PCWL22, War23, YP23, dCCM^{+}22]. packages [DG22]. Padé [AY21b]. paired [SCHS22]. pairs [AGA20, CP22]. Pairwise [BDL23, Lev21]. panel [AY21a, AÖ20b, DP22, DNWG22, FZM21, Nad21]. panels [Wu22]. Panichkitkosolkula [Kri22]. paper [Kri22]. paradigm [STH24]. Parallel [SCK22, Lui22]. Parameter [CLLX22, MFPP21, TH23, ZS23, ASZN21, AA22a, AA23a, AAA22a, Alg20, AHB23, BYK24, BS22b, BPN22, CCBJ21, CTCL22, CSX20, CS22b, CCM20, CS22c, DG23a, EE23, ENK^{+}22, FQKA23, FLPB20, GB22, GW22, Han21, Hua22, IRS20, KN22, KLM^{+}22, KLM24, KP21, Kur21c, LA23, LK21a, LF24, LS23, MCWL22, MJ20, MG24, MGS22, Nak22, NS22, ÖT20, ÖA23, OP22, PD22, PJL21, PMK23a, RAA23, RAA22, SAF21, Sha22b, TK22a, VG22, WQW^{+}22, ZK23, ZB22, Biç20]. parameter-free [TK22a]. parameters [AK23b, ASK23, BA22a, BPN22, BLXL24, BDL23, CKT^{+}20, CTK^{+}22, EE23, GZY22, GSG22, HCTZ21, JD23a, JHR21, KHK22, KB21, LTR21, Mag20, MRM22, MK21, MM20, MMdOC23, MZZF21, MMMCS22b, NBC23, SK24, SU22, SCK20, Sun23, TIK24, TY22, YCD23, YHGA21, YWL^{+}23]. Parametric [Bar20b, GZZ24, MSS20, Akd22, AHB23, ANP20, BCR^{+}21, DRB20, DBS23, HB23, Hut22, dALNPdOF21, QL21, RftADNI22, Sad24, SB20, Sol21, TF22a, YZ21]. parametrization [vO21]. parent [MKS22, NAJ22]. parent-generalized [MKS22]. Pareto [AK23b, BC23b, DA22, Gho23, JKH20, LSTE24, MV22b, WG23, ZWC^{+}23]. Parisian [CHL22b]. part [YZ21]. Partial [ZFG22, AG20, Che22, GZ22b, GZ22a, IMR21, KSAR21, Pol20, Seo22, WS20, WK23, YZ21, ZNLW20, ZGL21, ZHJ23]. Partially [KVJH23, AG20, Çet23, CL23, GCF^{+}20, HCZ21, JH23, LY24, LY23, Lu23, PLA22, RVV22, ZLF23, ZZWH20, ZXF22, ZWTY23]. Particle [IS20, ASM21, RY21, SI20, dOMD^{+}21]. particles [EHF23]. partition [LCC22]. partitioning [KJB23]. partitions [CQ22]. partly [PC22]. past [CP23]. patch [Cha23a]. path [PN21]. patient [DF22, SLME23]. patients [JBJ23]. pattern [CLL20, HCLF23]. patterns [Ger21, PK21, SN22a, SPS22]. payments [LC22]. PBIB [BGJ^{+}23, KS23b]. PBIBDs [SG21]. PBINAR [MPS21]. PCA [AMK^{+}21a]. PDF [Ras20]. Pearson [YWLD20]. Pegram [SM20]. PEIV [Zha22b]. Pena [KUA21]. Penalized [ABB22, CYS21, CYS23, CLJK21, LJC^{+}20, TY21, WSZ23, FF23, GPGMK22, IAG21, IGA24, JKLK24, KJB23, KLM^{+}22, MRB24, ONE22, YDHW21, YXYX22, ZWTY23]. Penalty [PLA22, VLvdW22]. perfect [AA22b]. Performance [AÇ23a, CWW23, CSE^{+}21, CdCG22, JuAH23, KR21, uAJHD22, OBN22, ZSM23, AG21, BS21, CS22a, Cos23, EAJ20, FS23, GABJ23, HCTZ21, HZZ^{+}24, MCJ22, NN20, ÖK21, Pol20, SRN^{+}22, SJ24, SZS20, SCK23, UEA20, Wan23a, WCRM23, WC23b, WCKC23, YLX^{+}22, ZYSS22]. Performances [Kur21a, Kur21b, AGI23, SNKA24]. period [MY23]. Periodic [MB23d, BA22c, BA22b, LG21, MB23c, MB23b, OB21, SB20, SB22, ZKW24]. periodogram [Wan23b]. periods [ARA23]. Permutation [MBDK21]. persistence [Aga23]. personalized [KS22b, SK23]. perspective [DuA23, MY23]. perturbed [LG21, TH23, Wan23b]. PH [JD23b, SI20]. phase [ASS22, CD23, MD22, PV20a, SSGH22, SIG23, Yad22, YHGA21, FAT23, HA24, KA24, KN22, YHGA21]. Phase-I [KN22]. phenomena [MD22, TK22a]. phylogenetic [Jhw24]. Piecewise [BBM24, MD22, Ors24]. PINAR [KL21, MB24]. Pitman [AR22, MR22, REMM22]. plan [AMA^{+}23, ABJ21a, ABJ21b, ARKA22, El 24, Hu23, KJB22, MS20, uASS22, PMK20, PMK23b, RAJ21, Spa22]. planar [Ger20]. planning [HB22a]. plans [BNA23, CLY21, DR22, NN21, RGA21, SBTP20, ZLSC20]. play [SS21b]. player [HCHK23]. Pliable [AWZ24]. plot [GGS23]. plots [Wij23]. PM [GZ21b]. Point [MG21, BSNT21, CCBJ21, CMM23, CHC20, EFD21, FSGMM23, HZ22, LZQ23, LY20, LD23, Mil21, Pic21, QYC22, SJ24, TY24, Ver21, WTN22, Wu22, ZHJ22]. points [BDD^{+}22, EHF23, LLLY23b, ZT20]. Poisson [AK23a, AK22b, ASM22, AAA22b, ABB22, BBM24, CFM21, DL20, GZY22, HA24, HZ23, KUA23, KR22, LKHC23, LRF20, LD23, MÖ23, MBBS22, MV22a, NM21, OY23, PS23, PASB21, SSZ23, Sun23, SKJ20, TIK24, TW24, TL23b, WZ21b]. Poisson-Modification [TW24]. policy [Bal22, LWS23, MY23, WMJ21]. Pollard [SRO20]. polling [SBB24]. polynomial [CBRL22, ZHW21]. polynomial-exponential [CBRL22]. polynomials [DY22a, LJC^{+}20]. polytomous [dSLHMB23]. population [ASA24, BPP20, BKPS23, BRA20, BF23, CWWT24, CK20, CD23, HUK23, IR21, KGF20, KK22b, MZ22b, MMVBE23, MYVD21, MSB23, NS20, NS23, PD21, QTH24, SNSE22, SZ20, SKV20b, SKV20a, SBB22, SSK22, Wan23a]. populations [AKE23, AB22, Baz20, GKS24, GK23, Gök22, dAIIS20, JD23a, KG22, KTK22, RY21, SRN^{+}22, SI23a, SN23, TNN20]. Portfolio [NN22, AM22b, BM21, dXLY20a, dXGLY20, dXLY20b]. positive [Bap22, DBS23]. positives [BR21]. possible [MHW24]. post [PLA22, ZL22a, Bar20a]. post-matching [ZL22a]. Posterior [KLK21, Han20, Han22, ZBA23]. Potential [BBD^{+}24, Sul23]. Power [BYK24, BGM20, BK23, HM21, WMS22, WC23b, AMA22, AS22d, CYS21, CYS23, CS22c, DWLS23, KLM24, LMFMA23, LLLY23b, Lui22, Pan20c, POCC22, RL21, RSG21, RSG22, RK20, SSS20, UY23, Uya22]. power-divergence-type [AS22d]. power-law [CS22c]. power-normal [LMFMA23]. powerful [DC20]. Practical [TD23]. pre [SMO21]. pre-timed [SMO21]. precipitation [IRA21, RLP23]. Precision [AM22b, Ava21, GAS22, MSNN^{+}24, WY21, WZZY23]. predefined [Wan22]. Predicting [BKG22, BT20, BBK23, CLL23, LZQ23]. Prediction [KK22b, PBC23, BJ20, BC21, CNL21, GAS22, HI22, HvZ21, LCC22, Lon23, MWNSB23, MBBS22, ÖK20, SS21b, SLME23, SZJG21, TED24, WS20, WWT22, WL20, WDZ22]. Predictive [HK22, DA22, KÖ21a, LLL22, SPG^{+}22, YLX^{+}22, ZBZS20]. predictor [BH23a, SAD^{+}21]. predictors [AR22, CZ22, HAT^{+}22, KR23, KÖ21b, SMB23]. Preface [FP20]. preprocessing [ÖDS24]. presence [BKPS23, CYS23, KK22a, KA23, KB23, MA23a, Meg23, MS20, MMVBE23, MSB23, uAJHD22, PZ20, RAG21, RGA21, Sad24, SEMR23, SKV20a, SU21, SBB22, SBB23, SR23, SCK21, TRF^{+}22, WGlTL23, ZA23]. preserving [HD22]. presmoothing [SMMO22]. PRESS [IHS22]. pressure [BT20, KKO21]. pretest [PLA22]. Pretesting [FO22]. prevalence [Avc21]. prevention [KAVR20]. preventive [PM20]. Price [LWS23]. prices [DE22, SI23b]. Pricing [CHL22b, LLLY23a, MFS21, OO21, WZX23, GWZ23, LLLL22, SM24]. prime [MAB23, dSPdSB23]. Principal [FN20, Moh22, Kur21b, ZNKN22]. principle [ZL23c]. prior [AÇ23a, EE23, Han22, SK22a, dSS22, UKK20]. priors [ALVQ22, KLK21, OY23]. privacy [HD22, LKR^{+}23, SSK22]. privacy-preserving [HD22]. probabilistic [BB24b]. probabilities [CYS21, IRS20, KR21, Mil21, NZ21, SMMO22, ZBA23]. probability [Aub23, BRTW24, Bar20a, CBRL22, Han22, HB23, KTZ20, KGF20, LTN23, LWP24, LX24, ON21, RSS22, SN22a, TF22b, TAA22]. probit [Gao23, MEA24]. problem [AKG23, Gai21, HCHK23, Squ22]. problems [AA20d, Che21, CCM20, HBM^{+}20, KS22a, KFM23, KLM^{+}22, SKK21, WZY^{+}21, YÖT20]. Procedure [KB23, AW22, Aub23, DC20, OM21, SS20a, WWW24]. procedures [CCBJ21, GO22, HZ22, HE21, KB21, SM23b, TA24, WZ21b]. process [ANR^{+}23, AH20, AK22a, ACKL23, ABJ21b, ARKA22, BGM20, Bal22, BNA23, Biç20, CWWT24, CKL24, CSDT20, CKL22, CKT^{+}20, CS22c, DR23, GK22a, GK22b, GBHA20, GKJ23, HU20, HB22b, HEK22, HCTZ21, HE21, KKZ21, KST^{+}22, LK21a, LLLL22, LLLY23b, LZQ23, LJ22b, LZ22, MMMCS22a, MMMG21, MKSH22, MMMCS22b, NAJ22, uAAF23, PG22, PDO^{+}20, PKA22, PIA22, RAJ21, RSS22, SK24, SNKA24, SM24, STdL21, SP23, SM23b, SSZ23, SCK23, SBZG22, TH23, WZ21b, WWT22, WJB21, YH21a, Yao23, ZWY20, ZX22b]. process-yield-based [BNA23]. processes [ASA23, ASM22, BNK21, BCP21, BS23a, BS23b, Bey21, Bou23, CMM23, CS22a, DK22a, DRB20, DL20, EAJ20, FSGMM23, GKM20, GE23, GS23, HORA22, Hu23, JHR21, KN22, LLLY23a, LC22, LD23, PG23, PV20a, PAT22, Pic21, RdAIS^{+}20, TGS22, VWZ21, Ver21, WZX23, XXH22, YWL24, YL23, ZL23a]. Procrustes [GP20]. produce [Ors24]. product [LCC22, NDWE24, QTH24, SVK22, SYPV23, ZZ22b]. production [KST^{+}22, KS20a, SNKA24, YC22, ZHJ22]. products [CS22a, MSL21, WC23b]. profile [PASB21, RAG21, ZGC21, ZG22b]. profiles [ASS22, AK22a, BNK21, HBS22, HA24, KA24, KN22, MÖ23, MV22a, RAG21, REN22, YHGA21]. profiling [CSE^{+}21, SCE^{+}20]. programming [AES21, RBKA23, TWSW23]. progressive [ANR^{+}23, AK22b, AHG20, BR24, BB24b, CSL20, CKK21, DSS23, KB21, KG22, LTR21, LHG24, MWTR23, MK21, SBTP20, Seo22, Sha22b, Tei22, WLT22, WL22, WG23, WC23b, WCKC23, YBC22, YSG23]. progressively [AKE23, AZAZ23, BMG22, HB22a, LTB23, MG21, MJ20, Pak23]. projected [NAG21b]. projection [BV22, FBNRG21, ZWCY24]. Propagation [BCDA21]. propensity [CLZ23b, HAT^{+}22, KT23, OBN22]. Properties [AAHR20, Alt20b, Bek24, KS23d, ASZN21, BCC24, CYS23, CZ22, DKL22, Ega24, GMC23, Gho23, KLK21, KS20c, KH23, PKW22, PV22, SAK24, SB24, WLC^{+}22]. property [XXH22]. prophet [TPC^{+}23]. proportion [Bis22, CWWT24, NS20, QLG23, SK22a]. proportional [BBM24, CMA21, MW20, RHG22, SK22b, SR23, WFLS24, ZCG24, ZWW23]. proportions [SYPV23, UYMK22]. proposal [AA22a]. Proposition [AF22b, MAF23]. Proschan [GM22]. protection [SSK22]. providers [SCE^{+}20]. proxy [DZLS22]. Pseudo [TKF^{+}20]. psychology [MMR22]. psychometric [DKL22]. pure [Ham22, RdAIS^{+}20]. pure-birth [RdAIS^{+}20]. purely [MHW24]. pursuit [FBNRG21]. put [WZX23]. Python [YP23].

QMLE [BA22c, LL23]. QMLE-based [LL23]. QSAR [AAK^{+}24]. quadrant [DBS23]. quadratic [KL22, SS23, YÖT20, YTL22]. quadrature [AA23a]. Quality [KS20a, BQCP23, ENA^{+}23, RSS22, SNA20, SN22b, ZK20a]. Quantile [MZ23, MSL21, MR22, SCK20, SCK21, Aki24, AA20b, AM22a, CKCG21, GGB20, Hut24, KM20, KHK22, KL23, KJB23, Kur23, LMP20, MRM23, SD22, SSSR23, SCA23, SHN23, TWTT20, TWTT21, WJZ^{+}22, ZZZ23, ZLSC20]. Quantile-based [MSL21, SCK21, SSSR23]. quantile-mean [MRM23]. quantiles [ABHF22, ACP21, DP20, ESES20, MCJ22]. quantitative [Jhw24, KVS22, SKV20b, SKV20a, SSK22, TPK22]. Quanto [LLLL22]. Quasi [DDK21, TW24, Bou23, CL22b, EL21, GNE21, JGL22, LPP21, LWC22, NT20, SS20b, YH21b, AK23c]. Quasi-Least [AK23c]. quasi-likelihood [JGL22, YH21b]. Quasi-maximum [DDK21, Bou23, LWC22]. quasi-Monte [CL22b]. quasi-random [EL21, NT20]. quasi-stationary [LPP21]. question [AE23a, NS20]. questionnaire [SY22]. queue [BC21, EAJ20, LJ22a]. queueing [BCDA21, JD23b]. queues [ALZ24]. queuing [BC21, BC23a].

R [AE23c, ATYK20, DG22, DG23b, FBNRG21, GGS23, JPH21, MdN22, MM22, MN17, War23, XW20, dCCM^{+}22]. R/S [XW20]. R/S-bootstrapping [XW20]. radiation [KKZ21]. Radon [JH23]. Rahim [PM20]. rainfall [LRF20]. random [AJ22, Bar20a, BJVV21, BKPS23, CK20, CGS21, CWLH23, CHL24, CKCG21, DY22b, EH22, EL21, FZZ20, GLT20, GZ21a, GKJ23, Gül22, HI22, HUK23, Hua20, KS22a, LHY^{+}23, Lev21, LZSC23, LJ22b, LZ22, MS22, MASA22, MYVD21, MSB23, MKS22, NT20, NZ21, PLS23, RY21, SNSE22, SJ23, SAA^{+}23, Sha22b, SN22a, SK22c, SBB22, SBB23, SLW20, TBK23, WJZ^{+}22, WSZ23, WZX23, Wan23b, YCD23, YHWW23, YGL22, YW20, ZQYW22, ZXZL22, Zhe24, ZL20]. random-effects [GLT20, Gül22]. randomization [JPH21]. Randomized [Arn23, NT20, Akd22, BV22, KVS22, NS20, SRN^{+}22, SKV20b, SKV20a, SSK22, SBB24, WKSeF22, XSS22, Yad22]. randomly [LHY^{+}23]. range [SA22, SS20b]. ranges [ACP21]. rank [AÖ20b, FS23, KE22, RSM22]. rank-based [RSM22]. Ranked [BC23b, AS22b, AOA23, AA22b, BRA20, KK22a, MZ22b, MLR^{+}22, MR22, Pan20b, RuA22, RTJ24, STdL21, SZ21, TdSCdSdCP22, VDN22, WCY24, YCD23]. ranking [AA22b, MK22, SDG22]. ranks [TP22a, ZHJ22]. Rao [XSS22]. rare [AÇ23a, ONE22]. rate [AMK21b, CLZ23a, KS23d, NG22, PR22, PK22, Sal23, SM24, XZJ^{+}21, ZS23]. rater [TDD20]. rates [AS23, BR21, Hor22, KF24, LAZ21, SKG22, Wan22]. rather [Sul22b]. Ratio [ESS22, KK22a, QYC22, BZ22b, CP22, Che21, CS22c, GGMM22, Gök22, HHH23, KGF20, Kri22, LZSC23, LTN23, MQ21, MKS22, ON21, PN21, QTH24, RHG22, SVK22, SPV22, SYPV23, WP22, WTN22, ZA23]. ratio-cum-regression [KGF20]. rational [Spa22]. ratios [ACP21]. Rayleigh [ABHF22, HORA22, Lon23, MWTR23, MK21, WC23b]. RCINAR [ZWY20]. re [JD23b, PP23, TF22a]. re-estimation [TF22a]. re-scaling [PP23]. re-servicing [JD23b]. real [GNE21, KF24, SPG^{+}22]. reap [BH23b]. reassessment [UKK20, ZWY21, ZWMY23]. receiver [AESY22, GM24a]. reciprocal [AM22a, Alh23]. recognition [ZfSRJ22]. Recognizing [Sul22b, Sul23]. recommended [AE23a]. reconstructions [OP20]. record [AR22, TS22, ZWC^{+}23]. record-based [TS22]. records [KR23, SJ23]. recovered [OP20]. recovery [KBB24]. rectifying [AMA^{+}23, RGA21]. Recurrence [LB22, KF24]. recurrent [FJ24, LLLY23b, PRH21, SL21]. recurrent-event [LLLY23b]. recursive [KJB23]. red [Özk21]. reduced [Büy23, GCF^{+}20]. reduced-bias [GCF^{+}20]. reducing [Daw23]. reduction [GPS23, KM20, WZY22, WN23, ZPW21, Zha24]. redundancies [LZ24]. reference [GABJ23, ZlBVGW^{+}23]. refinements [MVLMA22]. regarding [ZL20]. regenerative [Min22]. region [LLL22]. regional [SLW20]. regions [Bey21, dCCM^{+}22]. Regression [AÖ20a, GGMM22, HHY21, KKK23, MZ22a, WS20, XZS22, YZ23, AF22a, AT23, AA22a, AAA22a, AAF23, AJIAA22, AJI24, AES21, AK22a, AA20a, Alg20, AA20b, AM22a, AK21, AKS^{+}21, Ali22, AHB23, Ali24, AÇ23a, ARAR24, AQYA22, AAA22b, AQAN22, AAM23, AJ22, ARJ21, AK23c, ABB22, ASK23, AY21b, BYK24, BSNT21, BS22b, Bil23, BF23, CWW23, CR21, Cha23b, CFM21, CLJK21, CKCG21, DLRZ21, DYC22, DCY22, DA21, DF22, ESS22, EH22, ENK^{+}22, FQKA23, FLPB20, GI20, GWZC22, GZ22c, GZZ22, Gan22, GAS22, GY23, GGB20, GJK21, EJKL22, EJKL23, HMSRM23, HA24, IHS22, IRS20, KA24, KUA21, KGF20, KUA23, KM20, KHK22, KL23, KJB23, KEE20, KA23, KBB24, Kur21a, Kur21b, Kur23, Lee20, LB23a, LLL22, LY23, LFZ21, LAKA22, MB23a, MZ23, Mac24, MAAA23, MYS22, MAB23, Meg23]. regression [MVLMA22, Mil21, MRB24, MSNN^{+}24, Mur23, NS22, NIN20, dALNPdOF21, Nis21, ONE22, OBN22, OM21, ÖA22, PCWL22, PLS23, PB22, PV20a, PS23, Pho24, PJL21, PRH21, PB23, Pol20, POCC22, POC23, RRA20, ROL^{+}20, RAA23, RAA22, Ric21, RVV22, RBKA23, SSCT23, SGG22, SZ21, SM23a, SANH^{+}22, SL21, SR23, SMB23, SAD^{+}21, SCK20, SCK21, SCA23, TED24, Tak23, TW24, TWTT21, TÇA22, UOA23, UYMK22, ULN21, dOVQP21, VW21, VPOC24, VH23, WLC^{+}22, WSZ23, Wan23b, Wan23c, WZW23, WCY24, WCF21, XS22, YYZ23, YDLX20, YXYX22, YL22a, Yoo23, YL22b, YZ21, ZA23, ZK23, ZZ21, ZCG24, ZG21, ZG22b, ZQYW22, ZRC23, ZZZ23, ZHJ23, ZCW23, Zha23, Zhe21, Zhe24, ZLSC20, ZZ22b, dSdSTC21, dSPdSB23]. regression-type [ASK23]. regressions [CNL21, GPGMK22, dSC23]. regressors [AA22a, Gan22]. regular [HAA^{+}23, JAP21, LAZ21]. regularization [CZP21, HvZ21, ÖA23, ZWTY23]. regularized [EJKL22, EJKL23, LH20]. regularly [dXLY20b]. Rejection [AS23, SDG22]. related [CSG22, DG22]. relation [SRF^{+}23]. Relations [AZAZ23, Gül22, LB22]. relationship [AZH^{+}22, ZZ21]. relationships [Büy23, FS23, WC23a]. relative [CL23, GZ22c, KKO21, PM23, YZ23, ZZ22b]. Reliability [Bar20b, BMG22, CS22a, MWTR23, TKR24, WCKC23, ZX22b, ZYSS22, AS22b, AMK21b, Çet23, Han20, KBYS22, LDA23, LS23, MK21, MSL21, RY22, SMWJ20, TY22, WWT22]. reliable [WCRM23]. Remaining [LCC22]. Removal [Ano22]. removals [Sha22b]. Removing [PFM21]. renewal [LJ22b, PKA22, Yao23]. repair [Ami21, KCS24]. repairable [CS22c]. repeated [AATD21, ARA23, Akd22, KBR^{+}23, RBT^{+}22, dOVQP21]. repetitive [CLS22, KJB22]. replacement [Aub23, PM20]. replicated [LAZ21, Tür20]. replicates [CII22, Gho20]. REPPlab [FBNRG21]. Representation [Xia22, ZL22b]. Representativeness [VDN22]. representing [PK21, SJ23]. resample [WY21]. Resampling [CNL21, Fig23, RM23, NT20]. Resampling-based [CNL21, RM23]. Rescaled [MMVBE23]. Rescaling [BRA20]. research [MMMCS23, ZB22]. reserves [GHW20, WDZ22]. reserving [SIK24]. reservoir [CL22c]. residual [CP23, FL23, KL21, MÖ23, PZ20, RJJ23, Ric21, SJ23, SZJG21]. residual-based [FL23, MÖ23]. Residuals [ZZFG20, AG20, HHvR22, IRA21]. resistant [BGJ^{+}23]. respect [CCM20]. response [AS23, AAK^{+}24, Arn23, AAA20, Brz20, ÇKK23, DLRZ21, GAS22, HAA^{+}23, JBJ23, KS22a, KFM23, KS22b, KVS22, KB23, LZZZ23, Mag20, MASA22, MSB23, NN20, NS20, NS23, Sad24, SBR^{+}23, SKV20a, SU21, SK22c, SBB22, SBB24, Tür20, UM22, VJV^{+}23, WKSeF22, XSS22, ZLXT23, ZZL23, dSLHMB23, dCCM^{+}22]. response-adaptive [JBJ23]. responses [Esk21, GZ21a, LFMMRH22, PG22, PMM23, RSG21, RS24, TW24]. restricted [AK21, Baz22, Gha22, KÖ21b, LW22b, Pol22, TÇA22, XZJ^{+}21]. restriction [Baz20]. restrictions [YA21, YRAB22]. resubmitted [Bal22]. results [LB23a, NS22, TKF^{+}20]. retrieval [Pol22]. returns [IT21]. revenue [LJ22a]. Reverse [XZJ^{+}21]. reversed [BD23, HM21, SK22b]. reversed-hazard-based [BD23]. review [KGR21, MMMCS23]. revisited [GPGMK22, Kal22]. Revisiting [CMMV21]. rewards [LJ22b, Yao23]. reweighted [Zha22b, Zhe21]. rho [TK22a]. Ridge [Gan22, MRB23, RAA23, YRAB22, AF22a, AT23, Alg20, AK21, AKS^{+}21, AQYA22, AQAN22, AAM23, ABB22, AA22c, CWW23, ENK^{+}22, FLPB20, IS20, KUA21, KUA23, KEE20, KÖ21a, Kur21c, LAKA22, LAK24, MAAA23, MÖ23, NS22, ÖA23, PJL21, RAA22, SGG22, SCK20, SCK21, SCA23, VLvdW22, YA21]. ridge-type [AF22a, ABB22, LAKA22]. Riesz [KTZ20]. right [CKK21, KJB23, PV20a, SLSZ22, SVN24, TY22, Wan22, WLZ23, ZWW23]. right-censored [KJB23, PV20a, TY22, WLZ23, ZWW23]. Risk [ÖK21, BDG23, GHW20, GCF^{+}20, Han20, Han22, HB23, HZ22, KOT20, LZ24, LC22, LG21, RAN23, SD22, SHN23, TYLX22, TAA22, XXH22, dXGLY20, dXLY20b, YZ23, ZWMY23, ZC23, Alt20a]. risk-adjusted [RAN23, TAA22]. risk-averse [LZ24]. risks [AK22c, AHG20, CSL20, CSA23, LZQ23, LY24, LTB23, LHG24, MN17, PR22, WLT22, WL22]. RMCriteria [MM22]. RNA [YZF22]. RNA-seq [YZF22]. Robust [AG20, DA21, KS20b, LWZ^{+}23, Lu23, MKK22, Nak21, OM21, PB22, RVV22, SS23, SB21a, dCSF24, TÇA22, WJZ^{+}22, YXYX22, YZF22, ASS22, AKS^{+}21, ASM21, BZ22b, Bul24, CKK21, KN22, KA23, KA22, MAAA23, MA23b, NIN20, ÖA22, PM23, PKW22, Pol20, SM23a, SFG21, SANH^{+}22, Spa22, SCK21, VVNNT23, Wil22, YYZ23]. Robustification [RYA20]. Robustness [HAA^{+}23, OY24, SRF^{+}23, AA23b, DMP23, FO22, SPG^{+}22]. ROC [ANP20, HAT^{+}22, YZF22]. Roger [SF23]. role [ÖT20]. room [ASM21]. root [ABC^{+}23, FPAA21, FL23, GE23, GG22, Hep21b, LS22a]. rooted [JAP21]. rotatable [VSJ^{+}24, VJV^{+}23]. rotated [IT21]. rotation [Cha23a, SPS22]. rounded [ZB20]. rounding [UU23]. roundoff [WN21a].

RRT [KZG21, LKR^{+}23]. RSM [REN22]. ruin [HB23, HCHK23]. rules [Cha23b, CKL22, KA24, KTK22, SMM22]. Run [PAS23, VSJ^{+}24, CKL22, CTK^{+}22, HCTZ21, MKSH22, WCRM23]. runs [CKL22, CKT^{+}20, GK22a, GK22b, Ger21, KA24, KST^{+}22, MKSK23, SKM^{+}23, SNKA24, SMM22]. runs-rules [SMM22].

S [AE23c, MA23b, SKS21]. S-bootstrapping [XW20]. S-Gini [MA23b, SKS21]. Sahalia [ZS23]. Sample [CFM21, GK22a, HHOC23, KS23c, YL22a, Yoo23, BK23, Bil23, CZP21, Che21, CMMV21, DKL22, Ere20, Gai21, GBHA20, GBHA21, GK20, GWW^{+}23, Ham22, HHH23, HZY22, KM22, KJ20, LHG24, MK21, PKW22, PV20b, SS20b, SCK23, SF23, TF22a, TCJ23, WZY^{+}21, WWTM22, Wan23a, WJB21, WY23, WS21b, ZPW21]. sampled [Mos22]. sampler [ALZ24, AA20c]. samples [ASZN21, BRTW24, BQCP23, BKM23, BRA20, CLY21, CW23, DC24, LBL22, LWP24, LY20, MG21, Min23, SMM22, Wu21, WS22, ZB24]. sampling [AS22b, AKG23, AOA23, AÇ23a, ASA24, AMA^{+}23, AA22b, ARJ20, ABPK21, ABJ21a, ABJ21b, ARKA22, Aub23, BJU20, BNA23, BC23b, BKPS23, BMG22, BFS22, CK20, CD23, CLY21, CLS22, CL22c, DR22, El 24, FLXY22, HUK23, KJB22, KM22, KS22a, KGF20, KK22a, LK21a, LKHC23, LTJC23, LB23b, MNL21, MZ22b, MS22, MMMG21, MKSH22, MR22, MS20, MMMCS22b, MMMCS23, MYVD21, MSB23, MKS22, NS23, NN21, NT20, uASS22, OKO21, PP23, Pan20b, PN21, PDK23, PMK20, PMK23b, RSM22, RAJ21, RGA21, RuA22, RTJ24, SNSE22, SSGH22, STdL21, SZ21, SA22, SIG23, SAA^{+}23, SB21b, SKK21, SK22c, SBB22, SBB23, SYPV23, TdSCdSdCP22, TBK23, VDN22, WCY24, WWW24, WCKC23, XSS22, YCD23, YL23, ZD22, ZZL23, ZB22]. sampling/importance [NT20]. samplings [Min22]. Sankaran [Ano24]. SARIMA [AT21]. SAS [GGS23]. satisfying [Kaz23]. saturated [BDK^{+}23]. Saunders [BA22a, CLZ23a, KJB22, LMFMA23, LBH23, Sha23]. Scalable [WY23]. scale [CS22c, FBL22, GK20, GSG22, HW24, KAVdSFG22, MG21, MCWL22, MB21, PKW22, RM23, TJ20, VG22, Wan23a, Wil22]. scaled [DG23a, QTH24]. scaled-uniform [DG23a]. scales [BBD^{+}24, BH20, ZlBVGW^{+}23]. scaling [CSG22, PP23]. scan [dCdAB23]. scatter [Kal21]. scattering [KKZ21]. scenario [KSP23]. scheme [ASA24, BJU20, DSS23, GK23, MMMCS22b, PAS23, PMK23b, Seo22, Tei22, TP22a, YSG23]. schemes [BB24b, CKL22, MLR^{+}22, Mis20, MR22, MMMCS23, RSM22, RuA22, RTJ24, SMM22, SYPV23, Wan22, WG23, Yas22, ZL23a]. Score [BH23a, GSG22, TIK24, BAL24, GNE21, KT23, Nag21a, OBN22, VDN22, DS24]. Score-adjusted [TIK24]. score-driven [BAL24]. scores [CLZ23b, HAT^{+}22, SZS20]. Scott [NAJ22]. scramble [SKV20a]. scrambled [MSB23, NS23]. scrambling [SSGH22]. screening [CLC22, Fan22, SX22, SLSZ22]. sea [TF22b]. Second [MHW24, dXLY20b, Che21, DRB20, LJ22a, Zha23]. Second-order [MHW24, dXLY20b, Zha23]. seemingly [ENK^{+}22]. seismic [Lin22]. select [dALNPdOF21, Wes20]. Selecting [Fuk22, Spa22, SY22, SLJ23]. Selection [Ami21, AM24, AES21, Alg20, ALVQ22, Aub23, AM22b, AY21b, BDK^{+}23, Cha23a, CGS21, CLL20, CCBN22, Ciu22, DC23, DS24, EGVdR21, FW23, FPAA21, GPGMK22, EJKL22, IGA24, KS23a, KS22b, KÖ21a, Kur23, LYY21, Nak22, NS22, NN22, ÖA23, REMO22, RTJ24, RBKA23, SK23, ULN21, WJZ^{+}22, WZW23, WKcIC24, WFLS24, WLZ23, XH21, YOD^{+}23, ZRC23, ZZ22a, ZWCY24, ZCH23]. selector [Som22]. self [BS23b, CCH20, DuA23, HYSC21, LZQ23]. self-exciting [BS23b, LZQ23]. self-organizing [CCH20]. self-starting [DuA23]. self-updating [HYSC21]. seller [MY23]. selling [DE22]. semantic [MMR22]. Semi [HB23, HMSRM23, HZY22, Hut22, RftADNI22]. semi-Markov [HMSRM23]. Semi-parametric [HB23, Hut22, RftADNI22]. semi-varying [HZY22]. Semidefinite [TWSW23]. Semiparametric [FZZ20, Wan23b, ZT20, ZLTP23, AG20, ALB22, BS22a, BS23b, CST21, DY22a, Esk21, LW22a, LW22b, PCWL22, POCC22, POC23, SB21a, SL21, WP22, WLC^{+}22, WSZ23, WLZ23, ZWCY24]. sensitive [AE23a, DS24, GK22a, GK22b, KZG21, KVS22, MKSK23, MMMCS22b, NS23, SKM^{+}23, SKV20b, SKV20a, SSK22, TPK22, WKSeF22]. Sensitivity [ALVQ22, CLY23, FLXY22, FPAA21, NS20, RSG21, UU23]. separation [MRB24, ZZ21, ZNKN22]. seq [YZF22]. sequence [SN22a]. sequences [GDWY22, LZSC23]. Sequential [CCBJ21, FTC21, LTN23, ON21, AAA20, FS23, LDA23, LWX23, Lui22, MHW24, PDK23, VSJ^{+}24, Wu22, YW24, ZB22]. serially [LES21, SM20]. series [ADK21, AS21b, AKR23, Bap20, Bap22, BA22c, CL21, Gho20, GTW22, HCLF23, HE24, dAIIS20, KKO21, KuIK20, KL23, LDA23, LL23, LW23, LZ22, LY23, Mac24, MPS21, OLOK22, RL21, RSG21, RSG22, RK20, SNA20, SI23b, SLE20, TED24, WS21a, YH21b, dSdSTC21]. server [ALZ24, BCDA21, BC21]. service [EAJ20, LJ22a]. servicing [JD23b]. set [AS22b, AOA23, AA22b, BC23b, BRA20, KK22a, MZ22b, MLR^{+}22, MR22, RuA22, RTJ24, STdL21, SZ21, TdSCdSdCP22, VDN22, WCY24, YCD23]. SETAR} [HK22]. SETINAR} [MB23b, MB23c]. sets [EFD21]. setting [AFJ23, CH23, WWTM22, XH21, YWT22]. settings [Ega24, YWL^{+}23]. seven [Uya22]. Several [TII22, AE21, GKS24, Gök22, KLM24, MJ20, PD21, SB24, SS20b, SS20a, TNN20, WMS22]. Shannon [CQ22, LK21b]. shape [DY22a, MMR22, TIK24, TJ20]. shaped [BMG22]. shapes [SCK22]. shared [KKK23, PBLM22]. Shewhart [GM24b, SMM22]. Shewhart-type [SMM22]. shift [RYA20, WZ21a]. shifts [ACKL23, GK22a, GK22b, MMMCS22a]. Shiryaev [LPP21]. shock [BE20, ZX22b]. short [CKL22, KST^{+}22, SNKA24, TPC^{+}23]. short-term [TPC^{+}23]. shortfall [BDG23, SHN23]. Shrinkage [Alg20, LMP20, VG22, ABB22, BB24a, FLPB20, LW21, LXMT22, ÖK21, PLA22, dSS22, ZBB23]. Shrunken [SWG24]. Side [SKM^{+}23, GK22a, GK22b, MKSK23, MMMCS22b]. side-sensitive [MKSK23, MMMCS22b]. sided [HQZ^{+}23, RAJ21]. Sigma [KR21]. Sigma-based [KR21]. sigmoidal [Hut24]. sign [CLS22, GM24b, SB21b]. Signal [BAL24, TP22b, BLXL24]. signals [Lin22, SMO21]. signature [SZ23]. signed [CS22c, KE22]. signed-rank [KE22]. significance [Bil23, PLBP24]. significant [Gan22, XS22]. similarity [FSGMM23, REMM22]. simple [BB24b, BKPS23, Bil23, BCR^{+}21, HBS22, KA24, KN22, RAG21, REN22, SNSE22, TK23, WCY24, YCD23]. Simplex [CR21]. Simpson [SZJG21]. Simulating [TF22b, Wan22, Pan20c]. Simulation [Ano24, AE23c, FPAA21, Kri22, LLZ22, NAKK20, SDG22, AT23, Ali22, Ali24, AAHR20, AQAN22, AS21b, AK23c, ASM21, BBD^{+}24, Baz22, Bil23, BCR^{+}21, CMMV21, CLZ23b, Dog22, FQKA23, GZY22, GABJ23, GKZ^{+}22, IM20, KAVdSFG22, KKZ21, KOT20, MG24, MN17, ÖA23, PR22, Pol20, SEMR23, SLP22, TL23b, TEO20, UU23, Wan23a, WMS22, YLX^{+}22, ZLG23]. simulation-based [Ali24, CMMV21, IM20]. simulation/computation [KAVdSFG22]. simulations [GNE21, Som22]. Simultaneous [Par20, RAG21, RLP23, ZK20b, KVS22, RuA22, RM23, UIC23, WKcIC24, ZLTP23]. simultaneously [SK24]. Singhasomboona [Kri22]. Single [GZ22c, WWW24, ZXZL22, ALZ24, AA20b, BC21, BH23a, GZ22a, JKLK24, JGL22, Kur23, MNN21, MS20, Nad21, Pap20, PHS21, RGA21, Sha20, ZNLW20, ZGL21, ZXF22]. Single-index [GZ22c, ZXZL22, AA20b, GZ22a, JKLK24, JGL22, Kur23, Nad21, ZNLW20, ZGL21, ZXF22]. single-item [BH23a]. single-mediator [PHS21]. single-sampling [MS20]. single-server [ALZ24]. Single-stage [WWW24]. SingleCross [Tel22]. SingleCross-clustering [Tel22]. singular [AMP23, RM20]. SIR [JAP21]. situations [SK22c]. Six [KR21]. size [BQCP23, CFM21, CZP21, DKL22, HHOC23, Har21, KM22, KS23c, TF22a, TL23a, YL22a, Yoo23]. sizes [ARA23, Bil23, PLBP24, WWTM22]. skedastic [AS23]. Skew [FZM21, CWW23, CWWT24, CSDT20, FBL22, GZ21a, HR22, HK20, IT21, KAVR20, KAVdSFG22, MCWL22, OR23, ON21, PHN22, RY22, TJ20, WTN22, WWTM22, WZW23, XS23, YWL^{+}23]. skew-elliptical [RY22]. Skew-normal [FZM21, CWW23, CSDT20, HR22, MCWL22, WZW23, YWL^{+}23]. skew-normal-Cauchy [KAVR20, KAVdSFG22]. skewed [AS23, BFAR20, RYA20, VH23]. Skewing [Nis21]. skewness [PFM21, Ric21]. skill [GKZ^{+}22]. skip [El 24]. skip-lot [El 24]. slash [WTN22]. Slice [ZZL23]. sliced [HGWZ22]. slope [AZH^{+}22, IRS20]. slowly [TPvdH23]. Small [FHM22, IC21, KJ20, ZPW21, AC23b, MW20, MSS20, SF23, WJB21]. small-sample [SF23]. Smirnov [BK23, YZF22, ZZFG20]. smooth [Hep21a, Hep21b, MYS22]. Smoothed [RJJ23, FM23]. smoother [AS21b]. smoothing [AALR23, BAL24, DL20]. Sobol [Sol21]. social [FAT23, MASA22, SCHS22, ZSM23]. Software [Bar20b, GGS23]. solution [BYS24, SCHS22]. solutions [GHW20, Xia22]. solvers [EJKL23]. Solving [Ric21]. Some [AAA20, EFD21, GK22b, KAVdSFG22, KS22a, LB23a, ÖT20, SKV20a, SKK21, SU23b, SR23, Squ22, WZ21b, AKS^{+}21, AALR23, Bar20a, BAEk21, BA22b, Büy23, FS23, GHW20, KS20c, Kur21a, Kur21b, Lue23, MMMCS23, NBC23, OBN22, ÖK21, ÖÇ23, PAT22, WDZ22, ZSM23]. Space [DSO^{+}23, JGGDMRLD21, MMR22, Nag21a]. Space-time [DSO^{+}23]. spacing [NDWE24]. spacings [Pak23]. Sparse [GW20, LS22b, PK21, WCF21, AS22d, KBB24, Kur21b, Mos22, PB22, SFG21]. Spatial [dCdAB23, dOMD^{+}21, AC23b, BJ20, DK22a, DE22, FZM21, FSGMM23, Gao23, GO22, HK20, LD23, LW22b, MWNSB23, PC22, SB21a, SMB22, Ver21, WHW21, ZLF23, ZWCY24]. spatial-temporal [SB21a, SMB22]. Spatiotemporal [GO22]. SPC [KBR21]. SPDE [Ver21]. Spearman [TK22a]. special [KS20a]. specific [ZlBVGW^{+}23]. specification [CNL21, GMC23, Hu23, SBZG22]. specified [Mil21, WWTM22]. spectra [Pic21]. spectral [SMB22]. spectrally [LC22]. spectrum [AMP23, RM20]. speed [SMB22]. Speeding [Zhe24]. sphere [VWZ21]. spherical [JKLK24, TA24]. spikes [MMVBE23]. spillover [CS20]. spiral [BCC24]. Spline [LY23, CLJK21, KSAR21, YZ21]. splines [JKLK24, Mur23, YDHW21]. split [SY22]. spotting [TBH22]. spread [LWC22]. spreading [HCLF23]. square [AS22d, Che22, Han21, KS23b, Zhe21]. squared [MSS20, ZHJ22]. squares [DRB20, KLM^{+}22, MT23, SP23, WY21, WLY22, WZZY23, ZA23, ZG22b, Zha22b, AK23c]. stability [BV22]. stabilizing [Nis21]. stable [Bee22, Lin22, Moh22, SU22, ZKW24]. Stage [Bar20b, ARKA22, DP23, EAJ20, HZ22, KM20, Mac24, NN20, OM21, SMG23, WWW24, Wu21]. Stan [dSLHMB23]. Standard [Bou23, AE23b, BRTW24, GABJ23, Han21, WN23]. Standardized [Gök22, MGS22, SI23a]. starting [DuA23]. state [AMA^{+}23, ARJ20, ABPK21, ABJ21a, ABJ21b, EN22, JGGDMRLD21, Nag21a, RAJ21, SS21b, SZ23, ZYSS22]. Statement [Ano22]. Static [BH23b]. stationary [DRB20, DE22, GE23, HI22, KKO21, LPP21, MWNSB23, PAS23, SI23b, Wan23b]. statistic [BGM20, Bul24, FL23, KUA21, KL21, MK22, SX22, TK22a]. Statistical [AHG20, Biç20, CHC20, DCY22, FK22, LW22b, LHG24, MB21, MASA22, PSL23, PIA22, WG23, BCC24, GO22, GKZ^{+}22, HE21, HLHC22, HLT23, KL21, LKHC23, Mur23, NKSY22, OP20, RAN23, REN22, SRN^{+}22, Squ22, ZlBVGW^{+}23, ZB20, ZWC^{+}23]. statistically [BLXL24]. Statistics [Ano24, AE23c, Kri22, AZAZ23, ATYK20, AALR23, ARJ21, AAGP23, ARKA22, BRTW24, BC23b, Ere20, GM22, GK20, HR22, LB22, MX24, ML22, NJ22, SZS20, Sun23, ZLG23, dCdAB23]. Statistics-Simulation [Ano24]. status [HHH23, WFLS24, XZS22]. steady [EN22]. steady-state [EN22]. Stein [AAA22b, LD23]. Step [SS20a, Bey21, CSL20, GAS20, HB22a, Mis20, RYA20]. step-stress [CSL20, HB22a]. Step-up [SS20a]. stereo [OP20]. Sterling [WZZY23]. stick [DK22a]. stick-breaking [DK22a]. still [Bia22]. stimuli [MMR22]. Stochastic [Gha22, JGGDMRLD21, KKO21, KOT20, KÖ21b, WC23a, AK21, CS22a, ESLK22, JAP21, KKZ21, LX22, LYP22, eLhS23, LZ24, MF22, RRA20, SF24, Squ22, TEM20, XS23, YA21, YRAB22, YC22]. stock [CL21, DPS22, LZQ23, SI23b]. stocks [BH23b]. Strategies [GAS22, FO22, MHW24, PLA22, SCK22, ZBB23]. strategy [eLhS23, LWX23, LG21, SPS22, TYLX22, ZC23]. Stratified [Pan20b, BFS22, CK20, CD23, CL22c, GY23, HUK23, MYVD21, MSB23, MKS22, QLG23, SA22, SIG23, SAA^{+}23, SBB23, TBK23]. streams [CL22c]. strength [AS22b, Çet23, HORA22, LS23, MWTR23, RY22, SU23a]. Stress [RY22, AS22b, CSL20, Çet23, GAS20, HB22a, LS23, MWTR23, NDWE24, SU23a]. Stress-strength [RY22, AS22b, Çet23, LS23, MWTR23]. stringent [KuIK20, uIKKH22]. strong [DC24, SB24]. strongly [AATD21, KBR^{+}23, RBT^{+}22]. structural [AZH^{+}22, De 20, Dog22, FW23, KS23c, dSC23]. structure [Bia22, CZP21, HAA^{+}23, JM20, PK21, RYA20, dCSF24, SAD^{+}21, SLJ23, UKK20, Wes20, ZL22a, ZLXT23, ZK20b]. Structured [GZ21b]. structures [GZ21a, OR23, ROL^{+}20, RH23]. student [ML22, SN22b, BA22a, DDK21, NIN20, XS23]. studies [KT23, Sha22a, TWTT21, XWL22, YOD^{+}23, ZL22a]. Study [FPAA21, Han22, Kol24, ASA23, AFJ23, Ali22, Ali24, AR22, ANP20, AK23c, AE21, BA22a, BCC24, BBD^{+}24, Bay21, Baz22, Bil23, BCR^{+}21, CKK21, CLZ23b, DMP23, Dog22, FSGMM23, GZY22, GABJ23, EJKL22, GKZ^{+}22, KF24, KOT20, KGR21, LWP24, MNL21, MPS21, NN20, NN22, ÖA23, PR22, Pol20, Ras20, SEMR23, SPG^{+}22, SU23a, SMB22, SW23, SMB23, SSS23, SVN24, TEO20, UU23, Wan23a, WMS22, YWL24, YLX^{+}22, YW24, ZWY20, ZLG23]. style [BDL23]. sub [KL22]. sub-bifractional [KL22]. Subgroup [ZA21, Spa22]. subgroups [YHGA21]. subject [KJK20, WKcIC24]. submodels [Ali24]. subsample [WMXL24]. subsampling [LLZ22]. subset [AGA20, CKCG21]. subspace [EJKL23, PK21]. success [ML22]. successive [KS22a, MSB23, NS23, SKK21, SK22c, SBB22, TPK22]. sufficient [WZY22]. sum [CKCG21, KS23b, LX22, MKSH22, TPK22, ZHJ22]. sums [LHY^{+}23, NZ21, ZQYW22]. superiority [HHOC23]. superposition [RdAIS^{+}20]. supersaturated [BDK^{+}23]. supplementary [MKS22, SMM22]. Support [Nak22, AWZ24, Ega24, FN20, KL23, KBB24, MZ22a, MT21, MT23, MM22, Nak21, Zhe21, Zhe24]. supported [WJB21]. SUR/neural [IRA21]. sure [KL22]. surface [AAA20, HAA^{+}23, VJV^{+}23, dCCM^{+}22]. surgery [ASM21]. surrogate [HZY22]. surveillance [GO22]. survey [BJ20, GGS23, KK22b, PHS21, RH23]. surveys [IC21, LFZ21, QTH24]. Survival [RY21, AS22a, AE23c, BYK24, BBD^{+}24, BCR^{+}21, CSBdES22, CST21, CLC22, FZZ20, HLHC22, KKH17, LK21b, LSTE24, LYZ^{+}21, MZ22a, NCC^{+}22, Sad24, SMG23, TAA22, Wan22, WSZ23, YBC22, YW24, SVN24]. surviving [HG22]. survivors [FZZ20, ROL^{+}20]. survsim [MN17]. Sv [Wij23]. Sv-plots [Wij23]. swaps [CHL22b]. swarm [ASM21, IS20, SI20, dOMD^{+}21]. switching [KHK22, OKO21, SEMR23, TMZL21]. symmetric [AGI23, Lin22, MZZA24, UEA20, VWZ21, ZK20b]. symmetrized [HQN22]. symmetry [AS22d, GZ22b, GZ22a, GWZC22, GZZ22, Hut22, Kal21, NJ22, RP23, dCSF24, ZG22a, ZGL22]. synthetic [NKSY22]. system [BC23a, BQCP23, KFB23, KBYS22, MM22, PY20, SZJG21, TKR24]. systematic [MNL21, MG24]. systems [Ami21, CS22c, KCS24, KS20a, LHY^{+}23, LZ24, PM20, RAN23, SNA20, SF24, ZYSS22].

t [SZ23, RAA23]. T-K [RAA23]. t-signature [SZ23]. table [HLT23, KJK20]. tables [AS22d, CH23, WK23]. Taguchi [Hua23]. tail [Hwa23, KR21, dXLY20a, dXLY20b]. tailed [AMD22, GI20, GWW^{+}23, QYC22, VH23]. tails [BGM20, GO22, dXLY20a, dXGLY20]. tapering [KA23]. target [SRN^{+}22]. taste [MMR22]. teaching [SN22b]. technique [BRA20, EE23, MSB23, SG21, SBB24, TPK22]. techniques [GBHA20, NS20]. temperature [AMP23, KKO21, TPC^{+}23]. tempering [SCK22]. temporal [SB21a, SMB22]. temporally [LW23]. tensor [MT23]. term [Bia22, FZZ20, IRS20, ROL^{+}20, TPC^{+}23]. terms [AK23c, Cav24]. Terry [SDG22]. Test [DY22a, YZ21, AS22a, AF22b, ABC^{+}23, ABPK21, BGM20, Bay21, Baz22, Bul24, CSL20, Çet23, CP22, CS22c, DZLS22, DBS23, DR22, DKL22, EHHS20, Ere20, FS22, FS23, FL23, GZY22, GABJ23, GAS20, Gök22, GG22, GSG22, Ham22, HB22a, HM21, Hep21b, Hu23, HZ23, KF24, KJB22, KuIK20, uIKKH22, KL23, KLM24, LSC23, LL23, LTN23, Lue23, MQ21, MX24, MK22, ON21, PHY20, Pho24, Sul22a, TK22a, WTN22, XW20, XS22, YZF22, YWL^{+}23, YL22b, ZLG23, ZLSC20, dSPdSB23]. Testing [EH22, FSGMM23, GZZ22, GKS24, Hep21a, Lar20, MEA24, Pap20, PD21, PB23, RP23, YSH23, ZGL22, ATYK20, Baz20, Bil23, CWLH23, CH23, DR23, DBS23, Ger20, GK20, HW24, LW23, LS22a, NJ22, Pak23, PFM21, SBTP20, Sul22b, TA24, UY23, WWYX21, War23, Wij23, WC23b, YV20, YW24, ZD22, ZK20b]. Tests [Bay21, AS23, ABHF22, Akd22, AG21, AF22b, AÖ20b, AA22b, AE21, BK23, BPN22, BKM23, CYS21, CYS23, Cav24, CH23, CL22b, CSA23, CMA21, FM23, FS23, FPAA21, GZZ24, GSG22, Hor22, HHH23, HS23, Hwa23, KLM24, KGR21, LBL22, MG21, Mil20, MAF23, NdSA20, NM21, OP22, PLBP24, Par20, PV20b, dCSF24, SN23, TCJ23, UU23, Uya22, WY23, WYY22, WMS22, ZZFG20, ZLG23]. Their [FO22, BR21, BE20, Gan22, Jhw24, KBR^{+}23, LX22, LB22, NYM20, RBT^{+}22, YHWW23]. theorems [Xia22]. theoretic [PMK20]. theory [Brz20, DP22, Mag20, SBR^{+}23, TF22a, UM22, ZZL23, dSLHMB23]. there [AM21]. thick [BGM20, VH23]. thick-tailed [VH23]. thinning [SM20, WWYX21]. third [AAA20, VSJ^{+}24]. Thompson [IMR21]. Three [DR22, ASZN21, Ham22, HLT23, OP22, PMM23, PMK23a, WH24]. three-arm [WH24]. three-category [HLT23]. three-component [PMM23]. three-parameter [ASZN21, OP22, PMK23a]. three-sample [Ham22]. threshold [AHB23, BS23b, LLZ22, MV22b, WWYX21, YH21b, ZC23]. thresholding [SWG24, SD22]. thyroid [HAT^{+}22]. tie [KK22a]. tied [BYK24]. tightened [BJU20]. tightened-normal-tightened [BJU20]. time [AHAMP23, AK20, ADK21, AS21b, AKR23, Bap20, Bap22, BBD^{+}24, BH23b, BA22c, BLXL24, BCR^{+}21, CMM23, CHC20, CL21, DuA23, DSO^{+}23, Gho20, GTW22, GZ21b, GWZ23, HHOC23, HCLF23, HQZ^{+}23, HE24, HCHK23, KKO21, KuIK20, KCS24, KL23, LHY^{+}23, LL23, LW23, LC22, LJ22b, LY23, Mac24, MAMA20, MZZF21, MPS21, NCC^{+}22, OLOK22, QTH24, RHG22, SKG22, SI23b, SS21b, SLME23, SLE20, TED24, TAS24, TBH22, TAA22, VBB23, WCC22, WS21a, Xia22, XZS22, YZ23, YH21b, ZZ22a, ZWW23, ZHJ22, ZBZS20, dSdSTC21]. time-constrained [ZBZS20]. time-course [VBB23]. time-dependent [BCR^{+}21, MAMA20]. time-ordered [CMM23]. time-scaled [QTH24]. time-scales [BBD^{+}24]. time-series [KKO21]. time-to-event [HHOC23, RHG22, VBB23]. time-varying [BLXL24, GZ21b, MZZF21, NCC^{+}22, ZZ22a]. timed [SMO21]. times [BYK24, CST21, CsSL22, GHW20, KKK23, LJ22b, NCC^{+}22, PIA22, SL21, Wan23c]. Tobit [ZLTP23]. Toeplitz [JM20]. tolerance [GWW^{+}23]. toll [ZfSRJ22]. Too [Mil20]. tool [BH23a]. tools [DG22]. Topp [AMK21b, BR24, LB23b]. TOPSIS [REN22]. total [ASA24, WY21, WLY22, WZZY23, Zha22b]. toxicity [YH21a, ZWMY23]. trace [Lin22, Ava21]. traditional [AF22b, CLZ23b]. traffic [BC23a, SMO21]. trait [Jhw24]. trajectory [FLXY22]. transaction [ZC23]. transcendental [ÖÇ23]. transfer [KKZ21]. transform [LLLY23a]. transformation [CST21, EHHS20, JH23, LK24, MEA24, PFM21, SSS20, SL21, WLZ23]. transformations [EH22]. Transformed [MGS23, AMA22, Gül22, Pan20c, Pap20, SAA21]. transition [Hep21a, KKK23, MYS22, SMMO22]. transitions [Hep21b]. translog [ÖÇ23]. transmission [KM20, SKG22]. Transmuted [AE23c, KKH17, TS22]. transplant [LBH23]. treatment [IM20, KS22b, LWZ^{+}23, MPD20, MHW24, SEMR23, SK23, Tak23, WGlTL23, WH24, ZA21]. treatments [KS22b]. tree [BVJ^{+}22, MES21, SWC^{+}20]. TreeNet [Cha23b]. trees [CQ22, JAP21]. trend [BGJ^{+}23, CMM23, Gho20, Ham22, JGGDMRLD21, NHZ22, PKA22]. trends [TPvdH23]. trial [Arn23, JBJ23, WH24, ZL20]. trials [HHOC23, Med22, SRN^{+}22, Yad22, ZL23c]. triangular [FS23]. triangulation [NAJ22]. Trimmed [Aki24]. trimming [LB23a]. triple [CKL24]. trivariate [dOMA21]. true [Bis22]. truncated [CHL24, HvZ21, KJB22, LTR21, MAMA20, MB21, ONS22, ZX22b]. Truncation [AY21b, BPN22, Gho23]. Tsallis [CP23]. TSLx [Alt20a]. Tukey [MFPP21]. Tukey-type [MFPP21]. tumor [RdAIS^{+}20]. tuning [KLM^{+}22, SLP22]. Turn [ALZ24]. Tweedie [AK23a, ZQY22]. Twenty20 [BT20]. twin [MZ22a]. Two [AAA22a, ARKA22, Bar20b, DP23, ENK^{+}22, GB22, TCJ23, AKE23, AA22a, AE23b, AB22, BK23, BR24, BPN22, CST21, CD23, CS22a, Che21, DNWG22, Ere20, FQKA23, Gül22, HZAF22, HHH23, HZ22, HLHC22, IM20, dAIIS20, IC21, KS22a, KM20, KJK20, KG22, Kri22, KTK22, KVS22, LA23, LAZ21, LF24, LWP24, LS23, MG21, Mac24, MJ20, MNN21, MF22, MD22, MAF23, OY23, OM21, OP22, PMM23, SNA20, SSGH22, SIG23, Sha22b, SKG22, SZ23, SKK21, SK22c, SPS22, SPV22, SMG23, SN23, TII22, UYMK22, WZY^{+}21, WY23, Wil22, WS22, YH21a, YSH23, YLX^{+}22, YGL22, ZK23, ZB22, Biç20]. two-agent [YH21a]. two-component [TII22]. two-dimensional [DNWG22]. two-factor [MF22]. two-level [LAZ21]. two-occasion [KS22a, SKK21, SK22c, SPS22]. Two-parameter [AAA22a, ENK^{+}22, AA22a, FQKA23, LA23, LS23, MJ20, OP22, Sha22b, ZB22, Biç20]. two-phase [CD23, MD22, SSGH22, SIG23]. Two-sample [TCJ23, BK23, Che21, WZY^{+}21, WY23]. Two-Stage [Bar20b, ARKA22, DP23, HZ22, KM20, Mac24, SMG23]. two-state [SZ23]. two-type [ZK23]. two-unit [SNA20]. two-way [AE23b, KJK20, YGL22]. twofold [TY22]. Type [AS22b, AF22a, ALMA23, AA20a, AS22d, AJ22, AB22, ABB22, AA22c, ASK23, BB24b, Bul23, CK20, CKK21, CW23, DKKA22, DSS23, ESS22, GPP21, GAS20, GM22, GK23, KB21, KG22, LSTE24, LTR21, LAKA22, LAK24, LS23, LHG24, MK21, MX24, MBBS22, MFPP21, PV20a, PK22, QTH24, SNSE22, SZ21, Seo22, Sha22b, SMM22, SBB22, TRF^{+}22, WG23, WKSeF22, WCRM23, WC23b, WCKC23, YBC22, Yas22, YWLD20, YSG23, ZK23, ZS23, AZAZ23, AMK21b, BR24, CYS21, CLY21, HB22a, KP21, Lon23, LS23, MG21, MWTR23, MJ20, NDWE24, PHY20, Pak23, PMK20, PMK23b, SBTP20, Tei22]. type-I [MBBS22, CLY21, HB22a, Lon23, PHY20, PMK20, Tei22]. type-II [AB22, DSS23, GK23, KB21, KG22, MK21, Seo22, WG23, YBC22, AZAZ23, AMK21b, BR24, KP21, MG21, MWTR23, NDWE24, Pak23, PMK23b, SBTP20]. types [WHW21].

U [ALZ24]. ultrahigh [Fan22, SLSZ22, ZWTY23]. ultrahigh-dimensional [Fan22]. Uludag [AE23a]. unbalanced [BB24a, Sha22a, YGL22, ZQY22]. unbiased [AK21, AQYA22, DRB20, HM21, Hut22, LA23, VW21]. unblinded [TF22a]. uncensored [POC23]. Uncertain [Yao23, LZZZ23, LY22, LJ22b, MV22b]. uncertainty [BCDA21, KCS24]. Unconstrained [vO21]. underlying [Kal21]. undue [ZWMY23]. unequal [Aub23, Cav24, KGF20, PLBP24, ZC22]. unidimensional [UM22]. unified [LLL22, LKR^{+}23]. Uniform [DC24, MMdOC23, DG23a, HQN22, LFPL24, OY24]. unifying [NSlRGL23]. unilateral [QLG23]. unit [BH20, EHF23, FPAA21, FL23, GE23, GG22, Hep21b, LS22a, SNA20, VJV^{+}23, AAGP23]. unit-weighted [BH20]. units [KK22b]. univariate [Lin22, Pan20c, ZX22a]. Universal [Med22, AHB23]. universities [SBR^{+}23]. unknown [ARAR24, HCTZ21, JD23a, MCWL22, MMMCS22a, Sha20, TCJ23, WH24, ZLF23]. unobserved [WGlTL23]. unrelated [ENK^{+}22, NS20]. unreliable [BCDA21]. unstable [MB24]. until [LC22]. updating [HYSC21, XH21]. Upper [ZWC^{+}23]. UpSet [GGS23].

Use [SF23, AS22c, LW23, SN22b, SLJ23, Wes20]. used [Lui22, Uya22, ZSM23]. useful [ANP20, Bia22, LCC22, OM21]. Using [JBJ23, Pol22, ZBB23, AMA22, AS21a, AMK^{+}21a, AJIAA22, AMP23, AOA23, AS22c, ARAB21, AM21, AAM23, ARJ21, ASA24, Arn23, AMK21b, ABPK21, ARKA22, AY21b, ASM21, BCDA21, BJ20, Bey21, BT20, BPP20, BCR^{+}21, BZ22b, Cha23a, CD23, CGS21, CLZ23b, EH22, EHHS20, EE23, ENA^{+}23, FN20, FBNRG21, GZZ22, Gho20, GGMM22, GKS24, GGS23, GK20, HAT^{+}22, HBS22, HZAF22, HvZ21, Hua23, Hut21, dAIIS20, JuAH20, JKLK24, JHR21, KR21, KS23b, KZG21, KCS24, KBR21, KK22b, KA22, Lee20, LJC^{+}20, LTJC23, LLLY23a, MT21, MX24, MRM23, MR22, MV22a, MKS22, NAJ22, NS20, NS22, NIN20, NCC^{+}22, NG22, uARS22, uAJHD22, uAA23, uAAF23, OBN22, OP20, PHY20, QTH24, RAA23, RftADNI22, RAJ21, RSS22, RM23, REN22, Sad24, SS23, SSS20, SSGH22, SZ21, SIG23, SI23b]. using [SLME23, SRO20, SRF^{+}23, SS20b, SKV20b, SVM20, SVK22, Sul23, SHN23, TAS24, TJ20, TK22b, TY22, TPK22, WZ21a, WZW23, WKSeF22, YZF22, YP23, YCD23, YOD^{+}23, YWLD20, ZB22, dSLHMB23, dOMD^{+}21]. utilization [Wes20]. utilizing [DCL^{+}24, SANH^{+}22].

vacations [LJ22a]. Valid [HLT23]. validation [Alg20, HZY22, LY22, MAF23, ZCH23, dSC23]. validity [FL23]. value [BGM20, BRTW24, BDG23, CW23, CS22b, CL22b, GCF^{+}20, KGR21, PKA22, SHN23, VBB23, dXGLY20, ZMNL23, Alt20a]. value-at-risk [GCF^{+}20, SHN23, dXGLY20, Alt20a]. valued [BA22c, BSAS22, LL23, SSZ23, WWYX21, OB21]. values [ABC^{+}23, Cav24, CBRL22, NSlRGL23, UY23, ZlBVGW^{+}23, ZWC^{+}23, ZMNL23]. VAR [Bey21, GTW22]. vari [Bal22]. variability [BDL23, CKL24, PK22, SB21b, ZBA23]. Variable [BDK^{+}23, EGVdR21, KS23a, Kur23, uASS22, REMO22, ULN21, WZW23, WFLS24, AY21a, AES21, AMA^{+}23, ARJ20, ASK23, BPP20, ÇKK23, CW24, Ciu22, GAS22, GPGMK22, EJKL22, HZY22, KM22, KJK20, LKHC23, LYY21, MKSH22, MMVBE23, uARS22, RAJ21, RBKA23, SKG22, SB21b, SKV20a, SSK22, dOVQP21, WZ21a, WJZ^{+}22, WKcIC24, WKSeF22, WCF21, WLZ23, YOD^{+}23, ZG22a, ZZWH20, ZWTY23, ZWCY24]. variables [Bar20a, CD23, CGS21, HK20, JL24, KZG21, KVS22, LHY^{+}23, Lue23, MAF23, MKS22, NN21, NZ21, SJ23, SIG23, WS20, WX21, Wil22, YHWW23, ZQYW22, ZLXT23, ZWTY23]. Variance [XS22, AM22b, BRA20, CK20, CCM20, DP23, Gho20, Gök22, GTW22, HUK23, Har21, HGWZ22, JuAH23, KM20, KBR21, LW23, MRM23, MMVBE23, Nis21, uAA23, uAAF23, QTH24, SNSE22, SANH^{+}22, SAA^{+}23, SK22c, VLvdW22, WZ21a, WS21b, XSS22, YGL22, YW20, dSC23]. variance-covariance [SANH^{+}22]. Variance-estimation-free [XS22]. variance-stabilizing [Nis21]. variances [Cav24, Cur23, GWW^{+}23, KTK22, LXMT22, LS22a, MNL21, SI23a, SS20b, SS20a, Wil22]. variant [SVK22]. Variation [DR23, CKL22, HZZ^{+}24, IR21, KJ20, KL22, LTJC23, RuA22, TNN20, WH24]. Variational [Zha22a, DP23, ESLK22]. variograms [DRB20]. various [Ali24, SU21, SBB23]. varying [BH23b, BLXL24, ESES20, GWZC22, GZ21b, Hor22, HCZ21, HZY22, LW22b, MZZF21, Mos22, MSNN^{+}24, NCC^{+}22, TEO20, WKT22, dXLY20b, YXYX22, ZZFG20, Zha22a, ZZWH20, ZZ22a, ZXF22, ZXZL22]. varying-coefficient [HCZ21, LW22b, YXYX22, Zha22a, ZXF22, ZXZL22]. Vasicek [SM24]. vector [AWZ24, Ega24, FN20, KL23, MZ22a, Med22, Nak21, Nak22, RAG21, RM20, Zhe21, Zhe24]. vectors [Baz20, Baz22, Moh22, Nag21a, TA24, WHS23, YSH23]. version [AK21, HBM^{+}20, PM20]. versus [GCF^{+}20, HMV^{+}22]. via [Alg20, AM24, Bee22, BDK^{+}23, CZP21, CDB^{+}21, DP22, EGVdR21, GMC23, HW24, HHY21, LW21, LYY21, LZQ23, LD23, Lu23, MZ22a, ML22, Mur23, NDWE24, ÖA23, RGA21, SD22, SX22, SLE20, TWTT20, WY23, Xia22, YDHW21, YXYX22, YC22]. Viability [SF24]. VIFs [Özk21]. VII [YWLD20]. Vijayan [Aub23]. virtual [CY24, NG22, OP20]. visualizations [NSlRGL23]. volatility [CS20, DZLS22, GE23, eLhS23, MF22, WC23a, XS23]. Volodin [Kri22]. volumes [DF22]. vs [OKO21, SVN24]. VSI [WZ21a]. vulnerable [LLLY23a, WZX23].

W [Kri22]. Wald [GSG22]. walks [RY21]. Waring [RC23]. warping [Hut21]. warranty [MY23]. Watson [Ali22, AHB23, ESES20, NdSA20]. wavelet [AA20d, DYC22, dALNPdOF21, PASB21, dSS22]. wavelet-based [AA20d]. wavelets [BYS24, OO21]. way [AE23b, BD23, Cav24, KJK20, YGL22]. Weak [ZQYW22, GBHA20, TEM20]. weakly [RBT^{+}22, ZMNL23]. Weibull [Ano24, AKE23, ARAB21, Ami21, ABPK21, AHG20, ASP22, BJU20, DA22, HHOC23, JS19, KNSA23, KR23, KGR21, KP21, NG22, uAAF23, PV20a, PDK23, PIA22, Seo22, SLME23, TS22, TY22, TRF^{+}22, YL22a, Yoo23]. Weibull-based [SLME23]. Weight [HGWZ22, DCL^{+}24, Pol20]. Weighted [HQN22, MSNN^{+}24, NAKK20, NYM20, PLS23, TWTT21, TDD20, AK20, ACKL23, BH20, CP23, CLS22, CTK^{+}22, CKCG21, DRB20, ESLK22, ESES20, FH22, FK22, GB22, GW22, HMV^{+}22, HZZ^{+}24, HHY21, HS23, Kal22, LHY^{+}23, Med22, uARS22, uASS22, uAA23, QTH24, TY24, WY21, WZY22, WLY22, WHS23, ZA23, ZQYW22, ZYSS22]. weighting [AÇ23a, BF23, MES21]. weights [GGS23]. well [RSM22]. well-known [RSM22]. where [Lav21]. WHIM [Lav21]. Whitney [MMMCS22a]. Widder [Bar20a]. width [Aki24]. Wiener [Hu23, Lev21, SBZG22, ZL23a]. wildfires [Ver21]. wildland [TBH22]. wind [SMB22]. window [FTC21]. windows [De 20]. Winsorization [LB23a]. within [AK23a, RAG21, SF23, TKR24, Wes20]. without [Aub23, Ric21, TP22b, ZWMY23, ZZ22b]. WOD [YHWW23]. Wong [PHY20]. Wood [CdCG22]. working [LJ22a, SLJ23, Wes20]. WQS [CKCG21]. wrapped [Bee22]. WS [Lin22]. Wynn [Bar20a]. Wynn- [Bar20a].

X [AS22b, LS23]. XGBoost [WSZ23]. XII [MCJ22, WCKC23].

yield [BNA23, El 24, NN21]. Youden [KS23b]. Youden- [KS23b]. Yule [KH23].

Zero [RC23, AKG23, Ali24, ASM22, BLR23, CFM21, CBRL22, GZY22, HA24, HZ23, KS23c, KR22, LX22, LTN23, LRF20, Pho24, RLP23, RSG22, RS24, SZ20, SSZ23, SN23, YL22a, ZBB23, ZQY22]. zero-inflated [Ali24, ASM22, BLR23, CFM21, GZY22, KR22, Pho24, RLP23, RSG22, RS24, SZ20, SN23, YL22a, ZBB23, ZQY22]. zero-modified [SSZ23]. zero-numerator [AKG23]. zero-sum [LX22]. zeros [TEO20]. ZIP [Ali24].

References

Algamal:2020:LTE

[AA20a] Zakariya Yahya Algamal and Yasin Asar. Liu-type estimator for the gamma regression model. Communications in Statistics: Simulation and Computation, 49(8):2035–2048, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2020:BSI

[AA20b] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Bayesian single-index quantile regression for ordinal data. Communications in Statistics: Simulation and Computation, 49(5):1306–1320, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2020:NGS

[AA20c] Rahim Alhamzawi and Haithem Taha Mohammad Ali. A new Gibbs sampler for Bayesian lasso. Communications in Statistics: Simulation and Computation, 49(7):1855–1871, 2020. CODEN CSSCDB. ISSN 0361-0918.

Aykroyd:2020:BAW

[AA20d] Robert G. Aykroyd and Hassan Aljohani. A Bayesian approach to wavelet-based modelling of discontinuous functions applied to inverse problems. Communications in Statistics: Simulation and Computation, 49(1):207–225, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2022:APA

[AA22a] Shakeel Ahmad and Muhammad Aslam. Another proposal about the new two-parameter estimator for linear regression model with correlated regressors. Communications in Statistics: Simulation and Computation, 51(6):3054–3072, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ashour:2022:NDF

[AA22b] Samir K. Ashour and Mohamed S. Abdallah. New distribution function estimators and tests of perfect ranking in concomitant-based ranked set sampling. Communications in Statistics: Simulation and Computation, 51(3):823–848, 2022. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2022:MLR

[AA22c] Muhammad Aslam and Shakeel Ahmad. The modified Liuridge-type estimator: a new class of biased estimators to address multicollinearity. Communications in Statistics: Simulation and Computation, 51(11):6591–6609, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ahmadabadi:2023:GQM

[AA23a] Majid Nili Ahmadabadi and Mohammad Nili Ahmadabadi. Gaussian quadrature method for GLD parameter estimation. Communications in Statistics: Simulation and Computation, 52(4):1699–1711, 2023. CODEN CSSCDB. ISSN 0361-0918.

Alshqaq:2023:RMC

[AA23b] Shokrya Alshqaq and Ali Abuzaid. On the robustness of Mallows’ C_{p} criterion. Communications in Statistics: Simulation and Computation, 52(3):1149–1163, 2023. CODEN CSSCDB. ISSN 0361-0918.

Arshad:2020:SST

[AAA20] Hafiz Muhammad Arshad, Tanvir Ahmad, and Munir Akhtar. Some sequential third order response surface designs. Communications in Statistics: Simulation and Computation, 49 (7):1872–1885, 2020. CODEN CSSCDB. ISSN 0361-0918.

Akram:2022:TPE

[AAA22a] Muhammad Nauman Akram, Muhammad Amin, and Muhammad Amanullah. Two-parameter estimator for the inverse Gaussian regression model. Communications in Statistics: Simulation and Computation, 51(10):6208–6226, 2022. CODEN CSSCDB. ISSN 0361-0918.

Amin:2022:JSE

[AAA22b] Muhammad Amin, Muhammad Nauman Akram, and Muhammad Amanullah. On the James–Stein estimator for the Poisson regression model. Communications in Statistics: Simulation and Computation, 51(10):5596–5608, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Akram:2023:GBE

[AAF23] Muhammad Nauman Akram, Muhammad Amin, and Muhammad Faisal. On the generalized biased estimators for the gamma regression model: methods and applications. Communications in Statistics: Simulation and Computation, 52(9):4087–4100, 2023. CODEN CSSCDB. ISSN 0361-0918.

Arshad:2023:BIU

[AAGP23] Mohd. Arshad, Qazi J. Azhad, Neetu Gupta, and Ashok Kumar Pathak. Bayesian inference of Unit Gompertz distribution based on dual generalized order statistics. Communications in Statistics: Simulation and Computation, 52(8):3657– 3675, 2023. CODEN CSSCDB. ISSN 0361-0918.

Alizadeh:2020:EEF

[AAHR20] Morad Alizadeh, Mahmoud Afshari, Bistoon Hosseini, and Thiago G. Ramires. Extended exp-G family of distributions: Properties, applications and simulation. Communications in Statistics: Simulation and Computation, 49(7):1730–1745, 2020. CODEN CSSCDB. ISSN 0361-0918.

Akram:2024:NIL

[AAK^{+}24] Muhammad Nauman Akram, Muhammad Amin, B. M. Golam Kibria, Mohammad Arashi, Adewale F. Lukman, and Nimra Afzal. A new improved Liu estimator for the QSAR model with inverse Gaussian response. Communications in Statistics: Simulation and Computation, 53(4):1873–1888, 2024. CODEN CSSCDB. ISSN 0361-0918.

Almanjahie:2023:CAN

[AALR23] Ibrahim M. Almanjahie, Wafa Mesfer Alahmari, Ali Laksaci, and Mustapha Rachdi. Computational aspects of the k NN local linear smoothing for some conditional models in high dimensional statistics. Communications in Statistics: Simulation and Computation, 52(7):2985–3005, 2023. CODEN CSSCDB. ISSN 0361-0918.

Amin:2023:EBR

[AAM23] Muhammad Amin, Muhammad Nauman Akram, and Abdul Majid. On the estimation of Bell regression model using ridge estimator. Communications in Statistics: Simulation and Computation, 52(3):854–867, 2023. CODEN CSSCDB. ISSN 0361-0918.

Abdullah:2021:MNC

[AATD21] Muhammad Abdullah, Rashid Ahmed, M. H. Tahir, and Muhammad Daniyal. Minimal non-circular balanced and strongly balanced repeated measurements designs. Communications in Statistics: Simulation and Computation, 50(10): 2837–2844, 2021. CODEN CSSCDB. ISSN 0361-0918.

Asar:2022:ITL

[AB22] Yasin Asar and R. Arabi Belaghi. Inference for two Lomax populations under joint type-II censoring. Communications in Statistics: Simulation and Computation, 51(11):6808–6825, 2022. CODEN CSSCDB. ISSN 0361-0918.

Asl:2022:PRT

[ABB22] Mehri Noori Asl, Hossein Bevrani, and Reza Arabi Belaghi. Penalized and ridge-type shrinkage estimators in Poisson regression model. Communications in Statistics: Simulation and Computation, 51(7):4039–4056, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ang:2023:EAC

[ABC^{+}23] Gregory Tai Xiang Ang, Zhidong Bai, Kwok Pui Choi, Yasunori Fujikoshi, and Jiang Hu. Exact and approximate computation of critical values of the largest root test in high dimension. Communications in Statistics: Simulation and Computation, 52(5):2177–2193, 2023. CODEN CSSCDB. ISSN 0361-0918.

AbdEl-Bar:2018:EGM

[Abd18] Ahmed M. T. Abd El-Bar. An extended Gompertz–Makeham distribution with application to lifetime data. Communications in Statistics: Simulation and Computation, 47(8):2454– 2475, 2018. CODEN CSSCDB. ISSN 0361-0918. See comments [CL22a].

Ahrari:2022:GFT

[ABHF22] V. Ahrari, S. Baratpour, A. Habibirad, and V. Fakoor. Goodness of fit tests for Rayleigh distribution based on quantiles. Communications in Statistics: Simulation and Computation, 51(2):341–357, 2022. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2021:DED

[ABJ21a] Muhammad Aslam, Saminathan Balamurali, and Chi-Hyuck Jun. Determination and economic design of a generalized multiple dependent state sampling plan. Communications in Statistics: Simulation and Computation, 50(11):3465–3482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2021:NMD

[ABJ21b] Muhammad Aslam, Saminathan Balamurali, and Chi-Hyuck Jun. A new multiple dependent state sampling plan based on the process capability index. Communications in Statistics: Simulation and Computation, 50(6):1711–1727, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Aslam:2021:DAC

[ABPK21] Muhammad Aslam, Saminathan Balamurali, Jeyadurga Periyasamypandian, and Nasrullah Khan. Designing of an attribute control chart based on modified multiple dependent state sampling using accelerated life test under Weibull distribution. Communications in Statistics: Simulation and Computation, 50(3):902–916, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alpay:2023:PPW

[AÇ23a] Olcay Alpay and Emel Çankaya. Performance of prior and weighting bias correction methods for rare event logistic regression under the influence of sampling bias. Communications in Statistics: Simulation and Computation, 52(3):993– 1014, 2023. CODEN CSSCDB. ISSN 0361-0918.

Anjoy:2023:HBE

[AC23b] Priyanka Anjoy and Hukum Chandra. Hierarchical Bayes estimation of small area means under a spatial nonstationary Fay-herriot model. Communications in Statistics: Simulation and Computation, 52(7):3043–3061, 2023. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2022:DMA

[ACKL22] Vasileios Alevizakos, Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. A double moving average control chart: Discussion. Communications in Statistics: Simulation and Computation, 51(10):6043–6057, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2023:GWM

[ACKL23] Vasileios Alevizakos, Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. A generally weighted moving average t control chart for monitoring shifts in the process mean. Communications in Statistics: Simulation and Computation, 52(6):2425–2438, 2023. CODEN CSSCDB. ISSN 0361-0918.

Arachchige:2021:IER

[ACP21] Chandima N. P. G. Arachchige, Maxwell Cairns, and Luke A. Prendergast. Interval estimators for ratios of independent quantiles and interquantile ranges. Communications in Statistics: Simulation and Computation, 50(12):3914–3930, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alqawba:2021:CDD

[ADK21] Mohammed Alqawba, Norou Diawara, and Jong-Min Kim. Copula directional dependence of discrete time series marginals. Communications in Statistics: Simulation and Computation, 50(11):3733–3750, 2021. CODEN CSSCDB. ISSN 0361-0918.

Azouagh:2021:DEC

[AE21] Nabil Azouagh and Said El Melhaoui. Detecting exponential component in autoregressive models: comparative study between several tests of nonlinearity. Communications in Statistics: Simulation and Computation, 50(11):3273–3285, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahmadian:2023:ESQ

[AE23a] Robab Ahmadian and Ilker Ercan. Evaluating the sensitive question methods; recommended Uludag Adjustment for the Crosswise Model. Communications in Statistics: Simulation and Computation, 52(12):5759–5772, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Almehrizi:2023:ASE

[AE23b] Rashid S. Almehrizi and Mahmoud Emam. Asymptotic standard errors of intraclass correlation coefficients for two-way model. Communications in Statistics: Simulation and Computation, 52(5):2073–2092, 2023. CODEN CSSCDB. ISSN 0361-0918.

Azimi:2023:CKM

[AE23c] Reza Azimi and Mahdy Esmailian. Correction to: Khan, M. S., King, R. and Hudson, I. L., (2017), Transmuted generalized exponential distribution: a generalization of the exponential distribution with applications to survival data, Communications in Statistics — Simulation and computation, 46:6, 4377–4398. Communications in Statistics: Simulation and Computation, 52(10):i–vi, 2023. CODEN CSSCDB. ISSN 0361-0918. See [KKH17].

Albassam:2021:MIN

[AES21] Mohammed Albassam, Ali El Hefnawy, and Emad E. A. Soliman. Mixed integer nonlinear goal programming approach to variable selection in linear regression. Communications in Statistics: Simulation and Computation, 50(12):4028–4040, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahmadian:2022:CBC

[AESY22] Robab Ahmadian, Ilker Ercan, Deniz Sigirli, and Abdulmecit Yildiz. Combining binary and continuous biomarkers by maximizing the area under the receiver operating characteristic curve. Communications in Statistics: Simulation and Computation, 51(8):4396–4409, 2022. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Fattah:2022:NCB

[AF22a] Mahmoud A. Abdel-Fattah. On a new class of binomial ridgetype regression estimators. Communications in Statistics: Simulation and Computation, 51(6):3272–3290, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Alves:2022:PNA

[AF22b] Henrique J. P. Alves and Daniel F. Ferreira. Proposition of new alternative tests adapted to the traditional T^{2} test. Communications in Statistics: Simulation and Computation, 51 (5):2287–2300, 2022. CODEN CSSCDB. ISSN 0361-0918.

Aleagobe:2023:MEE

[AFJ23] Hussain A. A. Aleagobe, Vahid Fakoor, and Sarah Jomhoori. Maximum entropy and empirical likelihood in length-biased setting: a comparison study. Communications in Statistics: Simulation and Computation, 52(6):2439–2452, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Abonazel:2020:RPR

[AG20] Mohamed Reda Abonazel and Ahmed Abd-Elfatah Gad. Robust partial residuals estimation in semiparametric partially linear model. Communications in Statistics: Simulation and Computation, 49(5):1223–1236, 2020. CODEN CSSCDB. ISSN 0361-0918.

Altunkaynak:2021:CPN

[AG21] Bulent Altunkaynak and Hamza Gamgam. Comparing the performance of nonparametric tests for equality of location against ordered alternatives. Communications in Statistics: Simulation and Computation, 50(1):63–84, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2020:CAP

[AGA20] Tanvir Ahmad, Steven G. Gilmour, and Hafiz Muhammad Arshad. Comparisons of augmented pairs designs and subset designs. Communications in Statistics: Simulation and Computation, 49(7):1898–1921, 2020. CODEN CSSCDB. ISSN 0361-0918.

Agami:2023:CPD

[Aga23] Sarit Agami. Comparison of persistence diagrams. Communications in Statistics: Simulation and Computation, 52(5): 1948–1961, 2023. CODEN CSSCDB. ISSN 0361-0918.

Al-Gounmeein:2023:CPS

[AGI23] Remal Shaher Al-Gounmeein and Mohd Tahir Ismail. Comparing the performances of symmetric and asymmetric generalized autoregressive conditionally heteroscedasticity models based on long-memory models under different distributions. Communications in Statistics: Simulation and Computation, 52(5):1878–1908, 2023. CODEN CSSCDB. ISSN 0361-0918.

Abbasi:2020:NAC

[AH20] Saba Abbasi and Abdul Haq. New adaptive CUSUM charts for process mean. Communications in Statistics: Simulation and Computation, 49(11):2944–2962, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Al-Hakeem:2023:GED

[AHAMP23] Hussein Ali Al-Hakeem, Jayanthi Arasan, Mohd Shafie Bin Mustafa, and Lim Fong Peng. Generalized exponential distribution with interval-censored data and time dependent covariate. Communications in Statistics: Simulation and Computation, 52(12):6149–6159, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ali:2023:EBP

[AHB23] Taha Hussein Ali, Heyam Abd Al-Majeed Hayawi, and Delshad Shaker Ismael Botani. Estimation of the bandwidth parameter in Nadaraya–Watson kernel non-parametric regression based on universal threshold level. Communications in Statistics: Simulation and Computation, 52(4):1476–1489, 2023. CODEN CSSCDB. ISSN 0361-0918.

Azizi:2020:SIC

[AHG20] Fariba Azizi, Firoozeh Haghighi, and Niloofar Tabibi Gilani. Statistical inference for competing risks model under progressive interval censored Weibull data. Communications in Statistics: Simulation and Computation, 49(7):1931–1944, 2020. CODEN CSSCDB. ISSN 0361-0918.

Aminnejad:2022:BOD

[AJ22] Minoo Aminnejad and Habib Jafari. Bayesian optimal designs for Cox regression model with random and nonrandom intercept based on type I censored data. Communications in Statistics: Simulation and Computation, 51(2):583–603, 2022. CODEN CSSCDB. ISSN 0361-0918.

Al-Jawarneh:2024:ALR

[AJI24] Abdullah S. Al-Jawarneh and Mohd. Tahir Ismail. The adaptive LASSO regression and empirical mode decomposition algorithm for enhancing modelling accuracy. Communications in Statistics: Simulation and Computation, 53(2):714–726, 2024. CODEN CSSCDB. ISSN 0361-0918.

Al-Jawarneh:2022:IAM

[AJIAA22] Abdullah S. Al-Jawarneh, Mohd Tahir Ismail, Ahmad M. Awajan, and Ahmed R. M. Alsayed. Improving accuracy models using elastic net regression approach based on empirical mode decomposition. Communications in Statistics: Simulation and Computation, 51(7):4006–4025, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Alevizakos:2020:DEW

[AK20] Vasileios Alevizakos and Christos Koukouvinos. A double exponentially weighted moving average chart for time between events. Communications in Statistics: Simulation and Computation, 49(10):2765–2784, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alheety:2021:NVU

[AK21] Mustafa I. Alheety and B. M. Golam Kibria. A new version of unbiased ridge regression estimator under the stochastic restricted linear regression model. Communications in Statistics: Simulation and Computation, 50(6):1589–1599, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2022:EPC

[AK22a] Vasileios Alevizakos and Christos Koukouvinos. Evaluation of process capability in gamma regression profiles. Communications in Statistics: Simulation and Computation, 51(9): 5174–5189, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2022:PMC

[AK22b] Vasileios Alevizakos and Christos Koukouvinos. A progressive mean control chart for COM-Poisson distribution. Communications in Statistics: Simulation and Computation, 51(3): 849–867, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alqallaf:2022:BIG

[AK22c] Fatemah A. Alqallaf and Debasis Kundu. A bivariate inverse generalized exponential distribution and its applications in dependent competing risks model. Communications in Statistics: Simulation and Computation, 51(12):7019–7036, 2022. CODEN CSSCDB. ISSN 0361-0918.

Abid:2023:CBW

[AK23a] Rahma Abid and Célestin C. Kokonendji. Choice between and within the classes of Poisson–Tweedie and Poisson– exponential–Tweedie count models. Communications in Statistics: Simulation and Computation, 52(5):2115–2129, 2023. CODEN CSSCDB. ISSN 0361-0918.

Amponsah:2023:CAE

[AK23b] Charles K. Amponsah and Tomasz J. Kozubowski. A computational approach to estimation of discrete Pareto parameters. Communications in Statistics: Simulation and Computation, 52(8):3692–3711, 2023. CODEN CSSCDB. ISSN 0361-0918.

Asar:2023:CGE

[AK23c] Erdogan Asar and Erdem Karabulut. Comparison of generalized estimating equations and Quasi-Least Squares regression methods in terms of efficiency with a simulation study. Communications in Statistics: Simulation and Computation, 52 (3):1015–1025, 2023. CODEN CSSCDB. ISSN 0361-0918.

Akdur:2022:CNP

[Akd22] Hatice Tul Kubra Akdur. Comparison of non-parametric tests of ordered alternatives for repeated measures in randomized blocks. Communications in Statistics: Simulation and Computation, 51(7):4146–4158, 2022. CODEN CSSCDB. ISSN 0361-0918.

Abo-Kasem:2023:ATW

[AKE23] Osama E. Abo-Kasem and Ahmed Elshahhat. Analysis of two Weibull populations under joint progressively hybrid censoring. Communications in Statistics: Simulation and Computation, 52(9):4469–4490, 2023. CODEN CSSCDB. ISSN 0361-0918.

Al-Kandari:2023:BAM

[AKG23] Noriah M. Al-Kandari and Paul H. Garthwaite. Bayesian analysis of misclassified binomial data: double-sampling and the zero-numerator problem. Communications in Statistics: Simulation and Computation, 52(2):334–348, 2023. CODEN CSSCDB. ISSN 0361-0918.

Akinshin:2024:THD

[Aki24] Andrey Akinshin. Trimmed Harrell–Davis quantile estimator based on the highest density interval of the given width. Communications in Statistics: Simulation and Computation, 53(3):1565–1575, 2024. CODEN CSSCDB. ISSN 0361-0918.

Awale:2023:FOI

[AKR23] Manik Awale, Akanksha S. Kashikar, and T. V. Ramanathan. Forecasting overdispersed INAR(1) count time series with negative binomial marginal. Communications in Statistics: Simulation and Computation, 52(6):2497–2517, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Ali:2021:CSN

[AKS^{+}21] Sajid Ali, Himmad Khan, Ismail Shah, Muhammad Moeen Butt, and Muhammad Suhail. A comparison of some new and old robust ridge regression estimators. Communications in Statistics: Simulation and Computation, 50(8):2213–2231, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alao:2022:ESM

[ALB22] Virgelio M. Alao, Joseph Ryan G. Lansangan, and Erniel B. Barrios. Estimation of semiparametric mixed analysis of covariance model. Communications in Statistics: Simulation and Computation, 51(5):2301–2317, 2022. CODEN CSSCDB. ISSN 0361-0918.

Algamal:2020:SPS

[Alg20] Zakariya Yahya Algamal. Shrinkage parameter selection via modified cross-validation approach for ridge regression model. Communications in Statistics: Simulation and Computation, 49(7):1922–1930, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2023:RBB

[Alh23] Rahim Alhamzawi. The reciprocal Bayesian bridge for leftcensored data. Communications in Statistics: Simulation and Computation, 52(8):3520–3528, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ali:2022:MAN

[Ali22] Taha Hussein Ali. Modification of the adaptive Nadaraya– Watson kernel method for nonparametric regression (simulation study). Communications in Statistics: Simulation and Computation, 51(2):391–403, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ali:2024:SBS

[Ali24] Essoham Ali. A simulation-based study of ZIP regression with various zero-inflated submodels. Communications in Statistics: Simulation and Computation, 53(2):642–657, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Akpan:2023:AEK

[ALMA23] Emmanuel Alphonsus Akpan, Kazeem Etitayo Lasisi, Imoh Udo Moffat, and Ubon Akpan Abasiekwere. Appraisal of excess kurtosis through outlier-modified GARCH-type models. Communications in Statistics: Simulation and Computation, 52(4):1523–1537, 2023. CODEN CSSCDB. ISSN 0361-0918.

Altun:2020:NAV

[Alt20a] Emrah Altun. A new approach to Value-at-Risk: GARCH-TSLx model with inference. Communications in Statistics: Simulation and Computation, 49(12):3134–3151, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Altun:2020:NGG

[Alt20b] Emrah Altun. A new generalization of geometric distribution with properties and applications. Communications in Statistics: Simulation and Computation, 49(3):793–807, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Ariyo:2022:MSB

[ALVQ22] Oludare Ariyo, Emmanuel Lesaffre, Geert Verbeke, and Adrian Quintero. Model selection for Bayesian linear mixed models with longitudinal data: Sensitivity to the choice of priors. Communications in Statistics: Simulation and Computation, 51(4):1591–1615, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alawamy:2024:BAS

[ALZ24] Eman Ahmed Alawamy, Yuanyuan Liu, and Yiqiang Q. Zhao. Bayesian analysis for single-server Markovian queues based on the No-U-Turn sampler. Communications in Statistics: Simulation and Computation, 53(2):658–670, 2024. CODEN CSSCDB. ISSN 0361-0918.

Allaj:2021:AAH

[AM21] Erindi Allaj and Maria Elvira Mancino. On asset-allocation and high-frequency data: are there financial gains from using different covariance estimators? Communications in Statistics: Simulation and Computation, 50(12):4413–4441, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2022:BRL

[AM22a] Rahim Alhamzawi and Himel Mallick. Bayesian reciprocal LASSO quantile regression. Communications in Statistics: Simulation and Computation, 51(11):6479–6494, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Avagyan:2022:PME

[AM22b] Vahe Avagyan and Xiaoling Mei. Precision matrix estimation under data contamination with an application to minimum variance portfolio selection. Communications in Statistics: Simulation and Computation, 51(4):1381–1400, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Anzarmou:2024:CHD

[AM24] Youssef Anzarmou and Abdallah Mkhadri. Classification in high dimension with an Alternative Feature Augmentation via Nonparametrics and Selection (AFANS). Communications in Statistics: Simulation and Computation, 53(1):476– 498, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2022:MIL

[AMA22] Zubair Ahmad, Eisa Mahmoudi, and Morad Alizadeh. Modelling insurance losses using a new beta power transformed family of distributions. Communications in Statistics: Simulation and Computation, 51(8):4470–4491, 2022. CODEN CSSCDB. ISSN 0361-0918.

Arshad:2023:CMV

[AMA^{+}23] Rabia Arshad, Yasar Mahmood, Muhammad Aslam, Hina Khan, Nasrullah Khan, and Naufil Sakran. Cost model of variable multiple dependent state sampling plan with rectifying inspection. Communications in Statistics: Simulation and Computation, 52(6):2349–2364, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2022:NFH

[AMD22] Zubair Ahmad, Eisa Mahmoudi, and Sanku Dey. A new family of heavy tailed distributions with an application to the heavy tailed insurance loss data. Communications in Statistics: Simulation and Computation, 51(8):4372–4395, 2022. CODEN CSSCDB. ISSN 0361-0918.

Amini:2021:SBM

[Ami21] Morteza Amini. Selection of the best minimal repair systems for Weibull lifetime distribution: a Bayesian approach. Communications in Statistics: Simulation and Computation, 50 (3):832–853, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahsan:2021:ODU

[AMK^{+}21a] Muhammad Ahsan, Muhammad Mashuri, Heri Kuswanto, Dedy Dwi Prastyo, and Hidayatul Khusna. Outlier detection using PCA mix based T^{2} control chart for continuous and categorical data. Communications in Statistics: Simulation and Computation, 50(5):1496–1523, 2021. CODEN CSSCDB. ISSN 0361-0918.

Arora:2021:BER

[AMK21b] Sangeeta Arora, Kalpana K. Mahajan, and Ritu Kumari. Bayes estimators for the reliability and hazard rate functions of Topp–Leone distribution using Type-II censored data. Communications in Statistics: Simulation and Computation, 50(8):2327–2344, 2021. CODEN CSSCDB. ISSN 0361-0918.

Al-Marhoobi:2023:ATH

[AMP23] Safia Al-Marhoobi and Andrey Pepelyshev. Analysis of temperature and humidity in Oman using singular spectrum analysis. Communications in Statistics: Simulation and Computation, 52(4):1403–1416, 2023. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2022:SR

[Ano22] Anonymous. Statement of removal. Communications in Statistics: Simulation and Computation, 51(11):(i), 2022. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2024:CJK

[Ano24] Anonymous. Correction to: Jayakumar, K., and Sankaran, K. K. (2019). Discrete Linnik Weibull distribution. Communications in Statistics-Simulation and Computation, 48 (10): 3092–3117. Communications in Statistics: Simulation and Computation, 53(2):1088–1089, 2024. CODEN CSSCDB. ISSN 0361-0918. See [JS19].

Alonso:2020:SIU

[ANP20] Rosa Alonso, Christos T. Nakas, and M. Carmen Pardo. A study of indices useful for the assessment of diagnostic markers in non-parametric ROC curve analysis. Communications in Statistics: Simulation and Computation, 49(8):2102–2113, 2020. CODEN CSSCDB. ISSN 0361-0918.

Abbas:2023:END

[ANR^{+}23] Zameer Abbas, Hafiz Zafar Nazir, Muhammad Riaz, Muhammad Abid, and Noureen Akhtar. An efficient nonparametric double progressive mean chart for monitoring of the process location. Communications in Statistics: Simulation and Computation, 52(6):2578–2591, 2023. CODEN CSSCDB. ISSN 0361-0918.

Acar:2020:RDM

[AÖ20a] Tugba Söküt Açar and M. Revan Özkale. Regression diagnostics methods for Liu estimator under the general linear regression model. Communications in Statistics: Simulation and Computation, 49(3):771–792, 2020. CODEN CSSCDB. ISSN 0361-0918.

Arsova:2020:ITC

[AÖ20b] Antonia Arsova and Deniz Dilan Karaman Örsal. Intersection tests for the cointegrating rank in dependent panel data. Communications in Statistics: Simulation and Computation, 49(4):918–941, 2020. CODEN CSSCDB. ISSN 0361-0918.

Al-Omari:2023:EDF

[AOA23] Amer I. Al-Omari and Mohamed S. Abdallah. Estimation of the distribution function using moving extreme and MiniMax ranked set sampling. Communications in Statistics: Simulation and Computation, 52(5):1909–1925, 2023. CODEN CSSCDB. ISSN 0361-0918.

Amin:2022:NRE

[AQAN22] Muhammad Amin, Muhammad Qasim, Saima Afzal, and Khalid Naveed. New ridge estimators in the inverse Gaussian regression: Monte Carlo simulation and application to chemical data. Communications in Statistics: Simulation and Computation, 51(10):6170–6187, 2022. CODEN CSSCDB. ISSN 0361-0918.

Amin:2022:AUR

[AQYA22] Muhammad Amin, Muhammad Qasim, Ahad Yasin, and Muhammad Amanullah. Almost unbiased ridge estimator in the gamma regression model. Communications in Statistics: Simulation and Computation, 51(7):3830–3850, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Almuzaini:2022:CSA

[AR22] Heba A. Almuzaini and Mohammad Z. Raqab. Comparative study for assessing the Pitman’s closeness of predictors based on exponential record data. Communications in Statistics: Simulation and Computation, 51(8):4427–4449, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Abdullah:2023:GCM

[ARA23] Muhammad Abdullah, H. M. Kashif Rasheed, and Rashid Ahmed. General construction of minimal balanced repeated measurements designs in non-circular periods of k different sizes. Communications in Statistics: Simulation and Computation, 52(2):523–532, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ali:2021:CHD

[ARAB21] Sajid Ali, SyedMuhammad Muslim Raza, Muhammad Aslam, and Muhammad Moeen Butt. CEV-Hybrid Dewma charts for censored data using Weibull distribution. Communications in Statistics: Simulation and Computation, 50(2):446–461, 2021. CODEN CSSCDB. ISSN 0361-0918.

Altaf:2024:NBF

[ARAR24] Saima Altaf, Fareeha Rashid, Muhammad Aslam, and Sadaf Riasat. A novel Bayesian framework to address unknown heteroscedasticity for the linear regression model. Communications in Statistics: Simulation and Computation, 53(3): 1298–1307, 2024. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2020:NVC

[ARJ20] Muhammad Aslam, Muhammad Ali Raza, and Chi-Hyuck Jun. A new variable control chart under generalized multiple dependent state sampling. Communications in Statistics: Simulation and Computation, 49(9):2321–2332, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Amiri:2021:NRU

[ARJ21] M. Amiri, R. Roozegar, and A. Jamalizadeh. Nonlinear regression using order statistics from the multivariate generalized hyperbolic distributions. Communications in Statistics: Simulation and Computation, 50(5):1348–1363, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Aslam:2022:TSS

[ARKA22] Muhammad Aslam, Gadde Srinivasa Rao, Nasrullah Khan, and Liaquat Ahmad. Two-stage sampling plan using process loss index under neutrosophic statistics. Communications in Statistics: Simulation and Computation, 51(6):2831–2841, 2022. CODEN CSSCDB. ISSN 0361-0918.

Arnab:2023:RRT

[Arn23] Raghunath Arnab. Randomized response trial using negative binomial distribution. Communications in Statistics: Simulation and Computation, 52(12):6002–6010, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2021:ECD

[AS21a] Aneel Ahmed and Javid Shabbir. On estimation of coefficient of dispersion using the auxiliary information. Communications in Statistics: Simulation and Computation, 50(11): 3590–3606, 2021. CODEN CSSCDB. ISSN 0361-0918.

Asai:2021:SSL

[AS21b] Manabu Asai and Mike K. P. So. A simulation smoother for long memory time series with correlated and heteroskedastic additive noise. Communications in Statistics: Simulation and Computation, 50(2):388–399, 2021. CODEN CSSCDB. ISSN 0361-0918.

Adhikari:2022:DGF

[AS22a] A. P. N. G. Adhikari and M. R. Sooriyarachchi. Developing a goodness of fit test for a joint model of clustered survival and count data. Communications in Statistics: Simulation and Computation, 51(12):7151–7168, 2022. CODEN CSSCDB. ISSN 0361-0918.

Akgul:2022:ISS

[AS22b] Fatma Gül Akgül and Birdal Senoglu. Inferences for stressstrength reliability of Burr Type X distributions based on ranked set sampling. Communications in Statistics: Simulation and Computation, 51(6):3324–3340, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alam:2022:CEM

[AS22c] Shameem Alam and Javid Shabbir. Calibration estimation of mean by using double use of auxiliary information. Communications in Statistics: Simulation and Computation, 51(8): 4769–4787, 2022. CODEN CSSCDB. ISSN 0361-0918.

Altun:2022:DMF

[AS22d] Gökçen Altun and Tülay Saraçbasi. Determination of model fitting with power-divergence-type measure of departure from symmetry for sparse and non-sparse square contingency tables. Communications in Statistics: Simulation and Computation, 51(7):4087–4111, 2022. CODEN CSSCDB. ISSN 0361-0918.

Adamec:2023:RRB

[AS23] Václav Adamec and Eric P. Smith. Rejection rates of bootstrapped and exact heteroskedasticity tests in response to skedastic function and normal or skewed disturbances. Communications in Statistics: Simulation and Computation, 52 (6):2765–2780, 2023. CODEN CSSCDB. ISSN 0361-0918.

Albassam:2023:ESB

[ASA23] Mohammed Albassam, Emad E. A. Soliman, and Sherif S. Ali. An effectiveness study of the Bayesian inference with multivariate autoregressive moving average processes. Communications in Statistics: Simulation and Computation, 52 (10):4773–4788, 2023. CODEN CSSCDB. ISSN 0361-0918.

Anjum:2024:MBE

[ASA24] Safora Anjum, Javid Shabbir, and Shakeel Ahmed. Modelbased estimation for population total under model misspecification using the balanced sampling scheme. Communications in Statistics: Simulation and Computation, 53(4):1678–1689, 2024. CODEN CSSCDB. ISSN 0361-0918.

Audu:2023:NRT

[ASK23] Ahmed Audu, Rajesh Singh, and Supriya Khare. New regression-type compromised imputation class of estimators with known parameters of auxiliary variable. Communications in Statistics: Simulation and Computation, 52(10): 4789–4801, 2023. CODEN CSSCDB. ISSN 0361-0918.

Azizi:2021:RSO

[ASM21] Mohammad Javad Azizi, Farshad Seifi, and Samira Moghadam. A robust simulation optimization algorithm using kriging and particle swarm optimization: Application to surgery room optimization. Communications in Statistics: Simulation and Computation, 50(7):2025–2041, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aly:2022:AEC

[ASM22] Aya A. Aly, Nesma A. Saleh, and Mahmoud A. Mahmoud. An adaptive EWMA control chart for monitoring zero-inflated Poisson processes. Communications in Statistics: Simulation and Computation, 51(4):1564–1577, 2022. CODEN CSSCDB. ISSN 0361-0918.

Azizi:2022:IAB

[ASP22] Afsaneh Azizi, Abdolreza Sayyareh, and Hanieh Panahi. Inference about the bivariate new extended Weibull distribution based on complete and censored data. Communications in Statistics: Simulation and Computation, 51(3):738–756, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ahmadi:2022:NRC

[ASS22] Mohammad Mahdi Ahmadi, Hamid Shahriari, and Yaser Samimi. A novel robust control chart for monitoring multiple linear profiles in phase II. Communications in Statistics: Simulation and Computation, 51(11):6257–6268, 2022. CODEN CSSCDB. ISSN 0361-0918.

Afify:2021:TPE

[ASZN21] Ahmed Z. Afify, Adriano K. Suzuki, Chunfang Zhang, and Mazen Nassar. On three-parameter exponential distribution: properties, Bayesian and non-Bayesian estimation based on complete and censored samples. Communications in Statistics: Simulation and Computation, 50(11):3799–3819, 2021. CODEN CSSCDB. ISSN 0361-0918.

Arlt:2021:ASM

[AT21] Josef Arlt and Peter Trcka. Automatic SARIMA modeling and forecast accuracy. Communications in Statistics: Simulation and Computation, 50(10):2949–2970, 2021. CODEN CSSCDB. ISSN 0361-0918.

Abonazel:2023:BRR

[AT23] Mohamed R. Abonazel and Ibrahim M. Taha. Beta ridge regression estimators: simulation and application. Communications in Statistics: Simulation and Computation, 52(9): 4280–4292, 2023. CODEN CSSCDB. ISSN 0361-0918.

Albayrak:2020:PKR

[ATYK20] Muammer Albayrak, Kemal Turhan, Yasemin Yavuz, and Zeliha Aydin Kasap. kaphom: an R package for testing the homogeneity of intra-class kappa statistics. Communications in Statistics: Simulation and Computation, 49(12):3283–3298, 2020. CODEN CSSCDB. ISSN 0361-0918.

Aubry:2023:CIH

[Aub23] Philippe Aubry. On the correct implementation of the Hanurav–Vijayan selection procedure for unequal probability sampling without replacement. Communications in Statistics: Simulation and Computation, 52(5):1849–1877, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Avagyan:2021:DTE

[Ava21] Vahe Avagyan. D-Trace estimation of a precision matrix with eigenvalue control. Communications in Statistics: Simulation and Computation, 50(4):1231–1247, 2021. CODEN CSSCDB. ISSN 0361-0918.

Avci:2021:BAM

[Avc21] Esin Avci. Bayesian approach to the meta-analysis of multicategory prevalence. Communications in Statistics: Simulation and Computation, 50(5):1541–1559, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Ananda:2022:NLC

[AW22] Malwane M. A. Ananda and Samaradasa Weerahandi. A note on the limitations of the CAT procedure with application to mixed-effects models. Communications in Statistics: Simulation and Computation, 51(6):3138–3148, 2022. CODEN CSSCDB. ISSN 0361-0918.

Asenso:2024:PLS

[AWZ24] Theophilus Quachie Asenso, Puyu Wang, and Hai Zhang. Pliable lasso for the support vector machine. Communications in Statistics: Simulation and Computation, 53(2):786–798, 2024. CODEN CSSCDB. ISSN 0361-0918.

Akay:2021:HCM

[AY21a] Özlem Akay and Güzin Yüksel. Hierarchical clustering of mixed variable panel data based on new distance. Communications in Statistics: Simulation and Computation, 50(6): 1695–1710, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aydin:2021:TLS

[AY21b] Dursun Aydin and Ersin Yilmaz. Truncation level selection in nonparametric regression using Padé approximation. Communications in Statistics: Simulation and Computation, 50 (3):744–763, 2021. CODEN CSSCDB. ISSN 0361-0918.

Al-Zahrani:2023:RMP

[AZAZ23] Bander Al-Zahrani and Areej M. Al-Zaydi. Relations for moments of progressively Type-II censored order statistics from the exponential-geometric distribution and associated inference. Communications in Statistics: Simulation and Computation, 52(5):2151–2166, 2023. CODEN CSSCDB. ISSN 0361-0918.

AlMamuno:2022:IIO

[AZH^{+}22] Abu Sayed Md. Al Mamuno, Yong Zulina Zubairi, Abdul Ghapor Hussin, A. H. M. Rahmatullah Imon, Sohel Rana, and Jalmar Carrasco. Identification of influential observation in linear structural relationship model with known slope. Communications in Statistics: Simulation and Computation, 51(1):72–83, 2022. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2022:MLE

[BA22a] Narayanaswamy Balakrishnan and Farouq Mohammad A. Alam. Maximum likelihood estimation of the parameters of Student’s t Birnbaum–Saunders distribution: a comparative study. Communications in Statistics: Simulation and Computation, 51(3):793–822, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2022:SPI

[BA22b] Mohamed Bentarzi and Nawel Aries. On some periodic INARMA(p, q) models. Communications in Statistics: Simulation and Computation, 51(10):5773–5793, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2022:QPI

[BA22c] Mohamed Bentarzi and Nawel Aries. QMLE of periodic integer-valued time series models. Communications in Statistics: Simulation and Computation, 51(9):4973–4999, 2022. CODEN CSSCDB. ISSN 0361-0918.

Barakat:2021:SGF

[BAEk21] H. M. Barakat, A. W. Aboutahoun, and N. N. El-kadar. On some generalized families arising from mixture normal distribution with applications. Communications in Statistics: Simulation and Computation, 50(1):198–216, 2021. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2022:CAV

[Bal22] S. Balamurali. Combined attri-vari inspection policy for resubmitted lots based on the process capability index. Communications in Statistics: Simulation and Computation, 51 (9):5406–5425, 2022. CODEN CSSCDB. ISSN 0361-0918.

Blazsek:2024:SSS

[BAL24] Szabolcs Blazsek, Astrid Ayala, and Adrian Licht. Signal smoothing for score-driven models: a linear approach. Communications in Statistics: Simulation and Computation, 53 (2):829–852, 2024. CODEN CSSCDB. ISSN 0361-0918.

Bapat:2020:NCB

[Bap20] Sudeep R. Bapat. A new correlation for bivariate time series with a higher order of integration. Communications in Statistics: Simulation and Computation, 49(10):2546–2558, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bapat:2022:AMP

[Bap22] Sudeep R. Bapat. An alternative measure of positive correlation for bivariate time series. Communications in Statistics: Simulation and Computation, 51(6):3252–3258, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Barabesi:2020:CPD

[Bar20a] Lucio Barabesi. The computation of the probability density and distribution functions for some families of random variables by means of the Wynn-ρ accelerated Post–Widder formula. Communications in Statistics: Simulation and Computation, 49(5):1333–1351, 2020. CODEN CSSCDB. ISSN 0361-0918.

Barghout:2020:TSN

[Bar20b] May Barghout. A two-stage non-parametric software reliability model. Communications in Statistics: Simulation and Computation, 49(5):1159–1180, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bayoud:2021:TNN

[Bay21] Husam Awni Bayoud. Tests of normality: new test and comparative study. Communications in Statistics: Simulation and Computation, 50(12):4442–4463, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Bazyari:2020:GTO

[Baz20] Abouzar Bazyari. A general testing for order restriction on mean vectors of multivariate normal populations. Communications in Statistics: Simulation and Computation, 49(8): 1982–1998, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bazyari:2022:CTO

[Baz22] Abouzar Bazyari. On the conditional test of order restricted multivariate normal mean vectors: simulation study and application. Communications in Statistics: Simulation and Computation, 51(12):7037–7056, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bao:2024:MMS

[BB24a] Yihan Bao and James G. Booth. Mixed models and shrinkage estimation for balanced and unbalanced designs. Communications in Statistics: Simulation and Computation, 53 (1):398–408, 2024. CODEN CSSCDB. ISSN 0361-0918.

Bhattacharya:2024:MTS

[BB24b] Ritwik Bhattacharya and Narayanaswamy Balakrishnan. A MCMC-type simple probabilistic approach for determining optimal progressive censoring schemes. Communications in Statistics: Simulation and Computation, 53(4):1758–1767, 2024. CODEN CSSCDB. ISSN 0361-0918.

Batyrbekova:2024:PBI

[BBD^{+}24] Nurgul Batyrbekova, Hannah Bower, Paul W. Dickman, Robert Szulkin, Paul C. Lambert, and Therese M.-L. Andersson. Potential bias introduced by not including multiple time-scales in survival analysis: a simulation study. Communications in Statistics: Simulation and Computation, 53(2): 993–1006, 2024. CODEN CSSCDB. ISSN 0361-0918.

Barboza:2023:NMA

[BBK23] Flavio Barboza, Leonardo Fernando Cruz Basso, and Herbert Kimura. New metrics and approaches for predicting bankruptcy. Communications in Statistics: Simulation and Computation, 52(6):2615–2632, 2023. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2024:PLA

[BBM24] N. Balakrishnan, S. Barui, and F. S. Milienos. Piecewise linear approximations of baseline under proportional hazards based COM–Poisson cure models. Communications in Statistics: Simulation and Computation, 53(3):1135–1160, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Basak:2021:BIP

[BC21] Arpita Basak and Amit Choudhury. Bayesian inference and prediction in single server M/M/1 queuing model based on queue length. Communications in Statistics: Simulation and Computation, 50(6):1576–1588, 2021. CODEN CSSCDB. ISSN 0361-0918.

Basak:2023:CBI

[BC23a] Arpita Basak and Amit Choudhury. Classical and Bayesian inference on traffic intensity of multiserver Markovian queuing system. Communications in Statistics: Simulation and Computation, 52(5):2044–2057, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bhoj:2023:RSS

[BC23b] Dinesh S. Bhoj and Girish Chandra. Ranked set sampling with lowest order statistics for Pareto distribution. Communications in Statistics: Simulation and Computation, 52(6): 2327–2335, 2023. CODEN CSSCDB. ISSN 0361-0918.

Banerjee:2024:IEB

[BCC24] Prasenjit Banerjee, Tanuka Chattopadhyay, and Asis Kumar Chattopadhyay. Investigation of the effect of bars on the properties of spiral galaxies: a multivariate statistical study. Communications in Statistics: Simulation and Computation, 53 (3):1216–1246, 2024. CODEN CSSCDB. ISSN 0361-0918.

Bachi:2021:PEU

[BCDA21] Katia Bachi, Cédric Chauvière, Hacène Djellout, and Karim Abbas. Propagation of epistemic uncertainty in queueing models with unreliable server using chaos expansions. Communications in Statistics: Simulation and Computation, 50 (4):1039–1061, 2021. CODEN CSSCDB. ISSN 0361-0918.

Barbeito:2021:BCI

[BCP21] Inés Barbeito, Ricardo Cao, and Dimitris Politis. Bootstrap confidence intervals for conditional density function in Markov processes. Communications in Statistics: Simulation and Computation, 50(12):4315–4337, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Bower:2021:CSC

[BCR^{+}21] Hannah Bower, Michael J. Crowther, Mark J. Rutherford, Therese M.-L. Andersson, Mark Clements, Xing-Rong Liu, Paul W. Dickman, and Paul C. Lambert. Capturing simple and complex time-dependent effects using flexible parametric survival models: a simulation study. Communications in Statistics: Simulation and Computation, 50(11):3777–3793, 2021. CODEN CSSCDB. ISSN 0361-0918.

Baratnia:2023:RHB

[BD23] Mohammad Baratnia and Mahdi Doostparast. A reversedhazard-based nonlinear model for one-way classification. Communications in Statistics: Simulation and Computation, 52(9):4378–4391, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bodevan:2022:MOA

[BDD^{+}22] Emerson C. Bodevan, Luiz H. Duczmal, Anderson R. Duarte, Pedro H. L. Silva, and Gladston J. P. Moreira. Multi-objective approach for multiple clusters detection in data points events. Communications in Statistics: Simulation and Computation, 51(3):1313–1332, 2022. CODEN CSSCDB. ISSN 0361-0918.

Barczy:2023:AVR

[BDG23] Mátyás Barczy, Ádám Dudás, and József Gáll. On approximations of value at risk and expected shortfall involving kurtosis. Communications in Statistics: Simulation and Computation, 52(3):770–794, 2023. CODEN CSSCDB. ISSN 0361-

Buccini:2023:VSS

[BDK^{+}23] Alessandro Buccini, Omar De la Cruz Cabrera, Christos Koukouvinos, Marilena Mitrouli, and Lothar Reichel. Variable selection in saturated and supersaturated designs via ℓ_{p}– ℓ_{q} minimization. Communications in Statistics: Simulation and Computation, 52(9):4326–4347, 2023. CODEN CSSCDB. ISSN 0361-0918.

Boos:2023:PCL

[BDL23] Dennis D. Boos, Kaiyuan Duan, and Xiaoni Liu. Pairwise comparisons for levene-style variability parameters. Communications in Statistics: Simulation and Computation, 52(4): 1562–1576, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bozbulut:2020:GES

[BE20] Ali Riza Bozbulut and Serkan Eryilmaz. Generalized extreme shock models and their applications. Communications in Statistics: Simulation and Computation, 49(1):110–120, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bee:2022:EWS

[Bee22] Marco Bee. Estimating the wrapped stable distribution via indirect inference. Communications in Statistics: Simulation and Computation, 51(11):6371–6387, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Bekele:2023:AMA

[Bek23] Alemu Bekele. Alex mean (AlM) location estimator for measure of center of data. Communications in Statistics: Simulation and Computation, 52(6):2488–2496, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bekrizadeh:2024:GFC

[Bek24] Hakim Bekrizadeh. Generalized FGM copulas: Properties and applications. Communications in Statistics: Simulation and Computation, 53(2):744–755, 2024. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2021:CMS

[Bey21] Beste H. Beyaztas. Construction of multi-step forecast regions of VAR processes using ordered block bootstrap. Communications in Statistics: Simulation and Computation, 50 (7):2107–2125, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bruch:2023:AMR

[BF23] Christian Bruch and Barbara Felderer. Applying multilevel regression weighting when only population margins are available. Communications in Statistics: Simulation and Computation, 52(11):5401–5422, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bekker:2020:CMA

[BFAR20] A. Bekker, J. T. Ferreira, M. Arashi, and B. W. Rowland. Computational methods applied to a skewed generalized normal family. Communications in Statistics: Simulation and Computation, 49(11):2930–2943, 2020. CODEN CSSCDB. ISSN 0361-0918.

Brito:2022:GAA

[BFS22] José Brito, Augusto Fadel, and Gustavo Semaan. A genetic algorithm applied to optimal allocation in stratified sampling. Communications in Statistics: Simulation and Computation, 51(7):3714–3732, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhowmik:2023:CTR

[BGJ^{+}23] Arpan Bhowmik, Rahul Kumar Gupta, Seema Jaggi, Eldho Varghese, Mohd. Harun, Cini Varghese, and Anindita Datta. On the construction of trend resistant PBIB designs. Communications in Statistics: Simulation and Computation, 52 (9):4052–4064, 2023. CODEN CSSCDB. ISSN 0361-0918.

Balagyozyan:2020:PTT

[BGM20] Aram Balagyozyan, Christos Giannikos, and Barry K. Ma. Power and thick tails: an ARCH process example with extreme value as test statistic. Communications in Statistics: Simulation and Computation, 49(2):556–564, 2020. CODEN CSSCDB. ISSN 0361-0918.

Beauducel:2020:FMI

[BH20] André Beauducel and Norbert Hilger. On the fit of models implied by unit-weighted scales. Communications in Statistics: Simulation and Computation, 49(11):3054–3064, 2020. CODEN CSSCDB. ISSN 0361-0918.

Beauducel:2023:SPF

[BH23a] André Beauducel and Norbert Hilger. Score predictor factor analysis as a tool for the identification of single-item indicators. Communications in Statistics: Simulation and Computation, 52(2):453–465, 2023. CODEN CSSCDB. ISSN 0361-0918.

Benlagha:2023:DII

[BH23b] Noureddine Benlagha and Wael Hemrit. Does investment in insurance stocks reap diversification benefits? Static and time varying copula modeling. Communications in Statistics: Simulation and Computation, 52(4):1384–1402, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bianchi:2022:MFG

[Bia22] Michele Leonardo Bianchi. Are multi-factor Gaussian term structure models still useful? An empirical analysis on Italian BTPs. Communications in Statistics: Simulation and Computation, 51(7):3685–3713, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bicer:2020:SIG

[Biç20] Hayrinisa Demirci Biçer. Statistical inference for geometric process with the Two-Parameter Lindley Distribution. Communications in Statistics: Simulation and Computation, 49 (11):2979–3000, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bilon:2023:NST

[Bil23] Xavier Javines Bilon. Normality and significance testing in simple linear regression model for large sample sizes: a simulation study. Communications in Statistics: Simulation and Computation, 52(6):2781–2797, 2023. CODEN CSSCDB. ISSN 0361-0918.

Biswas:2022:EPT

[Bis22] Aniket Biswas. Estimating the proportion of true null hypotheses with application in microarray data. Communications in Statistics: Simulation and Computation, 51(11): 6294–6308, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bakar:2020:APS

[BJ20] K. Shuvo Bakar and Huidong Jin. Areal prediction of survey data using Bayesian spatial generalised linear models. Communications in Statistics: Simulation and Computation, 49 (11):2963–2978, 2020. CODEN CSSCDB. ISSN 0361-0918.

BenSlimen:2022:CCB

[BJA22] Yosra Ben Slimen, Julien Jacques, and Sylvain Allio. Coclustering for binary and functional data. Communications in Statistics: Simulation and Computation, 51(9):4845–4866, 2022. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2020:DTN

[BJU20] S. Balamurali, P. Jeyadurga, and M. Usha. Designing of tightened-normal-tightened sampling scheme under Weibull and gamma distributions for mean life assurance. Communications in Statistics: Simulation and Computation, 49(7): 1704–1729, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bhowmik:2021:NOD

[BJVV21] Arpan Bhowmik, Seema Jaggi, Eldho Varghese, and Cini Varghese. A note on optimal directional neighbour designs with random block effect. Communications in Statistics: Simulation and Computation, 50(3):865–880, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Baumgartner:2023:PCK

[BK23] Daniel Baumgartner and John Kolassa. Power considerations for Kolmogorov–Smirnov and Anderson–Darling two-sample tests. Communications in Statistics: Simulation and Computation, 52(7):3137–3145, 2023. CODEN CSSCDB. ISSN 0361-0918.

Barakat:2022:PFL

[BKG22] H. M. Barakat, O. M. Khaled, and Hadeer A. Ghonem. Predicting future lifetime for mixture exponential distribution. Communications in Statistics: Simulation and Computation, 51(7):3533–3552, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhatti:2023:EAT

[BKIR23] Sajjad Haider Bhatti, Faizan Wajid Khan, Muhammad Irfan, and Muhammad Ali Raza. An effective approach towards efficient estimation of general linear model in case of heteroscedastic errors. Communications in Statistics: Simulation and Computation, 52(2):392–403, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Bhattacharyya:2023:NFT

[BKM23] Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. A new family of tests for DMTTF alternatives under complete and censored samples. Communications in Statistics: Simulation and Computation, 52(10):4603–4620, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bhushan:2023:EPM

[BKPS23] Shashi Bhushan, Anoop Kumar, Abhay Pratap Pandey, and Saurabh Singh. Estimation of population mean in presence of missing data under simple random sampling. Communications in Statistics: Simulation and Computation, 52(12): 6048–6069, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bar-Lev:2023:EDM

[BLR23] Shaul K. Bar-Lev and Ad Ridder. Exponential dispersion models for overdispersed zero-inflated count data. Communications in Statistics: Simulation and Computation, 52(7): 3286–3304, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bian:2024:SEA

[BLXL24] Jiawen Bian, Zhihui Liu, Jing Xing, and Hongwei Li. A statistically efficient algorithm for estimating the parameters of a chirp signal model with time-varying amplitude. Communications in Statistics: Simulation and Computation, 53(3): 1423–1443, 2024. CODEN CSSCDB. ISSN 0361-0918.

Bnouachir:2021:ECB

[BM21] Najla Bnouachir and Abdallah Mkhadri. Efficient clusterbased portfolio optimization. Communications in Statistics: Simulation and Computation, 50(11):3241–3255, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Bi:2022:REB

[BMG22] Qixuan Bi, Yanbin Ma, and Wenhao Gui. Reliability estimation for the bathtub-shaped distribution based on progressively first-failure censoring sampling. Communications in Statistics: Simulation and Computation, 51(8):4564–4580, 2022. CODEN CSSCDB. ISSN 0361-0918.

Banihashemi:2023:DPY

[BNA23] Atefe Banihashemi, Mohammad Saber Fallah Nezhad, and Amirhossein Amiri. Developing process-yield-based acceptance sampling plans for AR(1) auto-correlated process. Communications in Statistics: Simulation and Computation, 52 (9):4230–4251, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bahrami:2021:MMP

[BNK21] Hassan Bahrami, Seyed Taghi Akhavan Niaki, and Majid Khedmati. Monitoring multivariate profiles in multistage processes. Communications in Statistics: Simulation and Computation, 50(11):3436–3464, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bakar:2022:FDH

[BNN22] S. A. Abu Bakar, S. Nadarajah, and N. Ngataman. A family of density-hazard distributions for insurance losses. Communications in Statistics: Simulation and Computation, 51(10): 5857–5875, 2022. CODEN CSSCDB. ISSN 0361-0918.

Boularouk:2023:SLQ

[Bou23] Yakoub Boularouk. Standard Laplace quasi-maximum likelihood estimator for GARCH processes. Communications in Statistics: Simulation and Computation, 52(4):1490–1503, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bhatt:2022:ATP

[BPN22] Milind B. Bhatt, Shantilal R. Patel, and Pratik B. Nandy. Asymptotic tests for parameters of two one-truncation parameter family of distributions. Communications in Statistics: Simulation and Computation, 51(4):1616–1636, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhushan:2020:OIM

[BPP20] Shashi Bhushan, Abhay Pratap Pandey, and Arvind Pandey. On optimality of imputation methods for estimation of population mean using higher order moment of an auxiliary variable. Communications in Statistics: Simulation and Computation, 49(6):1560–1574, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bessegato:2023:AOQ

[BQCP23] Lupércio F. Bessegato, Roberto C. Quinino, Frederico R. B. Cruz, and Augusto R. Pereira. An analysis of online quality control by attributes with an imperfect classification system and inspections with samples of size n. Communications in Statistics: Simulation and Computation, 52(7):2941–2966, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bickel:2021:CFD

[BR21] David R. Bickel and Abbas Rahal. Correcting false discovery rates for their bias toward false positives. Communications in Statistics: Simulation and Computation, 50(11):3699–3713, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bayoud:2024:CBI

[BR24] Husam A. Bayoud and Mohammad Z. Raqab. Classical and Bayesian inferences for two Topp–Leone models under joint progressive Type-II censoring. Communications in Statistics: Simulation and Computation, 53(1):427–445, 2024. CODEN CSSCDB. ISSN 0361-0918.

Biswas:2020:RBT

[BRA20] Ankur Biswas, Anil Rai, and Tauqueer Ahmad. Rescaling bootstrap technique for variance estimation for ranked set samples in finite population. Communications in Statistics: Simulation and Computation, 49(10):2704–2718, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2024:AAE

[BRTW24] Narayanaswamy Balakrishnan, Jan Rychtár, Dewey Taylor, and Stephen D. Walter. Accurate approximation of the expected value, standard deviation, and probability density function of extreme order statistics from Gaussian samples. Communications in Statistics: Simulation and Computation, 53(2):869–878, 2024. CODEN CSSCDB. ISSN 0361-0918.

Brzezinska:2020:IRT

[Brz20] Justyna Brzezińska. Item response theory models in the measurement theory. Communications in Statistics: Simulation and Computation, 49(12):3299–3313, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Bulinskaya:2021:MAA

[BS21] Ekaterina V. Bulinskaya and Boris I. Shigida. Modeling and asymptotic analysis of insurance company performance. Communications in Statistics: Simulation and Computation, 50 (9):2743–2756, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2022:EEP

[BS22a] Mohamed Bentarzi and Mohamed Sadoun. Efficient estimation in (PINAR(1)) model: semiparametric case. Communications in Statistics: Simulation and Computation, 51(12): 7110–7132, 2022. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2022:CPE

[BS22b] Ufuk Beyaztas and Han Lin Shang. A comparison of parameter estimation in function-on-function regression. Communications in Statistics: Simulation and Computation, 51(8): 4607–4637, 2022. CODEN CSSCDB. ISSN 0361-0918.

Beaumont:2023:IEG

[BS23a] Paul M. Beaumont and Aaron D. Smallwood. Inference for estimators of generalized long memory processes. Communications in Statistics: Simulation and Computation, 52(12): 6096–6115, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2023:EES

[BS23b] Mohamed Bentarzi and Mohamed Sadoun. Efficient estimation in semiparametric self-exciting threshold INAR processes. Communications in Statistics: Simulation and Computation, 52(6):2592–2614, 2023. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2022:FLM

[BSAS22] Ufuk Beyaztas, Han Lin Shang, and Abdel-Salam G. Abdel-Salam. Functional linear models for interval-valued data. Communications in Statistics: Simulation and Computation, 51(7):3513–3532, 2022. CODEN CSSCDB. ISSN 0361-0918.

Basalamah:2021:MIC

[BSNT21] Doaa Basalamah, Khamis K. Said, Wei Ning, and Yubin Tian. Modified information criterion for linear regression changepoint model with its applications. Communications in Statistics: Simulation and Computation, 50(1):180–197, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Bhattacharjee:2020:POM

[BT20] Dibyojyoti Bhattacharjee and Priyanka Talukdar. Predicting outcome of matches using pressure index: evidence from Twenty20 cricket. Communications in Statistics: Simulation and Computation, 49(11):3028–3040, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Bulut:2023:IGL

[Bul23] Y. Murat Bulut. Inverse Gaussian Liu-type estimator. Communications in Statistics: Simulation and Computation, 52 (10):4864–4879, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bulut:2024:RTS

[Bul24] Hasan Bulut. A robust test statistic for independence in high dimensional data. Communications in Statistics: Simulation and Computation, 53(2):702–713, 2024. CODEN CSSCDB. ISSN 0361-0918.

Buyukkaya:2023:CRB

[Büy23] Melek Eris Büyükkaya. Characterizing relationships between BLUPs under linear mixed model and some associated reduced models. Communications in Statistics: Simulation and Computation, 52(8):3438–3451, 2023. CODEN CSSCDB. ISSN 0361-0918.

Bulgakova:2022:NSR

[BV22] T. E. Bulgakova and A. V. Voytishek. On numerical stability of randomized projection functional algorithms. Communications in Statistics: Simulation and Computation, 51(4): 1637–1646, 2022. CODEN CSSCDB. ISSN 0361-0918.

Birteeb:2022:ECT

[BVJ^{+}22] Peter T. Birteeb, Cini Varghese, Seema Jaggi, Eldho Varghese, and Mohd Harun. An efficient class of tree network balanced designs for agroforestry experimentation. Communications in Statistics: Simulation and Computation, 51(12): 7169–7179, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhowmik:2022:GFD

[BVJV22] Arpan Bhowmik, Eldho Varghese, Seema Jaggi, and Cini Varghese. On the generation of factorial designs with minimum level changes. Communications in Statistics: Simulation and Computation, 51(6):3400–3409, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Bahcecitapar:2024:PAA

[BYK24] Melike Bahçecitapar, Elif Yildirim, and Duru Karasoy. Power analysis of approximation methods for parameter estimation in Cox regression model with longitudinal covariate and tied survival times. Communications in Statistics: Simulation and Computation, 53(3):1089–1106, 2024. CODEN CSSCDB. ISSN 0361-0918.

Benghoul:2024:DHO

[BYS24] Maroua Benghoul, Berna Yazici, and Ahmet Sezer. Detection and handling outliers in longitudinal data: wavelets decomposition as a solution. Communications in Statistics: Simulation and Computation, 53(3):1472–1483, 2024. CODEN CSSCDB. ISSN 0361-0918.

Bedouhene:2022:NMB

[BZ22a] Kahina Bedouhene and Nabil Zougab. Nonparametric multiplicative bias correction for von Mises kernel circular density estimator. Communications in Statistics: Simulation and Computation, 51(3):774–792, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bulut:2022:ICR

[BZ22b] Hasan Bulut and Tolga Zaman. An improved class of robust ratio estimators by using the minimum covariance determinant estimation. Communications in Statistics: Simulation and Computation, 51(5):2457–2463, 2022. CODEN CSSCDB. ISSN 0361-0918.

Culos:2022:ABC

[CAA22] Anthony E. Culos, Jeffrey L. Andrews, and Hamid Afshari. An artificial bee colony algorithm for mixture model-based clustering. Communications in Statistics: Simulation and Computation, 51(10):5658–5669, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cavus:2024:COW

[Cav24] Mustafa Cavus. Comparison of one-way ANOVA tests under unequal variances in terms of median p-values. Communications in Statistics: Simulation and Computation, 53(4): 1619–1632, 2024. CODEN CSSCDB. ISSN 0361-0918.

Chesneau:2022:PED

[CBRL22] Christophe Chesneau, Hassan S. Bakouch, Pedro L. Ramos, and Francisco Louzada. The polynomial-exponential distribution: a continuous probability model allowing for occurrence of zero values. Communications in Statistics: Simulation and Computation, 51(8):4581–4606, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Chaturvedi:2021:SPE

[CCBJ21] Ajit Chaturvedi, Saibal Chattopadhyay, Sudeep R. Bapat, and Neeraj Joshi. Sequential point estimation procedures for the parameter of a family of distributions. Communications in Statistics: Simulation and Computation, 50(9):2678–2704, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chiew:2022:CDS

[CCBN22] Esther Chiew, Katherine Cauthen, Nathanael Brown, and Linda Nozick. Comparison of distribution selection methods. Communications in Statistics: Simulation and Computation, 51(4):1982–2005, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:SOC

[CCH20] Jen-Hao Chen, Yen-Chang Chang, and Wen-Liang Hung. A self-organizing clustering algorithm for functional data. Communications in Statistics: Simulation and Computation, 49 (5):1237–1263, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chowdhury:2020:COP

[CCM20] M. Chowdhury, M. Chen, and S. Mandal. A class of optimization problems on minimizing variance based criteria in respect of parameter estimators of a linear model. Communications in Statistics: Simulation and Computation, 49(10): 2719–2731, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chaudhary:2023:CEP

[CD23] Manoj K. Chaudhary and Tulika Dutta. Calibration estimator of population mean under stratified two-phase sampling using two auxiliary variables. Communications in Statistics: Simulation and Computation, 52(11):5423–5433, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Couto:2021:NDM

[CDB^{+}21] Ricardo Couto, Luiz H. Duczmal, Denise Burgarelli, Felipe Álvares, and Gladston J. P. Moreira. Nonparametric dependence modeling via cluster analysis: a financial contagion application. Communications in Statistics: Simulation and Computation, 50(2):537–556, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chiapella:2022:PED

[CdCG22] Luciana Carla Chiapella and María del Carmen Garcia. Performance evaluation of different computational methods to estimate Wood’s lactation curve by nonlinear mixed-effects models. Communications in Statistics: Simulation and Computation, 51(11):6414–6424, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cetinkaya:2023:SSR

[Çet23]  Çagatay Çetinkaya. The stress-strength reliability model

with component strength under partially accelerated life test. Communications in Statistics: Simulation and Computation, 52(10):4665–4684, 2023. CODEN CSSCDB. ISSN 0361-0918.

Channouf:2021:SSC

[CFM21] Nabil Channouf, Marc Fredette, and Brenda MacGibbon. Sample size calculations for hierarchical Poisson and zeroinflated Poisson regression models. Communications in Statistics: Simulation and Computation, 50(4):937–956, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chavent:2021:CCV

[CGS21] Marie Chavent, Robin Genuer, and Jérôme Saracco. Combining clustering of variables and feature selection using random forests. Communications in Statistics: Simulation and Computation, 50(2):426–445, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2023:CTA

[CH23] Huan Cheng and Jianghua He. Comparison of tests for association of 2×2 tables under multiple testing setting. Communications in Statistics: Simulation and Computation, 52(6): 2336–2348, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chang:2023:FDC

[Cha23a] Jing Chang. Functional data classification using the rotation forest method combined with the patch selection. Communications in Statistics: Simulation and Computation, 52(7): 3365–3378, 2023. CODEN CSSCDB. ISSN 0361-0918.

Changpetch:2023:LRM

[Cha23b] Pannapa Changpetch. Logistic regression model with TreeNet and association rules analysis: applications with medical datasets. Communications in Statistics: Simulation and Computation, 52(6):2633–2645, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:SIM

[CHC20] Jinbao Chen, Yawen Hou, and Zheng Chen. Statistical inference methods for cumulative incidence function curves at a fixed point in time. Communications in Statistics: Simulation and Computation, 49(1):79–94, 2020. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2021:ELR

[Che21] Conghua Cheng. Empirical likelihood ratio under infinite second moment for two-sample problems. Communications in Statistics: Simulation and Computation, 50(10):3069–3076, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2022:CNP

[Che22] Hao Cheng. A class of new partial least square algorithms for first and higher order models. Communications in Statistics: Simulation and Computation, 51(8):4349–4371, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2022:BEA

[CHL22a] Hao Chen, Lanshan Han, and Alvin Lim. Beyond the EM algorithm: constrained optimization methods for latent class model. Communications in Statistics: Simulation and Computation, 51(9):5222–5244, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:PCD

[CHL22b] Wenting Chen, Xin-Jiang He, and Sha Lin. Pricing credit default swaps with Parisian and Parasian default mechanics. Communications in Statistics: Simulation and Computation, 51(2):421–431, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2024:ELM

[CHL24] Hao Chen, Lanshan Han, and Alvin Lim. Estimating linear mixed effects models with truncated normally distributed random effects. Communications in Statistics: Simulation and Computation, 53(4):2050–2070, 2024. CODEN CSSCDB. ISSN 0361-0918.

Cinar:2022:ENA

[CII22] Ozan Cinar, Cem Iyigun, and Ozlem Ilk. An evaluation of a novel approach for clustering genes with dissimilar replicates. Communications in Statistics: Simulation and Computation, 51(12):7458–7471, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ciuperca:2022:AVS

[Ciu22] Gabriela Ciuperca. Automatic variable selection in a linear model on massive data. Communications in Statistics: Simulation and Computation, 51(9):4937–4956, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cekim:2020:TEP

[CK20] Hatice Oncel Cekim and Cem Kadilar. In-type estimators for the population variance in stratified random sampling. Communications in Statistics: Simulation and Computation, 49(7):1665–1677, 2020. CODEN CSSCDB. ISSN 0361-0918.

Curtin:2021:RSI

[CKCG21] Paul Curtin, Joshua Kellogg, Nadja Cech, and Chris Gennings. A random subset implementation of weighted quantile sum (WQS_{RS}) regression for analysis of high-dimensional mixtures. Communications in Statistics: Simulation and Computation, 50(4):1119–1134, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chaturvedi:2021:NSR

[CKK21] Ajit Chaturvedi, Taruna Kumari, and Narendra Kumar. Numerical study of robust Bayesian analysis of generalized inverted family of distributions based on progressive type II right censoring. Communications in Statistics: Simulation and Computation, 50(11):3207–3240, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Cetinkaya:2023:DFO

[ÇKK23] Merve Kandemir Çetinkaya, Selahattin Kaçiranlar, and Fikriye Kurtoglu. Developed first-order approximated estimators for the gamma distributed response variable. Communications in Statistics: Simulation and Computation, 52 (8):3919–3938, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chew:2022:ERR

[CKL22] XinYing Chew, Khai Wah Khaw, and Ming Ha Lee. The efficiency of run rules schemes for the multivariate coefficient of variation in short runs process. Communications in Statistics: Simulation and Computation, 51(6):2942–2962, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Chatterjee:2024:MPM

[CKL24] Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. Monitoring process mean and variability with one triple EWMA chart. Communications in Statistics: Simulation and Computation, 53(2):611–641, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Chong:2020:ODM

[CKT^{+}20] Zhi Lin Chong, Michael Boon Chong Khoo, Wei Lin Teoh, Wai Chung Yeong, and Sok Li Lim. Optimal design of the modified group runs (MGR) X̄ chart when process parameters are estimated. Communications in Statistics: Simulation and Computation, 49(1):244–260, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chin:2021:NHF

[CL21] Wen Cheong Chin and Min Cherng Lee. Nonlinear highfrequency stock market time series: Modeling and combine forecast evaluations. Communications in Statistics: Simulation and Computation, 50(7):2126–2144, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Castellares:2022:CEG

[CL22a] Fredy Castellares and Artur J. Lemonte. Comments on “An extended Gompertz–Makeham distribution with application to lifetime data”. Communications in Statistics: Simulation and Computation, 51(5):2769–2777, 2022. CODEN CSSCDB. ISSN 0361-0918. See [Abd18].

Chiu:2022:IVA

[CL22b] Sung Nok Chiu and Kwong Ip Liu. Improving p-value approximation and level accuracy of Monte Carlo tests by quasi-Monte Carlo methods. Communications in Statistics: Simulation and Computation, 51(3):1272–1288, 2022. CODEN CSSCDB. ISSN 0361-0918.

Collins:2022:SRS

[CL22c] David Collins and Yan Lu. A stratified reservoir sampling algorithm in streams and large datasets. Communications in Statistics: Simulation and Computation, 51(4):1767–1782, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2023:NRE

[CL23] Yinjun Chen and Huilan Liu. A new relative error estimation for partially linear multiplicative model. Communications in Statistics: Simulation and Computation, 52(10):4962–4980, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:MFS

[CLC22] Xiaolin Chen, Wei Liu, and Xiaojing Chen. Model-free survival conditional feature screening. Communications in Statistics: Simulation and Computation, 51(10):5690–5708, 2022. CODEN CSSCDB. ISSN 0361-0918.

Choi:2021:PSM

[CLJK21] Junsouk Choi, JungJun Lee, Jae-Hwan Jhong, and Ja-Yong Koo. Penalized I-spline monotone regression estimation. Communications in Statistics: Simulation and Computation, 50(11):3714–3732, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:CML

[CLL20] Nan Chen, Meijuan Li, and Hongyun Liu. Comparison of maximum likelihood approach, Diggle–Kenward selection model, pattern mixture model with MAR and MNAR dropout data. Communications in Statistics: Simulation and Computation, 49(7):1746–1767, 2020. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2023:NMM

[CLL23] Maolin Cheng, Yun Liu, and Jianuo Li. A new modeling method of gray GM(1, N) model and its application to predicting China’s clean energy consumption. Communications in Statistics: Simulation and Computation, 52(8):3712–3723, 2023. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2022:PEM

[CLLX22] Maolin Cheng, Jiano Li, Yun Liu, and Minyin Xiang. Parameter estimation of modified gray model GM(1, N) and model application. Communications in Statistics: Simulation and Computation, 51(9):5447–5464, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:NGW

[CLS22] Jen-Hsiang Chen, Shin-Li Lu, and Shey-Huei Sheu. A nonparametric generally weighted moving average sign chart based on repetitive sampling. Communications in Statistics: Simulation and Computation, 51(3):1137–1156, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2021:OCB

[CLY21] Lee-Shen Chen, TaChen Liang, and Ming-Chung Yang. Optimal curtailed Bayesian sampling plans for exponential distributions with Type-I hybrid censored samples. Communications in Statistics: Simulation and Computation, 50(3): 764–777, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cao:2023:SAA

[CLY23] Wentao Cao, Yaling Li, and Qingzhao Yu. Sensitivity analysis for assumptions of general mediation analysis. Communications in Statistics: Simulation and Computation, 52(6): 2453–2470, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chekkal:2023:GBS

[CLZ23a] Sylia Chekkal, Karima Lagha, and Nabil Zougab. Generalized Birnbaum–Saunders kernel for hazard rate function estimation. Communications in Statistics: Simulation and Computation, 52(4):1546–1561, 2023. CODEN CSSCDB. ISSN 0361-0918.

Collier:2023:EPS

[CLZ23b] Zachary K. Collier, Walter L. Leite, and Haobai Zhang. Estimating propensity scores using neural networks and traditional methods: a comparative simulation study. Communications in Statistics: Simulation and Computation, 52(9): 4545–4560, 2023. CODEN CSSCDB. ISSN 0361-0918.

Coolen-Maturi:2022:OFO

[CM22] Tahani Coolen-Maturi. The ordering of future observations from multiple groups. Communications in Statistics: Simulation and Computation, 51(12):7526–7543, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cockeran:2021:GFT

[CMA21] Marike Cockeran, Simos George Meintanis, and James S. Allison. Goodness-of-fit tests in the Cox proportional hazards model. Communications in Statistics: Simulation and Computation, 50(12):4132–4143, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chaudhuri:2023:TDT

[CMM23] Somnath Chaudhuri, Mehdi Moradi, and Jorge Mateu. On the trend detection of time-ordered intensity images of point processes on linear networks. Communications in Statistics: Simulation and Computation, 52(4):1318–1330, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Chiodini:2021:RSA

[CMMV21] Paola Maddalena Chiodini, Giancarlo Manzi, Bianca Maria Martelli, and Flavio Verrecchia. Revisiting sample allocation methods: a simulation-based comparison. Communications in Statistics: Simulation and Computation, 50(8):2197–2212, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cribari-Neto:2021:RBP

[CNL21] Francisco Cribari-Neto and Fábio P. Lima. Resampling-based prediction intervals in beta regressions under correct and incorrect model specification. Communications in Statistics: Simulation and Computation, 50(5):1398–1416, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Costa:2023:PCC

[Cos23] Antonio F. B. Costa. The performance of the Cpk chart. Communications in Statistics: Simulation and Computation, 52(8):3912–3918, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:FTA

[CP22] Chao Chen and Haiyan Pan. A fiducial test for assessing the non-inferiority of odds ratio in matched-pairs design. Communications in Statistics: Simulation and Computation, 51 (6):2978–2991, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chakraborty:2023:WCT

[CP23] Siddhartha Chakraborty and Biswabrata Pradhan. On weighted cumulative Tsallis residual and past entropy measures. Communications in Statistics: Simulation and Computation, 52(5):2058–2072, 2023. CODEN CSSCDB. ISSN 0361-0918.

Corredor:2022:SEP

[CQ22] J. S. Corredor and A. J. Quiroz. Shannon’s entropy of partitions determined by hierarchical clustering trees in asymmetry and dimension identification. Communications in Statistics: Simulation and Computation, 51(10):5954–5966, 2022. CODEN CSSCDB. ISSN 0361-0918.

Carrasco:2021:SRM

[CR21] Jalmar M. F. Carrasco and Nancy Reid. Simplex regression models with measurement error. Communications in Statistics: Simulation and Computation, 50(11):3420–3435, 2021. CODEN CSSCDB. ISSN 0361-0918.

Choi:2020:BVS

[CS20] Ji-Eun Choi and Dong Wan Shin. Bootstrapping volatility spillover index. Communications in Statistics: Simulation and Computation, 49(1):66–78, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:RAP

[CS22a] Xudan Chen and Xinli Sun. Reliability assessment for products with two performance characteristics based on marginal stochastic processes and copulas. Communications in Statistics: Simulation and Computation, 51(7):3621–3644, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2022:IMP

[CS22b] Maolin Cheng and Guojun Shi. Improved methods for parameter estimation of gray model GM(1,1) based on new background value optimization and model application. Communications in Statistics: Simulation and Computation, 51(2): 647–669, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chumnaul:2022:MSL

[CS22c] Jularat Chumnaul and Mohammad Sepehrifar. Modified signed log-likelihood ratio test for the scale parameter of the power-law process with applications to repairable systems. Communications in Statistics: Simulation and Computation, 51(8):4828–4844, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chouia:2023:DEG

[CSA23] Sana Chouia and Nacira Seddik-Ameur. Different EDF goodness-of-fit tests for competing risks models. Communications in Statistics: Simulation and Computation, 52(8): 3491–3501, 2023. CODEN CSSCDB. ISSN 0361-0918.

Cancho:2022:MSM

[CSBdES22] Vicente G. Cancho, Adriano K. Suzuki, Gladys D. C. Barriga, and Ana P. J. do Espirito Santo. A multivariate survival model induced by discrete frailty. Communications in Statistics: Simulation and Computation, 51(11):6572–6590, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:IGP

[CSDT20] Xudan Chen, Xinli Sun, Xiong Ding, and Jue Tang. The inverse Gaussian process with a skew-normal distribution as a degradation model. Communications in Statistics: Simulation and Computation, 49(11):2827–2843, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2021:PCP

[CSE^{+}21] Yanjun Chen, Damla Sentürk, Jason P. Estes, Luis F. Campos, Connie M. Rhee, Lorien S. Dalrymple, Kamyar Kalantar-Zadeh, and Danh V. Nguyen. Performance characteristics of profiling methods and the impact of inadequate case-mix adjustment. Communications in Statistics: Simulation and Computation, 50(6):1854–1871, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cuadras:2022:CMF

[CSG22] Carles M. Cuadras and Sonia Salvo-Garrido. Comparing multiple factor analysis and related metric scaling. Communications in Statistics: Simulation and Computation, 51(5):2343– 2362, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cai:2020:BAD

[CSL20] Jing Cai, Yimin Shi, and Bin Liu. Bayesian analysis for dependent competing risks model with masked causes of failure in step-stress accelerated life test under progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 49(9):2302–2320, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:GFM

[CsSL22] Chyong-Mei Chen, Pao sheng Shen, and Ting-Hsuan Lee. A gamma-frailty model for interval-censored data with dependent examination times: a computationally efficient approach. Communications in Statistics: Simulation and Computation, 51(10):6071–6082, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chang:2021:CID

[CST21] Yu-Mei Chang, Pao-Sheng Shen, and Yu-Hsin Tang. Confidence interval for the difference between two median survival times with semiparametric transformation models. Communications in Statistics: Simulation and Computation, 50(3): 633–649, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2020:IPE

[CSX20] Maolin Cheng, Guojun Shi, and Mingyin Xiang. On the improvement of the parameter estimation of the grey model GM(1,1) and model application. Communications in Statistics: Simulation and Computation, 49(5):1367–1384, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:OCI

[CTCL22] Shun-Yi Chen, Chih-Chun Tsai, Hubert J. Chen, and Su-Hao Lee. Optimal confidence interval for the largest exponential location parameter. Communications in Statistics: Simulation and Computation, 51(7):4057–4072, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Chong:2022:ODE

[CTK^{+}22] Z. L. Chong, K. L. Tan, Michael B. C. Khoo, W. L. Teoh, and P. Castagliola. Optimal designs of the exponentially weighted moving average (EWMA) median chart for known and estimated parameters based on median run length. Communications in Statistics: Simulation and Computation, 51(7): 3660–3684, 2022. CODEN CSSCDB. ISSN 0361-0918.

Curto:2023:CIM

[Cur23] José Dias Curto. Confidence intervals for means and variances of nonnormal distributions. Communications in Statistics: Simulation and Computation, 52(9):4414–4430, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2023:LBE

[CW23] Tao Chen and Lichun Wang. Linear Bayes estimator of the extreme value distribution based on type II censored samples. Communications in Statistics: Simulation and Computation, 52(9):4532–4544, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chen:2024:MCC

[CW24] Pei-Le Chen and Mei-Han Wang. The multivariate control chart with variable dimension for monitoring both location and dispersion. Communications in Statistics: Simulation and Computation, 53(3):1534–1547, 2024. CODEN CSSCDB. ISSN 0361-0918.

Chen:2023:ETM

[CWLH23] Jin Chen, Dehui Wang, Cong Li, and Jingwen Huang. Estimation and testing of multivariate random coefficient autoregressive model based on empirical likelihood. Communications in Statistics: Simulation and Computation, 52(2): 291–308, 2023. CODEN CSSCDB. ISSN 0361-0918.

Cao:2023:PRE

[CWW23] Xingyun Cao, Danlu Wang, and Liucang Wu. Performance of ridge estimator in skew-normal mode regression model. Communications in Statistics: Simulation and Computation, 52 (3):1164–1177, 2023. CODEN CSSCDB. ISSN 0361-0918.

Cao:2024:AEP

[CWWT24] Lixia Cao, Cong Wang, Tonghui Wang, and David Trafimow. The APP for estimating population proportion based on skew normal approximations and the Beta–Bernoulli process. Communications in Statistics: Simulation and Computation, 53 (1):167–177, 2024. CODEN CSSCDB. ISSN 0361-0918.

Chou:2024:VML

[CY24] Elizabeth P. Chou and Shan-Ping Yang. A virtual multi-label approach to imbalanced data classification. Communications in Statistics: Simulation and Computation, 53(3):1461–1471, 2024. CODEN CSSCDB. ISSN 0361-0918.

Cavus:2021:PPA

[CYS21] Mustafa Cavus, Berna Yazici, and Ahmet Sezer. Penalized power approach to compare the power of the tests when Type I error probabilities are different. Communications in Statistics: Simulation and Computation, 50(7):1912–1926, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cavus:2023:PPP

[CYS23] Mustafa Cavus, Berna Yazici, and Ahmet Sezer. Penalized power properties of the normality tests in the presence of outliers. Communications in Statistics: Simulation and Computation, 52(8):3568–3580, 2023. CODEN CSSCDB. ISSN 0361-0918.

Chwila:2022:PEB

[CZ22] Adam Chwila and Tomasz Zadło. On properties of empirical best predictors. Communications in Statistics: Simulation and Computation, 51(1):220–253, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2021:CSR

[CZP21] Chen Chen, Jie Zhou, and Jianxin Pan. Correlation structure regularization via entropy loss function for high-dimension and low-sample-size data. Communications in Statistics: Simulation and Computation, 50(4):993–1008, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cui:2021:MDE

[CZS21] Lihong Cui, Qianhui Zhang, and Jianjun Sun. Multiwavelet density estimation for biased data. Communications in Statistics: Simulation and Computation, 50(1):234–253, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Dogru:2021:RMR

[DA21] Fatma Zehra Dogru and Olcay Arslan. Robust mixture regression modeling based on the generalized M (GM)-estimation method. Communications in Statistics: Simulation and Computation, 50(9):2643–2665, 2021. CODEN CSSCDB. ISSN 0361-0918.

Deng:2022:BPA

[DA22] M. Deng and M. S. Aminzadeh. Bayesian predictive analysis for Weibull–Pareto composite model with an application to insurance data. Communications in Statistics: Simulation and Computation, 51(5):2683–2709, 2022. CODEN CSSCDB. ISSN 0361-0918.

Inacio:2020:CTP

[dAIIS20] Marco Henrique de Almeida Inácio, Rafael Izbicki, and Luis Ernesto Salasar. Comparing two populations using Bayesian Fourier series density estimation. Communications in Statistics: Simulation and Computation, 49(1):261–282, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lima-Filho:2021:KCC

[dALFB21] Luiz Medeiros de Araujo Lima-Filho and Fábio Mariano Bayer. Kumaraswamy control chart for monitoring double bounded environmental data. Communications in Statistics: Simulation and Computation, 50(9):2513–2528, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Neto:2021:WSP

[dALNPdOF21] Eufrásio de Andrade Lima Neto, Aluísio Pinheiro, and Adenice Gomes de Oliveira Ferreira. On wavelet to select the parametric form of a regression model. Communications in Statistics: Simulation and Computation, 50(9):2619–2642, 2021. CODEN CSSCDB. ISSN 0361-0918.

Dawoud:2023:NIE

[Daw23] Issam Dawoud. A new improved estimator for reducing the multicollinearity effects. Communications in Statistics: Simulation and Computation, 52(8):3581–3592, 2023. CODEN CSSCDB. ISSN 0361-0918.

Dewan:2023:NNP

[DBS23] Isha Dewan, Deepesh Bhati, and K. K. Sudheesh. A new non-parametric test for testing positive quadrant dependence. Communications in Statistics: Simulation and Computation, 52(10):5090–5098, 2023. CODEN CSSCDB. ISSN 0361-0918.

Davidson:2020:MPF

[DC20] Heather Davidson and Robert A. Cribbie. A more powerful familywise error control procedure for evaluating mean equivalence. Communications in Statistics: Simulation and Computation, 49(11):2914–2929, 2020. CODEN CSSCDB. ISSN 0361-0918.

DeCanditiis:2023:MSI

[DC23] Daniela De Canditiis and Silvia Cirulli. Model selection for inferring Gaussian graphical models. Communications in Statistics: Simulation and Computation, 52(12):6084–6095, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ding:2024:USC

[DC24] Liwang Ding and Ping Chen. Uniform strong consistency of histogram density estimation for ϕ-mixing samples. Communications in Statistics: Simulation and Computation, 53(2): 1037–1047, 2024. CODEN CSSCDB. ISSN 0361-0918.

delCastillo:2022:CRL

[dCCM^{+}22] Enrique del Castillo, Peng Chen, Adam Meyers, John Hunt, and James Rapkin. Confidence regions for the location of response surface optima: the R package OptimaRegion. Communications in Statistics: Simulation and Computation, 51 (12):7074–7094, 2022. CODEN CSSCDB. ISSN 0361-0918.

deCarvalho:2023:SSS

[dCdAB23] Daniel Matos de Carvalho, Getúlio José Amorim do Amaral, and Fernanda De Bastiani. Spatial scan statistics based on empirical likelihood. Communications in Statistics: Simulation and Computation, 52(8):3897–3911, 2023. CODEN CSSCDB. ISSN 0361-0918.

Du:2024:WJA

[DCL^{+}24] Ruofei Du, Ye Jin Choi, Ji-Hyun Lee, Songthip Ounpraseuth, and Zhuopei Hu. A weight jackknife approach utilizing linear model based-estimators for clustered data. Communications in Statistics: Simulation and Computation, 53(2):1048–1067, 2024. CODEN CSSCDB. ISSN 0361-0918.

Silva:2024:RCI

[dCSF24] Bruna da Costa Silva and Daniel Furtado Ferreira. Robust computationally intensive and asymptotic tests for compound symmetry structure. Communications in Statistics: Simulation and Computation, 53(4):1645–1656, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Ding:2022:SIH

[DCY22] Liwang Ding, Ping Chen, and Li Yongming. Statistical inference for a heteroscedastic regression model with ϕ-mixing errors. Communications in Statistics: Simulation and Computation, 51(8):4658–4676, 2022. CODEN CSSCDB. ISSN 0361-0918.

Djeddour-Djaballah:2021:QML

[DDK21] K. Djeddour-Djaballah and L. Kerar. Quasi-maximum likelihood estimation of GARCH with Student distributed noise. Communications in Statistics: Simulation and Computation, 50(5):1249–1271, 2021. CODEN CSSCDB. ISSN 0361-0918.

DeGaetano:2020:FGM

[De 20] Davide De Gaetano. Forecasting with GARCH models under structural breaks: an approach based on combinations across estimation windows. Communications in Statistics: Simulation and Computation, 49(10):2559–2582, 2020. CODEN CSSCDB. ISSN 0361-0918.

DeOliveira:2022:NSN

[DE22] Victor De Oliveira and Mark D. Ecker. A non-stationary non-Gaussian hedonic spatial model for house selling prices. Communications in Statistics: Simulation and Computation, 51(6):2888–2905, 2022. CODEN CSSCDB. ISSN 0361-0918.

Duran:2022:MRA

[DF22] Ahmet Duran and Mohammed Farrukh. Multiple regression analysis for dynamics of patient volumes. Communications in Statistics: Simulation and Computation, 51(6):2906–2923, 2022. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2022:MDG

[DG22] H. Demirtas and R. Gao. Mixed data generation packages and related computational tools in R. Communications in Statistics: Simulation and Computation, 51(8):4520–4563, 2022. CODEN CSSCDB. ISSN 0361-0918.

DeSantis:2023:CIP

[DG23a] Fulvio De Santis and Stefania Gubbiotti. Confidence intervals for the parameter of the scaled-uniform model. Communications in Statistics: Simulation and Computation, 52(12): 6235–6248, 2023. CODEN CSSCDB. ISSN 0361-0918.

DiNardo:2023:GNB

[DG23b] Elvira Di Nardo and Giuseppe Guarino. On the generation of necklaces and bracelets in R. Communications in Statistics: Simulation and Computation, 52(11):5727–5737, 2023. CODEN CSSCDB. ISSN 0361-0918.

Dahdouh:2022:GSS

[DK22a] Omar Dahdouh and Majid Jafari Khaledi. Generalized spatial stick-breaking processes. Communications in Statistics: Simulation and Computation, 51(8):4638–4657, 2022. CODEN CSSCDB. ISSN 0361-0918.

Deebani:2022:MCE

[DK22b] Wejdan Deebani and Nezamoddin N. Kachouie. Monte Carlo ensemble correlation coefficient for association detection. Communications in Statistics: Simulation and Computation, 51(12):7095–7109, 2022. CODEN CSSCDB. ISSN 0361-0918.

Dag:2022:DCE

[DKKA22] Osman Dag, Merve Kasikci, Erdem Karabulut, and Reha Alpar. Diverse classifiers ensemble based on GMDH-type neural network algorithm for binary classification. Communications in Statistics: Simulation and Computation, 51(5):2440–2456, 2022. CODEN CSSCDB. ISSN 0361-0918.

Draxler:2022:GTF

[DKL22] Clemens Draxler, Andreas Kurz, and Artur J. Lemonte. The gradient test and its finite sample size properties in a conditional maximum likelihood and psychometric modeling context. Communications in Statistics: Simulation and Computation, 51(6):3185–3203, 2022. CODEN CSSCDB. ISSN 0361-0918.

Duan:2020:FSF

[DL20] Qihong Duan and Junrong Liu. Filtering and smoothing formulas of AR(p)-modulated Poisson processes. Communications in Statistics: Simulation and Computation, 49(6):1575– 1591, 2020. CODEN CSSCDB. ISSN 0361-0918.

Dey:2021:NRM

[DLRZ21] Dipak K. Dey, Junfeng Liu, Nalini Ravishanker, and Edwards Qiang Zhang. A note on response mean confidence band for linear regression models. Communications in Statistics: Simulation and Computation, 50(3):778–785, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Demni:2023:DBD

[DMP23] Houyem Demni, Amor Messaoud, and Giovanni C. Porzio. Distance-based directional depth classifiers: a robustness study. Communications in Statistics: Simulation and Computation, 52(11):5695–5713, 2023. CODEN CSSCDB. ISSN 0361-0918.

Du:2022:ETD

[DNWG22] Yujing Du, Hon Keung Tony Ng, Jun Wang, and Wei Gao. On estimation of two-dimensional dynamic panel model with confounders. Communications in Statistics: Simulation and Computation, 51(2):505–517, 2022. CODEN CSSCDB. ISSN 0361-0918.

Dogan:2022:SSC

[Dog22] Ilkay Dogan. A simulation study comparing model fit measures of structural equation modeling with multivariate contaminated normal distribution. Communications in Statistics: Simulation and Computation, 51(5):2526–2536, 2022. CODEN CSSCDB. ISSN 0361-0918.

deOliveira:2021:GBD

[dOMA21] Ricardo Puziol de Oliveira, Josmar Mazucheli, and Jorge Alberto Achcar. A generalization of Basu–Dhar’s bivariate geometric distribution to the trivariate case. Communications in Statistics: Simulation and Computation, 50(12):4388–4412, 2021. CODEN CSSCDB. ISSN 0361-0918.

deOliveira:2021:SCA

[dOMD^{+}21] Dênis Ricardo Xavier de Oliveira, Gladston Moreira, Anderson Ribeiro Duarte, André Cançado, and Eduardo Luz. Spatial cluster analysis using particle swarm optimization and dispersion function. Communications in Statistics: Simulation and Computation, 50(8):2368–2385, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Valadares:2021:NCR

[dOVQP21] Danilo Gilberto de Oliveira Valadares, Roberto Costa Quinino, and Magda Carvalho Pires. The need to conduct repeated classifications in a logistic regression model with misclassification in the dependent variable. Communications in Statistics: Simulation and Computation, 50(5):1459–1472, 2021. CODEN CSSCDB. ISSN 0361-0918.

Dedduwakumara:2020:CIQ

[DP20] Dilanka S. Dedduwakumara and Luke A. Prendergast. Confidence intervals for quantiles from histograms and other grouped data. Communications in Statistics: Simulation and Computation, 49(6):1546–1559, 2020. CODEN CSSCDB. ISSN 0361-0918.

Darkhovsky:2022:MFC

[DP22] Boris Darkhovsky and Alexandra Piryatinska. Model-free classification of panel data via the ϵ-complexity theory. Communications in Statistics: Simulation and Computation, 51 (7):4073–4086, 2022. CODEN CSSCDB. ISSN 0361-0918.

Dhanushya:2023:TSV

[DP23] S. Dhanushya and T. Palanisamy. Two-stage variational mode decomposition approach to enhance the estimates of variance function. Communications in Statistics: Simulation and Computation, 52(11):5605–5614, 2023. CODEN CSSCDB. ISSN 0361-0918.

Dufitinema:2022:MLE

[DPS22] Josephine Dufitinema, Seppo Pynnönen, and Tommi Sottinen. Maximum likelihood estimators from discrete data modeled by mixed fractional Brownian motion with application to the Nordic stock markets. Communications in Statistics: Simulation and Computation, 51(9):5264–5287, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Divecha:2022:TEL

[DR22] Jyoti Divecha and D. P. Raykundaliya. Three economical life test acceptance sampling plans. Communications in Statistics: Simulation and Computation, 51(6):3305–3323, 2022. CODEN CSSCDB. ISSN 0361-0918.

Danielius:2023:VCP

[DR23] Tadas Danielius and Alfredas Rackauskas. p-variation of CUSUM process and testing change in the mean. Communications in Statistics: Simulation and Computation, 52(1): 43–55, 2023. CODEN CSSCDB. ISSN 0361-0918.

Das:2020:AUW

[DRB20] Sourav Das, Tata Subba Rao, and Georgi N. Boshnakov. An asymptotically unbiased weighted least squares estimation criterion for parametric variograms of second order stationary geostatistical processes. Communications in Statistics: Simulation and Computation, 49(7):1839–1854, 2020. CODEN CSSCDB. ISSN 0361-0918.

Demirarslan:2024:OAS

[DS24] Mert Demirarslan and Asli Suner. OCtS: an alternative of the t-Score method sensitive to outliers and correlation in feature selection. Communications in Statistics: Simulation and Computation, 53(3):1409–1422, 2024. CODEN CSSCDB. ISSN 0361-0918.

dosSantos:2023:CAV

[dSC23] Patricia Mendes dos Santos and Marcelo Ângelo Cirillo. Construction of the average variance extracted index for construct validation in structural equation models with adaptive regressions. Communications in Statistics: Simulation and Computation, 52(4):1639–1650, 2023. CODEN CSSCDB. ISSN 0361-0918.

da-Silva:2021:DMA

[dSdSTC21] Cibele Q. da Silva, Paulo H. D. da Silva, Osiris Turnes, and Leandro T. Correia. Dynamic model averaging adapted to dynamic regression models for time series of counts. Communications in Statistics: Simulation and Computation, 50(2): 501–524, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ding:2020:MMB

[DSL20] Hao Ding, Ziwei Su, and Xiaoqian Liu. A modified multinomial baseline logit model with logit functions having different covariates. Communications in Statistics: Simulation and Computation, 49(11):2861–2875, 2020. CODEN CSSCDB. ISSN 0361-0918.

daSilva:2023:BEM

[dSLHMB23] Marcelo A. da Silva, Ren Liu, Anne Corinne Huggins-Manley, and Jorge L. Bazán. Bayesian estimation of multidimensional polytomous item response theory models with Q-matrices using Stan. Communications in Statistics: Simulation and Computation, 52(11):5178–5194, 2023. CODEN CSSCDB. ISSN 0361-0918.

Duarte:2023:STB

[DSO^{+}23] A. R. Duarte, S. B. Silva, F. L. P. Oliveira, A. C. L. Almeida, and L. H. Duczmal. Space-time border analysis to evaluate and detect clusters. Communications in Statistics: Simulation and Computation, 52(6):2275–2289, 2023. CODEN CSS-CDB. ISSN 0361-0918.

dosSantos:2023:MTB

[dSPdSB23] Kleber Henrique dos Santos, Tarciana Liberal Pereira, Tatiene Correia de Souza, and Marcelo Bourguignon. A misspecification test for beta prime regression models. Communications in Statistics: Simulation and Computation, 52(10): 4561–4574, 2023. CODEN CSSCDB. ISSN 0361-0918.

deSilva:2022:MAF

[dSRC22] Galappaththige S. R. de Silva, Lasitha N. Rathnayake, and Pankaj K. Choudhary. Modeling and analysis of functional method comparison data. Communications in Statistics: Simulation and Computation, 51(12):7298–7318, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Sousa:2022:BWS

[dSS22] Alex Rodrigo dos Santos Sousa. Bayesian wavelet shrinkage with logistic prior. Communications in Statistics: Simulation and Computation, 51(8):4700–4714, 2022. CODEN CSSCDB. ISSN 0361-0918.

Deepmala:2023:IBB

[DSS23] Deepmala, Sanjay Kumar Singh, and Umesh Singh. Inferences based on a balanced joint progressive type-II censoring scheme for Lindley distributed lifetimes. Communications in Statistics: Simulation and Computation, 52(9):4119–4133, 2023. CODEN CSSCDB. ISSN 0361-0918.

Dogu:2023:MED

[DuA23] Eralp Dogu and Muhammad Noor ul Amin. Monitoring exponentially distributed time between events data: self-starting perspective. Communications in Statistics: Simulation and Computation, 52(3):1104–1118, 2023. CODEN CSSCDB. ISSN 0361-0918.

Du:2023:EBH

[DWLS23] Kai Du, Min Wang, Tom Lu, and Xiaoqian Sun. Estimation based on hybrid censored data from the power Lindley distribution. Communications in Statistics: Simulation and Computation, 52(8):3939–3957, 2023. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:AVR

[dXGLY20] Guo dong Xing, Xiaoli Gan, Xiaohu Li, and Shanchao Yang. On the asymptotics of value-at-risk for portfolio loss under bivariate Eyraud–Farlie–Gumbel–Morgenstern copula and heavy tails. Communications in Statistics: Simulation and Computation, 49(9):2462–2471, 2020. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:ATC

[dXLY20a] Guo dong Xing, Xiaohu Li, and Shanchao Yang. On the asymptotics of tail conditional expectation for portfolio loss under bivariate Eyraud–Farlie–Gumbel–Morgenstern copula and heavy tails. Communications in Statistics: Simulation and Computation, 49(8):2049–2058, 2020. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:SOA

[dXLY20b] Guo dong Xing, Xiaohu Li, and Shanchao Yang. Second-order asymptotics of tail distortion risk measure for portfolio loss in the multivariate regularly varying model. Communications in Statistics: Simulation and Computation, 49(2):491–503, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ding:2022:TSC

[DY22a] Jianhua Ding and Ping Yu. Test shape constraints in semiparametric model with Bernstein polynomials. Communications in Statistics: Simulation and Computation, 51(5):2667– 2682, 2022. CODEN CSSCDB. ISSN 0361-0918.

DOttaviano:2022:MRE

[DY22b] Fabio D’Ottaviano and Wenzhao Yang. On missing random effects in machine learning. Communications in Statistics: Simulation and Computation, 51(11):6320–6331, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ding:2022:BEB

[DYC22] Liwang Ding, Shengwei Yao, and Ping Chen. On Berry– Esseen bound of wavelet estimators in nonparametric regression model under asymptotically negatively associated assumptions. Communications in Statistics: Simulation and Computation, 51(3):924–940, 2022. CODEN CSSCDB. ISSN 0361-0918.

Deng:2022:TVP

[DZLS22] Chunliang Deng, Xingfa Zhang, Yuan Li, and Zefang Song. On the test of the volatility proxy model. Communications in Statistics: Simulation and Computation, 51(12):7390–7403, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ebadi:2020:CCM

[EAJ20] Mohsen Ebadi and Amir Ahmadi-Javid. Control charts for monitoring multi-stage service processes with optimal queue performance. Communications in Statistics: Simulation and Computation, 49(9):2472–2484, 2020. CODEN CSSCDB. ISSN 0361-0918.

Eidous:2022:KMO

[EAT22] Omar M. Eidous and Saif Al-Deen A. Al-Talafha. Kernel method for overlapping coefficients estimation. Communications in Statistics: Simulation and Computation, 51(9):5139– 5156, 2022. CODEN CSSCDB. ISSN 0361-0918.

Erkoc:2023:NPE

[EE23] Ali Erkoç and M. Aydin Erar. A new parameter estimation technique using prior information of parameters in nonlinear model with multicollinear data. Communications in Statistics: Simulation and Computation, 52(5):1926–1936, 2023. CODEN CSSCDB. ISSN 0361-0918.

Elsawah:2021:SIB

[EFD21] A. M. Elsawah, Kai-Tai Fang, and Yu Hui Deng. Some interesting behaviors of good lattice point sets. Communications in Statistics: Simulation and Computation, 50(11):3650–3668, 2021. CODEN CSSCDB. ISSN 0361-0918.

Egashira:2024:APM

[Ega24] Kento Egashira. Asymptotic properties of multiclass support vector machine under high dimensional settings. Communications in Statistics: Simulation and Computation, 53(4): 1991–2005, 2024. CODEN CSSCDB. ISSN 0361-0918.

Epprecht:2021:VSF

[EGVdR21] Camila Epprecht, Dominique Guégan, Álvaro Veiga, and Joel Correa da Rosa. Variable selection and forecasting via automated methods for linear models: LASSO/adaLASSO and Autometrics. Communications in Statistics: Simulation and Computation, 50(1):103–122, 2021. CODEN CSSCDB. ISSN 0361-0918.

El-Horbaty:2022:TAR

[EH22] Yahia S. El-Horbaty. Testing the absence of random effects in the nested-error regression model using orthogonal transformations. Communications in Statistics: Simulation and Computation, 51(5):2736–2746, 2022. CODEN CSSCDB. ISSN 0361-0918.

Elsawah:2023:NAG

[EHF23] A. M. Elsawah, Li Meng Hua, and Kai-Tai Fang. A novel algorithm for generating minimum energy points from identically charged particles in 1D, 2D and 3D unit hypercubes. Communications in Statistics: Simulation and Computation, 52(8):3529–3552, 2023. CODEN CSSCDB. ISSN 0361-0918.

Enomoto:2020:MNT

[EHHS20] Rie Enomoto, Zofia Hanusz, Ayako Hara, and Takashi Seo. Multivariate normality test using normalizing transformation for Mardia’s multivariate kurtosis. Communications in Statistics: Simulation and Computation, 49(3):684–698, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Guide:2022:CSR

[EJKL22] M. El Guide, K. Jbilou, C. Koukouvinos, and A. Lappa. Comparative study of L_{1} regularized logistic regression methods for variable selection. Communications in Statistics: Simulation and Computation, 51(9):4957–4972, 2022. CODEN CSSCDB. ISSN 0361-0918.

Guide:2023:KSS

[EJKL23] M. El Guide, K. Jbilou, C. Koukouvinos, and A. Lappa. Krylov subspace solvers for ℓ_{1} regularized logistic regression method. Communications in Statistics: Simulation and Computation, 52(6):2738–2751, 2023. CODEN CSSCDB. ISSN 0361-0918.

ElFarme:2024:OSL

[El 24] Nabil El Farme. Optimal skip-lot sampling plan based on the yield index for auto-correlated data. Communications in Statistics: Simulation and Computation, 53(2):814–828, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ermakov:2021:DMQ

[EL21] Sergej M. Ermakov and Svetlana N. Leora. Decrease of the mean of the quasi-random integration error. Communications in Statistics: Simulation and Computation, 50(11): 3581–3589, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:MIS

[eLhS23] Cheng en Li and Jian hua Shi. MCMC interweaving strategy for estimating stochastic volatility model and its application. Communications in Statistics: Simulation and Computation, 52(2):557–568, 2023. CODEN CSSCDB. ISSN 0361-0918.

Elsawah:2024:NLC

[Els24] A. M. Elsawah. A novel low complexity fast algorithm for effectively designing optimal mixed-level experiments. Communications in Statistics: Simulation and Computation, 53 (1):315–343, 2024. CODEN CSSCDB. ISSN 0361-0918.

Elkimakh:2022:HMM

[EN22] Karima Elkimakh and Abdelaziz Nasroallah. Hidden Markov model steady-state estimation. Communications in Statistics: Simulation and Computation, 51(11):6792–6807, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ershadi:2023:MDQ

[ENA^{+}23] Mohammad Javad Ershadi, Seyed Taghi Akhavan Niaki, Amir Azizi, Aynaz Ashtarian Esfahani, and Reza Edris Abadi. Monitoring data quality using Hoteling T^{2} multivariate control chart. Communications in Statistics: Simulation and Computation, 52(4):1591–1606, 2023. CODEN CSSCDB. ISSN 0361-0918.

Esfanjani:2022:TPR

[ENK^{+}22] Robab Mehdizadeh Esfanjani, Dariush Najarzadeh, Hossein Jabbari Khamnei, Farshin Hormozinejad, and Mahnaz Talebi. Two-parameter ridge estimation in seemingly unrelated regression models. Communications in Statistics: Simulation and Computation, 51(9):4904–4918, 2022. CODEN CSSCDB. ISSN 0361-0918.

Erem:2020:BTS

[Ere20] Aysegul Erem. Bivariate two sample test based on exceedance statistics. Communications in Statistics: Simulation and Computation, 49(9):2389–2401, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ermakov:2022:GGO

[ES22] Sergej M. Ermakov and Dmitriy N. Semenchikov. Genetic global optimization algorithms. Communications in Statistics: Simulation and Computation, 51(4):1503–1512, 2022. CODEN CSSCDB. ISSN 0361-0918.

ElShekhAhmed:2020:AWN

[ESES20] Hazem I. El Shekh Ahmed, Raid B. Salha, and Hossam O. El-Sayed. Adaptive weighted Nadaraya–Watson estimation of the conditional quantiles by varying bandwidth. Communications in Statistics: Simulation and Computation, 49(5): 1105–1117, 2020. CODEN CSSCDB. ISSN 0361-0918.

Eskandari:2021:AES

[Esk21] Farzad Eskandari. Asymptotic efficient semiparametric empirical Bayes estimation of multinomial responses. Communications in Statistics: Simulation and Computation, 50(1): 123–141, 2021. CODEN CSSCDB. ISSN 0361-0918.

ElHaj:2022:EBS

[ESLK22] Abir El Haj, Yousri Slaoui, Pierre-Yves Louis, and Zaher Khraibani. Estimation in a binomial stochastic blockmodel for a weighted graph by a variational expectation maximization algorithm. Communications in Statistics: Simulation and Computation, 51(8):4450–4469, 2022. CODEN CSSCDB. ISSN 0361-0918.

Eappen:2022:RRT

[ESS22] Christin Variathu Eappen, Stephen A. Sedory, and Sarjinder Singh. Ratio and regression type estimators of a new measure of coefficient of dispersion. Communications in Statistics: Simulation and Computation, 51(4):1899–1920, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Fang:2022:NIF

[Fan22] Jianglin Fang. Nonparametric independence feature screening for ultrahigh-dimensional missing data. Communications in Statistics: Simulation and Computation, 51(10):5670–5689, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fotuhi:2023:PIM

[FAT23] Hatef Fotuhi, Amirhossein Amiri, and Ali Reza Taheriyoun. Phase II monitoring of autocorrelated attributed social networks based on generalized estimating equations. Communications in Statistics: Simulation and Computation, 52(4): 1504–1522, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ferreira:2022:LMM

[FBL22] Clécio S. Ferreira, Heleno Bolfarine, and Víctor H. Lachos. Linear mixed models based on skew scale mixtures of normal distributions. Communications in Statistics: Simulation and Computation, 51(12):7194–7214, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fischer:2021:PRR

[FBNRG21] Daniel Fischer, Alain Berro, Klaus Nordhausen, and Anne Ruiz-Gazen. REPPlab: an R package for detecting clusters and outliers using exploratory projection pursuit. Communications in Statistics: Simulation and Computation, 50(11): 3397–3419, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fan:2023:DAL

[FF23] Ye Fan and Suning Fan. Distributed adaptive lasso penalized generalized linear models for big data. Communications in Statistics: Simulation and Computation, 52(4):1679–1698, 2023. CODEN CSSCDB. ISSN 0361-0918.

Faisal:2022:BIW

[FH22] Shahla Faisal and Christian Heumann. Bootstrap inference for weighted nearest neighbors imputation. Communications in Statistics: Simulation and Computation, 51(6):2842–2857, 2022. CODEN CSSCDB. ISSN 0361-0918.

Faltys:2022:SAE

[FHM22] Ondrej Faltys, Tomás Hobza, and Domingo Morales. Small area estimation under area-level generalized linear mixed models. Communications in Statistics: Simulation and Computation, 51(12):7404–7426, 2022. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2023:RMA

[Fig23] Adelaide Figueiredo. Resampling methods in ANOVA for data from the von Mises–Fisher distribution. Communications in Statistics: Simulation and Computation, 52(10): 4999–5013, 2023. CODEN CSSCDB. ISSN 0361-0918.

Fangyuan:2024:MEM

[FJ24] Kang Fangyuan and Guo Jiang. A mixed effect model for clustered recurrent event data. Communications in Statistics: Simulation and Computation, 53(1):190–202, 2024. CODEN CSSCDB. ISSN 0361-0918.

Fallah:2022:SIG

[FK22] Afshin Fallah and Ramin Kazemi. Statistical inference for the generalized weighted exponential distribution. Communications in Statistics: Simulation and Computation, 51(8): 4296–4313, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fu:2023:VRB

[FL23] Ke-Ang Fu and Jie Li. On the validity of the residual-based bootstrap for the unit root test statistic with long memory observations. Communications in Statistics: Simulation and Computation, 52(2):309–319, 2023. CODEN CSSCDB. ISSN 0361-0918.

Firinguetti-Limone:2020:BES

[FLPB20] Luis Firinguetti-Limone and Manuel Pereira-Barahona. Bayesian estimation of the shrinkage parameter in ridge regression. Communications in Statistics: Simulation and Computation, 49(12):3314–3327, 2020. CODEN CSSCDB. ISSN 0361-0918.

Feng:2022:ETS

[FLXY22] Kaixuan Feng, Zhenzhou Lu, Sinan Xiao, and WanYing Yun. An efficient trajectory sampling design method for elementary effect based global sensitivity analysis. Communications in Statistics: Simulation and Computation, 51(12):6989–7004, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fauzi:2023:BFK

[FM23] Rizky Reza Fauzi and Yoshihiko Maesono. Boundary-free kernel-smoothed goodness-of-fit tests for data on general interval. Communications in Statistics: Simulation and Computation, 52(5):1962–1978, 2023. CODEN CSSCDB. ISSN 0361-0918.

Farokhnia:2020:PCA

[FN20] Moez Farokhnia and S. T. A. Niaki. Principal component analysis-based control charts using support vector machines for multivariate non-normal distributions. Communications in Statistics: Simulation and Computation, 49(7):1815–1838, 2020. CODEN CSSCDB. ISSN 0361-0918.

Flores:2022:PSH

[FO22] Pablo Flores and Jordi Ocaña. Pretesting strategies for homoscedasticity when comparing means. their robustness facing non-normality. Communications in Statistics: Simulation and Computation, 51(1):280–292, 2022. CODEN CSSCDB. ISSN 0361-0918.

Filipiak:2020:PAL

[FP20] Katarzyna Filipiak and Simo Puntanen. Preface — advances on linear models and inference: computational aspects. Communications in Statistics: Simulation and Computation, 49 (3):565–566, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ferrer-Perez:2021:SUR

[FPAA21] Hugo Ferrer-Pérez, María-Isabel Ayuda, and Antonio Aznar. The sensitivity of unit root tests to the initial condition and to the lag length selection: a Monte Carlo simulation study. Communications in Statistics: Simulation and Computation, 50(4):1062–1072, 2021. CODEN CSSCDB. ISSN 0361-0918.

Farghali:2023:GTP

[FQKA23] Rasha A. Farghali, Muhammad Qasim, B. M. Golam Kibria, and Mohamed R. Abonazel. Generalized two-parameter estimators in the multinomial logit regression model: methods, simulation and application. Communications in Statistics: Simulation and Computation, 52(7):3327–3342, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Fernando:2022:DGF

[FS22] Gayara Fernando and Roshini Sooriyarachchi. The development of a goodness-of-fit test for high level binary multilevel models. Communications in Statistics: Simulation and Computation, 51(5):2710–2730, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fernando:2023:PLR

[FS23] S. I. R. Fernando and M. R. Sooriyarachchi. The performance of the log rank test for different relationships between the hazards: the case of the triangular test in comparison with some other sequential tests. Communications in Statistics: Simulation and Computation, 52(11):5498–5521, 2023. CODEN CSSCDB. ISSN 0361-0918.

Fuentes-Santos:2023:TSB

[FSGMM23] I. Fuentes-Santos, W. González-Manteiga, and J. Mateu. Testing similarity between first-order intensities of spatial point processes. A comparative study. Communications in Statistics: Simulation and Computation, 52(5):2210–2230, 2023. CODEN CSSCDB. ISSN 0361-0918.

Su:2021:DFD

[fSLC21] Zhi fang Su, Ken-Yu Lin, and Mei-Yuan Chen. Dual-frailty default intensity model: estimations and an application. Communications in Statistics: Simulation and Computation, 50(5):1315–1332, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fang:2021:SDE

[FTC21] Yongfei Fang, Yubin Tian, and Xia Cai. Sequential designs for estimating the operating window. Communications in Statistics: Simulation and Computation, 50(8):2386–2398, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fukuda:2022:SAA

[Fuk22] Kosei Fukuda. Selecting from among 12 alternative distributions of financial data. Communications in Statistics: Simulation and Computation, 51(7):3943–3954, 2022. CODEN CSSCDB. ISSN 0361-0918.

Fan:2023:BMS

[FW23] Tsai-Hung Fan and Yi-Fu Wang. Bayesian model selection for structural equation models for myopia data. Communications in Statistics: Simulation and Computation, 52(11): 5680–5694, 2023. CODEN CSSCDB. ISSN 0361-0918.

Fan:2021:ACE

[FY21] Qingliang Fan and Wanlu Yu. Adaptive k-class estimation in high-dimensional linear models. Communications in Statistics: Simulation and Computation, 50(12):3885–3913, 2021. CODEN CSSCDB. ISSN 0361-0918.

Farzammehr:2021:SNG

[FZM21] Mohadeseh Alsadat Farzammehr, Mohammad Reza Zadkarami, and Geoffrey J. McLachlan. Skew-normal generalized spatial panel data model. Communications in Statistics: Simulation and Computation, 50(11):3286–3314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Feng:2024:ADE

[FZY24] Zhenghui Feng, Jun Zhang, and Baojun Yang. Average derivation estimation with multiplicative distortion measurement errors. Communications in Statistics: Simulation and Computation, 53(1):30–61, 2024. CODEN CSSCDB. ISSN 0361-0918.

Feng:2020:SRC

[FZZ20] Yan Feng, Xiaobing Zhao, and Xian Zhou. Semiparametric random censorship models for survival data with long-term survivors. Communications in Statistics: Simulation and Computation, 49(11):2876–2896, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghitany:2022:GLD

[GABG22] Mohamed E. Ghitany, Suja M. Aboukhamseen, Abrar A. Baqer, and Ramesh C. Gupta. Gompertz–Lindley distribution and associated inference. Communications in Statistics: Simulation and Computation, 51(5):2599–2618, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ghaemmaghami:2023:CAB

[GABJ23] Parvin Ghaemmaghami, S. Mohammad Taghi Ayatollahi, Zahra Bagheri, and S. Reza Jafarzadeh. Covariate-adjusted Bayesian estimation of the performance of a continuous diagnostic test with a limit of detection in the absence of a reference standard: a simulation study. Communications in Statistics: Simulation and Computation, 52(4):1331–1343, 2023. CODEN CSSCDB. ISSN 0361-0918.

Gaigall:2021:NAM

[Gai21] Daniel Gaigall. On a new approach to the multi-sample goodness-of-fit problem. Communications in Statistics: Simulation and Computation, 50(10):2971–2989, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gana:2022:RRL

[Gan22] Rajaram Gana. Ridge regression and the lasso: how do they do as finders of significant regressors and their multipliers? Communications in Statistics: Simulation and Computation, 51(10):5738–5772, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gao:2023:BAM

[Gao23] Ping Gao. Bayesian analysis of a multivariate spatial ordered probit model. Communications in Statistics: Simulation and Computation, 52(4):1300–1317, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Ghaly:2020:ACA

[GAS20] A. A. Abdel Ghaly, H. M. Aly, and R. N. Salah. Applying the copula approach on step stress accelerated life test under type II censoring. Communications in Statistics: Simulation and Computation, 49(1):159–177, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gangsei:2022:LRB

[GAS22] Lars Erik Gangsei, Trygve Almøy, and Solve Sæbø. Linear regression with bivariate response variable containing missing data. strategies to increase prediction precision. Communications in Statistics: Simulation and Computation, 51(2): 527–538, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghapani:2022:TPW

[GB22] F. Ghapani and B. Babadi. Two parameter weighted mixed estimator in linear measurement error models. Communications in Statistics: Simulation and Computation, 51(12): 6936–6946, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Barrios:2020:DAC

[GBHA20] José María González-Barrios and Ricardo Hoyos-Argüelles. Distributions associated to the counting techniques of the dsample copula of order m and weak convergence of the sample process. Communications in Statistics: Simulation and Computation, 49(10):2505–2532, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Barrios:2021:ECA

[GBHA21] José M. González-Barrios and Ricardo Hoyos-Argüelles. Estimating checkerboard approximations with sample d-copulas. Communications in Statistics: Simulation and Computation, 50(12):3992–4027, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2020:CHV

[GCF^{+}20] M. Ivette Gomes, Frederico Caeiro, Fernanda Figueiredo, Lgia Henriques-Rodrigues, and Dinis Pestana. Corrected-Hill versus partially reduced-bias value-at-risk estimation. Communications in Statistics: Simulation and Computation, 49 (4):867–885, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gao:2022:ADC

[GDWY22] Min Gao, Saisai Ding, Shipeng Wu, and Wenzhi Yang. The asymptotic distribution of CUSUM estimator based on αmixing sequences. Communications in Statistics: Simulation and Computation, 51(10):6101–6113, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Gogebakan:2023:BUR

[GE23] Kemal Çaglar Gögebakan and Burak Alparslan Eroglu. Bounded unit root processes with non-stationary volatility. Communications in Statistics: Simulation and Computation, 52(4):1245–1263, 2023. CODEN CSSCDB. ISSN 0361-0918.

Gera:2020:ATP

[Ger20] Amos E. Gera. Acceptance testing for a planar array. Communications in Statistics: Simulation and Computation, 49 (5):1198–1209, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gera:2021:RP

[Ger21] Amos E. Gera. From runs to patterns. Communications in Statistics: Simulation and Computation, 50(12):4300–4314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Guris:2022:GDN

[GG22] Selahattin Güris and Burak Güris. GLS detrending in nonlinear unit root test. Communications in Statistics: Simulation and Computation, 51(3):1096–1102, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghasemzadeh:2020:BQR

[GGB20] Siamak Ghasemzadeh, Mojtaba Ganjali, and Taban Baghfalaki. Bayesian quantile regression for joint modeling of longitudinal mixed ordinal and continuous data. Communications in Statistics: Simulation and Computation, 49(2):375– 395, 2020. CODEN CSSCDB. ISSN 0361-0918.

Giancristofaro:2022:RAC

[GGMM22] R. Arboretti Giancristofaro, M. Gastaldi, L. Martinello, and C. Meneguzzer. Regression analysis with compositional data using orthogonal log-ratio coordinates. Communications in Statistics: Simulation and Computation, 51(4):1932–1945, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gomez:2023:CUP

[GGS23] Camilo Gomez, Alexander V. Goponenko, and Julia N. Soulakova. Constructing UpSet plot for survey data with weights using SAS and R software. Communications in Statistics: Simulation and Computation, 52(6):2320–2326, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ghapani:2022:SRL

[Gha22] Fatemeh Ghapani. Stochastic restricted Liu estimator in linear mixed measurement error models. Communications in Statistics: Simulation and Computation, 51(3):1220–1233, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2020:NUE

[Gho20] Sucharita Ghosh. A note on using the empirical moment generating function to estimate the variance of nonparametric trend estimates from independent time series replicates. Communications in Statistics: Simulation and Computation, 49 (9):2287–2301, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2023:HTM

[Gho23] Indranil Ghosh. Hidden truncation in multivariate Pareto (II) data: properties and inference. Communications in Statistics: Simulation and Computation, 52(5):2093–2114, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Gao:2020:ESS

[GHW20] Zhongqin Gao, Jingmin He, and Bingbing Wang. Exact solutions of some exit times for the diffusion risk model with liquid reserves, credit and debit interest. Communications in Statistics: Simulation and Computation, 49(10):2693–2703, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gad:2020:ALR

[GI20] Ahmed M. Gad and Wafaa I. M. Ibrahim. An adaptive linear regression approach for modeling heavy-tailed longitudinal data. Communications in Statistics: Simulation and Computation, 49(5):1181–1197, 2020. CODEN CSSCDB. ISSN 0361-0918.

Goudarzi:2021:NBO

[GJK21] Manizheh Goudarzi, Habib Jafari, and Soleiman Khazaei. Nonparametric Bayesian optimal designs for exponential regression model. Communications in Statistics: Simulation and Computation, 50(6):1809–1819, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gao:2023:MLA

[GJQ23] Jia-Xing Gao, Da-Quan Jiang, and Min-Ping Qian. Multivariate locally adaptive kernel density estimation. Communications in Statistics: Simulation and Computation, 52(9): 4431–4444, 2023. CODEN CSSCDB. ISSN 0361-0918.

Goyal:2020:NMS

[GK20] Manish Goyal and Narinder Kumar. Nonparametric multiple sample scale testing using U-statistics. Communications in Statistics: Simulation and Computation, 49(11):3019–3027, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2022:SBS

[GK22a] M. P. Gadre and V. C. Kakade. Sample based side sensitive group runs control chart to detect shifts in the process mean. Communications in Statistics: Simulation and Computation, 51(1):190–200, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2022:SSS

[GK22b] Mukund P. Gadre and Vikas C. Kakade. Some side sensitive group runs based control charts to detect shifts in the process median. Communications in Statistics: Simulation and Computation, 51(4):1474–1487, 2022. CODEN CSSCDB. ISSN 0361-0918.

Goel:2023:LBI

[GK23] Rajni Goel and Hare Krishna. Likelihood and Bayesian inference for k Lindley populations under joint type-II censoring scheme. Communications in Statistics: Simulation and Computation, 52(8):3475–3490, 2023. CODEN CSSCDB. ISSN 0361-0918.

Goudarzi:2023:ODL

[GKJ23] Manizheh Goudarzi, Soleiman Khazaei, and Habib Jafari. Doptimal designs for linear mixed model with random effects of Dirichlet process. Communications in Statistics: Simulation and Computation, 52(11):5371–5380, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ghasami:2020:APG

[GKM20] Safdar Ghasami, Zahra Khodadadi, and Mohsen Maleki. Autoregressive processes with generalized hyperbolic innovations. Communications in Statistics: Simulation and Computation, 49(12):3080–3092, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gill:2024:THS

[GKS24] Amar Nath Gill, Jatesh Kumar, and Parminder Singh. Testing homogeneity of several exponential populations using censored data. Communications in Statistics: Simulation and Computation, 53(1):344–356, 2024. CODEN CSSCDB. ISSN 0361-0918.

Guo:2022:DLS

[GKZ^{+}22] Zhaoyu Guo, Irwin Khuu, Kevin Zhu, Jeffrey S. Rosenthal, and Frederic P. Schoenberg. Distinguishing luck from skill through statistical simulation: a case study. Communications in Statistics: Simulation and Computation, 51(5):2537–2559, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gan:2020:CCE

[GLT20] Shengjun Gan, Changli Lu, and Yongge Tian. Computation and comparison of estimators under different linear randomeffects models. Communications in Statistics: Simulation and Computation, 49(5):1210–1222, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2022:EDG

[GM22] Shyamal Ghosh and Murari Mitra. On the exact distribution of generalized Hollander–Proschan type statistics. Communications in Statistics: Simulation and Computation, 51(9): 5051–5067, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghebremichael:2024:CBL

[GM24a] Musie Ghebremichael and Haben Michael. Comparison of the binormal and Lehman receiver operating characteristic curves. Communications in Statistics: Simulation and Computation, 53(2):772–785, 2024. CODEN CSSCDB. ISSN 0361-

Godase:2024:CSC

[GM24b] Dadasaheb G. Godase and Shashibhushan B. Mahadik. The combined Shewhart–CUSUM sign charts. Communications in Statistics: Simulation and Computation, 53(1):357–366, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2023:BGD

[GMC23] Indranil Ghosh, Filipe Marques, and Subrata Chakraborty. A bivariate geometric distribution via conditional specification: properties and applications. Communications in Statistics: Simulation and Computation, 52(12):5925–5945, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Gammoudi:2021:GQM

[GNE21] Imed Gammoudi, Asma Nani, and Mohamed El Ghourabi. Generalized quasi maximum likelihood estimation for generalized autoregressive score models: simulations and real applications. Communications in Statistics: Simulation and Computation, 50(11):3338–3363, 2021. CODEN CSSCDB. ISSN 0361-0918.

Garthoff:2022:SPS

[GO22] Robert Garthoff and Philipp Otto. Spatiotemporal procedures for the statistical surveillance of spatial autoregressive models with heavy tails. Communications in Statistics: Simulation and Computation, 51(10):5709–5737, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gokpinar:2022:SLR

[Gök22] Esra Gökpinar. Standardized likelihood ratio test for homogeneity of variance of several normal populations. Communications in Statistics: Simulation and Computation, 51(11): 6309–6319, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gorecki:2020:GPC

[GP20] Tomasz Górecki and Paweł Piasecki. Generalization of the Procrustes coefficient to functional data. Communications in Statistics: Simulation and Computation, 49(3):808–816, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gegout-Petit:2022:RKM

[GPGMK22] Anne Gégout-Petit, Aurélie Gueudin-Muller, and Clémence Karmann. The revisited knockoffs method for variable selection in L_{1}-penalized regressions. Communications in Statistics: Simulation and Computation, 51(10):5582–5595, 2022. CODEN CSSCDB. ISSN 0361-0918.

Georghiou:2021:MNB

[GPP21] Costas Georghiou, Andreas N. Philippou, and Zaharias M. Psillakis. On the modes of the negative binomial distribution of order k, type I. Communications in Statistics: Simulation and Computation, 50(4):1217–1230, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gioia:2023:MBR

[GPS23] V. Gioia, E. C. Kenne Pagui, and A. Salvan. Median bias reduction in cumulative link models. Communications in Statistics: Simulation and Computation, 52(3):795–802, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Giraudo:2024:CMN

[GRR24] Maria Teresa Giraudo, Fulvio Ricceri, and Elena Rosso. Comment on Moriña and Navarro (2017). Communications in Statistics: Simulation and Computation, 53(2):1082–1087, 2024. CODEN CSSCDB. ISSN 0361-0918. See [MN17].

Grenier:2023:DNN

[GS23] Isabelle Grenier and Bruno Sansó. Distributed nearestneighbor Gaussian processes. Communications in Statistics: Simulation and Computation, 52(7):2886–2898, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Guven:2022:SWT

[GSG22] Gamze Güven, Hatice Samkar, and Fikri Gökpinar. Score and Wald tests for the homogeneity of inverse Gaussian scale parameters based on computational approach test. Communications in Statistics: Simulation and Computation, 51(11): 6649–6666, 2022. CODEN CSSCDB. ISSN 0361-0918.

Grzesiek:2022:CCB

[GTW22] Aleksandra Grzesiek, Marek Teuerle, and Agnieszka Wyłomańska. Cross-codifference for bidimensional VAR(1) time series with infinite variance. Communications in Statistics: Simulation and Computation, 51(3):1355–1380, 2022. CODEN CSSCDB. ISSN 0361-0918.

Guler:2022:RBB

[Gül22] Nesrin Güler. On relations between BLUPs under two transformed linear random-effects models. Communications in Statistics: Simulation and Computation, 51(9):5099–5125, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gu:2020:SCC

[GW20] Xiaolan Gu and Qiusheng Wang. Sparse canonical correlation analysis algorithm with alternating direction method of multipliers. Communications in Statistics: Simulation and Computation, 49(9):2372–2388, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghitany:2022:NPA

[GW22] M. E. Ghitany and Shaochen Wang. A note on parameter asymptotics for weighted Lindley distribution. Communications in Statistics: Simulation and Computation, 51(4):1513– 1524, 2022. CODEN CSSCDB. ISSN 0361-0918.

Guo:2023:EET

[GWW^{+}23] Baocai Guo, Wei Wang, Zhichao Wu, Xixiang He, and Yumei Zhu. Exact equal-tailed β-expectation tolerance intervals for sample variances. Communications in Statistics: Simulation and Computation, 52(11):5309–5326, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Guo:2023:OPU

[GWZ23] Zhidong Guo, Xianhong Wang, and Yunliang Zhang. Option pricing under time interval driven model. Communications in Statistics: Simulation and Computation, 52(4):1538–1545, 2023. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:MSM

[GWZC22] Yujie Gai, Yusheng Wei, Jun Zhang, and Aixian Chen. Measuring the symmetry of model errors for varying coefficient regression models based on correlation coefficient. Communications in Statistics: Simulation and Computation, 51(5): 2235–2251, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gao:2023:JEH

[GY23] Yimiao Gao and Yuehan Yang. A joint estimation for the high-dimensional regression modeling on stratified data.

Communications in Statistics: Simulation and Computation, 52(12):6129–6140, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ghalani:2021:ICS

[GZ21a] Mohammad Reza Ghalani and Mohammad Reza Zadkarami. Investigation of covariance structures in modelling longitudinal ordinal responses with skew normal random effect. Communications in Statistics: Simulation and Computation, 50 (1):254–269, 2021. CODEN CSSCDB. ISSN 0361-0918.

Guo:2021:SLT

[GZ21b] Xiao Guo and Hai Zhang. Structured learning of time-varying networks with application to PM_{2.5} data. Communications in Statistics: Simulation and Computation, 50(5):1364–1382, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:DSM

[GZ22a] Yujie Gai and Jun Zhang. Detection of the symmetry of model errors for partial linear single-index models. Communications in Statistics: Simulation and Computation, 51(6):3410–3427, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:DSA

[GZ22b] Yujie Gai and Jun Zhang. Detection the symmetry or asymmetry of model errors in partial linear models. Communications in Statistics: Simulation and Computation, 51(5):2217– 2234, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:SIR

[GZ22c] Yujie Gai and Jun Zhang. Single-index relative error regression models. Communications in Statistics: Simulation and Computation, 51(5):2252–2267, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gao:2022:IMN

[GZY22] Siyu Gao, Qianqian Zhang, and Jing Yu. The impact of misspecification of nuisance parameters on test for homogeneity in zero-inflated Poisson model: a simulation study. Communications in Statistics: Simulation and Computation, 51(1): 84–98, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:TSM

[GZZ22] Yujie Gai, Xuehu Zhu, and Jun Zhang. Testing symmetry of model errors for nonparametric regression models by using correlation coefficient^{1}. Communications in Statistics: Simulation and Computation, 51(4):1436–1453, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2024:PHT

[GZZ24] Yujie Gai, Jun Zhang, and Yue Zhou. Parametric hypothesis tests for exponentiality under multiplicative distortion measurement errors data. Communications in Statistics: Simulation and Computation, 53(3):1594–1617, 2024. CODEN CSSCDB. ISSN 0361-0918.

Hassan:2024:PIM

[HA24] Anwaar M. Hassan and Aya A. Aly. Phase II monitoring of zero inflated Poisson regression profiles. Communications in Statistics: Simulation and Computation, 53(3):1519–1533, 2024. CODEN CSSCDB. ISSN 0361-0918.

Hayat:2023:RMO

[HAA^{+}23] Humera Hayat, Atif Akbar, Tanvir Ahmad, Sajjad Haider Bhatti, and Muhammad Imdad Ullah. Robustness to missing observation and optimalities of response surface designs with regular and complex structure. Communications in Statistics: Simulation and Computation, 52(11):5213–5230, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hamdy:2022:DFT

[Ham22] Mohamed Ismail Hamdy. Distribution-free three-sample test for detecting trend in the pure location model. Communications in Statistics: Simulation and Computation, 51(7):4127– 4145, 2022. CODEN CSSCDB. ISSN 0361-0918.

Han:2020:BER

[Han20] Ming Han. E-Bayesian estimations of the reliability and its eposterior risk under different loss functions. Communications in Statistics: Simulation and Computation, 49(6):1527–1545, 2020. CODEN CSSCDB. ISSN 0361-0918.

Han:2021:BEP

[Han21] Ming Han. E-Bayesian estimations of parameter and its evaluation standard: E-MSE (expected mean square error) under different loss functions. Communications in Statistics: Simulation and Computation, 50(7):1971–1988, 2021. CODEN CSSCDB. ISSN 0361-0918.

Han:2022:SED

[Han22] Ming Han. Study on the effect of the different prior distributions on e-Bayesian estimation of failure probability and its eposterior risk. Communications in Statistics: Simulation and Computation, 51(11):6441–6455, 2022. CODEN CSSCDB. ISSN 0361-0918.

Harwell:2021:ESV

[Har21] Michael Harwell. An effect size for variance heterogeneity in meta-analysis. Communications in Statistics: Simulation and Computation, 50(7):1955–1970, 2021. CODEN CSSCDB. ISSN 0361-0918.

Han:2022:RAU

[HAT^{+}22] Seungbong Han, Adin-Cristian Andrei, Kam-Wah Tsui, Sung-Cheol Yun, and Jong Ho Yoon. ROC analysis using covariate balancing propensity scores with an application to biochemical predictors for thyroid cancer. Communications in Statistics: Simulation and Computation, 51(2):374–390, 2022. CODEN CSSCDB. ISSN 0361-0918.

Han:2022:OPP

[HB22a] David Han and Tianyu Bai. Optimal planning of progressively Type-I censored step-stress accelerated life test under interval inspection. Communications in Statistics: Simulation and Computation, 51(12):7565–7586, 2022. CODEN CSSCDB. ISSN 0361-0918.

Haq:2022:DCC

[HB22b] Abdul Haq and Lubna Bibi. The dual CUSUM charts with auxiliary information for process mean. Communications in Statistics: Simulation and Computation, 51(1):164–189, 2022. CODEN CSSCDB. ISSN 0361-0918.

Harfouche:2023:SPA

[HB23] Zineb Harfouche and Aicha Bareche. Semi-parametric approach for approximating the ruin probability of classical risk models with large claims. Communications in Statistics: Simulation and Computation, 52(11):5585–5604, 2023. CODEN CSSCDB. ISSN 0361-0918.

Holling:2020:ENV

[HBM^{+}20] Heinz Holling, Walailuck Böhning, Ehsan Masoudi, Dankmar Böhning, and Patarawan Sangnawakij. Evaluation of a new version of I^{2} with emphasis on diagnostic problems. Communications in Statistics: Simulation and Computation, 49(4): 942–972, 2020. CODEN CSSCDB. ISSN 0361-0918.

Haq:2022:NAM

[HBS22] Abdul Haq, Mehwish Bibi, and Burhan Ali Shah. A novel approach to monitor simple linear profiles using individual observations. Communications in Statistics: Simulation and Computation, 51(11):6269–6282, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2023:RTP

[HCHK23] Abid Hussain, Salman A. Cheema, Summaira Haroon, and Tanveer Kifayat. The ruin time for 3-player gambler’s problem: an approximate formula. Communications in Statistics: Simulation and Computation, 52(4):1651–1659, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Hu:2023:BDE

[HCLF23] Jie Hu, Zirui Chen, Wenwei Lin, and Xiaodan Fan. Bayesian detection of event spreading pattern from multivariate binary time series. Communications in Statistics: Simulation and Computation, 52(4):1225–1244, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hu:2021:GCP

[HCTZ21] XueLong Hu, Philippe Castagliola, Anan Tang, and Jianlan Zhong. Guaranteed conditional performance of the median run length based EWMA X̄ chart with unknown process parameters. Communications in Statistics: Simulation and Computation, 50(12):4280–4299, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hu:2021:FIC

[HCZ21] Guozhi Hu, Weihu Cheng, and Jie Zeng. Focused information criterion and model averaging for varying-coefficient partially linear models with longitudinal data. Communications in Statistics: Simulation and Computation, 50(8):2399–2417, 2021. CODEN CSSCDB. ISSN 0361-0918.

Han:2020:MLE

[HD20] Zifei Han and Victor De Oliveira. Maximum likelihood estimation of Gaussian copula models for geostatistical count data. Communications in Statistics: Simulation and Computation, 49(8):1957–1981, 2020. CODEN CSSCDB. ISSN 0361-0918.

He:2022:MAP

[HD22] Baihua He and Fangli Dong. Model averaging with privacypreserving. Communications in Statistics: Simulation and Computation, 51(4):1401–1414, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2021:MDP

[HE21] Xin-Wei Huang and Takeshi Emura. Model diagnostic procedures for copula-based Markov chain models for statistical process control. Communications in Statistics: Simulation and Computation, 50(8):2345–2367, 2021. CODEN CSSCDB. ISSN 0361-0918.

Huang:2024:CMC

[HE24] Xin-Wei Huang and Takeshi Emura. Computational methods for a copula-based Markov chain model with a binomial time series. Communications in Statistics: Simulation and Computation, 53(4):1973–1990, 2024. CODEN CSSCDB. ISSN 0361-0918.

Haq:2022:NDE

[HEK22] Abdul Haq, Sana Ejaz, and Michael B. C. Khoo. A new double EWMA-t chart for process mean. Communications in Statistics: Simulation and Computation, 51(11):6556–6571, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hepsag:2021:TCN

[Hep21a] Aycan Hepsag. Testing for cointegration in nonlinear asymmetric smooth transition error correction models. Communications in Statistics: Simulation and Computation, 50(2): 400–412, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hepsag:2021:URT

[Hep21b] Aycan Hepsag. A unit root test based on smooth transitions and nonlinear adjustment. Communications in Statistics: Simulation and Computation, 50(3):625–632, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Hamdeni:2022:MOG

[HG22] Tasnime Hamdeni and Soufiane Gasmi. The Marshall–Olkin generalized defective Gompertz distribution for surviving fraction modeling. Communications in Statistics: Simulation and Computation, 51(11):6511–6524, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Hu:2022:WFF

[HGWZ22] Wenjuan Hu, Jiaxian Guo, Guochang Wang, and Baoxue Zhang. Weight fused functional sliced average variance estimation. Communications in Statistics: Simulation and Computation, 51(9):5000–5008, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hsieh:2023:ELR

[HHH23] Jin-Jian Hsieh, Tai-Yu Hsu, and Ke-Yu Hou. Empirical likelihood ratio tests for two sample comparison under current status data. Communications in Statistics: Simulation and Computation, 52(2):425–441, 2023. CODEN CSSCDB. ISSN 0361-0918.

Han:2023:SSD

[HHOC23] Dong Han, Yawen Hou, Chunquan Ou, and Zheng Chen. Sample size determination in superiority or non-inferiority clinical trials with time-to-event data under exponential, Weibull and Gompertz distributions. Communications in Statistics: Simulation and Computation, 52(4):1212–1224, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hamid:2022:GAR

[HHvR22] Jemila S. Hamid, Wei Liang Huang, and Dietrich von Rosen. Graphical analysis of residuals in multivariate growth curve models and applications in the analysis of longitudinal data. Communications in Statistics: Simulation and Computation, 51(10):5556–5581, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2021:REI

[HHY21] Mian Huang, Kang He, and Weixin Yao. Regression estimation via information-weighted composite models with different dimensions. Communications in Statistics: Simulation and Computation, 50(6):1613–1621, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hamaz:2022:OPS

[HI22] Abdelghani Hamaz and Mohamed Ibazizen. An optimal prediction in stationary random fields based on a new interpolation approach. Communications in Statistics: Simulation and Computation, 51(4):1746–1758, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hosseini:2020:ALI

[HK20] Fatemeh Hosseini and Omid Karimi. Approximate likelihood inference in spatial generalized linear mixed models with closed skew normal latent variables. Communications in Statistics: Simulation and Computation, 49(1):121–134, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hamdi:2022:PDC

[HK22] Fayçal Hamdi and Abderaouf Khalfi. Predictive density criterion for SETAR models. Communications in Statistics: Simulation and Computation, 51(2):443–459, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Huang:2022:NSM

[HLHC22] Xinghui Huang, Jingjing Lyu, Yawen Hou, and Zheng Chen. A nonparametric statistical method for two crossing survival curves. Communications in Statistics: Simulation and Computation, 51(9):5041–5050, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2023:VSI

[HLT23] Xifen Huang, Jiajuan Liang, and Guo-Liang Tian. Valid statistical inference methods for incomplete contingency table with three-category missing data. Communications in Statistics: Simulation and Computation, 52(11):5195–5212, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hanyuda:2021:PCU

[HM21] Bunto Hanyuda and Hidetoshi Murakami. Power comparisons of the unbiased Berk–Jones test and the unbiased reversed Berk–Jones test. Communications in Statistics: Simulation and Computation, 50(4):1009–1024, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2022:NAD

[HMRN22] Shahid Hussain, Tahir Mahmood, Muhammad Riaz, and Hafiz Zafar Nazir. A new approach to design median control charts for location monitoring. Communications in Statistics: Simulation and Computation, 51(7):3553–3577, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Haji-Maghsoudi:2023:MHS

[HMSRM23] Saiedeh Haji-Maghsoudi, Majid Sadeghifar, Ghodratollah Roshanaei, and Hossein Mahjub. Multivariate hidden semi-Markov models for longitudinal data: a dynamic regression modeling. Communications in Statistics: Simulation and Computation, 52(12):5830–5849, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hermans:2022:OWE

[HMV^{+}22] Lisa Hermans, Geert Molenberghs, Geert Verbeke, Michael G. Kenward, Pavlos Mamouris, and Bert Vaes. Optimal weighted estimation versus Cochran–Mantel–Haenszel. Communications in Statistics: Simulation and Computation, 51(7):3645– 3659, 2022. CODEN CSSCDB. ISSN 0361-0918.

Homburg:2020:CEA

[Hom20] Annika Homburg. Criteria for evaluating approximations of count distributions. Communications in Statistics: Simulation and Computation, 49(12):3152–3170, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hormozinejad:2022:OTA

[Hor22] Farshin Hormozinejad. Optimal tests for arbitrarily varying object and the familywise error rates. Communications in Statistics: Simulation and Computation, 51(3):993–1001, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2022:DNC

[HORA22] M. Pear Hossain, M. Hafidz Omar, Muhammad Riaz, and Sheikh Y. Arafat. On designing a new control chart for Rayleigh distributed processes with an application to monitor glass fiber strength. Communications in Statistics: Simulation and Computation, 51(6):3168–3184, 2022. CODEN CSSCDB. ISSN 0361-0918.

Han:2021:EJD

[HPZ^{+}21] Sung Won Han, Sunghoon Park, Hua Zhong, Eun-Seok Ryu, Pei Wang, Sehee Jung, Jayeon Lim, Jeewhan Yoon, and SungHwan Kim. Estimation of joint directed acyclic graphs with lasso family for gene networks. Communications in Statistics: Simulation and Computation, 50(9):2793–2807, 2021. CODEN CSSCDB. ISSN 0361-0918.

He:2022:WSC

[HQN22] Lanlan He, Hong Qin, and Jianhui Ning. Weighted symmetrized centered discrepancy for uniform design. Communications in Statistics: Simulation and Computation, 51(8): 4509–4519, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hu:2023:MOS

[HQZ^{+}23] XueLong Hu, YuLong Qiao, PanPan Zhou, JianLan Zhong, and Shu Wu. Modified one-sided EWMA charts for monitoring time between events. Communications in Statistics: Simulation and Computation, 52(3):1041–1056, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hasanalipour:2022:ISN

[HR22] P. Hasanalipour and M. Razmkhah. Inference on skew-normal distribution based on Fisher information in order statistics. Communications in Statistics: Simulation and Computation, 51(4):1525–1541, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hudaverdi:2023:WTI

[HS23] Burcu Hudaverdi and Selim Orhun Susam. On the weighted tests of independence based on Bernstein empirical copula. Communications in Statistics: Simulation and Computation, 52(2):404–424, 2023. CODEN CSSCDB. ISSN 0361-0918.

Haq:2020:EGC

[HU20] Abdul Haq and Zain Ul Abidin. An enhanced GWMA chart for process mean. Communications in Statistics: Simulation and Computation, 49(4):847–866, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hu:2023:ODT

[Hu23] Cheng-Hung Hu. Optimum degradation test plan under model mis-specification for Wiener and gamma processes. Communications in Statistics: Simulation and Computation, 52(5):1719–1732, 2023. CODEN CSSCDB. ISSN 0361-0918.

Huang:2020:CFR

[Hua20] Bai Huang. Combined fixed and random effects estimators. Communications in Statistics: Simulation and Computation, 49(8):1945–1956, 2020. CODEN CSSCDB. ISSN 0361-0918.

Huang:2022:EPE

[Hua22] Wei-Heng Huang. The effect of parameter estimation on X̄ control charts for the lognormal distribution. Communications in Statistics: Simulation and Computation, 51(11): 6715–6728, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2023:EDM

[Hua23] Ching Chun Huang. Economic design of max charts using Taguchi’s loss function. Communications in Statistics: Simulation and Computation, 52(12):6178–6192, 2023. CODEN CSSCDB. ISSN 0361-0918.

Haq:2023:EFP

[HUK23] Abdul Haq, Muhammad Usman, and Manzoor Khan. Estimation of finite population variance under stratified random sampling. Communications in Statistics: Simulation and Computation, 52(12):6193–6209, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hutson:2021:IEL

[Hut21] Alan D. Hutson. Inference for L-estimators of location using a bootstrap warping approach. Communications in Statistics: Simulation and Computation, 50(7):2145–2150, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Hutson:2022:SPB

[Hut22] Alan D. Hutson. A semi-parametric bootstrap-based best linear unbiased estimator of location under symmetry. Communications in Statistics: Simulation and Computation, 51(9): 5323–5332, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hutson:2024:GSQ

[Hut24] Alan D. Hutson. The generalized sigmoidal quantile function. Communications in Statistics: Simulation and Computation, 53(2):799–813, 2024. CODEN CSSCDB. ISSN 0361-0918.

Hartog:2021:NBL

[HvZ21] Jarno Hartog and Harry van Zanten. Nonparametric Bayesian label prediction on a large graph using truncated Laplacian regularization. Communications in Statistics: Simulation and Computation, 50(12):3821–3838, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hou:2024:LSM

[HW24] Zhiqiang Hou and Pengfei Wang. Large-scale multiple testing via multivariate hidden Markov models. Communications in Statistics: Simulation and Computation, 53(4):1932–1951, 2024. CODEN CSSCDB. ISSN 0361-0918.

Hwang:2023:CBT

[Hwa23] Eunju Hwang. A class of bootstrap tests on the tail index. Communications in Statistics: Simulation and Computation, 52(5):1786–1804, 2023. CODEN CSSCDB. ISSN 0361-0918.

Hung:2021:MSU

[HYSC21] Wen-Liang Hung, Jenn-Hwai Yang, I-Wen Song, and Yen-Chang Chang. A modified self-updating clustering algorithm for application to dengue gene expression data. Communications in Statistics: Simulation and Computation, 50(2):483– 500, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hu:2022:BCM

[HZ22] Jun Hu and Yan Zhuang. A broader class of modified twostage minimum risk point estimation procedures for a normal mean. Communications in Statistics: Simulation and Computation, 51(12):7587–7601, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2023:ATZ

[HZ23] Jie Huang and Fukang Zhu. An alternative test for zero modification in the INAR(1) model with Poisson innovations. Communications in Statistics: Simulation and Computation, 52(3):803–816, 2023. CODEN CSSCDB. ISSN 0361-0918.

Harfouche:2022:DMA

[HZAF22] L. Harfouche, N. Zougab, S. Adjabi, and B. Funke. Discrete multivariate associated kernel estimators using two multiplicative bias correction methods. Communications in Statistics: Simulation and Computation, 51(2):404–420, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Huang:2022:ESV

[HZY22] Zhensheng Huang, Riquan Zhang, and Wenwen Yin. Estimation of semi-varying coefficient error-in-variable models with surrogate data and validation sample. Communications in Statistics: Simulation and Computation, 51(8):4159–4185, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hu:2024:PDE

[HZZ^{+}24] XueLong Hu, SuYing Zhang, XiaoJian Zhou, AnAn Tang, and FuPeng Xie. The performance of double exponentially weighted moving average control charts for monitoring the coefficient of variation. Communications in Statistics: Simulation and Computation, 53(4):1779–1798, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ijaz:2021:EPL

[IAG21] Musarrat Ijaz, Zahid Asghar, and Asma Gul. Ensemble of penalized logistic models for classification of high-dimensional data. Communications in Statistics: Simulation and Computation, 50(7):2072–2088, 2021. CODEN CSSCDB. ISSN 0361-0918.

Islam:2021:SAE

[IC21] Sadikul Islam and Hukum Chandra. Small area estimation combining data from two surveys. Communications in Statistics: Simulation and Computation, 50(6):1728–1749, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ijaz:2024:FSM

[IGA24] Musarrat Ijaz, Asma Gul, and Zahid Asghar. A feature selection method for classification based on ensemble of penalized logistic models. Communications in Statistics: Simulation and Computation, 53(3):1285–1297, 2024. CODEN CSSCDB. ISSN 0361-0918.

Imon:2022:PDB

[IHS22] A. H. M. Rahmatullah Imon, Soroar Hossain, and Noora N. Saleh. On PRESS and deletion bootstraps in linear regression. Communications in Statistics: Simulation and Computation, 51(12):7319–7333, 2022. CODEN CSSCDB. ISSN 0361-0918.

Iachina:2020:SBC

[IM20] Maria Iachina and Peter Morling. A simulation-based comparison of two methods for determining the treatment effect in children diagnosed with hyperkinetic disorder. Communications in Statistics: Simulation and Computation, 49(6): 1385–1396, 2020. CODEN CSSCDB. ISSN 0361-0918.

Izaguirre:2021:HTE

[IMR21] Alejandro Izaguirre and Gabriel Montes-Rojas. Horvitz– Thompson estimator under partial information with an application to network degree distribution. Communications in Statistics: Simulation and Computation, 50(2):343–366, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ivkovic:2021:BCI

[IR21] Ivana Ivković and Vesna Rajić. Better confidence intervals for the population coefficient of variation. Communications in Statistics: Simulation and Computation, 50(12):4215–4262, 2021. CODEN CSSCDB. ISSN 0361-0918.

Iriany:2021:CGM

[IRA21] Atiek Iriany, Diana Rosyida, and Arifin Arifin. A comparison of GSTAR-SUR models and a hybrid GSTAR-SUR/neural network model on residuals of precipitation forecasting. Communications in Statistics: Simulation and Computation, 50 (9):2782–2792, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ivkovic:2020:CPC

[IRS20] Ivana Ivković, Vesna Rajić, and Jelena Stanojević. Coverage probabilities of confidence intervals for the slope parameter of linear regression model when the error term is not normally distributed. Communications in Statistics: Simulation and Computation, 49(1):147–158, 2020. CODEN CSSCDB. ISSN 0361-0918.

Inan:2020:PSO

[IS20] Deniz Inan and Nuriye Sancar. Particle swarm optimization based ridge logistic estimator. Communications in Statistics: Simulation and Computation, 49(3):669–683, 2020. CODEN CSSCDB. ISSN 0361-0918.

Iqbal:2021:BIM

[IT21] Farhat Iqbal and Kostas Triantafyllopoulos. Bayesian inference of multivariate rotated GARCH models with skew returns. Communications in Statistics: Simulation and Computation, 50(10):3105–3123, 2021. CODEN CSSCDB. ISSN 0361-0918.

Tsukada:2023:DAC

[iTS23] Shin ichi Tsukada and Takatoshi Sugiyama. Distribution approximation of covariance matrix eigenvalues. Communications in Statistics: Simulation and Computation, 52(9):4313– 4325, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jamshidi:2021:DNI

[JAP21] Babak Jamshidi, Sayed Mohammad Reza Alavi, and Gholam Ali Parham. The distribution of the number of the infected individuals in a stochastic SIR model on regular rooted trees. Communications in Statistics: Simulation and Computation, 50(5):1427–1444, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jackson:2023:UBA

[JBJ23] H. Jackson, S. Bowen, and T. Jaki. Using biomarkers to allocate patients in a response-adaptive clinical trial. Communications in Statistics: Simulation and Computation, 52(12): 5946–5965, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jana:2023:COM

[JD23a] Nabakumar Jana and Santanu Dey. Classification of observations into von Mises–Fisher populations with unknown parameters. Communications in Statistics: Simulation and Computation, 52(9):4392–4413, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jose:2023:MPQ

[JD23b] Joby K. Jose and V. Deepthi. M/PH/1 queueing model with re-servicing. Communications in Statistics: Simulation and Computation, 52(7):2865–2885, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jorge-Gonzalez:2021:SDT

[JGGDMRLD21] Elisa Jorge-González, Enrique González-Dávila, Raquel Martín-Rivero, and Domingo Lorenzo-Díaz. Stochastic and deterministic trend in state space models. Communications in Statistics: Simulation and Computation, 50(10):2809–2822, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2022:CQL

[JGL22] Rong Jiang, Meng-Fan Guo, and Xin Liu. Composite quasilikelihood for single-index models with massive datasets. Communications in Statistics: Simulation and Computation, 51(9):5024–5040, 2022. CODEN CSSCDB. ISSN 0361-0918.

Jeng:2023:CPM

[JH23] Shuen-Lin Jeng and Ming-Jie Hsu. Clustering partially masked images with modified Radon transformation. Communications in Statistics: Simulation and Computation, 52 (11):5231–5255, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2021:EMP

[JHR21] Wenjun Jiang, H. P. Hong, and Jiandong Ren. Estimation of model parameters of dependent processes constructed using Lévy Copulas. Communications in Statistics: Simulation and Computation, 50(3):691–707, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jhwueng:2024:CPQ

[Jhw24] Dwueng-Chwuan Jhwueng. On the covariance of phylogenetic quantitative trait evolution models and their matrix condition. Communications in Statistics: Simulation and Computation, 53(2):952–971, 2024. CODEN CSSCDB. ISSN 0361-0918.

Jayakumar:2020:NGP

[JKH20] K. Jayakumar, Bindu Krishnan, and G. G. Hamedani. On a new generalization of Pareto distribution and its applications. Communications in Statistics: Simulation and Computation, 49(5):1264–1284, 2020. CODEN CSSCDB. ISSN 0361-0918.

Jhong:2024:LPS

[JKLK24] Jae-Hwan Jhong, Jae-Young Kim, Jae-Deok Lee, and Ja-Yong Koo. Locally penalized single-index model using Bsplines and spherical coordinates. Communications in Statistics: Simulation and Computation, 53(1):273–287, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Jin:2024:NMC

[JL24] Jiayun Jin and Geert Loosveldt. Nonparametric multivariate control chart for numerical and categorical variables. Communications in Statistics: Simulation and Computation, 53 (1):457–475, 2024. CODEN CSSCDB. ISSN 0361-0918.

John:2020:CEB

[JM20] Mateusz John and Adam Mieldzioc. The comparison of the estimators of banded Toeplitz covariance structure under the high-dimensional multivariate model. Communications in Statistics: Simulation and Computation, 49(3):734–752, 2020. CODEN CSSCDB. ISSN 0361-0918.

Jodra:2023:NCG

[Jod23] P. Jodrá. A note on the computer generation of the binomial exponential distribution and generalizations. Communications in Statistics: Simulation and Computation, 52(1): 260–266, 2023. CODEN CSSCDB. ISSN 0361-0918.

Jin:2021:AMR

[JPH21] Man Jin, Adam Polis, and Jonathan Hartzel. Algorithms for minimization randomization and the implementation with an R package. Communications in Statistics: Simulation and Computation, 50(10):3077–3087, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jayakumar:2019:DLW

[JS19] K. Jayakumar and K. K. Sankaran. Discrete Linnik Weibull distribution. Communications in Statistics: Simulation and Computation, 48(10):3092–3117, 2019. CODEN CSSCDB. ISSN 0361-0918. See correction [Ano24].

Javaid:2020:NMH

[JuAH20] Amjad Javaid, Muhammad Noor ul Amin, and Muhammad Hanif. A new Max-HEWMA control chart using auxiliary information. Communications in Statistics: Simulation and Computation, 49(5):1285–1305, 2020. CODEN CSSCDB. ISSN 0361-0918.

Javaid:2023:PME

[JuAH23] Amjad Javaid, Muhammad Noor ul Amin, and Muhammad Hanif. Performance of Max-EWMA control chart for joint monitoring of mean and variance with measurement error. Communications in Statistics: Simulation and Computation, 52(1):1–26, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kunjunni:2022:MOD

[KA22] Sajana O. Kunjunni and Sajesh T. Abraham. Multidimensional outlier detection and robust estimation using S_{n} covariance. Communications in Statistics: Simulation and Computation, 51(7):3912–3922, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kucuk:2023:NHR

[KA23] Serenay Kucuk and Baris Asikgil. A novel hybrid robust tapering approach for nonlinear regression in the presence of autocorrelation and outliers. Communications in Statistics: Simulation and Computation, 52(11):5550–5566, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Karavigh:2024:MBC

[KA24] Mohammad Hassan Ahmadi Karavigh and Amirhossein Amiri. MEWMA based control charts with runs rules for monitoring multivariate simple linear regression profiles in Phase II. Communications in Statistics: Simulation and Computation, 53(3):1107–1134, 2024. CODEN CSSCDB. ISSN 0361-0918.

Kalina:2021:CME

[Kal21] Jan Kalina. Common multivariate estimators of location and scatter capture the symmetry of the underlying distribution. Communications in Statistics: Simulation and Computation, 50(10):2845–2857, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kalina:2022:MWC

[Kal22] Jan Kalina. The minimum weighted covariance determinant estimator revisited. Communications in Statistics: Simulation and Computation, 51(7):3888–3900, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Kahrari:2022:SSC

[KAVdSFG22] Fereshte Kahrari, Reinaldo Boris Arellano-Valle, Clecio da Silva Ferreira, and Diego Gallardo. Some simulation/computation in multivariate linear models of scale mixtures of skew-normal-Cauchy distributions. Communications in Statistics: Simulation and Computation, 51(11):6425– 6440, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kahrari:2020:BPM

[KAVR20] Fereshte Kahrari, Reinaldo Boris Arellano-Valle, and Majid Rezaei. Bias prevention of maximum likelihood estimates for skew-normal-Cauchy distribution. Communications in Statistics: Simulation and Computation, 49(1):1–15, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kazemitabar:2023:GFC

[Kaz23] Javad Kazemitabar. A general framework for constructing distributions satisfying Benford’s law. Communications in Statistics: Simulation and Computation, 52(12):6160–6167, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kohansal:2021:EPK

[KB21] Akram Kohansal and Hassan S. Bakouch. Estimation procedures for Kumaraswamy distribution parameters under adaptive type-II hybrid progressive censoring. Communications in Statistics: Simulation and Computation, 50(12):4059–4078, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2023:PIM

[KB23] Raju Kumar and Lal Mohan Bhar. Procedure for the identification of multiple influential observations in block design for incomplete multi-response experiments in presence of masking. Communications in Statistics: Simulation and Computation, 52(5):2167–2176, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2024:NCS

[KBB24] Ankit Kumar, Sharmodeep Bhattacharyya, and Kristofer Bouchard. Numerical characterization of support recovery in sparse regression with correlated design. Communications in Statistics: Simulation and Computation, 53(3):1504–1518, 2024. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:CCM

[KBR21] Jong-Min Kim, Jaiwook Baik, and Mitch Reller. Control charts of mean and variance using copula Markov SPC and conditional distribution by copula. Communications in Statistics: Simulation and Computation, 50(1):85–102, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Khan:2023:CMC

[KBR^{+}23] Abid Khan, Zahid Bashir, H. M. Kashif Rasheed, Mahmood Ulhassan, and Rashid Ahmed. Construction of minimal circular nearly strongly balanced repeated measurements designs and their conversion. Communications in Statistics: Simulation and Computation, 52(12):5749–5758, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2022:ADG

[KBYS22] M. Kumar, P. N. Bajeel, O. P. Yadav, and K. C. Siju. Application of degradation growth model in the estimation of Bayesian system reliability. Communications in Statistics: Simulation and Computation, 51(12):7246–7265, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Khati:2024:AMS

[KCS24] Amisha Khati, Neelam Chantola, and S. B. Singh. Availability modeling of systems incorporating uncertainty and multiple failure modes using critical repair time and copula. Communications in Statistics: Simulation and Computation, 53 (1):595–609, 2024. CODEN CSSCDB. ISSN 0361-0918.

Kwessi:2022:ANN

[KE22] Eddy Kwessi and Lloyd J. Edwards. Artificial neural networks with a signed-rank objective function and applications. Communications in Statistics: Simulation and Computation, 51(6):3363–3388, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kizilaslan:2020:IFR

[KEE20] Busenur Kizilaslan, Erol Egrioglu, and Atif Ahmet Evren. Intuitionistic fuzzy ridge regression functions. Communications in Statistics: Simulation and Computation, 49(3):699– 708, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kalemkerian:2024:ITB

[KF24] Juan Kalemkerian and Diego Fernández. An independence test based on recurrence rates. An empirical study and applications to real data. Communications in Statistics: Simulation and Computation, 53(2):912–932, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Kamranfar:2023:IGD

[KFB23] Hoda Kamranfar, Mitra Fouladirad, and Narayanaswamy Balakrishnan. Inference for a gradually deteriorating system with imperfect maintenance. Communications in Statistics: Simulation and Computation, 52(12):5811–5829, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Ke:2023:NNI

[KFEL23] Xiao Ke, Kai-Tai Fang, A. M. Elsawah, and Yuxuan Lin. New non-isomorphic detection methods for orthogonal designs. Communications in Statistics: Simulation and Computation, 52(1):27–42, 2023. CODEN CSSCDB. ISSN 0361-0918.

Khalili:2023:NEC

[KFM23] Sanaz Khalili, Javad Faradmal, and Hossein Mahjub. A new estimator to control collinearity problems in correlated binary response. Communications in Statistics: Simulation and Computation, 52(12):5991–6001, 2023. CODEN CSSCDB. ISSN 0361-0918.

Krishna:2022:ITL

[KG22] Hare Krishna and Rajni Goel. Inferences for two Lindley populations based on joint progressive type-II censored data. Communications in Statistics: Simulation and Computation, 51(9):4919–4936, 2022. CODEN CSSCDB. ISSN 0361-0918.

Khan:2020:RRE

[KGF20] Hina Khan, Sat Gupta, and Habiba Farhat. A ratio-cum-regression estimator of population mean in unequal probability sampling design. Communications in Statistics: Simulation and Computation, 49(2):484–490, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Krit:2021:GFT

[KGR21] Meryam Krit, Olivier Gaudoin, and Emmanuel Remy. Goodness-of-fit tests for the Weibull and extreme value distributions: A review and comparative study. Communications in Statistics: Simulation and Computation, 50(7):1888–1911, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2023:BGY

[KH23] C. Satheesh Kumar and S. Harisankar. Bivariate generalized Yule distribution: properties and applications. Communications in Statistics: Simulation and Computation, 52(7): 3233–3256, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kim:2022:DMS

[KHK22] Yunmi Kim, Lijuan Huo, and Tae-Hwan Kim. Dealing with Markov-switching parameters in quantile regression models. Communications in Statistics: Simulation and Computation, 51(11):6773–6791, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:DIC

[Kim21] Chanmin Kim. Deviance information criteria for mixtures of distributions. Communications in Statistics: Simulation and Computation, 50(10):2935–2948, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kazemi:2020:SSI

[KJ20] M. R. Kazemi and A. A. Jafari. Small sample inference for the common coefficient of variation. Communications in Statistics: Simulation and Computation, 49(1):226–243, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Kannan:2022:EDR

[KJB22] G. Kannan, P. Jeyadurga, and S. Balamurali. Economic design of repetitive group sampling plan based on truncated life test under Birnbaum–Saunders distribution. Communications in Statistics: Simulation and Computation, 51(12): 7334–7350, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kim:2023:MBR

[KJB23] Jaeoh Kim, Byoungwook Jang, and Sungwan Bang. Modelbased recursive partitioning algorithm to penalized noncrossing multiple quantile regression for the right-censored data. Communications in Statistics: Simulation and Computation, 52(8):3741–3757, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kim:2020:LLM

[KJK20] Seongyong Kim, Saebom Jeon, and Daeyoung Kim. On log– linear modeling for an incomplete two-way contingency table with one variable subject to nonresponse. Communications in Statistics: Simulation and Computation, 49(4):973–988, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kocyigit:2022:RER

[KK22a] Eda Gizem Koçyigit and Cem Kadilar. Ratio estimators for ranked set sampling in the presence of tie information. Communications in Statistics: Simulation and Computation, 51 (11):6826–6839, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kordasiabi:2022:PNU

[KK22b] S. Rahnamay Kordasiabi and S. Khazaei. Prediction of the nonsampled units in survey design with the finite population using Bayesian nonparametric mixture model. Communications in Statistics: Simulation and Computation, 51(6):3457– 3470, 2022. CODEN CSSCDB. ISSN 0361-0918.

Khan:2017:TGE

[KKH17] Muhammad Shuaib Khan, Robert King, and Irene Lena Hudson. Transmuted generalized exponential distribution: A generalization of the exponential distribution with applications to survival data. Communications in Statistics: Simulation and Computation, 46(6):4377–4398, 2017. CODEN CSSCDB. ISSN 0361-0918. See correction [AE23c].

Kim:2023:RAI

[KKK23] Jinheum Kim, Jayoun Kim, and Seong W. Kim. Regression analysis of the illness-death model with a shared frailty when all transition times are interval censored. Communications in Statistics: Simulation and Computation, 52(1):247–259, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kargapolova:2021:SMJ

[KKO21] N. A. Kargapolova, E. I. Khlebnikova, and V. A. Ogorodnikov. Stochastic models of joint non-stationary time-series of air temperature, relative humidity and atmospheric pressure. Communications in Statistics: Simulation and Computation, 50(12):3972–3983, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kargin:2021:MCS

[KKZ21] B. A. Kargin, E. G. Kablukova, and P. Zhen. Monte Carlo simulation of the optical radiation transfer process in stochastic scattering media. Communications in Statistics: Simulation and Computation, 50(12):3984–3991, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:RCS

[KL21] Hanwool Kim and Sangyeol Lee. On residual CUSUM statistic for PINAR(1) model in statistical design and diagnostic of control chart. Communications in Statistics: Simulation and Computation, 50(5):1290–1314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kuang:2022:BEB

[KL22] Nenghui Kuang and Ying Li. Berry–Esséen bounds and almost sure CLT for the quadratic variation of the subbifractional Brownian motion. Communications in Statistics:

Simulation and Computation, 51(8):4257–4275, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Kim:2023:CQC

[KL23] Chang Kyeom Kim and Sangyeol Lee. Conditional quantile change test for time series based on support vector regression. Communications in Statistics: Simulation and Computation, 52(11):5567–5584, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kang:2021:PPU

[KLK21] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Posterior properties under matching priors for generalized gamma distribution. Communications in Statistics: Simulation and Computation, 50(4):1202–1216, 2021. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2022:NME

[KLM^{+}22] Christos Koukouvinos, Angeliki Lappa, Marilena Mitrouli, Paraskevi Roupa, and Ondrej Turek. Numerical methods for estimating the tuning parameter in penalized least squares problems. Communications in Statistics: Simulation and Computation, 51(4):1542–1563, 2022. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2024:CIT

[KLM24] K. Krishnamoorthy, Shanshan Lv, and Md Monzur Murshed. Combining independent tests for a common parameter of several continuous distributions: a new test and power comparisons. Communications in Statistics: Simulation and Computation, 53(4):1837–1856, 2024. CODEN CSSCDB. ISSN 0361-0918.

Kim:2020:ITV

[KM20] Tae-Hwan Kim and Christophe Muller. Inconsistency transmission and variance reduction in two-stage quantile regression. Communications in Statistics: Simulation and Computation, 49(4):1044–1077, 2020. CODEN CSSCDB. ISSN 0361-0918.

Katebi:2022:DSM

[KM22] Mehdi Katebi and M. Bameni Moghadam. A double sampling multivariate T^{2} control chart with variable sample size and variable sampling interval. Communications in Statistics:

Simulation and Computation, 51(7):3578–3595, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Khedmati:2022:PRP

[KN22] Majid Khedmati and Seyed Taghi Akhavan Niaki. Phase-I robust parameter estimation of simple linear profiles in multistage processes. Communications in Statistics: Simulation and Computation, 51(2):460–485, 2022. CODEN CSSCDB. ISSN 0361-0918.

Khan:2023:MAE

[KNSA23] Nasrullah Khan, Muhammad S. Nawaz, Rehan A. K. Sherwani, and Muhammad Aslam. Moving average EWMA chart for the Weibull distribution. Communications in Statistics: Simulation and Computation, 52(5):2231–2240, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Kuran:2021:MRC

[KÖ21a] Özge Kuran and M. Revan Özkale. Marginal ridge conceptual predictive model selection criterion in linear mixed models. Communications in Statistics: Simulation and Computation, 50(2):581–607, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kuran:2021:SRL

[KÖ21b] Özge Kuran and M. Revan Özkale. Stochastic restricted Liu predictors in linear mixed models. Communications in Statistics: Simulation and Computation, 50(9):2561–2580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kolyukhin:2024:SAC

[Kol24] Dmitriy Kolyukhin. Study the accuracy of the correlation fractal dimension estimation. Communications in Statistics: Simulation and Computation, 53(1):219–233, 2024. CODEN CSSCDB. ISSN 0361-0918.

Khalfi:2020:SRA

[KOT20] Linda Khalfi and Megdouda Ourbih-Tari. Stochastic risk analysis in Monte Carlo simulation: a case study. Communications in Statistics: Simulation and Computation, 49(11): 3041–3053, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kumari:2021:BPE

[KP21] Neera Kumari and Ranjita Pandey. On Bayesian parameter estimation of beta log Weibull distribution under Type-II censoring. Communications in Statistics: Simulation and Computation, 50(3):650–668, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kalaivani:2021:PEE

[KR21] S. Kalaivani and J. Ravichandran. Performance evaluation of exponential distribution using Six Sigma-based tail probabilities. Communications in Statistics: Simulation and Computation, 50(11):3125–3145, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2022:ZIH

[KR22] C. Satheesh Kumar and Rakhi Ramachandran. On zeroinflated hyper-Poisson distribution and its applications. Communications in Statistics: Simulation and Computation, 51 (3):868–881, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kotb:2023:CEP

[KR23] Mohammed S. Kotb and Mohammad Z. Raqab. Comparison of estimators and predictors based on modified Weibull records: Bayesian and non-Bayesian approaches. Communications in Statistics: Simulation and Computation, 52(6): 2518–2541, 2023. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2022:NPS

[Kri22] Kalimuthu Krishnamoorthy. A note on the paper “Singhasomboona, L., Panichkitkosolkula, W. and Volodin, A. (2020). Confidence intervals for the ratio of medians of two independent log-normal distributions. Communications in Statistics — Simulation and Computation. https://doi.org/10.1080/03610918.2020.1812649”. Communications in Statistics: Simulation and Computation, 51(12): 7613–7614, 2022. CODEN CSSCDB. ISSN 0361-0918.

Korzenowski:2020:QMS

[KS20a] A. L. Korzenowski and W. L. Simões. Quality monitoring by special charts for highly customized production systems. Communications in Statistics: Simulation and Computation, 49(8):2181–2192, 2020. CODEN CSSCDB. ISSN 0361-0918.

Krzysko:2020:RMF

[KS20b] Mirosław Krzyśko and Lukasz Smaga. Robust multivariate functional discriminant coordinates. Communications in Statistics: Simulation and Computation, 49(3):717–733, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2020:SPA

[KS20c] C. Satheesh Kumar and S. Sreejakumari. On some properties and applications of intervened cluster negative binomial distribution. Communications in Statistics: Simulation and Computation, 49(9):2357–2371, 2020. CODEN CSSCDB. ISSN 0361-0918.

Khalid:2022:SIM

[KS22a] Mohd Khalid and Garib Nath Singh. Some imputation methods to deal with the issue of missing data problems due to random non-response in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 51 (12):7266–7286, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kulasekera:2022:MRB

[KS22b] K. B. Kulasekera and Chathura Siriwardhana. Multiresponse based personalized treatment selection with data from crossover designs for multiple treatments. Communications in Statistics: Simulation and Computation, 51(2):554– 569, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kabata:2023:VSD

[KS23a] Daijiro Kabata and Mototsugu Shintani. Variable selection in double/debiased machine learning for causal inference: an outcome-adaptive approach. Communications in Statistics: Simulation and Computation, 52(12):5880–5893, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Kaur:2023:CBS

[KS23b] Parneet Kaur and Kush Sharma. Constant block-sum Youden-m square and PBIB designs using Galois field. Communications in Statistics: Simulation and Computation, 52 (11):5396–5400, 2023. CODEN CSSCDB. ISSN 0361-0918.

King:2023:SSD

[KS23c] Clay King and James D. Stamey. Sample size determination for a Bayesian cost-effectiveness model with structural zero costs. Communications in Statistics: Simulation and Computation, 52(5):2241–2256, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2023:NCL

[KS23d] C. Satheesh Kumar and A. S. Satheenthar. A new class of lifetime distribution with decreasing failure rate: Properties and applications. Communications in Statistics: Simulation and Computation, 52(7):3343–3364, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kazemi:2021:EPL

[KSAR21] M. Kazemi, D. Shahsavani, M. Arashi, and P. C. Rodrigues. Estimation in partial linear model with spline modal function. Communications in Statistics: Simulation and Computation, 50(11):3256–3272, 2021. CODEN CSSCDB. ISSN 0361-0918.

Karunarasan:2023:CBM

[KSP23] Darshika Karunarasan, Roshini Sooriyarachchi, and Vimukthini Pinto. A comparison of Bayesian Markov chain Monte Carlo methods in a multilevel scenario. Communications in Statistics: Simulation and Computation, 52(10):4756–4772, 2023. CODEN CSSCDB. ISSN 0361-0918.

Khoo:2022:MCC

[KST^{+}22] Michael B. C. Khoo, Sajal Saha, Sin Yin Teh, Abdul Haq, and How Chinh Lee. The median control chart for process monitoring in short production runs. Communications in Statistics: Simulation and Computation, 51(10):5816–5831, 2022. CODEN CSSCDB. ISSN 0361-0918.

Koh:2023:HGP

[KT23] Woon Yuen Koh and Chunhao Tu. A hybrid generalized propensity score approach for observational studies. Communications in Statistics: Simulation and Computation, 52 (9):4459–4468, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2022:ACR

[KTK22] Pushkal Kumar, Manas Ranjan Tripathy, and Somesh Kumar. Alternative classification rules for two normal populations with a common mean and ordered variances. Communications in Statistics: Simulation and Computation, 51(11): 6881–6901, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kessentini:2020:RPD

[KTZ20] Sameh Kessentini, Mariem Tounsi, and Raoudha Zine. The Riesz probability distribution: Generation and EM algorithm.

Communications in Statistics: Simulation and Computation, 49(8):2114–2133, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kashif:2021:IDP

[KUA21] Muhammad Kashif, Muhammad Aman Ullah, and Muhammad Aslam. Influential diagnostics with Pena’s statistic for the modified ridge regression. Communications in Statistics: Simulation and Computation, 50(12):3870–3884, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Khan:2023:PRD

[KUA23] Aamna Khan, Muhammad Aman Ullah, and Muhammad Amin. Poisson regression diagnostics with ridge estimation. Communications in Statistics: Simulation and Computation, 52(9):4174–4192, 2023. CODEN CSSCDB. ISSN 0361-0918.

Khan:2020:MST

[KuIK20] Waqar Muhammad Khan and Asad ul Islam Khan. Most stringent test of independence for time series. Communications in Statistics: Simulation and Computation, 49(11): 2808–2826, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kurnaz:2021:PSHa

[Kur21a] Fatma Sevinç Kurnaz. Performances of some high dimensional regression methods. Communications in Statistics: Simulation and Computation, 50(6):1820–1836, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Kurnaz:2021:PSHb

[Kur21b] Fatma Sevinç Kurnaz. Performances of some high dimensional regression methods: sparse principal component regression. Communications in Statistics: Simulation and Computation, 50(9):2529–2543, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kurtoglu:2021:MRP

[Kur21c] Fikriye Kurtoglu. Modified ridge parameter estimators for log-gamma model: Monte Carlo evidence with a graphical investigation. Communications in Statistics: Simulation and Computation, 50(7):2168–2183, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kuruwita:2023:VSS

[Kur23] C. N. Kuruwita. Variable selection in the single-index quantile regression model with high-dimensional covariates. Communications in Statistics: Simulation and Computation, 52(3): 1119–1131, 2023. CODEN CSSCDB. ISSN 0361-0918.

Karmakar:2023:PBD

[KVJH23] Sayantani Karmakar, Cini Varghese, Seema Jaggi, and Mohd Harun. Partially balanced t-designs. Communications in Statistics: Simulation and Computation, 52(12):6141–6148, 2023. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2022:IRR

[KVS22] Amod Kumar, Gajendra K. Vishwakarma, and G. N. Singh. An improved randomized response model for simultaneous estimation of means of two quantitative sensitive variables. Communications in Statistics: Simulation and Computation, 51(10):5967–5987, 2022. CODEN CSSCDB. ISSN 0361-0918.

Khalil:2021:MES

[KZG21] Sadia Khalil, Qi Zhang, and Sat Gupta. Mean estimation of sensitive variables under measurement errors using optional RRT models. Communications in Statistics: Simulation and Computation, 50(5):1417–1426, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Lattef:2023:MUT

[LA23] Mustafa Nadhim Lattef and Mustafa I. Alheety. On modified unbiased two-parameter estimator. Communications in Statistics: Simulation and Computation, 52(7):2843–2864, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lukman:2024:IMG

[LAK24] Adewale F. Lukman, Muhammad Amin, and B. M. Golam Kibria. Influence measures in gamma modified ridge type estimator. Communications in Statistics: Simulation and Computation, 53(1):147–163, 2024. CODEN CSSCDB. ISSN 0361-0918.

Lukman:2022:MRT

[LAKA22] Adewale F. Lukman, Kayode Ayinde, B. M. Golam Kibria, and Emmanuel T. Adewuyi. Modified ridge-type estimator for the gamma regression model. Communications in Statistics:

Simulation and Computation, 51(9):5009–5023, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Larsson:2020:TIE

[Lar20] Rolf Larsson. Testing for INAR effects. Communications in Statistics: Simulation and Computation, 49(10):2745–2764, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lavine:2021:WFA

[Lav21] Michael Lavine. WHIM: function approximation where it matters. Communications in Statistics: Simulation and Computation, 50(12):3839–3869, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:CIE

[LAZ21] Pengfei Li, Oludotun J. Akinlawon, and Shengli Zhao. Controlling individual and experimentwise error rates in replicated regular two-level factorial experiments. Communications in Statistics: Simulation and Computation, 50(6):1770– 1790, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:RRM

[LB22] Kai Liu and N. Balakrishnan. Recurrence relations for moments of order statistics from half logistic-geometric distribution and their applications. Communications in Statistics: Simulation and Computation, 51(11):6537–6555, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Lien:2023:SRM

[LB23a] Donald Lien and N. Balakrishnan. Some results on multiple regression analysis with data cleaned by trimming and Winsorization. Communications in Statistics: Simulation and Computation, 52(10):5082–5089, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lin:2023:TLD

[LB23b] Chien-Tai Lin and N. Balakrishnan. Topp–Leone distribution with an application to binomial sampling. Communications in Statistics: Simulation and Computation, 52(9):4075–4086, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2023:GBS

[LBH23] Kai Liu, Narayanaswamy Balakrishnan, and Mu He. Generalized Birnbaum–Saunders mixture cure frailty model: inferential method and an application to bone marrow transplant data. Communications in Statistics: Simulation and Computation, 52(11):5655–5679, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lin:2022:ETO

[LBL22] Chien-Tai Lin, Narayanaswamy Balakrishnan, and Man Ho Ling. Exact tests for outliers in Laplace samples. Communications in Statistics: Simulation and Computation, 51(10): 5794–5815, 2022. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:DPU

[LC22] Zhang Liu and Ping Chen. Dividend payments until drawdown time for risk models driven by spectrally negative Lévy processes. Communications in Statistics: Simulation and Computation, 51(12):7226–7245, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lau:2022:RUL

[LCC22] John W. Lau, Edward Cripps, and Sally Cripps. Remaining useful life prediction: a multiple product partition approach. Communications in Statistics: Simulation and Computation, 51(9):5288–5307, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lu:2023:EIF

[LD23] Yujiao Lu and Wanyang Dai. Estimating intensity functions of spatial inhomogeneous Poisson point processes via a Stein estimator. Communications in Statistics: Simulation and Computation, 52(10):4816–4848, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lagha:2023:NSE

[LDA23] Karima Lagha, Lamia Djerroud, and Smail Adjabi. Nonparametric sequential estimation of the cumulative function by orthogonal series and application in reliability. Communications in Statistics: Simulation and Computation, 52(11): 5633–5654, 2023. CODEN CSSCDB. ISSN 0361-0918.

Levy:2023:FCD

[LE23] Roy Levy and Craig K. Enders. Full conditional distributions for Bayesian multilevel models with additive or interactive effects and missing data on covariates. Communications in Statistics: Simulation and Computation, 52(7):2899–2923, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lee:2020:ACL

[Lee20] Seung-Hwan Lee. Assessing the censored linear regression model using martingale approximation. Communications in Statistics: Simulation and Computation, 49(2):408–418, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lin:2021:EUC

[LES21] Wei-Cheng Lin, Takeshi Emura, and Li-Hsien Sun. Estimation under copula-based Markov normal mixture models for serially correlated data. Communications in Statistics: Simulation and Computation, 50(12):4483–4515, 2021. CODEN CSSCDB. ISSN 0361-0918.

Levada:2021:CFB

[Lev21] Alexandre L. M. Levada. Closed-form Bayesian image denoising: improving the adaptive Wiener filter through pairwise Gaussian–Markov random fields. Communications in Statistics: Simulation and Computation, 50(4):1094–1118, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2024:NAP

[LF24] Yinan Li and Kai-Tai Fang. A new approach to parameter estimation of mixture of two normal distributions. Communications in Statistics: Simulation and Computation, 53(3): 1161–1187, 2024. CODEN CSSCDB. ISSN 0361-0918.

Lopez-Fidalgo:2022:EOE

[LFMMRH22] Jesús López-Fidalgo, Raúl Martín-Martín, and Mercedes Rodríguez-Hernández. Estimators and D-optimal experimental designs for mixtures of binary responses. Communications in Statistics: Simulation and Computation, 51(4):1488–1502, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lai:2024:CUD

[LFPL24] Jianfa Lai, Kai-Tai Fang, Xiaoling Peng, and Yuxuan Lin. Construction of uniform designs over continuous domain in computer experiments. Communications in Statistics: Simulation and Computation, 53(1):130–146, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Lu:2021:NRE

[LFZ21] Yan Lu, Ye Fu, and Guoyi Zhang. Nonparametric regression estimators in dual frame surveys. Communications in Statistics: Simulation and Computation, 50(3):854–864, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Long:2021:PRM

[LG21] Yang Long and Deng Guohe. A perturbed risk model with constant interest and periodic barrier dividend strategy. Communications in Statistics: Simulation and Computation, 50 (8):2467–2481, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2020:LRG

[LH20] Baohua Li and Lixia Hu. Laplacian regularized generalized Dirichlet mixture distribution for data clustering. Communications in Statistics: Simulation and Computation, 49(1): 16–28, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lv:2024:SIG

[LHG24] Qi Lv, Rui Hua, and Wenhao Gui. Statistical inference of Gompertz distribution under general progressive type II censored competing risks sample. Communications in Statistics: Simulation and Computation, 53(2):682–701, 2024. CODEN CSSCDB. ISSN 0361-0918.

Lang:2023:CCR

[LHY^{+}23] Junjun Lang, Tianyun He, Zhiqiang Yu, Yi Wu, and Xuejun Wang. Complete convergence for randomly weighted sums of random variables and its application in linear-time-invariant systems. Communications in Statistics: Simulation and Computation, 52(3):868–883, 2023. CODEN CSSCDB. ISSN 0361-0918.

Linhares:2022:EWE

[Lin22] Raquel Romes Linhares. The ECF-WS estimator for univariate symmetric stable distributions with application in seismic trace signals. Communications in Statistics: Simulation and Computation, 51(12):7487–7499, 2022. CODEN CSSCDB. ISSN 0361-0918.

Laxmi:2022:CRA

[LJ22a] P. Vijaya Laxmi and K. Jyothsna. Cost and revenue analysis of an impatient customer queue with second optional service and working vacations. Communications in Statistics: Simulation and Computation, 51(8):4799–4814, 2022. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:FHT

[LJ22b] Zhe Liu and Lifen Jia. First hitting time for renewal process with uncertain interarrival times and random rewards. Communications in Statistics: Simulation and Computation, 51 (10):5539–5555, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lee:2020:PLD

[LJC^{+}20] JungJun Lee, Jae-Hwan Jhong, Young-Rae Cho, SungHwan Kim, and Ja-Yong Koo. Penalized log-density estimation using Legendre polynomials. Communications in Statistics: Simulation and Computation, 49(11):2844–2860, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:2021:DSN

[LK21a] Ming Ha Lee and Michael B. C. Khoo. Double sampling np chart with estimated process parameter. Communications in Statistics: Simulation and Computation, 50(8):2232–2250, 2021. CODEN CSSCDB. ISSN 0361-0918.

Levene:2021:ESJ

[LK21b] Mark Levene and Aleksejus Kononovicius. Empirical survival Jensen-Shannon divergence as a goodness-of-fit measure for maximum likelihood estimation and curve fitting. Communications in Statistics: Simulation and Computation, 50(11): 3751–3767, 2021. CODEN CSSCDB. ISSN 0361-0918.

Lun:2024:GAI

[LK24] Zhixin Lun and Ravindra Khattree. A general approach for imputation of non-normal continuous data based on copula transformation. Communications in Statistics: Simulation and Computation, 53(1):567–594, 2024. CODEN CSSCDB. ISSN 0361-0918.

Lee:2023:ESD

[LKHC23] Ming Ha Lee, Michael B. C. Khoo, Abdul Haq, and XinYing Chew. Economic-statistical design of the variable sampling interval Poisson EWMA chart. Communications in Statistics: Simulation and Computation, 52(5):2139–2150, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Lovig:2023:MBR

[LKR^{+}23] Maxwell Lovig, Sadia Khalil, Sumaita Rahman, Pujita Sapra, and Sat Gupta. A mixture binary RRT model with a unified measure of privacy and efficiency. Communications in Statistics: Simulation and Computation, 52(6):2727–2737, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lee:2023:EFQ

[LL23] Sangyeol Lee and Sangjo Lee. Exponential family QMLEbased CUSUM test for integer-valued time series. Communications in Statistics: Simulation and Computation, 52(5): 2022–2043, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:ELB

[LLL22] Xiaohui Liu, Yuzi Liu, and Fucai Lu. Empirical likelihoodbased unified confidence region for a predictive regression model. Communications in Statistics: Simulation and Computation, 51(5):2122–2139, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:QOP

[LLLL22] Wenhan Li, Lixia Liu, Cuixiang Li, and Guiwen Lv. Quanto option pricing with a jump diffusion process. Communications in Statistics: Simulation and Computation, 51(5):2095–2109, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:PVO

[LLLY23a] Cuixiang Li, Huili Liu, Lixia Liu, and Qiumei Yao. Pricing vulnerable options under jump diffusion processes using double Mellin transform. Communications in Statistics: Simulation and Computation, 52(3):703–716, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Li:2023:BCP

[LLLY23b] Qing Li, Lijie Liu, Tianqi Li, and Kehui Yao. Bayesian change-points detection assuming a power law process in the recurrent-event context. Communications in Statistics: Simulation and Computation, 52(12):6011–6033, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lethikim:2023:ACA

[LLV23] Ngoc Lethikim, Tuan Lehoang, and Tai Vovan. Automatic clustering algorithm for interval data based on overlap distance. Communications in Statistics: Simulation and Computation, 52(5):2194–2209, 2023. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:SAS

[LLZ22] Dong Li, Muyi Li, and Lianbin Zeng. Simulation and application of subsampling for threshold autoregressive movingaverage models. Communications in Statistics: Simulation and Computation, 51(5):2110–2121, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lemonte:2023:MBS

[LMFMA23] Artur J. Lemonte, Guillermo Martínez-Flórez, and Germán Moreno-Arenas. Multivariate Birnbaum–Saunders powernormal model and associated inference. Communications in Statistics: Simulation and Computation, 52(5):1769–1785, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2020:SEI

[LMP20] Huilan Liu, Junjie Ma, and Changgen Peng. Shrinkage estimation for identification of linear components in composite quantile additive models. Communications in Statistics: Simulation and Computation, 49(10):2678–2692, 2020. CODEN CSSCDB. ISSN 0361-0918.

Long:2023:EPR

[Lon23] Bing Long. Estimation and prediction for the Rayleigh distribution based on double Type-I hybrid censored data. Communications in Statistics: Simulation and Computation, 52 (8):3553–3567, 2023. CODEN CSSCDB. ISSN 0361-0918.

Le:2020:IMC

[LPNP20] Hoa Le, Uyen Pham, Phuong Nguyen, and The Bao Pham. Improvement on Monte Carlo estimation of HPD intervals. Communications in Statistics: Simulation and Computation, 49(8):2164–2180, 2020. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:AEF

[LPP21] Kexuan Li, Aleksey S. Polunchenko, and Andrey Pepelyshev. Analytic evaluation of the fractional moments for the quasistationary distribution of the Shiryaev martingale on an interval. Communications in Statistics: Simulation and Computation, 50(9):2705–2720, 2021. CODEN CSSCDB. ISSN 0361-0918.

Larsson:2023:ADF

[LR23] Rolf Larsson and Jesper Rydén. Applications of discrete factor analysis. Communications in Statistics: Simulation and Computation, 52(10):4592–4602, 2023. CODEN CSSCDB. ISSN 0361-0918.

Louzada:2020:EPD

[LRF20] Francisco Louzada, Pedro Luiz Ramos, and Paulo Henrique Ferreira. Exponential-Poisson distribution: estimation and applications to rainfall and aircraft data with zero occurrence. Communications in Statistics: Simulation and Computation, 49(4):1024–1043, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lin:2022:NMT

[LS22a] Fuming Lin and Daimin Shi. A new method of testing for a unit root in the INAR(1) model based on variances. Communications in Statistics: Simulation and Computation, 51(10): 5915–5932, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lu:2022:SLI

[LS22b] Jun Lu and Lei Shi. Sparse local influence analysis. Communications in Statistics: Simulation and Computation, 51(11): 6235–6245, 2022. CODEN CSSCDB. ISSN 0361-0918.

Luo:2023:ISS

[LS23] Li Luo and James G. Surles. Inference for the stress-strength reliability for the two-parameter Burr type X under Type I and Type II censoring. Communications in Statistics: Simulation and Computation, 52(5):1713–1718, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lee:2023:CBA

[LSC23] J. Y. Lee, P. S. Shen, and K. F. Cheng. A comparison of bootstrap approaches for homogeneity test based on dichotomous outcome in meta-analysis. Communications in Statistics: Simulation and Computation, 52(9):4218–4229, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lin:2024:PTJ

[LSTE24] Yuan-Hsin Lin, Li-Hsien Sun, Yi-Ju Tseng, and Takeshi Emura. The Pareto type I joint frailty-copula model for clustered bivariate survival data. Communications in Statistics: Simulation and Computation, 53(4):2006–2030, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Lodhi:2023:PCC

[LTB23] Chandrakant Lodhi, Yogesh Mani Tripathi, and Ritwik Bhattacharya. On a progressively censored competing risks data from Gompertz distribution. Communications in Statistics: Simulation and Computation, 52(4):1278–1299, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:2023:MCV

[LTJC23] Pei-Hsi Lee, Chau-Chen Torng, Hao-Ren Jhong, and Chao-Yu Chou. Monitoring the coefficient of variation using a doublesampling control chart. Communications in Statistics: Simulation and Computation, 52(10):4849–4863, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2023:SPR

[LTN23] Tingting Liu, Weizhong Tian, and Wei Ning. Sequential probability ratio test for zero inflation in counting data. Communications in Statistics: Simulation and Computation, 52(4): 1344–1360, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lodhi:2021:EPT

[LTR21] Chandrakant Lodhi, Yogesh Mani Tripathi, and Manoj Kumar Rastogi. Estimating the parameters of a truncated normal distribution under progressive type II censoring. Communications in Statistics: Simulation and Computation, 50 (9):2757–2781, 2021. CODEN CSSCDB. ISSN 0361-0918.

Lu:2023:REM

[Lu23] Fei Lu. Robust estimation via modified Cholesky decomposition for modal partially nonlinear models with longitudinal data. Communications in Statistics: Simulation and Computation, 52(10):5110–5123, 2023. CODEN CSSCDB. ISSN 0361-0918.

Luepsen:2023:ABV

[Lue23] Haiko Luepsen. ANOVA with binary variables: the F -test and some alternatives. Communications in Statistics: Simulation and Computation, 52(3):745–769, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Lui:2022:NPC

[Lui22] Kung-Jong Lui. Notes on power comparison between the sequential parallel comparison and other commonly-used designs. Communications in Statistics: Simulation and Computation, 51(4):1783–1800, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:IMM

[LW21] Zhouping Li and Hui Wang. Improving K-means method via shrinkage estimation and LVQ algorithm. Communications in Statistics: Simulation and Computation, 50(11):3166–3181, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:CFE

[LW22a] Changbiao Liu and Jun Wang. A computationally fast estimator for semiparametric multinomial choice model. Communications in Statistics: Simulation and Computation, 51 (6):3355–3362, 2022. CODEN CSSCDB. ISSN 0361-0918.

Luo:2022:SIS

[LW22b] Guowang Luo and Mixia Wu. Statistical inference for semiparametric varying-coefficient spatial autoregressive models under restricted conditions. Communications in Statistics: Simulation and Computation, 51(5):2268–2286, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:2023:UTA

[LW23] Bu Hyoung Lee and William W. S. Wei. The use of temporally aggregated data in modeling and testing a variance change in a time series. Communications in Statistics: Simulation and Computation, 52(7):3183–3200, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:CFQ

[LWC22] Weiyi Liu, Die Wan, and Wei Chen. A closed-form quasimaximum likelihood estimator of bid–ask spread. Communications in Statistics: Simulation and Computation, 51(3): 1066–1079, 2022. CODEN CSSCDB. ISSN 0361-0918.

Liu:2024:CEN

[LWP24] Zhan Liu, Lin Wang, and Yingli Pan. Calibration estimation for non-probability samples under two distance functions: a comparative study. Communications in Statistics: Simulation and Computation, 53(3):1548–1564, 2024. CODEN CSSCDB. ISSN 0361-0918.

Lu:2023:PBB

[LWS23] Min Lu, Xing Wang, and Rosalie Speeckaert. Price bubbles in Beijing carbon market and environmental policy announcement. Communications in Statistics: Simulation and Computation, 52(3):884–897, 2023. CODEN CSSCDB. ISSN 0361-0918.

Lu:2022:DEC

[LWW22] Weixue Lu, Hecheng Wu, and Liyang Wan. Driving effect of CO_{2} emissions on economic growth-application of empirical likelihood for generalized method of moments. Communications in Statistics: Simulation and Computation, 51(12): 7500–7512, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:SDS

[LWX23] Xinmin Li, Xiao Wang, and Shifeng Xiong. A sequential design strategy for integrating low-accuracy and high-accuracy computer experiments. Communications in Statistics: Simulation and Computation, 52(3):817–824, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Li:2023:REH

[LWZ^{+}23] Ruohong Li, Honglang Wang, Yi Zhao, Jing Su, and Wanzhu Tu. Robust estimation of heterogeneous treatment effects: an algorithm-based approach. Communications in Statistics: Simulation and Computation, 52(10):4981–4998, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Li:2022:ABS

[LX22] Shengnan Li and Xiaoming Xu. Anticipated backward stochastic differential equations and their applications to zero-sum stochastic differential games. Communications in Statistics: Simulation and Computation, 51(5):2363–2380, 2022. CODEN CSSCDB. ISSN 0361-0918.

Long:2023:CLA

[LX23] Yuqi Long and Xingzhong Xu. Classification by likelihood accordance functions. Communications in Statistics: Simulation and Computation, 52(9):4101–4118, 2023. CODEN CSSCDB. ISSN 0361-0918.

Liu:2024:MEN

[LX24] Jiali Liu and Fuyuan Xiao. On the maximum extropy negation of a probability distribution. Communications in Statistics: Simulation and Computation, 53(1):234–246, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Li:2022:FFS

[LXMT22] Yu Li, Min Xiao, Ruixing Ming, and Dongsheng Tu. A family of flexible shrinkage estimators for the variances of highdimensional gene expressions. Communications in Statistics: Simulation and Computation, 51(11):6755–6772, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Lu:2020:ELD

[LY20] Weixue Lu and Shijuan Yang. Empirical likelihood of the distribution function in the finite point under φ-mixing samples. Communications in Statistics: Simulation and Computation, 49(10):2533–2545, 2020. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:CVU

[LY22] Zhe Liu and Xiangfeng Yang. Cross validation for uncertain autoregressive model. Communications in Statistics: Simulation and Computation, 51(8):4715–4726, 2022. CODEN CSSCDB. ISSN 0361-0918.

Liu:2023:SEP

[LY23] Yanping Liu and Juliang Yin. Spline estimation of partially linear regression models for time series with correlated errors. Communications in Statistics: Simulation and Computation, 52(11):5522–5536, 2023. CODEN CSSCDB. ISSN 0361-0918.

Li:2024:APA

[LY24] Yue Li and Jimin Ye. Analysis for partially accelerated dependent competing risks model with masked data based on copula function. Communications in Statistics: Simulation and Computation, 53(2):1020–1036, 2024. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:EFL

[LYP22] Bin Li, Qingzhao Yu, and Lu Peng. Ensemble of fast learning stochastic gradient boosting. Communications in Statistics: Simulation and Computation, 51(1):40–52, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:NEV

[LYY21] Ning Li, Hu Yang, and Jing Yang. Nonnegative estimation and variable selection via adaptive elastic-net for highdimensional data. Communications in Statistics: Simulation and Computation, 50(12):4263–4279, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Liya:2021:GEE

[LYZ^{+}21] Fu Liya, Zhuoran Yang, Jun Zhang, Anle Long, and Yan Zhou. Generalized estimating equations for analyzing multivariate survival data. Communications in Statistics: Simulation and Computation, 50(10):3060–3068, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:FOR

[LZ22] Jie Liu and Haixiang Zhang. First-order random coefficient INAR process with dependent counting series. Communications in Statistics: Simulation and Computation, 51(6):3341– 3354, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2024:RAS

[LZ24] Xiaohu Li and Hongyi Zhua. A risk-averse stochastic approximation of the optimal allocation of active redundancies to coherent systems. Communications in Statistics: Simulation and Computation, 53(4):1633–1644, 2024. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:MPC

[LZQ23] Lu Li, Weiwei Zhuang, and Guoxin Qiu. Modeling and predicting Chinese stock downside risks via Gaussian mixture models and marked self-exciting point process. Communications in Statistics: Simulation and Computation, 52(12): 6249–6267, 2023. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:AAR

[LZSC23] Xiaoqin Li, Lei Zhang, Yan Shen, and Zhiyong Chen. Asymptotic approximations of random ratio model based on AANA sequences. Communications in Statistics: Simulation and Computation, 52(8):3796–3819, 2023. CODEN CSSCDB. ISSN 0361-0918.

Li:2023:MDM

[LZZZ23] Weixia Li, Zhibin Zhu, Xiaoyuan Zhu, and Chongqi Zhang. Mixture design model with uncertain response and its application. Communications in Statistics: Simulation and Computation, 52(6):2310–2319, 2023. CODEN CSSCDB. ISSN 0361-0918.

Majid:2023:AED

[MA23a] Abdul Majid and Muhammad Aslam. The Almon Mestimator for the distributed lag model in the presence of outliers. Communications in Statistics: Simulation and Computation, 52(7):3273–3285, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mohamed:2023:REG

[MA23b] Laidi Mohamed and Rassoul Abdelaziz. A robust estimator of the S-Gini index for massive data. Communications in Statistics: Simulation and Computation, 52(8):3502–3519, 2023. CODEN CSSCDB. ISSN 0361-0918.

Majid:2021:ADL

[MAAA21] Abdul Majid, Muhammad Aslam, Saima Altaf, and Muhammad Amanullah. Addressing the distributed lag models with heteroscedastic errors. Communications in Statistics: Simulation and Computation, 50(12):4464–4482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Majid:2023:NRR

[MAAA23] Abdul Majid, Muhammad Amin, Muhammad Aslam, and Shakeel Ahmad. New robust ridge estimators for the linear regression model with outliers. Communications in Statistics: Simulation and Computation, 52(10):4717–4738, 2023. CODEN CSSCDB. ISSN 0361-0918.

Medeiros:2023:IEB

[MAB23] Francisco M. C. Medeiros, Mariana C. Araújo, and Marcelo Bourguignon. Improved estimators in beta prime regression models. Communications in Statistics: Simulation and Computation, 52(11):5125–5138, 2023. CODEN CSSCDB. ISSN 0361-0918.

Macedo:2024:TSM

[Mac24] Pedro Macedo. A two-stage maximum entropy approach for time series regression. Communications in Statistics: Simulation and Computation, 53(1):518–528, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Miranda:2023:PVM

[MAF23] Vânia F. L. Miranda, Henrique J. P. Alves, and Daniel F. Ferreira. Proposition and validation of multivariate tests of independence between two groups of variables. Communications in Statistics: Simulation and Computation, 52(7):2799–2810, 2023. CODEN CSSCDB. ISSN 0361-0918.

Magis:2020:NCC

[Mag20] David Magis. A note on converting calibrated parameters of item response theory models. Communications in Statistics: Simulation and Computation, 49(12):3065–3079, 2020. CODEN CSSCDB. ISSN 0361-0918.

Manoharan:2020:IML

[MAMA20] Thirunanthini Manoharan, Jayanthi Arasan, Habshah Midi, and Mohd Bakri Adam. Influential measures on log-normal model for left-truncated and case-k interval censored data with time-dependent covariate. Communications in Statistics: Simulation and Computation, 49(6):1445–1466, 2020. CODEN CSSCDB. ISSN 0361-0918.

Mogouie:2022:SMB

[MASA22] Hamed Mogouie, Gholam Ali Raissi Ardali, Ehsan Bahrami Samani, and Amirhossein Amiri. Statistical monitoring of binary response attributed social networks considering random effects. Communications in Statistics: Simulation and Computation, 51(3):973–992, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mitra:2021:SIB

[MB21] Debanjan Mitra and Narayanaswamy Balakrishnan. Statistical inference based on left truncated and interval censored data from log-location-scale family of distributions. Communications in Statistics: Simulation and Computation, 50(4): 1073–1093, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ma:2023:BAL

[MB23a] Wenkai Ma and W. John Braun. Bias assessment in local regression. Communications in Statistics: Simulation and Computation, 52(6):2376–2387, 2023. CODEN CSSCDB. ISSN 0361-0918.

Manaa:2023:PSM

[MB23b] Abderrahmen Manaa and Mohamed Bentarzi. On a periodic SETINAR model. Communications in Statistics: Simulation and Computation, 52(3):596–620, 2023. CODEN CSSCDB. ISSN 0361-0918.

Manaa:2023:PNBa

[MB23c] Abderrahmen Manaa and Mohamed Bentarzi. On a periodic negative binomial SETINAR model. Communications in Statistics: Simulation and Computation, 52(3):569–595, 2023. CODEN CSSCDB. ISSN 0361-0918.

Manaa:2023:PNBb

[MB23d] Abderrahmen Manaa and Mohamed Bentarzi. Periodic negative binomial INGARCH(1, 1) model. Communications in Statistics: Simulation and Computation, 52(11):5139–5162, 2023. CODEN CSSCDB. ISSN 0361-0918.

Manaa:2024:ADC

[MB24] Abderrahmen Manaa and Wissam Bentarzi. Asymptotic distribution of CLS estimators in the nearly unstable and unstable PINAR(1) model. Communications in Statistics: Simulation and Computation, 53(1):1–29, 2024. CODEN CSSCDB. ISSN 0361-0918.

Monfared:2022:EPB

[MBBS22] M. Mohammadi Monfared, Reza Arabi Belaghi, M. H. Behzadi, and Sukhdev Singh. Estimation and prediction based on type-I hybrid censored data from the Poisson–Exponential distribution. Communications in Statistics: Simulation and Computation, 51(5):2560–2585, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mefleh:2021:PBB

[MBDK21] Aline Mefleh, Romain Biard, Clément Dombry, and Zaher Khraibani. Permutation bootstrap and the block maxima method. Communications in Statistics: Simulation and Computation, 50(1):295–311, 2021. CODEN CSSCDB. ISSN 0361-0918.

Miljkovic:2022:APC

[MCJ22] Tatjana Miljkovic, Ryan Causey, and Milan Jovanović. Assessing the performance of confidence intervals for high quantiles of Burr XII and Inverse Burr mixtures. Communications in Statistics: Simulation and Computation, 51(8):4677–4699, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ma:2022:ILP

[MCWL22] Ziwei Ma, Ying-Ju Chen, Tonghui Wang, and Jing Liu. Inferences on location parameter in multivariate skew-normal family with unknown scale parameter. Communications in Statistics: Simulation and Computation, 51(9):5465–5481, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mir:2022:PMT

[MD22] Youness Mir and François Dubeau. A piecewise model for two-phase growth phenomena. Communications in Statistics: Simulation and Computation, 51(1):99–115, 2022. CODEN CSSCDB. ISSN 0361-0918.

MeSilva:2022:PMR

[MdN22] W. V. M.e Silva and F. F. do Nascimento. MCMC4Extremes: an R package for Bayesian inference for extremes and its extensions. Communications in Statistics: Simulation and Computation, 51(2):432–442, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mohammadi:2024:TBI

[MEA24] M. Mohammadi, M. Emadi, and M. Amini. Testing bivariate independence based on α-divergence by improved probit transformation method for copula density estimation. Communications in Statistics: Simulation and Computation, 53 (1):499–517, 2024. CODEN CSSCDB. ISSN 0361-0918.

Medvedev:2022:UMV

[Med22] Ilia N. Medvedev. Universal modification of vector weighted method of dependent trials. Communications in Statistics: Simulation and Computation, 51(2):518–526, 2022. CODEN CSSCDB. ISSN 0361-0918.

Megheib:2023:BAN

[Meg23] Mohamed Megheib. A Bayesian approach for nonparametric regression in the presence of correlated errors. Communications in Statistics: Simulation and Computation, 52(12): 6168–6177, 2023. CODEN CSSCDB. ISSN 0361-0918.

Momeni:2021:ENM

[MES21] Reza Momeni, Javad Etminan, and M. Khanjari Sadegh. Estimation of normal means in the tree order model by the weighting methods. Communications in Statistics: Simulation and Computation, 50(1):282–294, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Mehrdoust:2022:CFT

[MF22] Farshid Mehrdoust and Somayeh Fallah. On the calibration of fractional two-factor stochastic volatility model with non-Lipschitz diffusions. Communications in Statistics: Simulation and Computation, 51(11):6332–6351, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mostel:2021:PET

[MFPP21] Linda Möstel, Matthias Fischer, Fabian Pfälzner, and Marius Pfeuffer. Parameter estimation of Tukey-type distributions: a comparative analysis. Communications in Statistics: Simulation and Computation, 50(4):957–992, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mehrdoust:2021:PMA

[MFS21] Farshid Mehrdoust, Somayeh Fallah, and Oldouz Samimi. Pricing multi-asset American option under Heston–CIR diffusion model with jumps. Communications in Statistics: Simulation and Computation, 50(11):3182–3193, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ma:2021:PET

[MG21] Yanbin Ma and Wenhao Gui. Point estimation and two new goodness of fit tests for the scale family based on general progressively Type-II censored samples. Communications in Statistics: Simulation and Computation, 50(12):4187–4214, 2021. CODEN CSSCDB. ISSN 0361-0918.

Marrelec:2024:ECP

[MG24] Guillaume Marrelec and Alain Giron. Estimating the concentration parameter of a von Mises distribution: a systematic simulation benchmark. Communications in Statistics: Simulation and Computation, 53(1):117–129, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Melas:2022:SMC

[MGS22] Viatcheslav B. Melas, Roman Guchenko, and Vladislav Strashko. Standardized maximin criterion for discrimination and parameter estimation of nested models. Communications in Statistics: Simulation and Computation, 51(8):4314–4325, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mosaferi:2023:TFH

[MGS23] Sepideh Mosaferi, Malay Ghosh, and Rebecca C. Steorts. Transformed Fay–Herriot model with measurement error in covariates. Communications in Statistics: Simulation and Computation, 52(5):2257–2274, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mukhopadhyay:2024:SOA

[MHW24] Nitis Mukhopadhyay, Jun Hu, and Zhe Wang. Second-order asymptotics for comparing treatment means from purely sequential estimation strategies under possible outlying observations. Communications in Statistics: Simulation and Computation, 53(3):1330–1355, 2024. CODEN CSSCDB. ISSN 0361-0918.

Milosevic:2020:AEG

[Mil20] Bojana Milosević. Asymptotic efficiency of goodness-of-fit tests based on Too–Lin characterization. Communications in Statistics: Simulation and Computation, 49(8):2082–2101, 2020. CODEN CSSCDB. ISSN 0361-0918.

Miller:2021:MPP

[Mil21] Curtis Miller. Marginal probabilities and point estimation for conditionally specified logistic regression. Communications in Statistics: Simulation and Computation, 50(12):4338–4363, 2021. CODEN CSSCDB. ISSN 0361-0918.

Minh:2022:BNR

[Min22] Do Le Paul Minh. Bayesian networks: regenerative Gibbs samplings. Communications in Statistics: Simulation and Computation, 51(12):7554–7564, 2022. CODEN CSSCDB. ISSN 0361-0918.

Minh:2023:BNG

[Min23] Do Le Paul Minh. Bayesian networks: generating independent samples. Communications in Statistics: Simulation and Computation, 52(12):5739–5748, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mishra:2020:OOS

[Mis20] Nutan Mishra. Optimal one-step censoring schemes under entropy criterion. Communications in Statistics: Simulation and Computation, 49(8):2068–2081, 2020. CODEN CSSCDB. ISSN 0361-0918.

Malekzadeh:2020:IEM

[MJ20] A. Malekzadeh and A. A. Jafari. Inference on the equality means of several two-parameter exponential distributions under progressively Type II censoring. Communications in Statistics: Simulation and Computation, 49(12):3196–3211, 2020. CODEN CSSCDB. ISSN 0361-0918.

Maiti:2021:EPR

[MK21] Kousik Maiti and Suchandan Kayal. Estimation of parameters and reliability characteristics for a generalized Rayleigh distribution under progressive type-II censored sample. Communications in Statistics: Simulation and Computation, 50 (11):3669–3698, 2021. CODEN CSSCDB. ISSN 0361-0918.

Murakami:2022:TSR

[MK22] Hidetoshi Murakami and Soshi Kawada. A test statistic with a ranking method based on the Jeffreys divergence measure. Communications in Statistics: Simulation and Computation, 51(1):266–279, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2022:RND

[MKK22] Hamdy F. F. Mahmoud, Byung-Jun Kim, and Inyoung Kim. Robust nonparametric derivative estimator. Communications in Statistics: Simulation and Computation, 51(7):3809–3829, 2022. CODEN CSSCDB. ISSN 0361-0918.

Muneer:2022:PGF

[MKS22] Siraj Muneer, Alamgir Khalil, and Javid Shabbir. A parentgeneralized family of chain ratio exponential estimators in stratified random sampling using supplementary variables. Communications in Statistics: Simulation and Computation, 51(8):4727–4748, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mim:2022:NRS

[MKSH22] Faijun Nahar Mim, Michael B. C. Khoo, Sajal Saha, and Abdul Haq. New run sum t charts with variable sampling intervals for process mean. Communications in Statistics: Simulation and Computation, 51(9):5350–5372, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mim:2023:SSG

[MKSK23] Faijun Nahar Mim, Michael B. C. Khoo, Sajal Saha, and Khai Wah Khaw. A side-sensitive group runs median control chart with measurement errors. Communications in Statistics: Simulation and Computation, 52(4):1660–1678, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mitra:2022:ECB

[ML22] Sinjini Mitra and Kenny Le. The effect of cognitive and behavioral factors on student success in a bottleneck business statistics course via deeper analytics. Communications in Statistics: Simulation and Computation, 51(6):2779–2808, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mehmood:2022:HCC

[MLR^{+}22] Rashid Mehmood, Muhammed Hisyam Lee, Muhammad Riaz, Baber Zaman, and Iftikhar Ali. Hotelling T^{2} control chart based on bivariate ranked set schemes. Communications in Statistics: Simulation and Computation, 51(1):1–28, 2022. CODEN CSSCDB. ISSN 0361-0918.

Menezes:2020:IML

[MM20] André Felipe Berdusco Menezes and Josmar Mazucheli. Improved maximum likelihood estimators for the parameters of the Johnson S_{B} distribution. Communications in Statistics: Simulation and Computation, 49(6):1511–1526, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

MeloAlbuquerque:2022:PRD

[MM22] Pedro Henrique Melo Albuquerque and Gustavo Monteiro. RMCriteria: a decision making support system package for R. Communications in Statistics: Simulation and Computation, 51(6):3259–3271, 2022. CODEN CSSCDB. ISSN 0361-0918.

Menezes:2023:IML

[MMdOC23] A. F. B. Menezes, J. Mazucheli, R. P. de Oliveira, and S. Chakraborty. Improved maximum likelihood estimation of the parameters of the gamma-uniform distribution with bias-corrections. Communications in Statistics: Simulation and Computation, 52(9):4023–4035, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mabude:2022:DFM

[MMMCS22a] K. Mabude, J.-C. Malela-Majika, P. Castagliola, and S. C. Shongwe. Distribution-free mixed GWMA-CUSUM and CUSUM-GWMA Mann–Whitney charts to monitor unknown shifts in the process location. Communications in Statistics: Simulation and Computation, 51(11):6667–6690, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Motsepa:2022:SSD

[MMMCS22b] Collen M. Motsepa, Jean-Claude Malela-Majika, Philippe Castagliola, and Sandile C. Shongwe. A side-sensitive double sampling X̄ monitoring scheme with estimated process parameters. Communications in Statistics: Simulation and Computation, 51(7):3772–3808, 2022. CODEN CSSCDB. ISSN 0361-0918.

Motsepa:2023:DSM

[MMMCS23] Collen M. Motsepa, Jean-Claude Malela-Majika, Philippe Castagliola, and Sandile C. Shongwe. Double sampling monitoring schemes: a literature review and some future research ideas. Communications in Statistics: Simulation and Computation, 52(8):3391–3419, 2023. CODEN CSSCDB. ISSN 0361-0918.

Malela-Majika:2021:NDS

[MMMG21] J. C. Malela-Majika, C. M. Motsepa, and M. A. Graham. A new double sampling X̄ control chart for monitoring an abrupt change in the process location. Communications in Statistics: Simulation and Computation, 50(3):917–935, 2021. CODEN CSSCDB. ISSN 0361-0918.

Marozzi:2022:CIE

[MMR22] Marco Marozzi and Fernando Marmolejo-Ramos. Composite indicators in experimental psychology. an example with the semantic space of taste and shape stimuli stimuli. Communications in Statistics: Simulation and Computation, 51(1): 154–163, 2022. CODEN CSSCDB. ISSN 0361-0918.

Moya:2023:RBC

[MMVBE23] P. J. Moya, Juan F. Munoz, Encarnación Álvarez Verdejo, and F. J. Blanco-Encomienda. Rescaled bootstrap confidence intervals for the population variance in the presence of outliers or spikes in the distribution of a variable of interest. Communications in Statistics: Simulation and Computation, 52(2): 466–481, 2023. CODEN CSSCDB. ISSN 0361-0918.

Morina:2017:CRS

[MN17] D. Moriña and A. Navarro. Competing risks simulation with the survsim R package. Communications in Statistics: Simulation and Computation, 46(7):5712–5722, 2017. CODEN CSSCDB. ISSN 0361-0918. See [GRR24].

Magnussen:2021:DCM

[MNL21] Steen Magnussen and Thomas Nord-Larsen. Designconsistent model-based variances with systematic sampling: a case study with the Danish national forest inventory. Communications in Statistics: Simulation and Computation, 50 (1):38–48, 2021. CODEN CSSCDB. ISSN 0361-0918.

Maneerat:2021:BCI

[MNN21] Patcharee Maneerat, Sa-Aat Niwitpong, and Suparat Niwitpong. Bayesian confidence intervals for a single mean and the difference between two means of delta-lognormal distributions. Communications in Statistics: Simulation and Computation, 50(10):2906–2934, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mammadova:2023:CDR

[MÖ23] Ulduz Mammadova and M. Revan Özkale. Comparison of deviance and ridge deviance residual-based control charts for monitoring Poisson profiles. Communications in Statistics: Simulation and Computation, 52(3):826–853, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mohammadi:2022:PCA

[Moh22] Mohammad Mohammadi. Principal component analysis for α-stable vectors. Communications in Statistics: Simulation and Computation, 51(9):5245–5263, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mostafaiy:2022:EVC

[Mos22] Behdad Mostafaiy. On estimation in varying coefficient models for sparse and irregularly sampled functional data. Communications in Statistics: Simulation and Computation, 51 (3):1030–1045, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2020:COB

[MPD20] B. N. Mandal, Rajender Parsad, and Sukanta Dash. Construction of A-optimal balanced treatment incomplete block designs: an algorithmic approach. Communications in Statistics: Simulation and Computation, 49(6):1653–1664, 2020. CODEN CSSCDB. ISSN 0361-0918.

Monteiro:2021:BMT

[MPS21] Magda Monteiro, Isabel Pereira, and Manuel G. Scotto. Bivariate models for time series of counts: A comparison study between PBINAR models and dynamic factor models. Communications in Statistics: Simulation and Computation, 50 (7):1873–1887, 2021. CODEN CSSCDB. ISSN 0361-0918.

Marange:2021:NEL

[MQ21] Chioneso Show Marange and Yongsong Qin. A new empirical likelihood ratio goodness of fit test for normality based on moment constraints. Communications in Statistics: Simulation and Computation, 50(6):1561–1575, 2021. CODEN CSSCDB. ISSN 0361-0918.

Morabbi:2022:QEB

[MR22] Hakimeh Morabbi and Mostafa Razmkhah. Quantile estimation based on modified ranked set sampling schemes using Pitman closeness. Communications in Statistics: Simulation and Computation, 51(12):6968–6988, 2022. CODEN CSSCDB. ISSN 0361-0918.

Maksaei:2023:IMO

[MRB23] Najmieh Maksaei, Abdolrahman Rasekh, and Babak Babadi. Influence measures and outliers detection in linear mixed measurement error models with ridge estimation. Communications in Statistics: Simulation and Computation, 52(7):3062– 3078, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mondol:2024:PLA

[MRB24] Momenul Haque Mondol, M. Shafiqur Rahman, and Wasimul Bari. A penalized likelihood approach for dealing with separation in count data regression model. Communications in Statistics: Simulation and Computation, 53(4):1799–1813, 2024. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2022:ELP

[MRM22] M. A. W. Mahmoud, Dina A. Ramadan, and M. M. M. Mansour. Estimation of lifetime parameters of the modified extended exponential distribution with application to a mechanical model. Communications in Statistics: Simulation and Computation, 51(12):7005–7018, 2022. CODEN CSSCDB. ISSN 0361-0918.

Montes-Rojas:2023:DEU

[MRM23] Gabriel Montes-Rojas and Andrés Sebastián Mena. Density estimation using bootstrap quantile variance and quantilemean covariance. Communications in Statistics: Simulation and Computation, 52(4):1449–1461, 2023. CODEN CSSCDB. ISSN 0361-0918.

Mostofi:2020:DSS

[MS20] Amin Ghalamfarsa Mostofi and Sheyda Shirkani. Designing a single-sampling plan for attributes in the presence of classification errors. Communications in Statistics: Simulation and Computation, 49(7):1768–1782, 2020. CODEN CSSCDB. ISSN 0361-0918.

Makalic:2022:EAS

[MS22] Enes Makalic and Daniel F. Schmidt. An efficient algorithm for sampling from sin^{k}(x) for generating random correlation matrices. Communications in Statistics: Simulation and Computation, 51(5):2731–2735, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mukhopadhyay:2023:GET

[MSB23] Partha Mukhopadhyay, Garib Nath Singh, and A. Bandyopadhyay. A general estimation technique of population mean under stratified successive sampling in presence of random scrambled response and non-response. Communications in Statistics: Simulation and Computation, 52(11):5288–5308, 2023. CODEN CSSCDB. ISSN 0361-0918.

Malekzadeh:2021:QBR

[MSL21] Ahad Malekzadeh and Kamyar Sabri-Laghaie. Quantilebased reliability comparison of products: Applied to lognormal distribution. Communications in Statistics: Simulation and Computation, 50(8):2251–2265, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Mota:2024:WLR

[MSNN^{+}24] Alex L. Mota, Manoel Santos-Neto, Milton Miranda Neto, Jeremias Leão, Vera L. D. Tomazella, and Francisco Louzada. Weighted Lindley regression model with varying precision: estimation, modeling and its diagnostics. Communications in Statistics: Simulation and Computation, 53(4):1690–1710, 2024. CODEN CSSCDB. ISSN 0361-0918.

Moretti:2020:PBM

[MSS20] Angelo Moretti, Natalie Shlomo, and Joseph W. Sakshaug. Parametric bootstrap mean squared error of a small area multivariate EBLUP. Communications in Statistics: Simulation and Computation, 49(6):1474–1486, 2020. CODEN CSSCDB. ISSN 0361-0918.

Maboudou-Tchao:2021:HDD

[MT21] Edgard M. Maboudou-Tchao. High-dimensional data monitoring using support machines. Communications in Statistics: Simulation and Computation, 50(7):1927–1942, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Maboudou-Tchao:2023:LSS

[MT23] Edgard M. Maboudou-Tchao. Least-squares support tensor data description. Communications in Statistics: Simulation and Computation, 52(7):3026–3042, 2023. CODEN CSSCDB. ISSN 0361-0918.

Murat:2023:ODS

[Mur23] Naci Murat. Outlier detection in statistical modeling via multivariate adaptive regression splines. Communications in Statistics: Simulation and Computation, 52(7):3379–3390, 2023. CODEN CSSCDB. ISSN 0361-0918.

Morales:2022:EAM

[MV22a] Victor Hugo Morales and Jose Alberto Vargas. The effect of aggregating multivariate count data using Poisson profiles. Communications in Statistics: Simulation and Computation, 51(5):2646–2666, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mutali:2022:CLP

[MV22b] Sezen Mutali and Raluca Vernic. On the composite Lognormal–Pareto distribution with uncertain threshold. Communications in Statistics: Simulation and Computation, 51(8):4492–4508, 2022. CODEN CSSCDB. ISSN 0361-0918.

Melo:2022:HOA

[MVLMA22] Tatiane F. N. Melo, Tiago M. Vargas, Artur J. Lemonte, and Germán Moreno-Arenas. Higher-order asymptotic refinements in the multivariate Dirichlet regression model. Communications in Statistics: Simulation and Computation, 51 (1):53–71, 2022. CODEN CSSCDB. ISSN 0361-0918.

Moretti:2020:ESA

[MW20] Angelo Moretti and Adam Whitworth. Evaluations of small area composite estimators based on the iterative proportional fitting algorithm. Communications in Statistics: Simulation and Computation, 49(12):3093–3110, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Mojiri:2023:NSS

[MWNSB23] A. Mojiri, Y. Waghei, H. R. Nili-Sani, and G. R. Mohtashami Borzadaran. Non-stationary spatial autoregressive modeling for the prediction of lattice data. Communications in Statistics: Simulation and Computation, 52(11):5714–5726, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ma:2023:RIS

[MWTR23] Jin’ge Ma, Liang Wang, Yogesh Mani Tripathi, and Manoj Kumar Rastogi. Reliability inference for stressstrength model based on inverted exponential Rayleigh distribution under progressive Type-II censored data. Communications in Statistics: Simulation and Computation, 52(6): 2388–2407, 2023. CODEN CSSCDB. ISSN 0361-0918.

Miao:2024:JTH

[MX24] Runsheng Miao and Kai Xu. Joint test for homogeneity of high-dimensional means and covariance matrices using maximum-type statistics. Communications in Statistics: Simulation and Computation, 53(2):972–992, 2024. CODEN CSSCDB. ISSN 0361-0918.

Mutlu:2023:DOW

[MY23] Hakan Tahiri Mutlu and Mehmet Selami Yildiz. Determination of the optimal warranty policy and period from the manufacturer’s/seller’s perspective. Communications in Statistics: Simulation and Computation, 52(8):3612–3634, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Maneejuk:2022:EIS

[MYS22] Paravee Maneejuk, Woraphon Yamaka, and Songsak Sriboonchitta. Entropy inference in smooth transition kink regression. Communications in Statistics: Simulation and Computation, 51(12):7366–7389, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mradula:2021:EEP

[MYVD21] Mradula, Subhash Kumar Yadav, Rahul Varshney, and Madhulika Dube. Efficient estimation of population mean under stratified random sampling with linear cost function. Communications in Statistics: Simulation and Computation, 50 (12):4364–4387, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ma:2022:RSD

[MZ22a] Guangzhi Ma and Xuejing Zhao. Regression of survival data via twin support vector regression. Communications in Statistics: Simulation and Computation, 51(9):5126–5138, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mahdizadeh:2022:IEP

[MZ22b] M. Mahdizadeh and Ehsan Zamanzade. On interval estimation of the population mean in ranked set sampling. Communications in Statistics: Simulation and Computation, 51(5): 2747–2768, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ma:2023:QRC

[MZ23] Xuejun Ma and Ping Zhang. Quantile regression for compositional covariates. Communications in Statistics: Simulation and Computation, 52(3):658–668, 2023. CODEN CSSCDB. ISSN 0361-0918.

Makhloufi:2024:FAK

[MZZA24] S. Makhloufi, N. Zougab, Y. Ziane, and S. Adjabi. A family of asymmetric kernels based on log-symmetric distributions. Communications in Statistics: Simulation and Computation, 53(1):380–397, 2024. CODEN CSSCDB. ISSN 0361-0918.

Meng:2021:BET

[MZZF21] Lixin Meng, Jiwei Zhang, Xue Zhang, and Guozhong Feng. Bayesian estimation of time-varying parameters in ordinary differential equation models with noisy time-varying covariates. Communications in Statistics: Simulation and Computation, 50(3):708–723, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nademi:2021:SIP

[Nad21] Arash Nademi. The single-index panel data models with heterogeneous link function: mixture approach. Communications in Statistics: Simulation and Computation, 50(8):2418–2431, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nagakura:2021:CES

[Nag21a] Daisuke Nagakura. Computing exact score vectors for linear Gaussian state space models. Communications in Statistics: Simulation and Computation, 50(8):2313–2326, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Nunez-Antonio:2021:MPG

[NAG21b] Gabriel Núñez-Antonio and Emiliano Geneyro. A multivariate projected gamma model for directional data. Communications in Statistics: Simulation and Computation, 50(9): 2721–2742, 2021. CODEN CSSCDB. ISSN 0361-0918.

Najari:2022:IPL

[NAJ22] Nader Najari, Mohammad Q. Vahidi Asl, and Abdollah Jalilian. Identifying parent locations in the Neyman–Scott process using Delaunay triangulation. Communications in Statistics: Simulation and Computation, 51(10):6058–6070, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Nakayama:2021:RSV

[Nak21] Yugo Nakayama. Robust support vector machine for highdimensional imbalanced data. Communications in Statistics: Simulation and Computation, 50(5):1524–1540, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Nakayama:2022:SVM

[Nak22] Yugo Nakayama. Support vector machine and optimal parameter selection for high-dimensional imbalanced data. Communications in Statistics: Simulation and Computation, 51(11): 6739–6754, 2022. CODEN CSSCDB. ISSN 0361-0918.

Najarzadegan:2020:WBG

[NAKK20] Hossein Najarzadegan, Mohammad Hossein Alamatsaz, Iraj Kazemi, and Debasis Kundu. Weighted bivariate geometric distribution: Simulation and estimation. Communications in Statistics: Simulation and Computation, 49(9):2419–2443, 2020. CODEN CSSCDB. ISSN 0361-0918.

Najera:2023:CSI

[NBC23] Edilberto Nájera and Addy Bolívar-Cimé. Comparison of some interval estimation methods for the parameters of the gamma distribution. Communications in Statistics: Simulation and Computation, 52(6):2408–2424, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Ngwa:2022:GST

[NCC^{+}22] Julius S. Ngwa, Howard J. Cabral, Debbie M. Cheng, David R. Gagnon, Michael P. LaValley, and L. Adrienne Cupples. Generating survival times with time-varying covariates using the Lambert W function. Communications in Statistics: Simulation and Computation, 51(1):135–153, 2022. CODEN CSSCDB. ISSN 0361-0918.

Nascimento:2020:DBH

[NdSA20] Abraão D. C. Nascimento, Raquel C. da Silva, and Getúlio J. A. Amaral. Distance-based hypothesis tests on the Watson distribution. Communications in Statistics: Simulation and Computation, 49(9):2225–2238, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nassar:2024:ELC

[NDWE24] Mazen Nassar, Sanku Dey, Liang Wang, and Ahmed Elshahhat. Estimation of Lindley constant-stress model via product of spacing with Type-II censored accelerated life data. Communications in Statistics: Simulation and Computation, 53 (1):288–314, 2024. CODEN CSSCDB. ISSN 0361-0918.

Nissas:2022:HDD

[NG22] Wassila Nissas and Soufiane Gasmi. A hybrid decision dependent maintenance model of failure rate and virtual age classes using modified Weibull intensity. Communications in Statistics: Simulation and Computation, 51(5):2397–2411, 2022. CODEN CSSCDB. ISSN 0361-0918.

Nemeth:2022:CTD

[NHZ22] László Németh, Zuzana Hübnerová, and András Zempléni. Comparison of trend detection methods in GEV models. Communications in Statistics: Simulation and Computation, 51(11):6398–6413, 2022. CODEN CSSCDB. ISSN 0361-0918.

Nduka:2020:EAR

[NIN20] Uchenna C. Nduka, Iheanyi S. Iwueze, and Elezar C. Nwogu. Efficient algorithms for robust estimation in autoregressive regression models using Student’s t distribution. Communications in Statistics: Simulation and Computation, 49(2): 355–374, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nishida:2021:SMV

[Nis21] Kiheiji Nishida. Skewing methods for variance-stabilizing local linear regression estimation. Communications in Statistics: Simulation and Computation, 50(7):2089–2106, 2021. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2022:EOS

[NJ22] Hadi Alizadeh Noughabi and Jalil Jarrahiferiz. Extropy of order statistics applied to testing symmetry. Communications in Statistics: Simulation and Computation, 51(6):3389–3399, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ng:2022:EES

[NKSY22] Peh Sang Ng, Michael B. C. Khoo, Sajal Saha, and Wai Chung Yeong. Economic and economic-statistical designs of auxiliary information based X̄, synthetic and EWMA charts. Communications in Statistics: Simulation and Computation, 51(3):1157–1185, 2022. CODEN CSSCDB. ISSN 0361-0918.

Novoa-Munoz:2021:GFT

[NM21] Francisco Novoa-Muñoz. Goodness-of-fit tests for the bivariate Poisson distribution. Communications in Statistics: Simulation and Computation, 50(7):1998–2014, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nandy:2020:SPS

[NN20] Karabi Nandy and Rajesh Ranjan Nandy. A study of the performance of 2-stage adaptive optimal designs in a logistic dose-response model. Communications in Statistics: Simulation and Computation, 49(5):1118–1141, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nezhad:2021:DVS

[NN21] Mohammad Saber Fallah Nezhad and Marziyeh Nesaee. Developing variables sampling plans based on EWMA yield index. Communications in Statistics: Simulation and Computation, 50(4):1135–1153, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ngerng:2022:PSA

[NN22] Miang Hong Ngerng and Sherilynn S. F. Ngerng. Portfolio selection algorithm under financial crisis: a case study with Bursa Malaysia. Communications in Statistics: Simulation and Computation, 51(5):2586–2598, 2022. CODEN CSSCDB. ISSN 0361-0918.

Narjis:2020:EPP

[NS20] Ghulam Narjis and Javid Shabbir. Estimation of population proportion and sensitivity level using optional unrelated question randomized response techniques. Communications in Statistics: Simulation and Computation, 49(12):3212–3226, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ndabashinze:2022:COR

[NS22] Barnabe Ndabashinze and Gülesen Üstündag Siray. Comparing ordinary ridge and generalized ridge regression results obtained using genetic algorithms for ridge parameter selection. Communications in Statistics: Simulation and Computation, 51(10):6159–6169, 2022. CODEN CSSCDB. ISSN 0361-0918.

Narjis:2023:ENS

[NS23] Ghulam Narjis and Javid Shabbir. An efficient new scrambled response model for estimating sensitive population mean in successive sampling. Communications in Statistics: Simulation and Computation, 52(11):5327–5344, 2023. CODEN CSSCDB. ISSN 0361-0918.

Nienkemper-Swanepoel:2023:GUV

[NSlRGL23] J. Nienkemper-Swanepoel, N. J. le Roux, and S. Gardner-Lubbe. GPAbin: unifying visualizations of multiple imputations for missing values. Communications in Statistics: Simulation and Computation, 52(6):2666–2685, 2023. CODEN CSSCDB. ISSN 0361-0918.

Nasirzadeh:2021:LFN

[NSN21] R. Nasirzadeh, A. R. Soltani, and A. R. Nematollahi. Longitudinal functional nonlinear marginal mixed effect models. Communications in Statistics: Simulation and Computation, 50(12):4079–4099, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ning:2020:RQR

[NT20] Jianhui Ning and Huiqiang Tao. Randomized quasi-random sampling/importance resampling. Communications in Statistics: Simulation and Computation, 49(12):3367–3379, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nourbakhsh:2020:WET

[NYM20] Mohammadreza Nourbakhsh, Gholamhossein Yari, and Yaser Mehrali. Weighted entropies and their estimations. Communications in Statistics: Simulation and Computation, 49(5): 1142–1158, 2020. CODEN CSSCDB. ISSN 0361-0918.

Noonan:2021:ABC

[NZ21] Jack Noonan and Anatoly Zhigljavsky. Approximations for the boundary crossing probabilities of moving sums of normal random variables. Communications in Statistics: Simulation and Computation, 50(11):3547–3568, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ozdemir:2022:CEL

[ÖA22] Senay Özdemir and Olcay Arslan. Combining empirical likelihood and robust estimation methods for linear regression models. Communications in Statistics: Simulation and Computation, 51(3):941–954, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2023:BSR

[ÖA23] M. Revan Özkale and Hüsniye Altuner. Bootstrap selection of ridge regularization parameter: a comparative study via a simulation study. Communications in Statistics: Simulation and Computation, 52(8):3820–3838, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ouzzani:2021:MPI

[OB21] Fares Ouzzani and Mohamed Bentarzi. On mixture periodic Integer-Valued ARCH models. Communications in Statistics: Simulation and Computation, 50(12):3931–3957, 2021. CODEN CSSCDB. ISSN 0361-0918.

Olmus:2022:PES

[OBN22] Hülya Olmus, Esra Bespinar, and Ezgi Nazman. Performance evaluation of some propensity score matching methods by using binary logistic regression model. Communications in Statistics: Simulation and Computation, 51(4):1647–1660, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ozel:2023:CSE

[ÖÇ23] Sibel Örk Özel and H. Altan Çabuk. Comparisons of the some estimators for the transcendental logarithmic (translog) model*. Communications in Statistics: Simulation and Computation, 52(9):4008–4022, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ozkan:2024:EDP

[ÖDS24] Yüksel Özkan, Mert Demirarslan, and Asli Suner. Effect of data preprocessing on ensemble learning for classification in disease diagnosis. Communications in Statistics: Simulation and Computation, 53(4):1657–1677, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2020:FPM

[ÖK20] M. Revan Özkale and Özge Kuran. A further prediction method in linear mixed models: Liu prediction. Communications in Statistics: Simulation and Computation, 49(12): 3171–3195, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ozbay:2021:RPS

[ÖK21] Nimet Özbay and Selahattin Kaçiranlar. Risk performance of some shrinkage estimators. Communications in Statistics: Simulation and Computation, 50(2):323–342, 2021. CODEN CSSCDB. ISSN 0361-0918.

Osmundsen:2021:MMS

[OKO21] Kjartan Kloster Osmundsen, Tore Selland Kleppe, and Atle Oglend. MCMC for Markov-switching models — Gibbs sampling vs. marginalized likelihood. Communications in Statistics: Simulation and Computation, 50(3):669–690, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Obeidat:2022:BMT

[OLOK22] Mohammed Obeidat, Juxin Liu, Nathaniel Osgood, and Geoff Klassen. Bayesian methods for time series of count data. Communications in Statistics: Simulation and Computation, 51(2):486–504, 2022. CODEN CSSCDB. ISSN 0361-0918.

Orenti:2021:RRA

[OM21] Annalisa Orenti and Ettore Marubini. Robust regression analysis: a useful two stage procedure. Communications in Statistics: Simulation and Computation, 50(1):16–37, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Opperman:2021:SPR

[ON21] Logan Opperman and Wei Ning. Sequential probability ratio test for skew normal distribution. Communications in Statistics: Simulation and Computation, 50(10):2823–2836, 2021. CODEN CSSCDB. ISSN 0361-0918.

Olmus:2022:CPL

[ONE22] Hülya Olmus, Ezgi Nazman, and Semra Erbas. Comparison of penalized logistic regression models for rare event case. Communications in Statistics: Simulation and Computation, 51(4):1578–1590, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ogawa:2022:FMT

[ONS22] Mitsunori Ogawa, Kazuki Nakamoto, and Tomonari Sei. On the fractional moments of a truncated centered multivariate normal distribution. Communications in Statistics: Simulation and Computation, 51(7):3923–3942, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Okorn:2021:LMG

[OO21] Rok Okorn and Matjaz Omladic. Lévy modeled GMWB: Pricing with wavelets. Communications in Statistics: Simulation and Computation, 50(2):413–425, 2021. CODEN CSSCDB. ISSN 0361-0918.

Osborne:2020:DSD

[OP20] Daniel E. Osborne and Vic Patrangenaru. Detecting statistical differences between average glaucoma indices of the eye recovered from virtual 3D reconstructions of the eye’s optic nerve head using 2D stereo images. Communications in Statistics: Simulation and Computation, 49(8):2193–2205, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ozonur:2022:GFT

[OP22] Deniz Ozonur and Sudhir Paul. Goodness of fit tests of the two-parameter gamma distribution against the threeparameter generalized gamma distribution. Communications in Statistics: Simulation and Computation, 51(3):687–697, 2022. CODEN CSSCDB. ISSN 0361-0918.

Opheim:2023:MAA

[OR23] Timothy Opheim and Anuradha Roy. Moving average and autoregressive correlation structures under multivariate skew normality. Communications in Statistics: Simulation and Computation, 52(1):84–97, 2023. CODEN CSSCDB. ISSN 0361-0918.

Orszulik:2024:CFM

[Ors24] Stefan T. Orszulik. Curve fitting: a method of joining piecewise functions to produce models of complex data. Communications in Statistics: Simulation and Computation, 53(4): 2071–2077, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ozbay:2020:EMM

[ÖT20] Nimet Özbay and Selma Toker. Efficiency of Mansson’s method: Some numerical findings about the role of biasing parameter in the estimation of distributed lag model. Communications in Statistics: Simulation and Computation, 49 (9):2333–2346, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ogura:2023:BEM

[OY23] Toru Ogura and Takemi Yanagimoto. Bayesian estimator of multiple Poisson means assuming two different priors. Communications in Statistics: Simulation and Computation, 52 (3):649–657, 2023. CODEN CSSCDB. ISSN 0361-0918.

Oladugba:2024:ROU

[OY24] Abimibola Victoria Oladugba and Brenda Mbouamba Yankam. Robustness of orthogonal uniform composite designs against outlier. Communications in Statistics: Simulation and Computation, 53(4):1729–1743, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2021:RIC

[Özk21] M. Revan Özkale. The red indicator and corrected VIFs in generalized linear models. Communications in Statistics: Simulation and Computation, 50(12):4144–4170, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Pakyari:2023:GFT

[Pak23] Reza Pakyari. Goodness-of-fit testing based on Gini Index of spacings for progressively Type-II censored data. Communications in Statistics: Simulation and Computation, 52(7): 3223–3232, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pan:2020:GCC

[Pan20a] Yingli Pan. Generalized case-cohort analysis for constrained estimation in the Cox’s model. Communications in Statistics: Simulation and Computation, 49(1):45–65, 2020. CODEN CSSCDB. ISSN 0361-0918.

Panahbehagh:2020:SRC

[Pan20b] Bardia Panahbehagh. Stratified and ranked composite sampling. Communications in Statistics: Simulation and Computation, 49(2):504–515, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pant:2020:TPM

[Pan20c] Mohan D. Pant. The t-transformed power method distributions for simulating univariate and multivariate non-normal distributions. Communications in Statistics: Simulation and Computation, 49(4):825–846, 2020. CODEN CSSCDB. ISSN 0361-0918.

Papanastassiou:2020:TSM

[Pap20] Demetrios Papanastassiou. Testing for a single mean with transformed data. Communications in Statistics: Simulation and Computation, 49(2):436–452, 2020. CODEN CSSCDB. ISSN 0361-0918.

Park:2020:SMT

[Par20] Hyo-Il Park. Simultaneous multivariate tests under the normality assumption. Communications in Statistics: Simulation and Computation, 49(7):1886–1897, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Phanthuna:2023:RLD

[PAS23] Piyatida Phanthuna, Yupaporn Areepong, and Saowanit Sukparungsee. Run length distribution for a modified EWMA scheme fitted with a stationary AR(p) model. Communications in Statistics: Simulation and Computation, 52(9):4260– 4279, 2023. CODEN CSSCDB. ISSN 0361-0918.

Piri:2021:WAP

[PASB21] Sepehr Piri, Abdel-Salam G. Abdel-Salam, and Edward L. Boone. A wavelet approach for profile monitoring of Poisson distribution with application. Communications in Statistics: Simulation and Computation, 50(2):525–536, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:DBS

[PAT22] Mustafa Hilmi Pekalp, Halil Aydogdu, and Kamil Feridun Türkman. Discriminating between some lifetime distributions in geometric counting processes. Communications in Statistics: Simulation and Computation, 51(3):715–737, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Pannu:2022:RSF

[PB22] Jasdeep Pannu and Nedret Billor. Robust sparse functional regression model. Communications in Statistics: Simulation and Computation, 51(9):4883–4903, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pokhriyal:2023:TME

[PB23] Himanshu Pokhriyal and N. Balakrishna. Testing for measurement error in regression with autoregressive innovations. Communications in Statistics: Simulation and Computation, 52(5):1834–1848, 2023. CODEN CSSCDB. ISSN 0361-0918.

Palm:2023:PIB

[PBC23] Bruna Gregory Palm, Fábio M. Bayer, and Renato J. Cintra. Prediction intervals in the beta autoregressive moving average model. Communications in Statistics: Simulation and Computation, 52(8):3635–3656, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pandey:2022:CAS

[PBLM22] Arvind Pandey, Shashi Bhushan, Ralte Lalpawimawha, and Praveen Kumar Misra. Comparison of additive shared frailty models under Lindley baseline distribution. Communications in Statistics: Simulation and Computation, 51(3):1121–1136, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pan:2022:BMS

[PC22] Chun Pan and Bo Cai. A Bayesian model for spatial partly interval-censored data. Communications in Statistics: Simulation and Computation, 51(12):7513–7525, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pan:2022:IPB

[PCWL22] Chun Pan, Bo Cai, Lianming Wang, and Xiaoyan Lin. ICBayes: a package of Bayesian semiparametric regression for intervel-censored data. Communications in Statistics: Simulation and Computation, 51(4):1852–1866, 2022. CODEN CSSCDB. ISSN 0361-0918.

Park:2021:THS

[PD21] Junyong Park and Andrei Draganescu. Testing homogeneity of several normal population means based on an interval hypothesis. Communications in Statistics: Simulation and Computation, 50(12):4114–4131, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pan:2022:GCC

[PD22] Yingli Pan and Lifeng Deng. Generalized case-cohort and inference for Cox’s model with parameter constraints. Communications in Statistics: Simulation and Computation, 51 (7):3487–3512, 2022. CODEN CSSCDB. ISSN 0361-0918.

Paul:2023:SSA

[PDK23] Biplab Paul, Shyamal K. De, and Debasis Kundu. A sequential sampling approach for discriminating log-normal, Weibull, and log-logistic distributions. Communications in Statistics: Simulation and Computation, 52(12):5857–5879, 2023. CODEN CSSCDB. ISSN 0361-0918.

Park:2020:IBC

[PDO^{+}20] Chanseok Park, Sanku Dey, Linhan Ouyang, Jai-Hyun Byun, and Mark Leeds. Improved bootstrap confidence intervals for the process capability index C_{pk}. Communications in Statistics: Simulation and Computation, 49(10):2583–2603, 2020. CODEN CSSCDB. ISSN 0361-0918.

Parra-Frutos:2021:RSK

[PFM21] Isabel Parra-Frutos and Lourdes Molera. Removing skewness and kurtosis by transformation when testing for mean equality. Communications in Statistics: Simulation and Computation, 50(9):2598–2618, 2021. CODEN CSSCDB. ISSN 0361-

Pal:2022:ECP

[PG22] Surajit Pal and Susanta Kumar Gauri. Evaluating capability of a process with ordinal responses. Communications in Statistics: Simulation and Computation, 51(6):2871–2887, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pal:2023:MPO

[PG23] Surajit Pal and Susanta Kumar Gauri. Monitoring processes with ordinal data: an area-based approach. Communications in Statistics: Simulation and Computation, 52(4):1361–1383, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pakbaz:2022:MAG

[PHN22] F. Pakbaz, F. Hosseini, and A. R. Nematollahi. Modeling additive genetic effects in animal models by closed skew normal distribution. Communications in Statistics: Simulation and Computation, 51(3):1186–1198, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pho:2024:GFT

[Pho24] Kim-Hung Pho. Goodness of fit test for a zero-inflated Bernoulli regression model. Communications in Statistics: Simulation and Computation, 53(2):756–771, 2024. CODEN CSSCDB. ISSN 0361-0918.

Pham:2021:DBS

[PHS21] Thanh Pham, Trung Ha, and Julia N. Soulakova. Designbased single-mediator approach for complex survey data. Communications in Statistics: Simulation and Computation, 50(3):822–831, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pakgohar:2020:GFT

[PHY20] A. Pakgohar, A. Habibirad, and F. Yousefzadeh. Goodness of fit test using Lin-Wong divergence based on Type-I censored data. Communications in Statistics: Simulation and Computation, 49(9):2485–2504, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:SID

[PIA22] Mustafa Hilmi Pekalp, Gültaç Eroglu Inan, and Halil Aydogdu. Statistical inference for doubly geometric process with Weibull interarrival times. Communications in Statistics: Simulation and Computation, 51(6):3428–3440, 2022. CODEN CSSCDB. ISSN 0361-0918.

Picinbono:2021:CFB

[Pic21] Bernard Picinbono. Coincidence functions and Bartlett spectra of point processes. Communications in Statistics: Simulation and Computation, 50(9):2581–2597, 2021. CODEN CSSCDB. ISSN 0361-0918.

Phrueksawatnon:2021:DOR

[PJL21] Piyada Phrueksawatnon, Jirawan Jitthavech, and Vichit Lorchirachoonkul. Determining the optimal ridge parameter in logistic regression. Communications in Statistics: Simulation and Computation, 50(11):3569–3580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Park:2021:SKS

[PK21] Heewon Park and Sadanori Konishi. Sparse kernel subspace method for classifying and representing patterns from data with complex structure. Communications in Statistics: Simulation and Computation, 50(10):3088–3104, 2021. CODEN CSSCDB. ISSN 0361-0918.

Park:2022:EBB

[PK22] Junhui Park and Seung-Ho Kang. Effects of between-batch variability on the type I error rate in biosimilar development. Communications in Statistics: Simulation and Computation, 51(1):323–340, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:EMV

[PKA22] Mustafa Hilmi Pekalp, Melike Özlem Karaduman, and Halil Aydogdu. Estimation of the mean value function for gamma trend renewal process. Communications in Statistics: Simulation and Computation, 51(6):3441–3456, 2022. CODEN CSSCDB. ISSN 0361-0918.

Park:2022:IFS

[PKW22] Chanseok Park, Haewon Kim, and Min Wang. Investigation of finite-sample properties of robust location and scale estimators. Communications in Statistics: Simulation and Computation, 51(5):2619–2645, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pandey:2022:CCF

[PL22] Arvind Pandey and Ralte Lalpawimawha. Comparison of correlated frailty models. Communications in Statistics: Simulation and Computation, 51(9):5387–5405, 2022. CODEN CSSCDB. ISSN 0361-0918.

Phukongtong:2022:PPP

[PLA22] Siwaporn Phukongtong, Supranee Lisawadi, and S. Ejaz Ahmed. Penalty, post pretest and shrinkage strategies in a partially linear model. Communications in Statistics: Simulation and Computation, 51(10):6004–6025, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pal:2024:ATS

[PLBP24] Subhamoy Pal, Rong Luo, Subhash Bagui, and Sudhir Paul. Alternative tests for the significance of the intraclass correlation coefficients under unequal family sizes. Communications in Statistics: Simulation and Computation, 53(1):409– 426, 2024. CODEN CSSCDB. ISSN 0361-0918.

Pan:2023:WER

[PLS23] Yingli Pan, Zhan Liu, and Guangyu Song. Weighted expectile regression with covariates missing at random. Communications in Statistics: Simulation and Computation, 52(3): 1057–1076, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pasha:2020:GVB

[PM20] M. A. Pasha and M. Bameni Moghadam. A generalized version of Ben–Daya–Rahim (2000) and Rahim–Banerjee (1993) cost models in economic design of X-control charts for systems with early replacement and preventive maintenance under decreasing integrated hazard. Communications in Statistics: Simulation and Computation, 49(1):178–193, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Padgett:2023:ERE

[PM23] R. Noah Padgett and Grant B. Morgan. Evaluating the relative efficiency among robust estimation methods for multilevel factor analysis with categorical data. Communications in Statistics: Simulation and Computation, 52(12):6070–6083, 2023. CODEN CSSCDB. ISSN 0361-0918.

Prajapati:2020:NDT

[PMK20] Deepak Prajapati, Sharmistha Mitra, and Debasis Kundu. A new decision theoretic sampling plan for exponential distribution under Type-I censoring. Communications in Statistics:

Simulation and Computation, 49(2):453–471, 2020. CODEN CSSCDB. ISSN 0361-0918.

Prajapat:2023:CME

[PMK23a] Kiran Prajapat, Sharmishtha Mitra, and Debasis Kundu. A consistent method of estimation for three-parameter generalized exponential distribution. Communications in Statistics: Simulation and Computation, 52(6):2471–2487, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Prajapati:2023:BSP

[PMK23b] Deepak Prajapati, Sharmishtha Mitra, and Debasis Kundu. Bayesian sampling plan for the exponential distribution with generalized Type-II hybrid censoring scheme. Communications in Statistics: Simulation and Computation, 52(2):533– 556, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pal:2023:ODE

[PMM23] Manisha Pal, Nripes Kumar Mandal (Retd.), and Hare Krishna Maity. D-optimal design for estimation of optimum mixture in a three-component mixture experiment with two responses. Communications in Statistics: Simulation and Computation, 52(5):2012–2021, 2023. CODEN CSSCDB. ISSN 0361-0918.

Patummasut:2021:REP

[PN21] Mena Patummasut and Sudarat Nidsunkid. A ratio estimator in path sampling. Communications in Statistics: Simulation and Computation, 50(2):376–387, 2021. CODEN CSSCDB. ISSN 0361-0918.

Prataviera:2023:EMS

[POC23] Fábio Prataviera, Edwin M. M. Ortega, and Gauss M. Cordeiro. An extended Maxwell semiparametric regression for censored and uncensored data. Communications in Statistics: Simulation and Computation, 52(7):3305–3326, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Prataviera:2022:EPE

[POCC22] Fábio Prataviera, Edwin M. M. Ortega, Gauss M. Cordeiro, and Vicente G. Cancho. The exponentiated power exponential semiparametric regression model. Communications in Statistics: Simulation and Computation, 51(10):5933–5953, 2022. CODEN CSSCDB. ISSN 0361-0918.

Polat:2020:EDW

[Pol20] Esra Polat. The effects of different weight functions on partial robust M-regression performance: a simulation study. Communications in Statistics: Simulation and Computation, 49 (4):1089–1104, 2020. CODEN CSSCDB. ISSN 0361-0918.

Polajnar:2022:UEN

[Pol22] Emil Polajnar. Using elastic net restricted kernel canonical correlation analysis for cross-language information retrieval. Communications in Statistics: Simulation and Computation, 51(6):2924–2941, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pal:2023:MRS

[PP23] Sanghamitra Pal and Dipika Patra. Modifications on rescaling bootstrap for adaptive sampling. Communications in Statistics: Simulation and Computation, 52(4):1607–1620, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pal:2022:NNL

[PR22] Suvra Pal and Souvik Roy. A new non-linear conjugate gradient algorithm for destructive cure rate model and a simulation study: illustration with negative binomial competing risks. Communications in Statistics: Simulation and Computation, 51(11):6866–6880, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pietrosanu:2021:HMB

[PRH21] Matthew Pietrosanu, Rhonda J. Rosychuk, and X. Joan Hu. Handling missing birthdates in marginal regression analysis with recurrent events. Communications in Statistics: Simulation and Computation, 50(1):142–152, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Perveen:2023:BLE

[PS23] Ismat Perveen and Muhammad Suhail. Bootstrap Liu estimators for Poisson regression model. Communications in Statistics: Simulation and Computation, 52(7):2811–2821, 2023. CODEN CSSCDB. ISSN 0361-0918.

Pan:2023:SIC

[PSL23] Yingli Pan, Guangyu Song, and Zhan Liu. Statistical inference for case-cohort design under the additive hazards model with covariate adjustment. Communications in Statistics:

Simulation and Computation, 52(10):4575–4591, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Panza:2020:MWR

[PV20a] Carlos Arturo Panza and Jose Alberto Vargas. Monitoring Weibull regression models with type I right-censored observations in phase II processes. Communications in Statistics: Simulation and Computation, 49(2):335–354, 2020. CODEN CSSCDB. ISSN 0361-0918.

Psaradakis:2020:NTD

[PV20b] Zacharias Psaradakis and Marián Vávra. Normality tests for dependent data: large-sample and bootstrap approaches. Communications in Statistics: Simulation and Computation, 49(2):283–304, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pathak:2022:BGL

[PV22] A. K. Pathak and P. Vellaisamy. A bivariate generalized linear exponential distribution: properties and estimation. Communications in Statistics: Simulation and Computation, 51(9): 5426–5446, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pan:2020:IAI

[PY20] Tianshu Pan and Yue Yin. Improving the accuracy of identifying the lognormal curve in the Johnson system. Communications in Statistics: Simulation and Computation, 49(5): 1321–1332, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pourjafar:2020:EMR

[PZ20] Hojjat Pourjafar and Vali Zardasht. Estimation of the mean residual life function in the presence of measurement errors. Communications in Statistics: Simulation and Computation, 49(2):532–555, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pei:2024:MLE

[PZ24] Jian Pei and Fukang Zhu. Marginal likelihood estimation for the negative binomial INGARCH model. Communications in Statistics: Simulation and Computation, 53(4):1814–1823, 2024. CODEN CSSCDB. ISSN 0361-0918.

Qu:2021:CDE

[QL21] Leming Qu and Yang Lu. Copula density estimation by finite mixture of parametric copula densities. Communications in Statistics: Simulation and Computation, 50(11):3315–3337, 2021. CODEN CSSCDB. ISSN 0361-0918.

Qiu:2023:CIP

[QLG23] Shi-Fang Qiu, Qing-Song Liu, and Yang Ge. Confidence intervals of proportion differences for stratified combined unilateral and bilateral data. Communications in Statistics: Simulation and Computation, 52(8):3839–3862, 2023. CODEN CSSCDB. ISSN 0361-0918.

Qureshi:2024:MTR

[QTH24] Muhammad Nouman Qureshi, Muhammad Umair Tariq, and Muhammad Hanif. Memory-type ratio and product estimators for population variance using exponentially weighted moving averages for time-scaled surveys. Communications in Statistics: Simulation and Computation, 53(3):1484–1493, 2024. CODEN CSSCDB. ISSN 0361-0918.

Qin:2022:RDC

[QYC22] Ruibing Qin, Xiaoqin Yang, and Zhanshou Chen. Ratio detections for change point in heavy tailed observations. Communications in Statistics: Simulation and Computation, 51 (5):2487–2510, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rashid:2022:BEB

[RAA22] Fareeha Rashid, Saima Altaf, and Muhammad Aslam. Bayesian estimation of the biasing parameter for ridge regression: a novel approach. Communications in Statistics: Simulation and Computation, 51(12):7215–7225, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ramzan:2023:RPE

[RAA23] Qasim Ramzan, Muhammad Nauman Akram, and Muhammad Amin. Ridge parameter estimation for the linear regression model under different loss functions using T-K approximation. Communications in Statistics: Simulation and Computation, 52(9):4361–4377, 2023. CODEN CSSCDB. ISSN 0361-0918.

Rahimi:2021:SMM

[RAG21] Sajede Baratian Rahimi, Amirhossein Amiri, and Reza Ghashghaei. Simultaneous monitoring of mean vector and covariance matrix of multivariate simple linear profiles in the presence of within profile autocorrelation. Communications in Statistics: Simulation and Computation, 50(6):1791–1808, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rao:2021:VSP

[RAJ21] G. Srinivasa Rao, Muhammad Aslam, and Chi-Hyuck Jun. A variable sampling plan using generalized multiple dependent state based on a one-sided process capability index. Communications in Statistics: Simulation and Computation, 50(9): 2666–2677, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rafiei:2023:MOD

[RAN23] Navid Rafiei, Shervin Asadzadeh, and Seyed Taghi Akhavan Niaki. Multi-objective design of risk-adjusted control chart in healthcare systems with economic and statistical considerations. Communications in Statistics: Simulation and Computation, 52(7):2967–2984, 2023. CODEN CSSCDB. ISSN 0361-0918.

Rasekhi:2020:SME

[Ras20] Mahdi Rasekhi. A study on methods for estimating the PDF and the CDF in the exponentiated gamma distribution. Communications in Statistics: Simulation and Computation, 49 (8):1999–2013, 2020. CODEN CSSCDB. ISSN 0361-0918.

Roozbeh:2023:NMI

[RBKA23] Mahdi Roozbeh, Saman Babaie-Kafaki, and Zohre Aminifard. A nonlinear mixed-integer programming approach for variable selection in linear regression model. Communications in Statistics: Simulation and Computation, 52(11):5434–5445, 2023. CODEN CSSCDB. ISSN 0361-0918.

Regis:2022:LMM

[RBN^{+}22] Marta Regis, Alberto Brini, Nazanin Nooraee, Reinder Haakma, and Edwin R. van den Heuvel. The t linear mixed model: model formulation, identifiability and estimation. Communications in Statistics: Simulation and Computation, 51(5):2318–2342, 2022. CODEN CSSCDB. ISSN 0361-0918.

Raqab:2021:IEH

[RBRAa21] Mohammad Z. Raqab, Omar M. Bdair, Manoj K. Rastogi, and Fahad M. Al-aboud. Inference for an exponentiated half logistic distribution with application to cancer hybrid censored data. Communications in Statistics: Simulation and Computation, 50(4):1178–1201, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rasheed:2022:EMC

[RBT^{+}22] Kashif Rasheed, Zahid Bashir, M. H. Tahir, Farrukh Shehzad, and Rashid Ahmed. Efficient minimal circular weakly balanced repeated measurements designs and their conversion into efficient minimal circular balanced and strongly balanced. Communications in Statistics: Simulation and Computation, 51(10):6197–6207, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rivas:2023:ZIW

[RC23] Luisa Rivas and Francisca Campos. Zero inflated Waring distribution. Communications in Statistics: Simulation and Computation, 52(8):3676–3691, 2023. CODEN CSSCDB. ISSN 0361-0918.

RodriguesLiska:2022:MLB

[RCSS22] Gilberto Rodrigues Liska, Marcelo Ângelo Cirillo, Fortunato Silva de Menezes, and Julio Silvio de Sousa Bueno Filho. Machine learning based on extended generalized linear model applied in mixture experiments. Communications in Statistics: Simulation and Computation, 51(5):2511–2525, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Rodrigues:2020:BSP

[RdAIS^{+}20] Josemar Rodrigues, Marco Henrique de Almeida Inacio, Adriano K. Suzuki, Fernando Raimundo da Silva, and Narayanaswamy Balakrishnan. Bayesian superposition of pure-birth destructive cure processes for tumor latency. Communications in Statistics: Simulation and Computation, 49 (12):3240–3253, 2020. CODEN CSSCDB. ISSN 0361-0918.

Reybod:2022:GPM

[REMM22] Arman Reybod, Javad Etminan, Rahim Moineddin, and Adel Mohammadpour. The generalized Pitman measure of similarity and hierarchical clustering. Communications in Statistics: Simulation and Computation, 51(9):5190–5201, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Rezazadeh:2022:VSF

[REMO22] H. Rezazadeh, F. Eskandari, M. Bameni Moghadam, and E. Ormoz. Variable selection in finite mixture of generalized estimating equations. Communications in Statistics: Simulation and Computation, 51(6):3237–3251, 2022. CODEN CSSCDB. ISSN 0361-0918.

Roshanbin:2022:MOE

[REN22] N. Roshanbin, M. J. Ershadi, and S. T. A. Niaki. Multiobjective economic-statistical design of simple linear profiles using a combination of NSGA-II, RSM, and TOPSIS. Communications in Statistics: Simulation and Computation, 51 (4):1704–1720, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rana:2022:ALO

[RftADNI22] Subrata Rana and for the Alzheimer’s Disease Neuroimaging Initiative. Analysis of longitudinal ordinal data using semi-parametric mixed model under missingness. Communications in Statistics: Simulation and Computation, 51(10): 5631–5642, 2022. CODEN CSSCDB. ISSN 0361-0918.

Razmkhah:2021:EDR

[RGA21] Mansooreh Razmkhah, Bahram Sadeghpour Gildeh, and Jafar Ahmadi. An economic design of rectifying single sampling plans via maxima nomination sampling in the presence of inspection errors. Communications in Statistics: Simulation and Computation, 50(1):217–233, 2021. CODEN CSSCDB. ISSN 0361-0918.

Razzak:2023:AIH

[RH23] Humera Razzak and Christian Heumann. Application of iterative hybrid MI approach to household survey data with complex dependence structures. Communications in Statistics: Simulation and Computation, 52(1):229–246, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Ran:2022:ECE

[RHG22] Liao Ran, Tianle Hu, and Sujuan Gao. Estimation of covariate effects in proportional cross-ratio model of bivariate time-to-event outcomes. Communications in Statistics: Simulation and Computation, 51(12):7472–7486, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ricker:2021:SLR

[Ric21] Martin Ricker. Solving linear regression without skewness of the residuals’ distribution. Communications in Statistics:

Simulation and Computation, 50(8):2482–2495, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Rajesh:2023:SKE

[RJJ23] G. Rajesh, Linto Joseph, and T. M. Jacob. Smoothed kernel estimation of bivariate residual entropy function. Communications in Statistics: Simulation and Computation, 52(9): 4065–4074, 2023. CODEN CSSCDB. ISSN 0361-0918.

Roozegar:2020:ACB

[RK20] Rasool Roozegar and Debasis Kundu. On absolutely continuous bivariate generalized exponential power series distribution. Communications in Statistics: Simulation and Computation, 49(7):1678–1703, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ramos:2021:NEG

[RL21] Pedro Luiz Ramos and Francisco Louzada. A note on the exponential geometric power series distribution. Communications in Statistics: Simulation and Computation, 50(11): 3794–3798, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ren:2023:SCI

[RLP23] Pengcheng Ren, Guanfu Liu, and Xiaolong Pu. Simultaneous confidence intervals for mean differences of multiple zeroinflated gamma distributions with applications to precipitation. Communications in Statistics: Simulation and Computation, 52(10):4705–4716, 2023. CODEN CSSCDB. ISSN 0361-0918.

Rodrigues:2020:NAV

[RM20] Paulo Canas Rodrigues and Rahim Mahmoudvand. A new approach for the vector forecast algorithm in singular spectrum analysis. Communications in Statistics: Simulation and Computation, 49(3):591–605, 2020. CODEN CSSCDB. ISSN 0361-0918.

Richter:2023:RBS

[RM23] Scott J. Richter and Melinda. H. McCann. Resampling-based simultaneous confidence intervals to compare scale using deviances. Communications in Statistics: Simulation and Computation, 52(7):3146–3155, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ramires:2020:FBM

[ROL^{+}20] Thiago G. Ramires, Edwin M. M. Ortega, Artur J. Lemonte, Niel Hens, and Gauss M. Cordeiro. A flexible bimodal model with long-term survivors and different regression structures. Communications in Statistics: Simulation and Computation, 49(10):2639–2660, 2020. CODEN CSSCDB. ISSN 0361-0918.

Riahi:2023:TCS

[RP23] S. Riahi and P. N. Patil. Testing for central symmetry and symmetry about an axis. Communications in Statistics: Simulation and Computation, 52(3):717–734, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Rai:2023:AMD

[RPT23] Himanshu Rai, M. S. Panwar, and Sanjeev K. Tomer. Analysis of masked data with Lindley failure model. Communications in Statistics: Simulation and Computation, 52(4): 1192–1211, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ramanathan:2020:SFR

[RRA20] T. V. Ramanathan, Neelabh Rohan, and Bovas Abraham. A stochastic frontier regression model with dynamic frontier. Communications in Statistics: Simulation and Computation, 49(6):1415–1428, 2020. CODEN CSSCDB. ISSN 0361-0918.

Rezaee:2024:GCJ

[RS24] Fateme Rezaee and Ehsan Bahrami Samani. Gaussian copula joint models for mixed longitudinal zero-inflated count and continuous responses. Communications in Statistics: Simulation and Computation, 53(2):853–868, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Razie:2021:AML

[RSG21] Farzaneh Razie, Ehsan Bahrami Samani, and Mojtaba Ganjali. Analysis of mixed longitudinal (k, l)-inflated power series, ordinal and continuous responses with sensitivity analysis to non-ignorable missing mechanism. Communications in Statistics: Simulation and Computation, 50(8):2286–2312, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rezaee:2022:GCB

[RSG22] Fatemeh Rezaee, Ehsan Bahrami Samani, and Mojtaba Ganjali. Gaussian copula-based zero-inflated power series joint models to analyze correlated count data. Communications in Statistics: Simulation and Computation, 51(10):6136–6148, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rad:2022:EII

[RSM22] Najmeh Nakhaei Rad, Mahdi Salehi, and Yaser Mehrali. Estimation of the inequality indices based on the well-known rank-based sampling schemes. Communications in Statistics: Simulation and Computation, 51(9):5308–5322, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Rasouli:2022:MBA

[RSS22] Mohammad Rasouli, Hamid Shahriari, and Yaser Samimi. Monitoring a bi-attribute high-quality process using a mixture probability distribution. Communications in Statistics: Simulation and Computation, 51(12):7427–7443, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Riffel:2024:DSU

[RTJ24] Vinicius Ricardo Riffel, Cesar Augusto Taconeli, and Paulo Justiniano Ribeiro Junior. On distribution selection under ranked set sampling schemes. Communications in Statistics: Simulation and Computation, 53(1):546–566, 2024. CODEN CSSCDB. ISSN 0361-0918.

Riaz:2022:ISM

[RuA22] Afshan Riaz and Muhammad Noor ul Amin. Improved simultaneous monitoring of mean and coefficient of variation under ranked set sampling schemes. Communications in Statistics: Simulation and Computation, 51(8):4410–4426, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ruzankin:2022:CNM

[Ruz22] Pavel S. Ruzankin. A class of nonparametric mode estimators. Communications in Statistics: Simulation and Computation, 51(6):3291–3304, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rodriguez:2022:REP

[RVV22] Daniela Rodríguez, Marina Valdora, and Pablo Vena. Robust estimation in partially linear regression models with monotonicity constraints. Communications in Statistics: Simulation and Computation, 51(4):2039–2052, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Rytova:2021:SAP

[RY21] Anastasiia Rytova and Elena Yarovaya. Survival analysis of particle populations in branching random walks. Communications in Statistics: Simulation and Computation, 50(10): 3031–3045, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rezaei:2022:SSR

[RY22] Amir Rezaei and Fatemeh Yousefzadeh. Stress-strength reliability of generalized skew-elliptical distributions and its Bayes estimation. Communications in Statistics: Simulation and Computation, 51(7):3471–3486, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rahman:2020:RCC

[RYA20] Ayu Abdul Rahman, Sharipah Soaad Syed Yahaya, and Abdu Mohammed Ali Atta. Robustification of CUSUM control structure for monitoring location shift of skewed distributions based on modified one-step M-estimator. Communications in Statistics: Simulation and Computation, 49(11):3001–3018, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shabbir:2022:EIR

[SA22] Javid Shabbir and Aneel Ahmed. Estimation of interquartile range in stratified sampling under non-linear cost function. Communications in Statistics: Simulation and Computation, 51(4):1891–1898, 2022. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2021:MEC

[SAA21] Aamir Saghir, Liaquat Ahmad, and Muhammad Aslam. Modified EWMA control chart for transformed gamma data. Communications in Statistics: Simulation and Computation, 50(10):3046–3059, 2021. CODEN CSSCDB. ISSN 0361-0918.

Shahzad:2023:NFV

[SAA^{+}23] Usman Shahzad, Ishfaq Ahmad, Ibrahim M. Almanjahie, Nadia H. Al-Noor, and Muhammad Hanif. A novel family of variance estimators based on L-moments and calibration approach under stratified random sampling. Communications in Statistics: Simulation and Computation, 52(8):3782–3795, 2023. CODEN CSSCDB. ISSN 0361-0918.

Stephenson:2021:DLP

[SAD^{+}21] Matthew Stephenson, R. Ayesha Ali, Gerarda A. Darlington, Flavio S. Schenkel, and E. James Squires. DSLRIG: Leveraging predictor structure in logistic regression. Communications in Statistics: Simulation and Computation, 50(6):1600–1612, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sadiq:2024:MSR

[Sad24] Maryam Sadiq. Modeling survival response using a parametric approach in the presence of multicollinearity. Communications in Statistics: Simulation and Computation, 53(4): 1889–1898, 2024. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2021:EMD

[SAF21] Aamir Saghir, Saddam Akber Abbasi, and Alireza Faraz. The exact method for designing the Maxwell chart with estimated parameter. Communications in Statistics: Simulation and Computation, 50(1):270–281, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sarhan:2024:NBL

[SAK24] Ammar M. Sarhan, Joseph Apaloo, and Debasis Kundu. A new bivariate lifetime distribution: properties, estimations and its extension. Communications in Statistics: Simulation and Computation, 53(2):879–896, 2024. CODEN CSSCDB. ISSN 0361-0918.

Salha:2023:LKE

[Sal23] Raid B. Salha. Lognormal kernel estimator of the hazard rate function. Communications in Statistics: Simulation and Computation, 52(9):3991–4007, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shahzad:2022:IBM

[SANH^{+}22] Usman Shahzad, Nadia H. Al-Noor, Muhammad Hanif, Irsa Sajjad, and Malik Muhammad Anas. Imputation based mean estimators in case of missing data utilizing robust regression and variance-covariance matrices. Communications in Statistics: Simulation and Computation, 51(8):4276–4295, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sadoun:2020:EEP

[SB20] Mohamed Sadoun and Mohamed Bentarzi. Efficient estimation in periodic INAR(1) model: parametric case. Communications in Statistics: Simulation and Computation, 49(8): 2014–2034, 2020. CODEN CSSCDB. ISSN 0361-0918.

Santos:2021:RIS

[SB21a] Julius Santos and Erniel Barrios. Robust inference in semiparametric spatial-temporal models. Communications in Statistics: Simulation and Computation, 50(8):2266–2285, 2021. CODEN CSSCDB. ISSN 0361-0918.

Shirke:2021:VSI

[SB21b] Digambar T. Shirke and Mahesh S. Barale. A variable sampling interval sign chart for variability based on deciles. Communications in Statistics: Simulation and Computation, 50 (5):1484–1495, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sadoun:2022:PEM

[SB22] Mohamed Sadoun and Mohamed Bentarzi. On periodic EGARCH models. Communications in Statistics: Simulation and Computation, 51(7):3733–3759, 2022. CODEN CSSCDB. ISSN 0361-0918.

Seba:2024:SLA

[SB24] Djillali Seba and Karima Belaide. On several local asymptotic properties for fractional autoregressive models with strong mixing noises. Communications in Statistics: Simulation and Computation, 53(4):1744–1757, 2024. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:FLT

[SBB22] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. Formulation of logarithmic type estimators to estimate population mean in successive sampling in presence of random non response and measurement errors. Communications in Statistics: Simulation and Computation, 51(3):901–923, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Singh:2023:GCC

[SBB23] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. A general class of calibration estimators under stratified random sampling in presence of various kinds of non-sampling errors. Communications in Statistics: Simulation and Computation, 52(2):320–333, 2023. CODEN CSSCDB. ISSN 0361-0918.

Singh:2024:NRR

[SBB24] Garib Nath Singh, Diya Bhattacharyya, and Arnab Bandyopadhyay. A new randomized response technique with application to election polling. Communications in Statistics: Simulation and Computation, 53(1):94–116, 2024. CODEN CSSCDB. ISSN 0361-0918.

Sagan:2023:IRT

[SBR^{+}23] Adam Sagan, Justyna Brzezińska, Aneta Rybicka, Mirosława Sztemberg-Lewandowska, and Marcin Pełka. Item response theory network analysis of European universities. Communications in Statistics: Simulation and Computation, 52(8): 3593–3611, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sen:2020:IOL

[SBTP20] Tanmay Sen, Ritwik Bhattacharya, Yogesh Mani Tripathi, and Biswabrata Pradhan. Inference and optimum life testing plans based on Type-II progressive hybrid censored generalized exponential data. Communications in Statistics: Simulation and Computation, 49(12):3254–3282, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sun:2022:MSA

[SBZG22] Li Sun, Narayanaswamy Balakrishnan, Fang-Chao Zhao, and Xiao-Hui Gu. Mis-specification analysis of the impact of covariates on the diffusion coefficient in Wiener degradation process. Communications in Statistics: Simulation and Computation, 51(6):3204–3222, 2022. CODEN CSSCDB. ISSN 0361-0918.

Suhail:2023:NQB

[SCA23] Muhammad Suhail, Sohail Chand, and Muhammad Aslam. New quantile based ridge M-estimator for linear regression models with multicollinearity and outliers. Communications in Statistics: Simulation and Computation, 52(4):1417–1434, 2023. CODEN CSSCDB. ISSN 0361-0918.

Senturk:2020:ICM

[SCE^{+}20] Damla Sentürk, Yanjun Chen, Jason P. Estes, Luis F. Campos, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Danh V. Nguyen. Impact of case-mix measurement error on estimation and inference in profiling of health care providers. Communications in Statistics: Simulation and Computation, 49(8): 2206–2224, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shah:2022:MDS

[SCHS22] Said Farooq Shah, Salman Arif Cheema, Zawar Hussain, and Ejaz Ali Shah. Masking data: a solution to social desirability bias in paired comparison experiments. Communications in Statistics: Simulation and Computation, 51(6):3149–3167, 2022. CODEN CSSCDB. ISSN 0361-0918.

Suhail:2020:QBE

[SCK20] Muhammad Suhail, Sohail Chand, and B. M. Golam Kibria. Quantile based estimation of biasing parameters in ridge regression model. Communications in Statistics: Simulation and Computation, 49(10):2732–2744, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Suhail:2021:QBR

[SCK21] Muhammad Suhail, Sohail Chand, and B. M. Golam Kibria. Quantile-based robust ridge M-estimator for linear regression model in presence of multicollinearity and outliers. Communications in Statistics: Simulation and Computation, 50(11): 3194–3206, 2021. CODEN CSSCDB. ISSN 0361-0918.

Strait:2022:PTS

[SCK22] Justin Strait, Oksana Chkrebtii, and Sebastian Kurtek. Parallel tempering strategies for model-based landmark detection on shapes. Communications in Statistics: Simulation and Computation, 51(4):1415–1435, 2022. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:2023:FSP

[SCK23] Aradhana Srivastava, Anoop Chaturvedi, and Nirpeksh Kumar. Finite sample performance of an estimator of process capability index C_{pm} for the autocorrelated data. Communications in Statistics: Simulation and Computation, 52(9): 4348–4360, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shirazi:2022:NEQ

[SD22] Esmaeil Shirazi and Hassan Doosti. Nonparametric estimation of a quantile density function under L_{p} risk via block thresholding method. Communications in Statistics: Simulation and Computation, 51(2):539–553, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Sawadogo:2022:SEM

[SDG22] Amadou Sawadogo and Simplice Dossou-Gbété. Simulation of an extension of Mallows–Bradley–Terry ranking model by acceptance-rejection method. Communications in Statistics: Simulation and Computation, 51(9):5530–5538, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Safari:2023:ETE

[SEMR23] Malihe Safari, Habib Esmaeili, Hossein Mahjub, and Ghodratollah Roshanaei. Estimation of treatment effect in presence of noncompliance with early or late switching: a simulation study. Communications in Statistics: Simulation and Computation, 52(10):4802–4815, 2023. CODEN CSSCDB. ISSN 0361-0918.

Seo:2022:OBA

[Seo22] Jung In Seo. Objective Bayesian analysis for the Weibull distribution with partial information under the generalized type-II progressive hybrid censoring scheme. Communications in Statistics: Simulation and Computation, 51(9):5157–5173, 2022. CODEN CSSCDB. ISSN 0361-0918.

Staggs:2023:UBW

[SF23] Vincent S. Staggs and Keith Feldman. Use of between-within degrees of freedom as an alternative to the Kenward–Roger method for small-sample inference in generalized linear mixed modeling of clustered count data. Communications in Statistics: Simulation and Computation, 52(10):5099–5109, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shi:2024:VIS

[SF24] Xuejun Shi and Qun Feng. Viability for Itô stochastic systems with non-Lipschitzian coefficients and its application. Communications in Statistics: Simulation and Computation, 53(1):446–456, 2024. CODEN CSSCDB. ISSN 0361-0918.

Shahriari:2021:RSL

[SFG21] Shirin Shahriari, Susana Faria, and A. Manuela Gonçalves. A robust sparse linear approach for contaminated data. Communications in Statistics: Simulation and Computation, 50 (6):1662–1678, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sharma:2021:NTC

[SG21] Kush Sharma and Davinder Kumar Garg. A new technique for comparing PBIBDs. Communications in Statistics: Simulation and Computation, 50(4):1154–1164, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sanchez:2022:CDR

[SGG22] Ainara Rodríguez Sánchez, Román Salmerón Gómez, and Catalina García. The coefficient of determination in the ridge regression. Communications in Statistics: Simulation and Computation, 51(1):201–219, 2022. CODEN CSSCDB. ISSN 0361-0918.

Shang:2020:EFS

[Sha20] Han Lin Shang. Estimation of a functional single index model with dependent errors and unknown error density. Communications in Statistics: Simulation and Computation, 49(12): 3111–3133, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shan:2022:EBC

[Sha22a] Guogen Shan. Estimation of bias-corrected intraclass correlation coefficient for unbalanced clustered studies with continuous outcomes. Communications in Statistics: Simulation and Computation, 51(11):6691–6700, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sharafi:2022:ITP

[Sha22b] Maryam Sharafi. Inference of the two-parameter Lindley distribution based on progressive type II censored data with random removals. Communications in Statistics: Simulation and Computation, 51(4):1967–1981, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sha:2023:CBI

[Sha23] Naijun Sha. Classical and Bayesian inference for an extended generalized Birnbaum–Saunders distribution. Communications in Statistics: Simulation and Computation, 52 (11):5615–5632, 2023. CODEN CSSCDB. ISSN 0361-0918.

Syuhada:2023:EBV

[SHN23] Khreshna Syuhada, Arief Hakim, and Risti Nur’aini. The expected-based value-at-risk and expected shortfall using quantile and expectile with application to electricity market data. Communications in Statistics: Simulation and Computation, 52(7):3104–3121, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sancar:2020:IIO

[SI20] Nuriye Sancar and Deniz Inan. Identification of influential observations based on binary particle swarm optimization in the Cox PH model. Communications in Statistics: Simulation and Computation, 49(3):567–590, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shahzadi:2023:NCI

[SI23a] Asifa Shahzadi and Maryam Ilyas. A new confidence interval for standardized generalized variances of k-multivariate normal populations. Communications in Statistics: Simulation and Computation, 52(10):5014–5023, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shetty:2023:FSP

[SI23b] Dileep Kumar Shetty and B. Ismail. Forecasting stock prices using hybrid non-stationary time series model with ERNN. Communications in Statistics: Simulation and Computation, 52(3):1026–1040, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shabbir:2023:DFU

[SIG23] Javid Shabbir, Zahida Iqbal, and Sat Gupta. On distribution function under two-phase stratified sampling using two auxiliary variables for mean estimation. Communications in Statistics: Simulation and Computation, 52(8):3758–3770, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shakoori:2024:CBB

[SIK24] Afrooz Shakoori, Muhyiddin Izadi, and Baha-Eldin Khaledi. Copula based Bayesian data analysis of loss reserving. Communications in Statistics: Simulation and Computation, 53 (2):727–743, 2024. CODEN CSSCDB. ISSN 0361-0918.

Sathar:2023:EBR

[SJ23] E. I. Abdul Sathar and Jitto Jose. Extropy based on records for random variables representing residual life. Communications in Statistics: Simulation and Computation, 52(1):196– 206, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shaabani:2024:ILP

[SJ24] J. Shaabani and A. A. Jafari. Inference on the lifetime performance index of gamma distribution: point and interval estimation. Communications in Statistics: Simulation and Computation, 53(3):1368–1386, 2024. CODEN CSSCDB. ISSN 0361-0918.

Seo:2022:NPE

[SK22a] Jung In Seo and Yongku Kim. Nonparametric prior elicitation for a binomial proportion. Communications in Statistics: Simulation and Computation, 51(6):2809–2821, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Seo:2022:NFP

[SK22b] Jung In Seo and Yongku Kim. Note on the family of proportional reversed hazard distributions. Communications in Statistics: Simulation and Computation, 51(10):5832–5844, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:CCE

[SK22c] Garib Nath Singh and Mohd Khalid. A composite class of estimators to deal with the issue of variance estimation under the situations of random non-response in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 51(4):1454–1473, 2022. CODEN CSSCDB. ISSN 0361-0918.

Siriwardhana:2023:OPT

[SK23] Chathura Siriwardhana and K. B. Kulasekera. Optimal personalized treatment selection with multivariate outcome measures in a multiple treatment case. Communications in Statistics: Simulation and Computation, 52(12):5773–5787, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sabahno:2024:MAC

[SK24] Hamed Sabahno and Michael B. C. Khoo. A multivariate adaptive control chart for simultaneously monitoring of the process parameters. Communications in Statistics: Simulation and Computation, 53(4):2031–2049, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Shende:2022:MMI

[SKG22] Subhash Shende, Mohan Kale, and Nikhil Gupte. Modeling maternal infant HIV transmission having variable hazard rates with two lag time distributions. Communications in Statistics: Simulation and Computation, 51(6):2858–2870, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sunecher:2020:BMC

[SKJ20] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. BINMA(1) model with COM-Poisson innovations: Estimation and application. Communications in Statistics: Simulation and Computation, 49(6):1631–1652, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2021:SIM

[SKK21] Garib Nath Singh, Mohd Khalid, and Jong-Min Kim. Some imputation methods to deal with the problems of missing data in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 50(2):557–580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Saha:2023:SSG

[SKM^{+}23] Sajal Saha, Michael B. C. Khoo, Faijun Nahar Mim, Theam Foo Ng, and Khai Wah Khaw. Side sensitive group runs t chart and its application in manufacturing. Communications in Statistics: Simulation and Computation, 52(9): 4036–4051, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sreelakshmi:2021:JEL

[SKS21] N. Sreelakshmi, Sudheesh K. Kattumannil, and Rituparna Sen. Jackknife empirical likelihood-based inference for S-Gini indices. Communications in Statistics: Simulation and Computation, 50(6):1645–1661, 2021. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:SAA

[SKV20a] G. N. Singh, Amod Kumar, and Gajendra K. Vishwakarma. Some alternative additive randomized response models for estimation of population mean of quantitative sensitive variable in the presence of scramble variable. Communications in Statistics: Simulation and Computation, 49(11):2785–2807, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:EPM

[SKV20b] Garib Nath Singh, Amod Kumar, and Gajendra K. Vishwakarma. Estimation of population mean of sensitive quantitative character using blank cards in randomized device.

Communications in Statistics: Simulation and Computation, 49(6):1603–1630, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shen:2021:MRR

[SL21] Pao-Sheng Shen and Yi Liu. Marginal regression of recurrent gap times based on semiparametric transformation cure model. Communications in Statistics: Simulation and Computation, 50(5):1383–1397, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sun:2020:BIT

[SLE20] Li-Hsien Sun, Chang-Shang Lee, and Takeshi Emura. A Bayesian inference for time series via copula-based Markov chain models. Communications in Statistics: Simulation and Computation, 49(11):2897–2913, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sultana:2023:CUM

[SLJ23] Abida Sultana, Nasrin Lipi, and Ajmery Jaman. A caution in the use of multiple criteria for selecting working correlation structure in generalized estimating equations. Communications in Statistics: Simulation and Computation, 52(3): 980–992, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shinohara:2023:DLP

[SLME23] Sayaka Shinohara, Yuan-Hsin Lin, Hirofumi Michimae, and Takeshi Emura. Dynamic lifetime prediction using a Weibullbased bivariate failure time model: a meta-analysis of individual-patient data. Communications in Statistics: Simulation and Computation, 52(2):349–368, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Seo:2022:IMT

[SLP22] Yun Am Seo, Youngsaeng Lee, and Jeong-Soo Park. Iterative method for tuning complex simulation code. Communications in Statistics: Simulation and Computation, 51(7):3975–3992, 2022. CODEN CSSCDB. ISSN 0361-0918.

Song:2022:MFF

[SLSZ22] Fengli Song, Peng Lai, Baohua Shen, and Lianhua Zhu. Model free feature screening for ultrahigh dimensional covariates with right censored outcomes. Communications in Statistics: Simulation and Computation, 51(8):4815–4827, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Song:2020:MRE

[SLW20] Jingwen Song, Zhenzhou Lu, and Pengfei Wei. Measuring regional effects of model inputs with random forest. Communications in Statistics: Simulation and Computation, 49(9): 2444–2461, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shirozhan:2020:IMB

[SM20] Masoumeh Shirozhan and Mehrnaz Mohammadpour. An INAR(1) model based on the Pegram and thinning operators with serially dependent innovation. Communications in Statistics: Simulation and Computation, 49(10):2617–2638, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sazak:2023:CRM

[SM23a] Hakan Savas Sazak and Nalan Mutlu. Comparison of the robust methods in the general linear regression model. Communications in Statistics: Simulation and Computation, 52 (7):3163–3182, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shamma:2023:APD

[SM23b] N. Shamma and M. Mohammadpour. Alternative procedures in dependent counting INAR process with application on COVID-19. Communications in Statistics: Simulation and Computation, 52(11):5381–5395, 2023. CODEN CSSCDB. ISSN 0361-0918.

Samimia:2024:VIR

[SM24] Oldouz Samimia and Farshid Mehrdoust. Vasicek interest rate model under Lévy process and pricing bond option. Communications in Statistics: Simulation and Computation, 53 (1):529–545, 2024. CODEN CSSCDB. ISSN 0361-0918.

Shahnavaz:2022:HSA

[SMB22] Ali Shahnavaz, Ali M. Mosammam, and Mohammad Hassan Behzadi. Half-spectral analysis of spatial-temporal data: The case study of Iranian daily wind speed data. Communications in Statistics: Simulation and Computation, 51(3):1103–1120, 2022. CODEN CSSCDB. ISSN 0361-0918.

Slanzi:2023:CSH

[SMB23] Debora Slanzi, Valentina Mameli, and Philip J. Brown. A comparative study on high-dimensional Bayesian regression with binary predictors. Communications in Statistics: Simulation and Computation, 52(5):1979–1999, 2023. CODEN CSSCDB. ISSN 0361-0918.

Singini:2023:DTS

[SMG23] I. L. Singini, H. G. Mwambi, and F. N. Gumedze. Diagnostics for a two-stage joint survival model. Communications in Statistics: Simulation and Computation, 52(11):5163–5177, 2023. CODEN CSSCDB. ISSN 0361-0918.

Shongwe:2022:STM

[SMM22] S. C. Shongwe and J.-C. Malela-Majika. Shewhart-type monitoring schemes with supplementary w-of-w runs-rules to monitor the mean of autocorrelated samples. Communications in Statistics: Simulation and Computation, 51(1):293–322, 2022. CODEN CSSCDB. ISSN 0361-0918.

Soutinho:2022:CPM

[SMMO22] Gustavo Soutinho, Luís Meira-Machado, and Pedro Oliveira. A comparison of presmoothing methods in the estimation of transition probabilities. Communications in Statistics: Simulation and Computation, 51(9):5202–5221, 2022. CODEN CSSCDB. ISSN 0361-0918.

Simoes:2021:ACM

[SMO21] Maria L. Simões and Paula Milheiro-Oliveira. An algorithm for the computation of mean delay of traffic at pre-timed control signals. Communications in Statistics: Simulation and Computation, 50(12):4100–4113, 2021. CODEN CSSCDB. ISSN 0361-0918.

Shi:2020:DDB

[SMWJ20] Jian Shi, Yixuan Meng, Shaoping Wang, and Zongxia Jiao. A dynamic delay-based reliability evaluation model for communication networks. Communications in Statistics: Simulation and Computation, 49(6):1397–1414, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sheng:2022:EPF

[SN22a] Ke-Ning Sheng and Joseph I. Naus. Exact probability of fixed patterns occurring in a random sequence. Communications in Statistics: Simulation and Computation, 51(9):4867–4882, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sivena:2022:IQH

[SN22b] Sofia Sivena and Yiannis Nikolaidis. Improving the quality of higher education teaching through the exploitation of student evaluations and the use of control charts. Communications in Statistics: Simulation and Computation, 51(3):1289–1312, 2022. CODEN CSSCDB. ISSN 0361-0918.

Stewart:2023:ELB

[SN23] Patrick Stewart and Wei Ning. Empirical-likelihood-based hypothesis tests for the means of two zero-inflated populations. Communications in Statistics: Simulation and Computation, 52(10):4933–4961, 2023. CODEN CSSCDB. ISSN 0361-0918.

Salmasnia:2020:IQM

[SNA20] Ali Salmasnia, Mohammadreza Namdar, and Behnam Abdzadeh. An integrated quality and maintenance model for two-unit series systems. Communications in Statistics: Simulation and Computation, 49(4):886–917, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Saha:2024:PEC

[SNKA24] Sajal Saha, Wei Chun Ng, Michael B. C. Khoo, and Sani Salihu Abubakar. On the performances of the X̄ and EWMA charts with auxiliary information for monitoring the process mean in short production runs. Communications in Statistics: Simulation and Computation, 53(2):1007–1019, 2024. CODEN CSSCDB. ISSN 0361-0918.

Sanaullah:2022:CHT

[SNSE22] Aamir Sanaullah, Iqra Niaz, Javid Shabbir, and Iqra Ehsan. A class of hybrid type estimators for variance of a finite population in simple random sampling. Communications in Statistics: Simulation and Computation, 51(10):5609–5619, 2022. CODEN CSSCDB. ISSN 0361-0918.

Solis:2021:NPE

[Sol21] Maikol Solís. Non-parametric estimation of the first-order Sobol indices with bootstrap bandwidth. Communications in Statistics: Simulation and Computation, 50(9):2497–2512, 2021. CODEN CSSCDB. ISSN 0361-0918.

Some:2022:BSA

[Som22] Sobom M. Somé. Bayesian selector of adaptive bandwidth for gamma kernel density estimator on [0, ∞): simulations and applications. Communications in Statistics: Simulation and Computation, 51(12):7287–7297, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sassi:2023:CLS

[SP23] Gilberto Pereira Sassi and Carolina Costa Mota Paraíba. Conditional least squares estimation for the SINAR(1, 1) process. Communications in Statistics: Simulation and Computation, 52(3):945–960, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sparks:2022:SRS

[Spa22] Ross Stewart Sparks. Selecting the rational subgroup is crucial in developing a robust monitoring plan for dispersion of normally distributed data. Communications in Statistics: Simulation and Computation, 51(12):6925–6935, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Salmaso:2022:DEM

[SPG^{+}22] Luigi Salmaso, Luca Pegoraro, Rosa Arboretti Giancristofaro, Riccardo Ceccato, Alberto Bianchi, Silvio Restello, and Davide Scarabottolo. Design of experiments and machine learning to improve robustness of predictive maintenance with application to a real case study. Communications in Statistics: Simulation and Computation, 51(2):570–582, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:GES

[SPS22] Garib Nath Singh, Awadhesh K. Pandey, and Chandraketu Singh. Generalized estimation strategy for mean estimation on current occasion in two-occasion rotation patterns. Communications in Statistics: Simulation and Computation, 51 (4):1661–1684, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singhasomboon:2022:CIR

[SPV22] Lapasrada Singhasomboon, Wararit Panichkitkosolkul, and Andrei Volodin. Confidence intervals for the ratio of medians of two independent log-normal distributions. Communications in Statistics: Simulation and Computation, 51(11): 6729–6738, 2022. CODEN CSSCDB. ISSN 0361-0918.

Squarcio:2022:SNE

[Squ22] Claudio R. S. Ávila Jr. Roberto M. F. Squarcio. Some new error estimates for statistical estimators obtained by Neumann– Monte Carlo methodology applied to the stochastic bending problem. Communications in Statistics: Simulation and Computation, 51(12):6947–6967, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sirohi:2023:SCP

[SR23] Anu Sirohi and Piyush Kant Rai. Some r–k class proportional hazard regression models in the presence of collinearity: an evidence from Indian infant mortality. Communications in Statistics: Simulation and Computation, 52(10):4880–4890, 2023. CODEN CSSCDB. ISSN 0361-0918.

Silva:2023:RIC

[SRF^{+}23] A. M. Silva, M. Resende, M. Facco, A. R. de Morais, and M. A. Cirillo. Robustness of interpretable components in relation to the effect of outliers using measures and circular distances. Communications in Statistics: Simulation and Computation, 52(5):1822–1833, 2023. CODEN CSSCDB. ISSN 0361-0918.

Schmid:2022:CPS

[SRN^{+}22] Ian Schmid, Kara E. Rudolph, Trang Quynh Nguyen, Hwanhee Hong, Marissa J. Seamans, Benjamin Ackerman, and Elizabeth A. Stuart. Comparing the performance of statistical methods that generalize effect estimates from randomized controlled trials to much larger target populations. Communications in Statistics: Simulation and Computation, 51(8): 4326–4348, 2022. CODEN CSSCDB. ISSN 0361-0918.

Silva:2020:IMU

[SRO20] Eliud Silva, Aram Ramos, and Karen G. Olvera. Inferences on mortality using the Heligman–Pollard model: the Mexican case. Communications in Statistics: Simulation and Computation, 49(12):3227–3239, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:SMC

[SS20a] Parminder Singh and Navdeep Singh. Step-up multiple comparison procedure for comparing several normal variances. Communications in Statistics: Simulation and Computation, 49(6):1592–1602, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:CSN

[SS20b] Rajvir Chauhan Singh and Parminder Singh. Comparing several normal variances with a control using sample quasi range. Communications in Statistics: Simulation and Computation, 49(2):396–407, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sarada:2021:BDA

[SS21a] Y. Sarada and R. Shenbagam. Bi-dimensional availability function and its application. Communications in Statistics: Simulation and Computation, 50(5):1333–1347, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Shi:2021:DTF

[SS21b] Jian Shi and Kai Song. A discrete-time and finite-state Markov chain based in-play prediction model for NBA basketball matches. Communications in Statistics: Simulation and Computation, 50(11):3768–3776, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sajana:2023:RQD

[SS23] O. K. Sajana and T. A. Sajesh. Robust quadratic discriminant analysis using S_{n} covariance. Communications in Statistics: Simulation and Computation, 52(3):735–744, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sasiwannapong:2022:ECB

[SSBA22] Sirasak Sasiwannapong, Saowanit Sukparungsee, Piyapatr Busababodhin, and Yupaporn Areepong. The efficiency of constructed bivariate copulas for MEWMA and Hotelling’s T^{2} control charts. Communications in Statistics: Simulation and Computation, 51(4):1837–1851, 2022. CODEN CSSCDB. ISSN 0361-0918.

Saglam:2023:AEA

[SSCT23] Fatih Saglam, Tuba Sanli, Mehmet Ali Cengiz, and Yüksel Terzi. Alternative expectation approaches for expectationmaximization missing data imputations in Cox regression. Communications in Statistics: Simulation and Computation, 52(12):5966–5974, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sanaullah:2022:MEG

[SSGH22] Aamir Sanaullah, Iram Saleem, Sat Gupta, and Muhammad Hanif. Mean estimation with generalized scrambling using two-phase sampling. Communications in Statistics: Simulation and Computation, 51(10):5643–5657, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:MRD

[SSK22] Garib Nath Singh, Chandraketu Singh, and Amod Kumar. A modified randomized device for estimation of population mean of quantitative sensitive variable with measure of privacy protection. Communications in Statistics: Simulation and Computation, 51(4):1867–1890, 2022. CODEN CSSCDB. ISSN 0361-0918.

Salinas:2020:CUP

[SSS20] Veronica I. Salinas, Stephen A. Sedory, and Sarjinder Singh. Calibration using power transformation. Communications in Statistics: Simulation and Computation, 49(9):2256–2286, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sohail:2023:CSD

[SSS23] Muhammad Umair Sohail, Fariha Sohil, and Javid Shabbir. Comparative study of different imputation methods. Communications in Statistics: Simulation and Computation, 52 (8):3452–3474, 2023. CODEN CSSCDB. ISSN 0361-0918.

Subhash:2023:NEQ

[SSSR23] Silpa Subhash, S. M. Sunoj, P. G. Sankaran, and G. Rajesh. Nonparametric estimation of quantile-based entropy function. Communications in Statistics: Simulation and Computation, 52(5):1805–1821, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sharafi:2023:FOI

[SSZ23] M. Sharafi, Z. Sajjadnia, and A. Zamani. A first-order integervalued autoregressive process with zero-modified Poisson– Lindley distributed innovations. Communications in Statistics: Simulation and Computation, 52(3):685–702, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Santore:2021:ACC

[STdL21] Fabiola Santore, Cesar Augusto Taconeli, and Idemauro Antonio Rodrigues de Lara. An adaptive control chart for the process location based on ranked set sampling. Communications in Statistics: Simulation and Computation, 50(11): 3364–3382, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sharifi:2024:BPL

[STH24] Ali Sharifi, Ali R. Taheriyoun, and Hamideh D. Hamedani. A Bayesian paradigm in a large class of Lévy-driven CARMA models for high frequency data. Communications in Statistics: Simulation and Computation, 53(4):1824–1836, 2024. CODEN CSSCDB. ISSN 0361-0918.

Singh:2021:ECV

[SU21] Garib Nath Singh and Mahamood Usman. Efficient combination of various estimators in the presence of non-response. Communications in Statistics: Simulation and Computation, 50(8):2432–2466, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sathe:2022:EPM

[SU22] Aastha M. Sathe and N. S. Upadhye. Estimation of the parameters of multivariate stable distributions. Communications in Statistics: Simulation and Computation, 51(10): 5897–5914, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sen:2023:BSI

[SU23a] Rijji Sen and S. K. Upadhyay. A Bayes study of inverse Gaussian based strength models with accelerating stress. Communications in Statistics: Simulation and Computation, 52(8): 3973–3990, 2023. CODEN CSSCDB. ISSN 0361-0918.

Singh:2023:SEE

[SU23b] G. N. Singh and M. Usman. Some efficient estimators in case of missing data. Communications in Statistics: Simulation and Computation, 52(3):1077–1103, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sulewski:2022:MLG

[Sul22a] Piotr Sulewski. Modified Lilliefors goodness-of-fit test for normality. Communications in Statistics: Simulation and Computation, 51(3):1199–1219, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sulewski:2022:RDR

[Sul22b] Piotr Sulewski. Recognizing distributions rather than goodness-of-fit testing. Communications in Statistics: Simulation and Computation, 51(11):6701–6714, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sulewski:2023:RDU

[Sul23] Piotr Sulewski. Recognizing distributions using method of potential functions. Communications in Statistics: Simulation and Computation, 52(6):2542–2558, 2023. CODEN CSSCDB. ISSN 0361-0918.

Sun:2023:IPP

[Sun23] Ping Sun. Identification of parameters of Poisson distributions by the extreme order statistics. Communications in Statistics: Simulation and Computation, 52(7):3156–3162, 2023. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:CEM

[SVK22] Neha Singh, Gajendra K. Vishwakarma, and Jong Min Kim. Computing the effect of measurement errors on efficient variant of the product and ratio estimators of mean using auxiliary information. Communications in Statistics: Simulation and Computation, 51(2):604–625, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:NEC

[SVM20] Rajesh Singh, Gautam Kumar Vishwakarma, and Prabhakar Mishra. A note on the estimators for coefficient of dispersion using auxiliary information. Communications in Statistics: Simulation and Computation, 49(9):2347–2356, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Srujana:2024:MLV

[SVN24] B. Srujana, Dhananjay Verma, and Sameen Naqvi. Machine Learning vs. Survival Analysis Models: a study on right censored heart failure data. Communications in Statistics: Simulation and Computation, 53(4):1899–1916, 2024. CODEN CSSCDB. ISSN 0361-0918.

Shang:2023:MEE

[SW23] Wenpeng Shang and Chunjie Wu. More effective estimation for additive hazards model in generalized case-cohort study. Communications in Statistics: Simulation and Computation, 52(11):5345–5370, 2023. CODEN CSSCDB. ISSN 0361-0918.

Speiser:2020:BTD

[SWC^{+}20] Jaime Lynn Speiser, Bethany J. Wolf, Dongjun Chung, Constantine J. Karvellas, David G. Koch, and Valerie L.

Durkalski. BiMM tree: a decision tree method for modeling clustered and longitudinal binary outcomes. Communications in Statistics: Simulation and Computation, 49(4):1004–1023, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sahtout:2024:DTM

[SWG24] Mohammad Omar Sahtout, Haiyan Wang, and Santosh Ghimire. Different thresholding methods on Nearest Shrunken Centroid algorithm. Communications in Statistics: Simulation and Computation, 53(3):1444–1460, 2024. CODEN CSSCDB. ISSN 0361-0918.

Song:2022:GFS

[SX22] Won Chul Song and Jun Xie. Group feature screening via the F statistic. Communications in Statistics: Simulation and Computation, 51(4):1921–1931, 2022. CODEN CSSCDB. ISSN 0361-0918.

Stuart:2022:CEM

[SY22] Matthew Stuart and Cindy Yu. A computationally efficient method for selecting a split questionnaire design. Communications in Statistics: Simulation and Computation, 51(5): 2464–2486, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sungboonchoo:2023:LCI

[SYPV23] Chanakan Sungboonchoo, Su-Fen Yang, Wararit Panichkitkosolkul, and Andrei Volodin. Logarithmic confidence intervals for the cross-product ratio of binomial proportions under different sampling schemes. Communications in Statistics: Simulation and Computation, 52(6):2686–2704, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Satter:2020:NIE

[SZ20] Faysal Satter and Yichuan Zhao. Nonparametric interval estimation for the mean of a zero-inflated population. Communications in Statistics: Simulation and Computation, 49(8): 2059–2067, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sazak:2021:MML

[SZ21] Hakan Savas Sazak and Melis Zeybek. The modified maximum likelihood regression type estimators using bivariate ranked set sampling. Communications in Statistics: Simulation and Computation, 50(11):3614–3649, 2021. CODEN CSSCDB. ISSN 0361-0918.

Siavashi:2023:ACS

[SZ23] Mohammad Siavashi and Somayeh Zarezadeh. An algorithm for computing the t-signature of two-state networks. Communications in Statistics: Simulation and Computation, 52(5): 2130–2138, 2023. CODEN CSSCDB. ISSN 0361-0918.

Song:2021:NMS

[SZJG21] Zhao-Li Song, Qian Zhao, Ping Jiang, and Bo Guo. A numerical method for system residual life prediction based on Simpson formula. Communications in Statistics: Simulation and Computation, 50(12):4171–4186, 2021. CODEN CSSCDB. ISSN 0361-0918.

Song:2020:MSP

[SZS20] Kai Song, Qingrong Zou, and Jian Shi. Modelling the scores and performance statistics of NBA basketball games. Communications in Statistics: Simulation and Computation, 49 (10):2604–2616, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tsagris:2024:IHT

[TA24] Michail Tsagris and Abdulaziz Alenazi. An investigation of hypothesis testing procedures for circular and spherical mean vectors. Communications in Statistics: Simulation and Computation, 53(3):1387–1408, 2024. CODEN CSSCDB. ISSN 0361-0918.

Tighkhorshid:2022:RAE

[TAA22] Elahe Tighkhorshid, Amirhossein Amiri, and Farzad Amirkhani. A risk-adjusted EWMA chart with dynamic probability control limits for monitoring survival time. Communications in Statistics: Simulation and Computation, 51(3):1333–1354, 2022. CODEN CSSCDB. ISSN 0361-0918.

Takahashi:2023:MIR

[Tak23] Masayoshi Takahashi. Multiple imputation regression discontinuity designs: Alternative to regression discontinuity designs to estimate the local average treatment effect at the cutoff. Communications in Statistics: Simulation and Computation, 52(9):4293–4312, 2023. CODEN CSSCDB. ISSN 0361-0918.

Talib:2024:MEC

[TAS24] Ayesha Talib, Sajid Ali, and Ismail Shah. Max-EWMA chart for time and magnitude monitoring using generalized exponential distribution. Communications in Statistics: Simulation and Computation, 53(4):1857–1872, 2024. CODEN CSS-CDB. ISSN 0361-0918.

Thomson:2022:TFS

[TBH22] Trevor J. Thomson, W. John Braun, and X. Joan Hu. On the time to first spotting in wildland fires. Communications in Statistics: Simulation and Computation, 51(5):2381–2396, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tiwari:2023:GCE

[TBK23] Kuldeep Kumar Tiwari, Sandeep Bhougal, and Sunil Kumar. A general class of estimators in stratified random sampling. Communications in Statistics: Simulation and Computation, 52(2):442–452, 2023. CODEN CSSCDB. ISSN 0361-0918.

Toka:2022:RER

[TÇA22] Onur Toka, Meral Çetin, and Olcay Arslan. Robust estimation in restricted linear regression. Communications in Statistics: Simulation and Computation, 51(3):1015–1029, 2022. CODEN CSSCDB. ISSN 0361-0918.

Thonghnunui:2023:TSM

[TCJ23] Nittaya Thonghnunui, Samruam Chongcharoen, and Knavoot Jiamwattanapong. Two-sample multivariate tests for highdimensional data when one covariance matrix is unknown. Communications in Statistics: Simulation and Computation, 52(3):669–684, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tarapara:2023:PEA

[TD23] Bharat Tarapara and Jyoti Divecha. Practical efficient alpha: Encompassing all alphas of central composite design. Communications in Statistics: Simulation and Computation, 52 (4):1621–1629, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tran:2020:WIR

[TDD20] Duyet Tran, Anil Dolgun, and Haydar Demirhan. Weighted inter-rater agreement measures for ordinal outcomes. Communications in Statistics: Simulation and Computation, 49 (4):989–1003, 2020. CODEN CSSCDB. ISSN 0361-0918.

Taconeli:2022:ECE

[TdSCdSdCP22] Cesar Augusto Taconeli, Angelo da Silva Cabral, José Luiz Padilha da Silva, and Anderson de Castro Peres. Evaluation of C_{pm} estimators in ranked set sampling designs. Communications in Statistics: Simulation and Computation, 51 (8):4749–4768, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tajmouati:2024:ARC

[TED24] Samya Tajmouati, Bouazza El Wahbi, and Mohamed Dakkon. Applying regression conformal prediction with nearest neighbors to time series data. Communications in Statistics: Simulation and Computation, 53(4):1768–1778, 2024. CODEN CSSCDB. ISSN 0361-0918.

Teimouri:2022:BCM

[Tei22] Mahdi Teimouri. Bias corrected maximum likelihood estimators under progressive Type-I interval censoring scheme. Communications in Statistics: Simulation and Computation, 51(11):6854–6865, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tellaroli:2022:SCA

[Tel22] P. Tellaroli. SingleCross-clustering: an algorithm for finding elongated clusters with automatic estimation of outliers and number of clusters. Communications in Statistics: Simulation and Computation, 51(5):2412–2428, 2022. CODEN CSSCDB. ISSN 0361-0918.

Taremi:2020:EWI

[TEM20] Mohammad Taremi, Farzad Eskandari, and Mohammad Bameni Moghadam. Effects of weak identification on the MD estimator in dynamic stochastic general equilibrium models. Communications in Statistics: Simulation and Computation, 49(10):2661–2677, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tuzen:2020:SSC

[TEO20] Fatih Tüzen, Semra Erbas, and Hülya Olmus. A simulation study for count data models under varying degrees of outliers and zeros. Communications in Statistics: Simulation and Computation, 49(4):1078–1088, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tarima:2022:DTF

[TF22a] Sergey Tarima and Nancy Flournoy. Distribution theory following blinded and unblinded sample size re-estimation under parametric models. Communications in Statistics: Simulation and Computation, 51(4):2053–2064, 2022. CODEN CSSCDB. ISSN 0361-0918.

Teng:2022:SFA

[TF22b] Huei-Wen Teng and Cheng-Der Fuh. Simulating false alarm probability in K-distributed sea clutter. Communications in Statistics: Simulation and Computation, 51(9):5081–5098, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tian:2022:MBG

[TGS22] Xinyu Tian, Yiran Gao, and Jian Shi. Modeling basketball games by inverse Gaussian processes. Communications in Statistics: Simulation and Computation, 51(11):6246–6256, 2022. CODEN CSSCDB. ISSN 0361-0918.

Trong:2023:PED

[TH23] Dang Duc Trong and Thai Phuc Hung. Parameter estimation for diffusion process from perturbed discrete observations. Communications in Statistics: Simulation and Computation, 52(3):925–944, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tajuddin:2022:STC

[TII22] Razik Ridzuan Mohd Tajuddin, Noriszura Ismail, and Kamarulzaman Ibrahim. Several two-component mixture distributions for count data. Communications in Statistics: Simulation and Computation, 51(7):3760–3771, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tamae:2024:SAM

[TIK24] Hiromasa Tamae, Kaoru Irie, and Tatsuya Kubokawa. Scoreadjusted methods for estimation of shape parameters in Gamma-Poisson and Beta-Binomial distributions. Communications in Statistics: Simulation and Computation, 53(3): 1247–1257, 2024. CODEN CSSCDB. ISSN 0361-0918.

Tamandi:2020:FMM

[TJ20] Mostafa Tamandi and Ahad Jamalizadeh. Finite mixture modeling using shape mixtures of the skew scale mixtures of normal distributions. Communications in Statistics: Simulation and Computation, 49(12):3345–3366, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tanil:2022:MSR

[TK22a] Halil Tanil and Agah Kozan. A modified Spearman’s rho parameter-free test statistic for early detection of newly emerging phenomena. Communications in Statistics: Simulation and Computation, 51(6):2822–2830, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tee:2022:EIC

[TK22b] Liivika Tee and Meelis Käärik. Estimating IBNR claim counts using different levels of data aggregation. Communications in Statistics: Simulation and Computation, 51(6):3013–3036, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tsay:2023:SAB

[TK23] Wen-Jen Tsay and Peng-Hsuan Ke. A simple approximation for the bivariate normal integral. Communications in Statistics: Simulation and Computation, 52(4):1462–1475, 2023. CODEN CSSCDB. ISSN 0361-0918.

Takam:2020:PML

[TKF^{+}20] Patrice Soh Takam, Eugene Kouassi, Renaud Fadonougbo, Jean Marcelin Bosson Brou, and Mbodja Mougoué. Pseudo maximum likelihood estimation and asymptotic results of the GARCH(1, 2) model under dependent innovations. Communications in Statistics: Simulation and Computation, 49(8): 2134–2163, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tyagi:2024:RAW

[TKR24] Subhi Tyagi, Akshay Kumar, and Mangey Ram. Reliability analysis of m-out-of–r-within-k-out-of-n system. Communications in Statistics: Simulation and Computation, 53(2): 671–681, 2024. CODEN CSSCDB. ISSN 0361-0918.

Tsai:2023:EME

[TL23a] Jia-Ren Tsai and Pen-Hwang Liau. Effect of measurement error size in linear heteroscedastic measurement error models. Communications in Statistics: Simulation and Computation, 52(10):5052–5081, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tsai:2023:NMA

[TL23b] Miao-Yu Tsai and Chao-Chun Lin. New model-averaged estimators of concordance correlation coefficients: simulation and application to longitudinal overdispersed Poisson data. Communications in Statistics: Simulation and Computation, 52(3):961–979, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tong:2021:NEC

[TMZL21] Jinying Tong, Qingting Meng, Zhenzhong Zhang, and Yunfang Lu. A note on ergodicity for CIR model with Markov switching. Communications in Statistics: Simulation and Computation, 50(5):1445–1458, 2021. CODEN CSSCDB. ISSN 0361-0918.

Thangjai:2020:AGC

[TNN20] Warisa Thangjai, Sa-Aat Niwitpong, and Suparat Niwitpong. Adjusted generalized confidence intervals for the common coefficient of variation of several normal populations. Communications in Statistics: Simulation and Computation, 49(1): 194–206, 2020. CODEN CSSCDB. ISSN 0361-0918.

Triantafyllou:2022:NDF

[TP22a] Ioannis S. Triantafyllou and Nikolaos I. Panayiotou. A new distribution-free monitoring scheme based on ranks. Communications in Statistics: Simulation and Computation, 51(11): 6456–6478, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tunno:2022:SDD

[TP22b] Ferebee Tunno and Miranda Perry. Signal discrimination without denoising. Communications in Statistics: Simulation and Computation, 51(2):626–646, 2022. CODEN CSSCDB. ISSN 0361-0918.

Toharudin:2023:ELS

[TPC^{+}23] Toni Toharudin, Resa Septiani Pontoh, Rezzy Eko Caraka, Solichatus Zahroh, Youngjo Lee, and Rung Ching Chen. Employing long short-term memory and Facebook prophet model in air temperature forecasting. Communications in Statistics: Simulation and Computation, 52(2):279–290, 2023. CODEN CSSCDB. ISSN 0361-0918.

Trisandhya:2022:AIS

[TPK22] Pidugu Trisandhya, Kumari Priyanka, and Ajay Kumar. Application of item sum technique for estimating quantitative sensitive mean on successive moves using auxiliary information. Communications in Statistics: Simulation and Computation, 51(7):3868–3887, 2022. CODEN CSSCDB. ISSN 0361-0918.

Teuling:2023:CMC

[TPvdH23] N. G. P. Den Teuling, S. C. Pauws, and E. R. van den Heuvel. A comparison of methods for clustering longitudinal data with slowly changing trends. Communications in Statistics: Simulation and Computation, 52(3):621–648, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Tomazella:2022:LTI

[TRF^{+}22] Vera L. D. Tomazella, Pedro L. Ramos, Paulo H. Ferreira, Alex L. Mota, and Francisco Louzada. The Lehmann type II inverse Weibull distribution in the presence of censored data. Communications in Statistics: Simulation and Computation, 51(12):7057–7073, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tanis:2022:RBT

[TS22] Caner Tanis and Bugra Saraçoglu. On the record-based transmuted model of Balakrishnan and He based on Weibull distribution. Communications in Statistics: Simulation and Computation, 51(8):4204–4224, 2022. CODEN CSSCDB. ISSN 0361-0918.

Turksen:2020:NML

[Tür20] Özlem Türksen. A nonlinear modeling with linear fuzzy numbers for replicated response measures. Communications in Statistics: Simulation and Computation, 49(3):629–646, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tharshan:2024:PMQ

[TW24] Ramajeyam Tharshan and Pushpakanthie Wijekoon. Poissonmodification of quasi Lindley regression model for overdispersed count responses. Communications in Statistics: Simulation and Computation, 53(3):1258–1273, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Tang:2023:SPB

[TWSW23] Fengqin Tang, Chunning Wang, Jinxia Su, and Yuanyuan Wang. Semidefinite programming based community detection for node-attributed networks and multiplex networks. Communications in Statistics: Simulation and Computation, 52 (1):68–83, 2023. CODEN CSSCDB. ISSN 0361-0918.

Tian:2020:LBQ

[TWTT20] Yuzhu Tian, Liyong Wang, Manlai Tang, and Maozai Tian. Likelihood-based quantile mixed effects models for longitudinal data with multiple features via MCEM algorithm. Communications in Statistics: Simulation and Computation, 49 (2):317–334, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tian:2021:WCQ

[TWTT21] Yuzhu Tian, Liyong Wang, Manlai Tang, and Maozai Tian. Weighted composite quantile regression for longitudinal mixed effects models with application to AIDS studies. Communications in Statistics: Simulation and Computation, 50(6):1837–1853, 2021. CODEN CSSCDB. ISSN 0361-0918.

Tan:2021:PEL

[TY21] Xiaoyan Tan and Li Yan. Penalized empirical likelihood for generalized linear models with longitudinal data. Communications in Statistics: Simulation and Computation, 50(2): 608–623, 2021. CODEN CSSCDB. ISSN 0361-0918.

Tekeli:2022:EPT

[TY22] Erkut Tekeli and Güzin Yüksel. Estimating the parameters of twofold Weibull mixture model in right-censored reliability data by using genetic algorithm. Communications in Statistics: Simulation and Computation, 51(11):6621–6634, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tian:2024:CPA

[TY24] Weizhong Tian and Yaoting Yang. Change point analysis for weighted exponential distribution. Communications in Statistics: Simulation and Computation, 53(1):367–379, 2024. CODEN CSSCDB. ISSN 0361-0918.

Tan:2022:CEO

[TYLX22] Jiyang Tan, Yang Yang, Shuren Liu, and Kainan Xiang. A consistent estimation of optimal dividend strategy in a risk model with delayed claims. Communications in Statistics: Simulation and Computation, 51(11):6840–6853, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2023:HWM

[uAA23] Muhammad Noor ul Amin and Asma Arshad. Homogeneously weighted moving average-variance control chart using auxiliary information. Communications in Statistics: Simulation and Computation, 52(10):4891–4908, 2023. CODEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2023:JMM

[uAAF23] Muhammad Noor ul Amin, Irfan Aslam, and Navid Feroze. Joint monitoring of mean and variance using Max-EWMA for Weibull process. Communications in Statistics: Simulation and Computation, 52(7):3257–3272, 2023. CODEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2022:PME

[uAJHD22] Muhammad Noor ul Amin, Amjad Javaid, Muhammad Hanif, and Eralp Dogu. Performance of maximum EWMA control chart in the presence of measurement error using auxiliary information. Communications in Statistics: Simulation and Computation, 51(9):5482–5506, 2022. CODEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2022:EWM

[uARS22] Muhammad Noor ul Amin, Afshan Riaz, and Amna Safeer. Exponentially weighted moving average control chart using auxiliary variable with measurement error. Communications in Statistics: Simulation and Computation, 51(3):1002–1014, 2022. CODEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2022:VAS

[uASS22] Muhammad Noor ul Amin, Amna Safeer, and Paryas Sharma. Variable acceptance sampling plan based on hybrid exponentially weighted moving averages. Communications in Statistics: Simulation and Computation, 51(12):7544–7553, 2022. CODEN CSSCDB. ISSN 0361-0918.

Uzabaci:2020:EOD

[UEA20] Ender Uzabaci, Ilker Ercan, and Ozlem Alpu. Evaluation of outlier detection method performance in symmetric multivariate distributions. Communications in Statistics: Simulation and Computation, 49(2):516–531, 2020. CODEN CSSCDB. ISSN 0361-0918.

Unsal:2022:CDC

[ÜFR22] Mehmet Güray Ünsal, Daniel Friesner, and Robert Rosenman. The curse of dimensionality (COD), misclassified DMUs, and Bayesian DEA. Communications in Statistics:

Simulation and Computation, 51(8):4186–4203, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Umelo-Ibemere:2023:ASC

[UIC23] Ngozi C. Umelo-Ibemere and Polycarp E. Chigbu. An algorithm for the simultaneous construction of D-optimal experimental designs. Communications in Statistics: Simulation and Computation, 52(9):4156–4173, 2023. CODEN CSSCDB. ISSN 0361-0918.

Khan:2022:MST

[uIKKH22] Asad ul Islam Khan, Waqar Muhammad Khan, and Mehmood Hussan. Most stringent test of null of cointegration: a Monte Carlo comparison. Communications in Statistics: Simulation and Computation, 51(4):2020–2038, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ulas:2020:IDM

[UKK20] Efehan Ulas, Filiz Karaman, and Tuba Koc. Impact of different model structure and prior distribution in continual reassessment method. Communications in Statistics: Simulation and Computation, 49(1):95–109, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Utazirubanda:2021:VSG

[ULN21] Jean Claude Utazirubanda, Tomás M. León, and Papa Ngom. Variable selection with group LASSO approach: Application to Cox regression with frailty model. Communications in Statistics: Simulation and Computation, 50(3):881–901, 2021. CODEN CSSCDB. ISSN 0361-0918.

UlHassan:2022:DUM

[UM22] Mahmood Ul Hassan and Frank Miller. Discrimination with unidimensional and multidimensional item response theory models for educational data. Communications in Statistics: Simulation and Computation, 51(6):2992–3012, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ugwuowo:2023:JKL

[UOA23] Fidelis Ifeanyi Ugwuowo, Henrietta Ebele Oranye, and Kingsley Chinedu Arum. On the jackknife Kibria–Lukman estimator for the linear regression model. Communications in Statistics: Simulation and Computation, 52(12):6116–6128, 2023. CODEN CSSCDB. ISSN 0361-0918.

UlIslam:2022:CCM

[UR22] Tanweer Ul Islam and Mahvish Rizwan. Comparison of correlation measures for nominal data. Communications in Statistics: Simulation and Computation, 51(3):698–714, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ushakov:2023:SET

[UU23] N. G. Ushakov and V. G. Ushakov. On sensitivity of exponentiality tests to data rounding: a Monte Carlo simulation study. Communications in Statistics: Simulation and Computation, 52(12):6225–6234, 2023. CODEN CSSCDB. ISSN 0361-0918.

Uhm:2023:CNT

[UY23] Taewoong Uhm and Seongbaek Yi. A comparison of normality testing methods by empirical power and distribution of P - values. Communications in Statistics: Simulation and Computation, 52(9):4445–4458, 2023. CODEN CSSCDB. ISSN 0361-0918.

Uyanto:2022:MCP

[Uya22] Stanislaus S. Uyanto. Monte Carlo power comparison of seven most commonly used heteroscedasticity tests. Communications in Statistics: Simulation and Computation, 51(4):2065– 2082, 2022. CODEN CSSCDB. ISSN 0361-0918.

Uozumi:2022:CID

[UYMK22] Ryuji Uozumi, Shinjo Yada, Kazushi Maruo, and Atsushi Kawaguchi. Confidence intervals for difference between two binomial proportions derived from logistic regression. Communications in Statistics: Simulation and Computation, 51 (6):3223–3236, 2022. CODEN CSSCDB. ISSN 0361-0918.

Vishwakarma:2023:HMV

[VBB23] Gajendra K. Vishwakarma, Atanu Bhattacharjee, and Souvik Banerjee. Handling missingness value on jointly measured time-course and time-to-event data. Communications in Statistics: Simulation and Computation, 52(1):126–141, 2023. CODEN CSSCDB. ISSN 0361-0918.

Verma:2022:RRS

[VDN22] Vivek Verma, Radhakanta Das, and Dilip C. Nath. Representativeness of ranked set sampling based on Bayesian score.

Communications in Statistics: Simulation and Computation, 51(3):1080–1095, 2022. CODEN CSSCDB. ISSN 0361-0918.

Verdoy:2021:ESM

[Ver21] Pablo Juan Verdoy. Enhancing the SPDE modeling of spatial point processes with INLA, applied to wildfires. Choosing the best mesh for each database. Communications in Statistics: Simulation and Computation, 50(10):2990–3030, 2021. CODEN CSSCDB. ISSN 0361-0918.

Vishwakarma:2022:SES

[VG22] Gajendra K. Vishwakarma and Shubham Gupta. Shrinkage estimator for scale parameter of gamma distribution. Communications in Statistics: Simulation and Computation, 51 (6):3073–3080, 2022. CODEN CSSCDB. ISSN 0361-0918.

Vijverberg:2023:GRL

[VH23] Wim Vijverberg and Takuya Hasebe. GTL regression: a linear model with skewed and thick-tailed disturbances. Communications in Statistics: Simulation and Computation, 52 (6):2290–2309, 2023. CODEN CSSCDB. ISSN 0361-0918.

Verma:2023:CML

[VJV^{+}23] Ankita Verma, Seema Jaggi, Eldho Varghese, Cini Varghese, Arpan Bhowmik, Anindita Datta, and Hemavathi M. On the construction of mixed-level rotatable response surface designs when experimental unit experiences overlap effects. Communications in Statistics: Simulation and Computation, 52(5): 1753–1768, 2023. CODEN CSSCDB. ISSN 0361-0918.

Veerman:2022:EVC

[VLvdW22] Jurre R. Veerman, Gwenaël G. R. Leday, and Mark A. van de Wiel. Estimation of variance components, heritability and the ridge penalty in high-dimensional generalized linear models. Communications in Statistics: Simulation and Computation, 51(1):116–134, 2022. CODEN CSSCDB. ISSN 0361-0918.

vanOest:2021:UCB

[vO21] Rutger van Oest. Unconstrained Cholesky-based parametrization of correlation matrices. Communications in Statistics: Simulation and Computation, 50(11):3607–3613, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Vasconcelos:2024:ELN

[VPOC24] J. C. S. Vasconcelos, F. Prataviera, E. M. M. Ortega, and G. M. Cordeiro. An extended logit-normal regression with application to human development index data. Communications in Statistics: Simulation and Computation, 53(3):1356–1367, 2024. CODEN CSSCDB. ISSN 0361-0918.

Vovan:2021:AGA

[VPT21] Tai Vovan, Dinh Phamtoan, and Dung Tranthituy. Automatic genetic algorithm in clustering for discrete elements. Communications in Statistics: Simulation and Computation, 50(6):1679–1694, 2021. CODEN CSSCDB. ISSN 0361-0918.

Varghese:2024:ROC

[VSJ^{+}24] Hemavathi M. Eldho Varghese, Shashi Shekhar, Seema Jaggi, Arpan Bhowmik, and T. V. Sathianandan. Run order consideration for sequential third order rotatable designs. Communications in Statistics: Simulation and Computation, 53(2): 1068–1081, 2024. CODEN CSSCDB. ISSN 0361-0918.

Vo-Van:2023:ERA

[VVNNT23] Tai Vo-Van, Lethikim Ngoc, and Thao Nguyen-Trang. An efficient robust automatic clustering algorithm for interval data. Communications in Statistics: Simulation and Computation, 52(10):4621–4635, 2023. CODEN CSSCDB. ISSN 0361-0918.

Varathan:2021:MAU

[VW21] Nagarajah Varathan and Pushpakanthie Wijekoon. Modified almost unbiased Liu estimator in logistic regression. Communications in Statistics: Simulation and Computation, 50(11): 3530–3546, 2021. CODEN CSSCDB. ISSN 0361-0918.

Vanlengenberg:2021:DGA

[VWZ21] Christopher D. Vanlengenberg, Wenshuang Wang, and Haimeng Zhang. Data generation for axially symmetric processes on the sphere. Communications in Statistics: Simulation and Computation, 50(6):1750–1769, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wan:2022:SSD

[Wan22] Fei Wan. Simulating survival data with predefined censoring rates under a mixture of non-informative right censoring schemes. Communications in Statistics: Simulation and Computation, 51(7):3851–3867, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:SSF

[Wan23a] Jin Wang. A simulation study of the finite-sample performance of the sample scale curve as an estimator of its population counterpart. Communications in Statistics: Simulation and Computation, 52(9):4522–4531, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:SNL

[Wan23b] Jing Wang. Semiparametric nonlinear log-periodogram regression estimation for perturbed stationary anisotropic long memory random fields. Communications in Statistics: Simulation and Computation, 52(12):6034–6047, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:TML

[Wan23c] Yanfeng Zhang Lichun Wang. K′ times k-means logistic regression algorithm for imbalanced classification. Communications in Statistics: Simulation and Computation, 52(9):4252– 4259, 2023. CODEN CSSCDB. ISSN 0361-0918.

Warasi:2023:GRP

[War23] Md S. Warasi. groupTesting: an R package for group testing estimation. Communications in Statistics: Simulation and Computation, 52(12):6210–6224, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:SMV

[WC23a] Joanna J. J. Wang and Jennifer S. K. Chan. Stochastic modelling of volatility and inter-relationships in the Australian electricity markets. Communications in Statistics: Simulation and Computation, 52(8):3877–3896, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Wu:2023:PCT

[WC23b] Shu-Fei Wu and Wei-Tsung Chang. Power comparison of the testing on the lifetime performance index for Rayleigh lifetime products under progressive type I interval censoring. Communications in Statistics: Simulation and Computation, 52(4):1435–1448, 2023. CODEN CSSCDB. ISSN 0361-0918.

Weng:2022:JAL

[WCC22] Shuo-Chun Weng, Yin-Chu Chang, and Chyong-Mei Chen. Joint analysis of longitudinal and interval-censored failure time data. Communications in Statistics: Simulation and Computation, 51(9):5333–5349, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:SLR

[WCF21] Yuh-Jenn Wu, Li-Hsueh Cheng, and Wei-Quan Fang. Sparse linear regression models of high dimensional covariates with non-Gaussian outliers and Berkson error-in-variable under heteroscedasticity. Communications in Statistics: Simulation and Computation, 50(11):3146–3165, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Wu:2023:RSD

[WCKC23] Shu-Fei Wu, Yi-Wun Cheng, Che-Wei Kang, and Wei-Tsung Chang. Reliability sampling design for the lifetime performance index of Burr XII lifetime distribution under progressive type I interval censoring. Communications in Statistics: Simulation and Computation, 52(11):5483–5497, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2023:DAE

[WCRM23] Shu Wu, Philippe Castagliola, Athanasios C. Rakitzis, and Petros E. Maravelakis. Design of attribute EWMA type control charts with reliable run length performance. Communications in Statistics: Simulation and Computation, 52(9): 4193–4209, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2024:FIR

[WCY24] Shuo Wang, Wangxue Chen, and Rui Yang. Fisher information in ranked set sampling from the simple linear regression model. Communications in Statistics: Simulation and Computation, 53(3):1274–1284, 2024. CODEN CSSCDB. ISSN 0361-0918.

Wolny-Dominiak:2022:BES

[WDZ22] Alicja Wolny-Dominiak and Tomasz Za̧dło. On bootstrap estimators of some prediction accuracy measures of loss reserves in a non-life insurance company. Communications in Statistics: Simulation and Computation, 51(8):4225–4240, 2022. CODEN CSSCDB. ISSN 0361-0918.

Westgate:2020:AUM

[Wes20] Philip M. Westgate. Approaches for the utilization of multiple criteria to select a working correlation structure for use within generalized estimating equations. Communications in Statistics: Simulation and Computation, 49(2):305–316, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wang:2024:VSM

[WFLS24] Wenshan Wang, Lijun Fang, Shuwei Li, and Jianguo Sun. Variable selection for misclassified current status data under the proportional hazards model. Communications in Statistics: Simulation and Computation, 53(3):1494–1503, 2024. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:SIE

[WG23] Kexin Wang and Wenhao Gui. Statistical inference of exponentiated Pareto distribution under adaptive type-II progressive censored schemes. Communications in Statistics: Simulation and Computation, 52(11):5256–5287, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:ETE

[WGlTL23] Jun Wang, Wei Gao, Man lai Tang, and Changbiao Liu. Estimating treatment effects in the presence of unobserved confounders. Communications in Statistics: Simulation and Computation, 52(10):4685–4704, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wu:2024:ANI

[WH24] Wei-Hwa Wu and Hsin-Neng Hsieh. Assessing the noninferiority of a new treatment in a three-arm trial with unknown coefficient of variation. Communications in Statistics: Simulation and Computation, 53(3):1576–1593, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2023:GEW

[WHS23] Binhui Wang, Zhifeng He, and Lianjie Shu. A generalized exponentially weighted moving average control chart for monitoring autocorrelated vectors. Communications in Statistics: Simulation and Computation, 52(6):2559–2577, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2021:FSA

[WHW21] Huiwen Wang, Tingting Huang, and Shanshan Wang. A flexible spatial autoregressive modelling framework for mixed covariates of multiple data types. Communications in Statistics: Simulation and Computation, 50(11):3498–3515, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wijesuriya:2023:SPI

[Wij23] Uditha Amarananda Wijesuriya. Sv-plots for identifying characteristics of the distribution and testing hypotheses. Communications in Statistics: Simulation and Computation, 52(1):207–228, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2022:CVR

[Wil22] Rand R. Wilcox. Comparing the variances or robust measures of scale of two dependent variables. Communications in Statistics: Simulation and Computation, 51(11):6610–6620, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2024:NCC

[Wil24] Rand R. Wilcox. A note on computing a confidence interval for the mean. Communications in Statistics: Simulation and Computation, 53(1):164–166, 2024. CODEN CSSCDB. ISSN 0361-0918.

Wehrhahn:2021:SSB

[WJB21] Claudia Wehrhahn, Alejandro Jara, and Andrés F. Barrientos. On the small sample behavior of Dirichlet process mixture models for data supported on compact intervals. Communications in Statistics: Simulation and Computation, 50 (3):786–810, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:RVS

[WJZ^{+}22] Yan Wang, Yunlu Jiang, Jiantao Zhang, Zhongran Chen, Baojian Xie, and Chengxiang Zhao. Robust variable selection based on the random quantile LASSO. Communications in Statistics: Simulation and Computation, 51(1):29–39, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wei:2023:PCA

[WK23] Zheng Wei and Daeyoung Kim. On partial and conditional association measures for ordinal contingency tables. Communications in Statistics: Simulation and Computation, 52(6): 2705–2726, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2024:ALL

[WKcIC24] Zhanfeng Wang, Amy M. Kwon, and Yuan chin Ivan Chang. Active learning with logistic models featuring simultaneous variable and subject selection. Communications in Statistics: Simulation and Computation, 53(2):897–911, 2024. CODEN CSSCDB. ISSN 0361-0918.

Waseem:2022:GCE

[WKSeF22] Zara Waseem, Hina Khan, Javid Shabbir, and Shan e Fatima. A generalized class of exponential type estimators for estimating the mean of the sensitive variable when using optional randomized response model. Communications in Statistics: Simulation and Computation, 51(12):7602–7612, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2022:MAG

[WKT22] Mingqiu Wang, Xiaoning Kang, and Guo-Liang Tian. Modified adaptive group lasso for high-dimensional varying coefficient models. Communications in Statistics: Simulation and Computation, 51(11):6495–6510, 2022. CODEN CSSCDB. ISSN 0361-0918.

Watters:2020:CFM

[WL20] Christine G. Watters and Lynn R. LaMotte. A comparison of four methods of inverse prediction. Communications in Statistics: Simulation and Computation, 49(3):709–716, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:IEC

[WL22] Liang Wang and Huanyu Li. Inference for exponential competing risks data under generalized progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 51(3):1255–1271, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:NAP

[WLC^{+}22] Yan Wang, Xiaoqin Li, Ling Chen, Minghui Wang, and Xuejun Wang. A note on the asymptotic properties of the estimators in a semiparametric regression model. Communications in Statistics: Simulation and Computation, 51(2):358– 373, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:IDC

[WLT22] Liang Wang, Mengyang Li, and Yogesh Mani Tripathi. Inference for dependent competing risks from bivariate Kumaraswamy distribution under generalized progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 51(6):3100–3123, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:JML

[WLY22] Leyang Wang, Zhiqiang Li, and Fengbin Yu. Jackknife method for the location of gross errors in weighted total least squares. Communications in Statistics: Simulation and Computation, 51(4):1946–1966, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2023:BLV

[WLZ23] Jingjing Wu, Xuewen Lu, and Wenyan Zhong. Bi-level variable selection in semiparametric transformation mixture cure models for right-censored data. Communications in Statistics: Simulation and Computation, 52(7):3006–3025, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:MID

[WMJ21] Pengyue Wu, Jing Ma, and Xiongfei Jiang. A model of innovation diffusion based on policy incentives. Communications in Statistics: Simulation and Computation, 50(9):2544–2560, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wijekularathna:2022:PAS

[WMS22] Danush K. Wijekularathna, Ananda B. W. Manage, and Stephen M. Scariano. Power analysis of several normality tests: a Monte Carlo simulation study. Communications in Statistics: Simulation and Computation, 51(3):757–773, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2024:ASE

[WMXL24] Xi Wu, Weiyan Mu, Shifeng Xiong, and Xinmin Li. Adaptive subsample estimation for multivariate normal distributions. Communications in Statistics: Simulation and Computation, 53(4):1711–1718, 2024. CODEN CSSCDB. ISSN 0361-0918.

Wiegand:2021:GMR

[WN21a] Martin Wiegand and Saralees Nadarajah. General moments of roundoff error. Communications in Statistics: Simulation and Computation, 50(10):2885–2905, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Wiegand:2021:MME

[WN21b] Martin Wiegand and Saralees Nadarajah. MEPDF: Multivariate empirical density functions. Communications in Statistics: Simulation and Computation, 50(2):367–375, 2021. CODEN CSSCDB. ISSN 0361-0918.

Withers:2023:BRS

[WN23] Christopher S. Withers and Saralees Nadarajah. Bias reduction for standard and extreme estimates. Communications in Statistics: Simulation and Computation, 52(4):1264–1277, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wan:2022:BDA

[WP22] Shuwen Wan and Kai Peng. A Bayesian discriminant analysis method under semiparametric density ratio models. Communications in Statistics: Simulation and Computation, 51(4): 1759–1766, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wen:2022:NCP

[WQW^{+}22] Luliang Wen, Yanjun Qiu, Minghui Wang, Juliang Yin, and Pingyan Chen. Numerical characteristics and parameter estimation of finite mixed generalized normal distribution. Communications in Statistics: Simulation and Computation, 51 (7):3596–3620, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2020:RPM

[WS20] Leyang Wang and Jianqiang Sun. Regression prediction method that is based on the partial errors-in-variables model. Communications in Statistics: Simulation and Computation, 49(12):3380–3395, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:DJF

[WS21a] Yanhua Wu and Yufeng Shi. Detection of jumps in financial time series. Communications in Statistics: Simulation and Computation, 50(2):313–322, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wywial:2021:VSM

[WS21b] Janusz L. Wywiał and Grzegorz Sitek. On variance of sample matrix eigenvalue. Communications in Statistics: Simulation and Computation, 50(7):1943–1954, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wu:2022:ELM

[WS22] Yanhua Wu and Yufeng Shi. Empirical likelihood for mean difference between two samples with missing data. Communications in Statistics: Simulation and Computation, 51(10): 6227–6234, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2024:DJF

[WS24] Yanhua Wu and Yufeng Shi. Detection of jumps in financial market. Communications in Statistics: Simulation and Computation, 53(1):247–258, 2024. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:PSC

[WSZ23] Yating Wang, Jinxia Su, and Xuejing Zhao. Penalized semiparametric Cox regression model on XGBoost and random survival forests. Communications in Statistics: Simulation and Computation, 52(7):3095–3103, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:LRT

[WTN22] Tianping Wang, Weizhong Tian, and Wei Ning. Likelihood ratio test change-point detection in the skew slash distribution. Communications in Statistics: Simulation and Computation, 51(9):5068–5080, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:OSM

[Wu21] Shu-Fei Wu. One stage multiple comparisons with the control for exponential mean lifetimes based on doubly censored samples under heteroscedasticity. Communications in Statistics: Simulation and Computation, 50(5):1473–1483, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2022:ICP

[Wu22] Yanhong Wu. Isolating changed panels and estimating common change point after sequential detection with FDR control. Communications in Statistics: Simulation and Computation, 51(10):6149–6158, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:NGM

[WW23] Hui Wang and Guizhi Wang. A novel gray model based on normalized beta function and its applications. Communications in Statistics: Simulation and Computation, 52(8):3724– 3740, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:ABP

[WWT22] Jun Wang, Dianpeng Wang, and Yubin Tian. Adaptive Bayesian prediction of reliability based on degradation process. Communications in Statistics: Simulation and Computation, 51(8):4788–4798, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:NSS

[WWTM22] Cong Wang, Tonghui Wang, David Trafimow, and Hunter A. Myüz. Necessary sample sizes for specified closeness and confidence of matched data under the skew normal setting. Communications in Statistics: Simulation and Computation, 51 (5):2083–2094, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wen:2024:SSS

[WWW24] Miin-Jye Wen, Chun-Che Wen, and Wei-Ming Wang. Singlestage sampling procedure for heteroscedasticity in multiple comparisons with a control. Communications in Statistics: Simulation and Computation, 53(4):1719–1728, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2021:ETI

[WWYX21] Xiaohong Wang, Dehui Wang, Kai Yang, and Da Xu. Estimation and testing for the integer-valued threshold autoregressive models based on negative binomial thinning. Communications in Statistics: Simulation and Computation, 50 (6):1622–1644, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:IHD

[WX21] Zhaoliang Wang and Liugen Xue. Inference for high dimensional linear models with error-in-variables. Communications in Statistics: Simulation and Computation, 50(1):164–179, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:JRM

[WY21] Leyang Wang and Fengbin Yu. Jackknife resample method for precision estimation of weighted total least squares. Communications in Statistics: Simulation and Computation, 50 (5):1272–1289, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wen:2023:SKT

[WY23] Qian Wen and Mingao Yuan. Scalable kernel two-sample tests via empirical likelihood and jackknife. Communications in Statistics: Simulation and Computation, 52(12):5975–5990, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wichitchan:2022:NCM

[WYY22] Supawadee Wichitchan, Weixin Yao, and Chun Yu. A new class of multivariate goodness of fit tests for multivariate normal mixtures. Communications in Statistics: Simulation and Computation, 51(11):6635–6648, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wan:2021:DSV

[WZ21a] Qiang Wan and Mei Zhu. Detecting a shift in variance using economically designed VSI control chart with combined attribute-variable inspection. Communications in Statistics: Simulation and Computation, 50(11):3483–3497, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2021:SEF

[WZ21b] Yu Wang and Haixiang Zhang. Some estimation and forecasting procedures in Poisson–Lindley INAR(1) process. Communications in Statistics: Simulation and Computation, 50(1): 49–62, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:VSF

[WZW23] Danlu Wang, Xin Zeng, and Liucang Wu. Variable selection in finite mixture of location and mean regression models using skew-normal distribution. Communications in Statistics: Simulation and Computation, 52(12):5788–5810, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2023:PVA

[WZX23] Shengan Wang, Qing Zhou, and Weilin Xiao. Pricing vulnerable American put options under jump-diffusion processes when corporate liabilities are random. Communications in Statistics: Simulation and Computation, 52(11):5462–5482, 2023. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:FIG

[WZY^{+}21] Xiao Wang, Chenchen Zou, Li Yi, Juan Wang, and Xinmin Li. Fiducial inference for gamma distributions: two-sample problems. Communications in Statistics: Simulation and Computation, 50(3):811–821, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:FSD

[WZY22] Bingcan Wang, Baoxue Zhang, and Haibo Yan. Functional sufficient dimension reduction based on weighted method. Communications in Statistics: Simulation and Computation, 51(11):6902–6923, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2023:SIM

[WZZY23] Leyang Wang, Yingwen Zhao, Chuanyi Zou, and Fengbin Yu. Sterling interpolation method for precision estimation of total least squares. Communications in Statistics: Simulation and Computation, 52(1):142–160, 2023. CODEN CSSCDB. ISSN 0361-0918.

Xue:2021:OUI

[XH21] Yishu Xue and Guanyu Hu. Online updating of information based model selection in the big data setting. Communications in Statistics: Simulation and Computation, 50(11): 3516–3529, 2021. CODEN CSSCDB. ISSN 0361-0918.

Xiao:2022:RTM

[Xia22] Lishun Xiao. Representation theorems of monotonicity generators for BSDEs via L^{p}(p > 1) solutions in general time intervals. Communications in Statistics: Simulation and Computation, 51(7):3993–4005, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xu:2022:VEF

[XS22] Kai Xu and Zhiling Shen. Variance-estimation-free test of significant covariates in high-dimensional regression. Communications in Statistics: Simulation and Computation, 51 (10):6083–6100, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xie:2023:BES

[XS23] Feng-Chang Xie and Ya-Yu Shen. Bayesian estimation for stochastic volatility model with jumps, leverage effect and generalized hyperbolic skew Student’s t-distribution. Communications in Statistics: Simulation and Computation, 52 (8):3420–3437, 2023. CODEN CSSCDB. ISSN 0361-0918.

Xu:2022:LCR

[XSS22] Tonghui Xu, Stephen A. Sedory, and Sarjinder Singh. Lowering the Cramér–Rao lower bounds of variance in randomized response sampling. Communications in Statistics: Simulation and Computation, 51(7):4112–4126, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:RBT

[XW20] Yixun Xing and Wayne A. Woodward. R/S-bootstrapping test for fractional integration. Communications in Statistics: Simulation and Computation, 49(6):1467–1473, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Xing:2022:MIE

[XWL22] Yanchun Xing, Ma Wenqing, and Chunhui Liang. A methodology for improving efficiency estimation based on conditional mix-GEE models in longitudinal studies. Communications in Statistics: Simulation and Computation, 51(1):254–265, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xie:2022:OCF

[XWZL22] Wanli Xie, Wen-Ze Wu, Tao Zhang, and Qi Li. An optimized conformable fractional non-homogeneous gray model and its application. Communications in Statistics: Simulation and Computation, 51(10):5988–6003, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xie:2022:LPR

[XXH22] Lin Xie, Hongmin Xiao, and Yan He. The limit property of a risk model based on entrance processes. Communications in Statistics: Simulation and Computation, 51(3):955–972, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xu:2021:RRM

[XZJ^{+}21] Qifa Xu, Xingxuan Zhuo, Cuixia Jiang, Fang Sun, and Xue Huang. Reverse restricted MIDAS model with application to US interest rate forecasts. Communications in Statistics: Simulation and Computation, 50(2):462–482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Xu:2022:RAD

[XZS22] Da Xu, Shishun Zhao, and Jianguo Sun. Regression analysis of dependent current status data with the accelerated failure time model. Communications in Statistics: Simulation and Computation, 51(10):6188–6196, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yavarizadeh:2021:NMR

[YA21] Bahareh Yavarizadeh and Syed Ejaz Ahmed. The new mixed ridge estimator in a linear mixed model with measurement error under stochastic linear mixed restrictions. Communications in Statistics: Simulation and Computation, 50(8):2185– 2196, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yada:2022:BAD

[Yad22] Shinjo Yada. A Bayesian adaptive design for addressing correlated late-onset outcomes in phase I/II randomized trials of drug combinations in oncology. Communications in Statistics: Simulation and Computation, 51(6):2963–2977, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yang:2023:BAI

[Yan23] Mingan Yang. A Bayesian analysis of the incomplete block crossover design. Communications in Statistics: Simulation and Computation, 52(10):4654–4664, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Yao:2023:URP

[Yao23] Kai Yao. Uncertain renewal process with general rewards. Communications in Statistics: Simulation and Computation, 52(1):56–67, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yashchin:2022:GAM

[Yas22] Emmanuel Yashchin. Gradient analysis of Markov-type control schemes and its applications. Communications in Statistics: Simulation and Computation, 51(4):1814–1836, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yadav:2022:BES

[YBC22] Abhimanyu Singh Yadav, Hassan S. Bakouch, and Christophe Chesneau. Bayesian estimation of the survival characteristics for Hjorth distribution under progressive type-II censoring.

Communications in Statistics: Simulation and Computation, 51(3):882–900, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yildirim:2022:MDB

[YC22] Emre Yildirim and Mehmet Ali Cengiz. Modeling dependency between industry production and energy market via stochastic copula approach. Communications in Statistics: Simulation and Computation, 51(4):2006–2019, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Yang:2023:LEE

[YCD23] Rui Yang, Wangxue Chen, and Yanfei Dong. Log-extended exponential–geometric parameters estimation using simple random sampling and moving extremes ranked set sampling. Communications in Statistics: Simulation and Computation, 52(1):267–277, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yang:2021:EFD

[YDHW21] Lianqiang Yang, Mengzhen Ding, Yongmiao Hong, and Xuejun Wang. Estimating functions and derivatives via adaptive penalized splines. Communications in Statistics: Simulation and Computation, 50(7):2054–2071, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yan:2020:GFI

[YDLX20] Liang Yan, Xiaofang Dong, Xuhua Liu, and Xingzhong Xu. Generalized fiducial inference in the multiple regression model with measurement errors. Communications in Statistics: Simulation and Computation, 49(9):2239–2255, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Ye:2022:BIV

[YGL22] Ren-Dao Ye, Wen-Ting Ge, and Kun Luo. Bootstrap inference on variance component functions in the unbalanced two-way random effects model. Communications in Statistics: Simulation and Computation, 51(9):5373–5386, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yada:2021:AGP

[YH21a] Shinjo Yada and Chikuma Hamada. Application of gamma process to two-agent combinations with delayed toxicity. Communications in Statistics: Simulation and Computation, 50(1):153–163, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yoon:2021:TIQ

[YH21b] Jae Eun Yoon and Sun Young Hwang. On the threshold innovation in quasi-likelihood for conditionally heteroscedastic time series. Communications in Statistics: Simulation and Computation, 50(7):2042–2053, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yazdi:2021:PIM

[YHGA21] Ahmad Ahmadi Yazdi, Ali Zeinal Hamadani, Mohammad Hosein Karimi Gavareshki, and Amirhossein Amiri. Phase II monitoring of multivariate profiles with estimated parameters and optimal phase I subgroups. Communications in Statistics: Simulation and Computation, 50(10):2858–2884, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yao:2023:LLN

[YHWW23] Chi Yao, Yongping He, Rui Wang, and Xuejun Wang. Laws of large numbers and complete convergence for WOD random variables and their applications. Communications in Statistics: Simulation and Computation, 52(1):175–195, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Yoo:2022:SSC

[YL22a] Hanna Yoo and Jae Won Lee. Sample size calculation based on discrete Weibull and zero-inflated discrete Weibull regression models. Communications in Statistics: Simulation and Computation, 51(12):7180–7193, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yu:2022:CTI

[YL22b] Qiqing Yu and Ruiqi Liu. A consistent test of independence and goodness-of-fit in linear regression models. Communications in Statistics: Simulation and Computation, 51(7):3955– 3974, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yue:2023:DSM

[YL23] Jin Yue and Liu Liu. A dynamic sampling for monitoring nonparametric multivariate processes. Communications in Statistics: Simulation and Computation, 52(8):3771–3781, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yang:2022:PPT

[YLX^{+}22] Jingyun Yang, Rong Li, Peirong Xiang, Jingyi Hu, Wenjin Lu, Zhongxin Ni, and Guoliang Cai. On the predictive performance of two Bayesian joint models: a simulation study. Communications in Statistics: Simulation and Computation, 51(11):6388–6397, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yimer:2023:BMA

[YOD^{+}23] Belay Birlie Yimer, Martin Otava, Teshome Degefa, Delenasaw Yewhalaw, and Ziv Shkedy. Bayesian model averaging in longitudinal studies using Bayesian variable selection methods. Communications in Statistics: Simulation and Computation, 52(6):2646–2665, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yoo:2023:SSC

[Yoo23] Hanna Yoo. Sample size for clustered count data based on discrete Weibull regression model. Communications in Statistics: Simulation and Computation, 52(12):5850–5856, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Yerlikaya-Ozkurt:2020:NCM

[YÖT20] Fatma Yerlikaya-Özkurt and Pakize Taylan. New computational methods for classification problems in the existence of outliers based on conic quadratic optimization. Communications in Statistics: Simulation and Computation, 49(3):753– 770, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yang:2023:JPP

[YP23] Xuerui Yang and Jianxin Pan. jmcm: a Python package for analyzing longitudinal data using joint mean-covariance models. Communications in Statistics: Simulation and Computation, 52(11):5446–5461, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yavarizadeh:2022:REL

[YRAB22] Bahareh Yavarizadeh, Abdolrahman Rasekh, Syed Ejaz Ahmed, and Babak Babadi. Ridge estimation in linear mixed measurement error models with stochastic linear mixed restrictions. Communications in Statistics: Simulation and Computation, 51(6):3037–3053, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yu:2023:EGI

[YSG23] Jiao Yu, Yuqi Shan, and Wenhao Gui. Estimation for the generalized inverted exponential distribution under adaptive progressive type II hybrid censoring scheme. Communications in Statistics: Simulation and Computation, 52(9):4134–4155, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yagi:2023:TET

[YSH23] Ayaka Yagi, Takashi Seo, and Zofia Hanusz. Testing equality of two mean vectors with monotone incomplete data. Communications in Statistics: Simulation and Computation, 52 (2):506–522, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yu:2022:NEI

[YTL22] Hengshi Yu, Guangyu Tong, and Fan Li. A note on the estimation and inference with quadratic inference functions for correlated outcomes. Communications in Statistics: Simulation and Computation, 51(11):6525–6536, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yau:2020:NHT

[YV20] Tat S. Yau and Julia Volaufova. A note on hypothesis testing in the envelope model. Communications in Statistics: Simulation and Computation, 49(3):606–628, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yin:2020:LAB

[YW20] Shanshan Yin and Lichun Wang. Linear approximate Bayes estimator for variance components in random effects model*. Communications in Statistics: Simulation and Computation, 49(5):1352–1366, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yin:2024:ETS

[YW24] Li Yin and Xiaoqin Wang. Estimating and testing sequential causal effects based on alternative G-formula: an observational study of the influence of early diagnosis on survival of cardia cancer. Communications in Statistics: Simulation and Computation, 53(4):1917–1931, 2024. CODEN CSSCDB. ISSN 0361-0918.

Ye:2023:BAH

[YWL^{+}23] Rendao Ye, Zhongchi Wang, Kun Luo, Ya Lin, Na An, and Weixiao Du. Bootstrap approaches for homogeneous test of location parameters under skew-normal settings. Communications in Statistics: Simulation and Computation, 52(10): 4739–4755, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yan:2024:SNP

[YWL24] Han Yan, Dehui Wang, and Cong Li. A study for the NM-BAR(1) processes. Communications in Statistics: Simulation and Computation, 53(3):1308–1329, 2024. CODEN CSSCDB. ISSN 0361-0918.

Yin:2020:NEM

[YWLD20] Junhui Yin, Liucang Wu, Hanchi Lu, and Lin Dai. New estimation in mixture of experts models using the Pearson type VII distribution. Communications in Statistics: Simulation and Computation, 49(2):472–483, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ye:2022:NEB

[YWT22] Gen Ye, Songjian Wang, and Niansheng Tang. Nonparametric empirical Bayesian method for noncontractual setting of customer-base analysis. Communications in Statistics: Simulation and Computation, 51(3):1234–1254, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yang:2022:REO

[YXYX22] Guangren Yang, Sijia Xiang, Weixin Yao, and Lin Xu. Robust estimation and outlier detection for varying-coefficient models via penalized regression. Communications in Statistics: Simulation and Computation, 51(10):5845–5856, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yali:2023:AED

[YYZ23] Fan Yali, Xiang Yayun, and Guo Zijun. Adaptive efficient and double-robust regression based on generalized empirical likelihood. Communications in Statistics: Simulation and Computation, 52(7):3079–3094, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yuan:2021:TPP

[YZ21] Mingao Yuan and Yue Zhang. Test for the parametric part in partial functional linear regression based on B-spline. Communications in Statistics: Simulation and Computation, 50 (1):1–15, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yan:2023:REM

[YZ23] Feifei Yan and Lin Zhu. Regression estimation of the marginal models with general relative risk form for multivariate failure time data. Communications in Statistics: Simulation and Computation, 52(8):3958–3972, 2023. CODEN CSSCDB. ISSN 0361-0918.

Yang:2022:RRS

[YZF22] Shengping Yang, Kun Zhang, and Zhide Fang. Robust RNAseq data analysis using an integrated method of ROC curve and Kolmogorov–Smirnov test. Communications in Statistics: Simulation and Computation, 51(12):7444–7457, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:SAC

[ZA21] Qing Zhang and Hongshik Ahn. Subgroup analysis of censored data on cancer treatment. Communications in Statistics: Simulation and Computation, 50(12):4041–4058, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zafar:2023:AWL

[ZA23] Zahra Zafar and Muhammad Aslam. An adaptive weighted least squares ratio approach for estimation of heteroscedastic linear regression model in the presence of outliers. Communications in Statistics: Simulation and Computation, 52(6): 2365–2375, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2020:BSI

[ZB20] Ningning Zhao and Zhidong Bai. Bayesian statistical inference based on rounded data. Communications in Statistics: Simulation and Computation, 49(1):135–146, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhuang:2022:CLT

[ZB22] Yan Zhuang and Sudeep R. Bapat. On comparing locations of two-parameter exponential distributions using sequential sampling with applications in cancer research. Communications in Statistics: Simulation and Computation, 51(10): 6114–6135, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2024:ELI

[ZB24] Xiaojun Zhu and Narayanaswamy Balakrishnan. Exact likelihood inference for Laplace distribution based on generalized hybrid censored samples. Communications in Statistics: Simulation and Computation, 53(1):259–272, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:AVP

[ZBA23] Yuchi Zhang, Ryan P. Browne, and Jeffrey L. Andrews. Assessing the variability of posterior probabilities in Gaussian model-based clustering. Communications in Statistics: Simulation and Computation, 52(5):1937–1947, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zandi:2023:USS

[ZBB23] Zahra Zandi, Hossein Bevrani, and Reza Arabi Belaghi. Using shrinkage strategies to estimate fixed effects in zero-inflated negative binomial mixed model. Communications in Statistics: Simulation and Computation, 52(7):3201–3222, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2020:EPL

[ZBZS20] Xiaojun Zhu, Narayanaswamy Balakrishnan, Yiliang Zhou, and Hon-Yiu So. Exact predictive likelihood inference for Laplace distribution based on a time-constrained experiment. Communications in Statistics: Simulation and Computation, 49(3):647–668, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2022:MLM

[ZC22] Qingle Zheng and Yuezhang Che. Multicategory large margin classification with unequal costs. Communications in Statistics: Simulation and Computation, 51(4):1801–1813, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zou:2023:MMR

[ZC23] Feng Zhen Zou and Xu Chen. A Markov-modulated risk model with transaction costs and threshold dividend strategy. Communications in Statistics: Simulation and Computation, 52(4):1577–1590, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zeng:2024:CIM

[ZCG24] Hui Zeng, Vernon M. Chinchilli, and Nasrollah Ghahraman. Causal inference with a mediated proportional hazards regression model. Communications in Statistics: Simulation and Computation, 53(1):203–218, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zhong:2023:NCV

[ZCH23] Yi Zhong, Prabhakar Chalise, and Jianghua He. Nested crossvalidation with ensemble feature selection and classification model for high-dimensional biological data. Communications in Statistics: Simulation and Computation, 52(1):110–125, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:KDE

[ZCW23] Jun Zhang, Aixian Chen, and Zhenghong Wei. Kernel density estimation for multiplicative distortion measurement regression models. Communications in Statistics: Simulation and Computation, 52(5):1733–1752, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:HTO

[ZD22] Haodong Zhang and Jieli Ding. Hypothesis testing in outcome-dependent sampling design under generalized linear models. Communications in Statistics: Simulation and Computation, 51(4):1721–1745, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:PIA

[ZFG22] Jun Zhang, Sanying Feng, and Yujie Gai. Partial index additive models with additive distortion measurement errors. Communications in Statistics: Simulation and Computation, 51(5):2191–2216, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:TER

[ZfSRJ22] Yong-He Zhao, Hui feng Shi, Xiao-Chen Ren, and Bing-Jie Jiao. A toll evasion recognition method based on Gaussian mixture clustering. Communications in Statistics: Simulation and Computation, 51(7):3901–3911, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:LRM

[ZG21] Jun Zhang and Yujie Gai. Linear regression models with general distortion measurement errors. Communications in Statistics: Simulation and Computation, 50(11):3383–3396, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:CCB

[ZG22a] Jun Zhang and Yujie Gai. Correlation coefficient-based measure for checking symmetry or asymmetry of a continuous variable with additive distortion. Communications in Statistics: Simulation and Computation, 51(5):2158–2190, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:NRM

[ZG22b] Jun Zhang and Yujie Gai. Nonlinear regression models with profile nonlinear least squares estimation. Communications in Statistics: Simulation and Computation, 51(5):2140–2157, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:EFM

[ZGC21] Caiya Zhang, Xiaolu Gu, and Yingyu Chen. Estimation for frailty measurement error Cox models based on profile likelihood and Bayes methods. Communications in Statistics: Simulation and Computation, 50(4):1025–1038, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2021:DMH

[ZGL21] Jun Zhang, Yujie Gai, and Bingqing Lin. Detection of marginal heteroscedasticity for partial linear single-index models. Communications in Statistics: Simulation and Computation, 50(3):724–743, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:TSA

[ZGL22] Jun Zhang, Yujie Gai, and Feng Li. Testing symmetry for additive distortion measurement errors data. Communications in Statistics: Simulation and Computation, 51(3):1046–1065, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:VIV

[Zha22a] Jiamin Zhang. Variational inference for varying-coefficient model. Communications in Statistics: Simulation and Computation, 51(2):670–685, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:IRT

[Zha22b] Jun Zhao. Iteratively reweighted total least squares for PEIV model. Communications in Statistics: Simulation and Computation, 51(7):4026–4038, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:NMD

[Zha23] Jun Zhang. Nonlinear multiplicative distortion regression models with second-order estimation. Communications in Statistics: Simulation and Computation, 52(12):5894–5924, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2024:NMD

[Zha24] Jun Zhang. Nonparametric multiplicative distortion measurement errors models with bias reduction. Communications in Statistics: Simulation and Computation, 53(4):1952–1972, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2021:IRL

[Zhe21] Songfeng Zheng. Iteratively reweighted least square for asymmetric L_{2}-loss support vector regression. Communications in Statistics: Simulation and Computation, 50(7):2151–2167, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2024:SLS

[Zhe24] Songfeng Zheng. Speeding up L_{2}-loss support vector regression by random Fourier features. Communications in Statistics: Simulation and Computation, 53(2):933–951, 2024. CO-DEN CSSCDB. ISSN 0361-0918.

Zhong:2022:CPC

[ZHJ22] Jianlan Zhong, Xulong Hu, and Tian Jiang. A change point control chart for monitoring the production lead time based on sum of squared ranks. Communications in Statistics: Simulation and Computation, 51(11):6283–6293, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:GPE

[ZHJ23] Jing Zhang, Zhensheng Huang, and Zhiqiang Jiang. Groupwise partial envelope model: efficient estimation in multivariate linear regression. Communications in Statistics: Simulation and Computation, 52(7):2924–2940, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:FPM

[ZHW21] Tao Zhang, Haifeng Huo, and Yanling Wan. Functional polynomial multiple-index model. Communications in Statistics: Simulation and Computation, 50(12):3958–3971, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Zeybek:2020:EGN

[ZK20a] Melis Zeybek and Onur Köksoy. The effects of gamma noise on quality improvement. Communications in Statistics: Simulation and Computation, 49(7):1783–1797, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zmyslony:2020:SHT

[ZK20b] Roman Zmyślony and Arkadiusz Kozioł. Simultaneous hypotheses testing in models with blocked compound symmetric covariance structure. Communications in Statistics: Simulation and Computation, 49(3):817–823, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Zeinal:2023:GNT

[ZK23] Amir Zeinal and Mohammad Reza Azmoun Zavie Kivi. The generalized new two-type parameter estimator in linear regression model. Communications in Statistics: Simulation and Computation, 52(1):98–109, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zulawinski:2024:ADM

[ZKW24] Wojciech Zuławiński, Piotr Kruczek, and Agnieszka Wyłomańska. Alternative dependency measures-based approach for estimation of the α-stable periodic autoregressive model. Communications in Statistics: Simulation and Computation, 53(3): 1188–1215, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2020:NMR

[ZL20] Lixia Zhu and Kung-Jong Lui. Notes on misspecifying the random effects distribution regarding analysis under the AB/ BA crossover trial in dichotomous data — a Monte Carlo evaluation. Communications in Statistics: Simulation and Computation, 49(2):419–435, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:AMS

[ZL22a] Yuyang Zhang and Bo Lu. Accounting for matching structure in post-matching analysis of observational studies. Communications in Statistics: Simulation and Computation, 51(6): 3081–3099, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2022:RFC

[ZL22b] Shiqiu Zheng and Shoumei Li. A representation for filtrationconsistent nonlinear expectations and its application. Communications in Statistics: Simulation and Computation, 51 (12):7133–7150, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhan:2023:CCS

[ZL23a] Mengmeng Zhan and Liping Liu. CUSUM control schemes for monitoring Wiener processes. Communications in Statistics: Simulation and Computation, 52(1):161–174, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:ECC

[ZL23b] Jun Zhang and Bingqing Lin. Estimation of correlation coefficient with general distortion measurement errors. Communications in Statistics: Simulation and Computation, 52(9): 4491–4521, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2023:ACP

[ZL23c] Hong Zhou and Huajiang Li. Applications of covering principle to clinical trials with multiple objectives. Communications in Statistics: Simulation and Computation, 52(6):2752–2764, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhan:2023:CSM

[ZlBVGW^{+}23] Z. Zhan, S. la Bastide-Van Gemert, M. Wiersum, K. R. Heineman, M. Hadders-Algra, and E. R. van den Heuvel. A comparison of statistical methods for age-specific reference values of discrete scales. Communications in Statistics: Simulation and Computation, 52(10):5024–5041, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:IPL

[ZLF23] Yuanqing Zhang, Hong Li, and Yaqin Feng. Inference for partially linear additive higher-order spatial autoregressive model with spatial autoregressive error and unknown heteroskedasticity. Communications in Statistics: Simulation and Computation, 52(3):898–924, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2023:SSC

[ZLG23] Wenjun Zheng, Dejian Lai, and K. Lance Gould. A simulation study of a class of nonparametric test statistics: a close look of empirical distribution function-based tests. Communications in Statistics: Simulation and Computation, 52(3):1132–1148, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2020:BOA

[ZLSC20] Yicheng Zhou, Zhenzhou Lu, Yan Shi, and Kai Cheng. Bayesian optimum accelerated life test plans based on quantile regression. Communications in Statistics: Simulation and Computation, 49(9):2402–2418, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2023:SEN

[ZLTP23] Xianbo Zhou, Lianhua Luo, Ying Tao, and Zhewen Pan. Semiparametric estimation of Nelson–Olson simultaneous Tobit model. Communications in Statistics: Simulation and Computation, 52(8):3863–3876, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:BMM

[ZLXT23] Jiwei Zhang, Jing Lu, Xin Xu, and Jian Tao. Bayesian multilevel multidimensional item response modeling approach for multiple latent variables in a hierarchical structure. Communications in Statistics: Simulation and Computation, 52(7): 2822–2842, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2023:BMC

[ZMNL23] Jiayan Zhu, Li Ma, Mengying Ni, and Zhengbang Li. A bootstrap method to calculate the p-value of Fisher’s combination for a large number of weakly dependent p-values. Communications in Statistics: Simulation and Computation, 52(9): 4210–4217, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:NCP

[ZNKN22] Kaimeng Zhang, Chi Tim Ng, Yong Man Kwon, and Myung Hwan Na. New concepts of principal component analysis based on maximum separation of clusters. Communications in Statistics: Simulation and Computation, 51(5):2429– 2439, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:EED

[ZNLW20] Jun Zhang, Cuizhen Niu, Tao Lu, and Zhenghong Wei. Estimation of the error distribution function for partial linear single-index models. Communications in Statistics: Simulation and Computation, 49(1):29–44, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:SSB

[ZPW21] Xuemao Zhang, Sudhir Paul, and You-Gan Wang. Small sample bias correction or bias reduction? Communications in Statistics: Simulation and Computation, 50(4):1165–1177, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2022:TGB

[ZQY22] He Zhou, Wei Qian, and Yi Yang. Tweedie gradient boosting for extremely unbalanced zero-inflated data. Communications in Statistics: Simulation and Computation, 51(9):5507–5529, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:WCW

[ZQYW22] Lu Zhang, Jibing Qi, Hairong Yang, and Xuejun Wang. Weak convergence for weighted sums of negatively associated random variables and its application in nonparametric regression models. Communications in Statistics: Simulation and Computation, 51(10):5876–5896, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:BCM

[ZRC23] Jingying Zhang, Ronald L. Rardin, and Justin R. Chimka. Budget constrained model selection for multiple linear regression. Communications in Statistics: Simulation and Computation, 52(11):5537–5549, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:PEG

[ZS23] Xuekang Zhang and Huisheng Shu. Parameter estimation for generalized Ait-Sahalia-type interest rate model. Communications in Statistics: Simulation and Computation, 52(4): 1630–1638, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zaki:2023:PCS

[ZSM23] Abeer A. Zaki, Nesma A. Saleh, and Mahmoud A. Mahmoud. Performance comparison of some centrality measures used in detecting anomalies in directed social networks. Communications in Statistics: Simulation and Computation, 52(7):3122– 3136, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:SMI

[ZT20] Shuxia Zhang and Boping Tian. Semiparametric method for identifying multiple change-points in financial market. Communications in Statistics: Simulation and Computation, 49 (6):1429–1444, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2023:URV

[ZWC^{+}23] Xu Zhao, Shaojie Wei, Weihu Cheng, Pengyue Zhang, Yang Zhang, and Qi Xu. Upper record values from the generalized Pareto distribution and associated statistical inference. Communications in Statistics: Simulation and Computation, 52 (2):369–391, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2024:OPB

[ZWCY24] Peixin Zhao, Hao Wu, Suli Cheng, and Yiping Yang. Orthogonal projection based variable selection for semiparametric spatial autoregressive models. Communications in Statistics: Simulation and Computation, 53(1):178–189, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:CRM

[ZWMY23] Weijia Zhang, Xikui Wang, Saman Muthukumarana, and Po Yang. A continual reassessment method without undue risk of toxicity. Communications in Statistics: Simulation and Computation, 52(4):1179–1191, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2023:DPR

[ZWTY23] Peixin Zhao, Junqi Wang, Xinrong Tang, and Weiming Yang. Double penalized regularization estimation for partially linear instrumental variable models with ultrahigh dimensional instrumental variables. Communications in Statistics: Simulation and Computation, 52(10):4636–4653, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2023:BEA

[ZWW23] Bo Zhao, Shuying Wang, and Chunjie Wang. Bayesian empirical analysis of the proportional hazards model for rightcensored failure time data. Communications in Statistics: Simulation and Computation, 52(10):5042–5051, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2020:SRP

[ZWY20] Jie Zhang, Dehui Wang, and Kai Yang. A study of RCINAR(1) process with generalized negative binomial marginals. Communications in Statistics: Simulation and Computation, 49(6):1487–1510, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:NDC

[ZWY21] Weijia Zhang, Xikui Wang, and Po Yang. A new design of the continual reassessment method. Communications in Statistics: Simulation and Computation, 50(7):2015–2024, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:IDC

[ZX22a] Guimei Zhao and Xingzhong Xu. The information domain confidence intervals in univariate linear calibration. Communications in Statistics: Simulation and Computation, 51(10): 5620–5630, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2022:RAD

[ZX22b] Huiling Zheng and Houbao Xu. Reliability analysis for degradation and shock process based on truncated normal distribution. Communications in Statistics: Simulation and Computation, 51(8):4241–4256, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:EPL

[ZXF22] Yang Zhao, Liugen Xue, and Sanying Feng. Estimation for a partially linear single-index varying-coefficient model. Communications in Statistics: Simulation and Computation, 51 (4):1685–1703, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:ALC

[ZXW23] Jun Zhang, Zhuoer Xu, and Zhenghong Wei. Absolute logarithmic calibration for correlation coefficient with multiplicative distortion. Communications in Statistics: Simulation and Computation, 52(2):482–505, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:SIV

[ZXZL22] Yang Zhao, Liugen Xue, Jinghua Zhang, and Juanfang Liu. Single-index varying-coefficient models with missing covariates at random. Communications in Statistics: Simulation and Computation, 51(12):7351–7365, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2024:ECC

[ZYF24] Jun Zhang, Baojun Yang, and Zhenghui Feng. Estimation of correlation coefficient under a linear multiplicative distortion measurement errors model. Communications in Statistics: Simulation and Computation, 53(1):62–93, 2024. CODEN CSSCDB. ISSN 0361-0918.

Zhuang:2022:RCE

[ZYSS22] Xinchen Zhuang, Tianxiang Yu, Zhongchao Sun, and Kunling Song. Reliability and capacity evaluation of multiperformance multi-state weighted K^{−}-out-of-n systems. Communications in Statistics: Simulation and Computation, 51(10):6026–6042, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zeng:2021:RBM

[ZZ21] Guoping Zeng and Emily Zeng. On the relationship between multicollinearity and separation in logistic regression. Communications in Statistics: Simulation and Computation, 50 (7):1989–1997, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:BBS

[ZZ22a] Jian-Qiang Zhao and Yan-Yong Zhao. Bootstrap bandwidth selection in time-varying coefficient models with jumps. Communications in Statistics: Simulation and Computation, 51 (6):3124–3137, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2022:GLP

[ZZ22b] Huili Zhou and Jun Zhang. General least product relative error estimation for multiplicative regression models with or without multiplicative distortion measurement errors. Communications in Statistics: Simulation and Computation, 51 (11):6352–6370, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:RBK

[ZZFG20] Jun Zhang, Jing Zhang, Zhenghui Feng, and Xu Guo. Residuals based Kolmogorov–Smirnov and Cramér–von Mises tests for varying coefficient models. Communications in Statistics: Simulation and Computation, 49(7):1798–1814, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2023:SSA

[ZZL23] Jiwei Zhang, Zhaoyuan Zhang, and Jing Lu. Slice sampling algorithm for estimating the item response theory model with ordinal response data. Communications in Statistics: Simulation and Computation, 52(10):4909–4932, 2023. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2020:NOE

[ZZWH20] Peixin Zhao, Xiaoshuang Zhou, Xiuli Wang, and Xingshou Huang. A new orthogonality empirical likelihood for varying coefficient partially linear instrumental variable models with longitudinal data. Communications in Statistics: Simulation and Computation, 49(12):3328–3344, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2023:BCQ

[ZZZ23] Feipeng Zhang, Shenglin Zheng, and Xiaoying Zhou. Bentcable quantile regression model. Communications in Statistics: Simulation and Computation, 52(5):2000–2011, 2023. CODEN CSSCDB. ISSN 0361-0918.